### v1

# Fully Dynamic Euclidean $k$-Means

Abstract
We consider the fundamental Euclidean $k$-means clustering problem in a dynamic setting, where the input $X\subseteq\mathbb{R}^{d}$ evolves over time via a sequence of point insertions/deletions. We have to explicitly maintain a solution (a set of $k$ centers) $S\subseteq\mathbb{R}^{d}$ throughout these updates, while minimizing the approximation ratio, the update time (time taken to handle a point insertion/deletion) and the recourse (number of changes made to the solution $S$) of the algorithm.
We present a dynamic algorithm for this problem with $\text{poly}(1/\epsilon)$-approximation ratio, $\tilde{O}(k^{\epsilon})$ update time and $\tilde{O}(1)$ recourse. In the general regime, where the dimension $d$ cannot be assumed to be a fixed constant, our algorithm has almost optimal guarantees across all these three parameters. Indeed, improving our update time or approximation ratio would imply beating the state-of-the-art static algorithm for this problem (which is widely believed to be the best possible), and the recourse of any dynamic algorithm must be $\Omega(1)$.
We obtain our result by building on top of the recent work of [Bhattacharya, Costa, Farokhnejad; STOC’25], which gave a near-optimal dynamic algorithm for $k$-means in general metric spaces (as opposed to in the Euclidean setting). Along the way, we design several novel geometric data structures that are of independent interest. Specifically, one of our main contributions is designing the first consistent hashing scheme [Czumaj, Jiang, Krauthgamer, Veselý, Yang; FOCS’22] that achieves $\text{poly}(d)$ running time per point evaluation with competitive parameters.
Part I Extended Abstract
1 Introduction
In Euclidean $k$-means clustering, the input consists of a set of $n$ points $X\subseteq\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space, where each point $x\in X$ has a nonnegative weight $w(x)\geq 0$, and a positive integer $k\in[n]$. We have to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$, so as to minimize the objective $\sum_{x\in X}w(x)\cdot\left(\mathrm{dist}(x,S)\right)^{2}$, where $\mathrm{dist}(x,S):=\min_{y\in S}\{\mathrm{dist}(x,y)\}$ and $d(x,y)$ denotes the Euclidean distance between the points $x$ and $y$. Our focus in this paper is on a general setting, where the parameters $k,d$ can not be treated as absolute constants. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84].
The study of Euclidean $k$-means dates back to 1950s, starting with the celebrated Lloyd’s algorithm [Llo82]. It is regarded as one of the most fundamental clustering problems, and serves as a basic primitive in unsupervised learning. The problem is APX-hard in Euclidean spaces of dimension $d=\Omega(\log n)$ [ACKS15]. In contrast, for every sufficiently small $\epsilon>0$, it admits a $\text{poly}(1/\epsilon)$-approximation algorithm that runs in $\tilde{O}(n^{1+\epsilon})$ time [lTS24, JJLL25]. [^1] We use $\tilde{O}(\cdot)$ notation to hide polylogarithmic factors in $n$ and the aspect ratio, and polynomial factors in $1/\epsilon$. Using standard coreset based sparsification [DSS24], the runtime can be further reduced to $\tilde{O}(n+k^{1+\epsilon})$. This tradeoff between $\mathrm{poly}(1/\epsilon)$-approximation and $\tilde{O}(n+k^{1+\epsilon})$ runtime is believed to be essentially the best possible.222 When $n=k-1$, $k$-means reduces to closest pair (CP) problem on $n$ points, whose current tradeoff is $\mathrm{poly}(1/\epsilon)$-approximation and $n^{1+\epsilon}$ time via locality sensitive hashing [AI06, AR15]. Whether this tradeoff for CP can be improved, i.e., achieving $\tilde{O}(n)$ time for $O(1)$-approximation, is a long-standing open question and is believed that a lower bound exists [Rub18].
Dynamic $k$-Means Clustering. In this paper, we consider the Euclidean $k$-means problem in a natural dynamic setting, where the input changes by a sequence of updates: Each update inserts/deletes a point in the current input $X\subseteq\mathbb{R}^{d}$. [^3] In case of an insertion, the update also specifies the weight of the concerned point. We wish to design a dynamic algorithm that can maintain an approximately optimal solution (i.e., a set of $k$ centers) $S\subseteq\mathbb{R}^{d}$ throughout these updates. The performance of such an algorithm is measured in terms of three parameters:
  • (i) Approximation ratio, which indicates the quality of the maintained solution.
  • (ii) Update time, which denotes the time taken by the algorithm to process an update, and measures the efficiency of the algorithm in handling dynamic inputs.
  • (iii) Recourse, which denotes the number of changes (i.e., point insertions/deletions) in the maintained solution $S$ per update, and indicates how stable the solution is over time.
A naive way to solve this problem would be to run the best possible static algorithm from scratch after each update [lTS24, DSS24, JJLL25]; leading to $\text{poly}(1/\epsilon)$-approximation, $\tilde{O}(n+k^{1+\epsilon})$ update time and $O(k)$ recourse. But can we do substantially better? In particular, can we design a dynamic algorithm for this problem with (almost) optimal performance guarantees? We answer this question in the affirmative. Our main result is summarized in the theorem below.
Theorem 1.1.
For every sufficiently small $\epsilon>0$, there is a randomized dynamic algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio, $\tilde{O}(k^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse. If the length of the update sequence is polynomially bounded by $n$ (the maximum number of input points at any time), then the approximation, update time and recourse guarantees of the algorithm hold with probability $1-1/\text{poly}(n)$ across the entire sequence of updates.
Remarks.
At the outset, we highlight several important aspects of our result.
  1. 1.
    The algorithm works against an oblivious adversary if we use the Johnson-Lindenstrauss transform [JL84] to reduce the dimension $d$ to $O(\log n)$. Otherwise, if we are not using the Johnson-Lindenstrauss transform, then our dynamic algorithm works against an adaptive adversary, at the cost of incurring an exponential in $d$ factor in the update time.
  2. 2.
    Our result seamlessly extends to the related $k$-median problem, where the goal is to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$ so as to minimize the objective $\sum_{x\in X}w(x)\cdot\mathrm{dist}(x,S)$.
  3. 3.
    In recent years, significant attention has been devoted towards designing dynamic algorithms for $k$-means and $k$-median objectives [LV17, CHP+19, HK20, FLNS21, BCLP23, BCG+24, BCF25a]. All these papers, however, focus on solving the problem in general metric spaces, where it is known that any dynamic algorithm has to incur an update time of $\Omega(k)$ [CS04]. In contrast, ours is the first work that designs an algorithm for this problem specifically tailored to the Euclidean setting, and accordingly bypasses the $\Omega(k)$ update time barrier.
  4. 4.
    Using our dynamic algorithm as a black-box subroutine, we can obtain a static algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(n+k^{1+\epsilon})$ runtime: Indeed, we can first reduce the number of input points to $\tilde{O}(k)$ by computing a coreset [DSS24] in $\tilde{O}(n)$ time, and then insert each of these (weighted) input points into our dynamic algorithm. Our algorithm would handle these $\tilde{O}(k)$ updates in $\tilde{O}(k^{1+\epsilon})$ total time, and in the end it will return a $\text{poly}(1/\epsilon)$-approximate solution to the original set of input points. This matches the current best tradeoff between approximation ratio and runtime in the static setting [lTS24, DSS24, JJLL25], which is also believed to be (almost) optimal (see Footnote 2). Furthermore, it is easy to show that any nontrivial dynamic $k$-means algorithm must necessarily incur $\Omega(1)$ recourse, whereas the recourse guarantee achieved by our algorithm is only $\tilde{O}(1)$. In summary, the approximation ratio, update time and recourse bounds of our algorithm are almost optimal.
Our Techniques. At a high level, our approach towards proving Theorem 1.1 consists of two stages. (i) Take an existing dynamic $k$-means algorithm on general metrics as a starting point. (ii) Speed up the update time of the concerned algorithm in Euclidean spaces, without significantly degrading its approximation ratio and recourse, by designing appropriate geometric data structures. At present, there are only two such dynamic algorithms with close to optimal guarantees (in terms of approximation ratio, update time and recourse) on general metrics [BCG+24, BCF25a].
Among them, the dynamic algorithm developed by [BCG+24] incurs $\Theta(k^{\epsilon})$ recourse. Furthermore, it seems that implementing this framework in Euclidean spaces would inherently incur an exponential in $1/\epsilon$ factor loss in the approximation ratio (see Appendix A for a more detailed discussion). Since we are aiming for $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(1)$ recourse, we cannot use the algorithm of [BCG+24] as the starting point of our approach.
Instead, we build on top of the dynamic algorithm of [BCF25a], which achieves $O(1)$ approximation ratio, $\tilde{O}(k)$ update time and $\tilde{O}(1)$ recourse on general metrics. To achieve our $k^{\epsilon}$ update time, we must break the $\tilde{O}(k)$ update time barrier of general metrics, and improve the framework of [BCF25a] to utilize the Euclidean structure. In Section 3, we review the basic framework of this algorithm, and outline the major challenges we need to overcome to implement a (modified) version of it in the geometric setting. This particularly requires us to implement several geometric data structures that work against adaptive adversary. Section 4 presents an overview of our geometric data structures. One of our key contributions here is the first consistent hashing scheme [CJK+22] that achieves $\mathrm{poly}(d)$ running time per point evaluation with competitive parameters. This new consistent hashing is crucial to ensure that our geometric data structures work against an adaptive adversary. Beyond clustering, our data structures generally address more fundamental tasks including dynamic range queries and maintenance of approximately the nearest assignment, which may be of independent interest. Finally, Section 5 outlines how we modify the framework of [BCF25a] and implement the resulting algorithm using our geometric data structures, leading us to Theorem 1.1. Parts II and III of the paper contains the full version of our algorithm and analysis.
Other Related Work. The Euclidean $k$-means problem has also been extensively studied in the special setting of low-dimensional spaces, where the dimension $d$ is assumed to be a constant. This problem in NP-hard even when $d=2$ [MNV12]. In contrast, it is known to admit a static $(1+\epsilon)$-approximation algorithm that runs in $\tilde{O}\left(2^{(1/\epsilon)^{O(d^{2})}}\cdot n\right)$ time [CFS21].
Finally, we note that an influential line of recent work has focused on two other canonical clustering problems in the dynamic setting: facility location [GHL18, CHP+19, BLP22, BGJ+24] and $k$-center [CGS18, GHL+21, BEF+23, CFG+24, LHG+24, BCF+25b, FS25].
2 Notations and Preliminaries
Parameters and Assumptions.
We fix the following parameters in the paper: $\epsilon$ is a sufficiently small constant, $n$ is the number of points in the data-set, $d$ is the dimension of the space, and $\Delta$ is the aspect ratio, i.e., the max distance divided by the smallest distance between distinct points. As a typical setup, we assume $\Delta\leq\mathrm{poly}(n)$ is an integer. Without loss of generality, the input points that we consider are from the discrete grid $[\Delta]^{d}:=\{(x_{1},\ldots,x_{d})\mid 1\leq x_{i}\leq\Delta\ \forall 1\leq i\leq d\}$, by a standard rounding and rescaling. Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84] (see the first paragraph in Section 5.1 for a detailed discussion).
Notations.
We now introduce some key notations that will be used in subsequent sections. For integer $n\geq 1$, denote $[n]:=\{1,\ldots,n\}$. Consider a generic function $f:U\to V$; for $T\subseteq U$ we write $f(T):=\{f(u):u\in T\}$, and for $v\in V$ we write $f^{-1}(v):=\{u\in U\mid f(u)=v\}$. For every two sets $A$ and $B$, we use $A-B$ instead of $A\setminus B$, and whenever $B=\{b\}$ is a singleton, we write $A-b$ instead of $A-\{b\}$ for simplicity. For $x\in[\Delta]^{d},r\geq 0$, let $\textsc{Ball}(x,r):=\{y\in[\Delta]^{d}:\mathrm{dist}(x,y)\leq r\}$ be the metric ball. For $T\subseteq[\Delta]^{d}$, let $\mathrm{diam}(T):=\max_{x,y\in T}\mathrm{dist}(x,y)$ be the diameter of $T$. For every weighted dataset $X\subseteq[\Delta]^{d}$ and center set $S\subseteq[\Delta]^{d}$, the cost of $S$ on $X$ is denoted by $\textsc{cost}(X,S)$ and is defined as $\textsc{cost}(X,S):=\sum_{x\in X}w(x)\cdot\mathrm{dist}(x,S)^{2}$. Accordingly, we define $\textsc{AverageCost}(X,S):=\textsc{cost}(X,S)/w(X)$. When the size of $S$ is one, i.e. $S=\{c\}$, we abbreviate $\textsc{cost}(X,\{c\})$ by $\textsc{cost}(X,c)$ and $\textsc{AverageCost}(X,\{c\})$ by $\textsc{AverageCost}(X,c)$. We denote by $\mathrm{OPT}_{k}(X)$ the cost of the optimum solution of size $k$. For every $W\subseteq[\Delta]^{d}$, we also denote by $\mathrm{OPT}_{k}^{W}(X)$ the cost of the optimum solution of size $k$ which is restricted to open centers only in $W$, i.e. $\mathrm{OPT}_{k}^{W}(X):=\min_{S\subseteq W,|S|=k}\textsc{cost}(X,S)$.
3 Overview of the Dynamic Algorithm of [BCF25a]
In this section, we explain the algorithm of [BCF25a] and its main ingredients that works for general metric spaces. Although their algorithm is primarily designed for $k$-median, it can be seamlessly extended to work for $k$-means as well. See Part III for a thorough analysis.
3.1 Robust Centers
Definition 3.1 (Robust Center, Definition 3.2 in the arxiv version of [BCF25a]).
Assume $X\subseteq[\Delta]^{d}$. Let $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points in $[\Delta]^{d}$, and let $B_{i}=\textsc{Ball}_{X}(x_{i},10^{i})$ for each $i\in[0,t]$. We refer to $(x_{0},x_{1},\ldots,x_{t})$ as a $t$-robust sequence w.r.t. $X$ iff for every $i\in[1,t]$:
$x_{i-1}=\begin{cases}x_{i}\quad&\text{if}\ \textsc{AverageCost}(B_{i},x_{i})\geq 10^{i}/5;\\ y_{i}\quad&\text{otherwise, where }y_{i}=\arg\min\limits_{y\in B_{i}+x_{i}}\textsc{cost}(B_{i},y).\end{cases}$
We say that a point $x\in[\Delta]^{d}$ is $t$-robust w.r.t. $X$ iff there exists a $t$-robust sequence $(x_{0},x_{1},\ldots,x_{t})$ w.r.t. $X$ such that $x_{0}=x$.
Definition 3.2 (Robust Solution, Definition 3.5 in the arxiv version of [BCF25a]).
We call a center set $S\subseteq[\Delta]^{d}$, robust w.r.t. $X$ iff each center $u\in S$ satisfies the following
$u\ \text{is}\ t\text{-robust w.r.t.\ $X$, where}\ t\ \text{is the smallest integer satisfying}\ 10^{t}\geq\mathrm{dist}(u,S-u)/200.$ (1)
3.2 Description of The Algorithm
Here, we explain the version of the algorithm that has an exponential running time. Later in Section 5, we provide our implementation of their algorithm in high-dimensional Euclidean spaces that has a small update time.
The algorithm works in epochs; each epoch lasts for some consecutive updates in $X$. Let $S\subseteq X$ denote the maintained solution (set of $k$ centers). We satisfy the following invariant.
Invariant A.
At the start of an epoch, the set $S$ is robust and $\textsc{cost}(S,X)\leq O(1)\cdot\mathrm{OPT}_{k}(X)$.
We now describe how the dynamic algorithm works in a given epoch, in four steps.
Step 1: Determining the length of the epoch. At the start of an epoch, the algorithm computes the maximum $\ell^{\star}\geq 0$ such that $\mathrm{OPT}_{k-\ell^{\star}}(X)\leq c\cdot\mathrm{OPT}_{k}(X)$, and sets $\ell\leftarrow\lfloor\ell^{\star}/\Theta(c)\rfloor$, where $c=O(1)$ is a large constant. The epoch will last for the next $\ell+1$ updates. [^4] Note that it is very well possible that $\ell=0$. From now on, consider the superscript $t\in[0,\ell+1]$ to denote the status of some object after the algorithm has finished processing the $t^{th}$ update in the epoch. For example, at the start of the epoch, $X=X^{(0)}$.
Step 2: Preprocessing at the start of the epoch. Let $S_{\textnormal{{init}}}\leftarrow S$ be the solution maintained by the algorithm after it finished processing the last update in the previous epoch. Before handling the very first update in the current epoch, the algorithm initializes the maintained solution by setting
$S^{(0)}\leftarrow\arg\min_{S^{\prime}\subseteq S_{\textnormal{{init}}}\,:\,|S^{\prime}|=k-\ell}\textsc{cost}(X^{(0)},S^{\prime}).$ (2)
Step 3: Handling the updates within the epoch. Consider the $t^{th}$ update in the epoch, for $t\in[1,\ell+1]$. The algorithm handles this update in a lazy manner, as follows. If the update involves the deletion of a point from $X$, then it does not change the maintained solution, and sets $S^{(t)}\leftarrow S^{(t-1)}$. In contrast, if the update involves the insertion of a point $p$ into $X$, then it sets $S^{(t)}\leftarrow S^{(t-1)}+p$.
Step 4: Post-processing at the end of the epoch. After the very last update in the epoch, the algorithm does some post-processing, and computes another set $S_{\textnormal{{final}}}\subseteq X$ of at most $k$ centers (i.e., $|S_{\textnormal{{final}}}|\leq k$) that satisfies Invariant A. Then the next epoch can be initiated with $S\leftarrow S_{\textnormal{{final}}}$ being the current solution. The post-processing is done as follows.
The algorithm adds $O(\ell+1)$ extra centers to the set $S_{\textnormal{{init}}}$, while minimizing the cost of the resulting solution w.r.t. $X^{(0)}$.
$A^{\star}\leftarrow\arg\min\limits_{\begin{subarray}{c}A\subseteq[\Delta]^{d}:|A|\leq O(\ell+1)\end{subarray}}\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+A),\text{ and }S^{\star}\leftarrow S_{\textnormal{{init}}}+A^{\star}.$ (3)
The algorithm then adds the newly inserted points within the epoch to the set of centers, so as to obtain the set $T^{\star}$. Next, the algorithm identifies the subset $W^{\star}\subseteq T^{\star}$ of $k$ centers that minimizes the $k$-median objective w.r.t. $X^{(\ell+1)}$.
$T^{\star}\leftarrow S^{\star}+\left(X^{(\ell+1)}-X^{(0)}\right),\text{ and }W^{\star}\leftarrow\arg\min\limits_{\begin{subarray}{c}W\subseteq T^{\star}\,:\,|W|=k\end{subarray}}\textsc{cost}(X^{(\ell+1)},W).$ (4)
Finally, it calls a subroutine referred to as Robustify on $W^{\star}$ and lets $S_{\textnormal{{final}}}$ be the set of $k$ centers returned by this subroutine. The goal of calling Robustify is to make the final solution $S_{\textnormal{{final}}}$, robust w.r.t. the current space $X=X^{(\ell+1)}$. We will provide our implementation of this subroutine in Section 5.3. Finally, before starting the next epoch, the algorithm sets $S\leftarrow S_{\textnormal{{final}}}$.
$S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}(W^{\star}),\ \text{and}\ S\leftarrow S_{\textnormal{{final}}}.$ (5)
It can be shown that the maintained set $S\subseteq[\Delta]^{d}$ is always of size at most $k$, and the solution $S=S_{\textnormal{{final}}}$ satisfies Invariant A at the end of Step 4, that validates the analysis of the next epoch.
Barriers to Implementation in Euclidean Spaces. To implement this algorithmic framework in general metric spaces, [BCF25a] give various different subroutines for each of these steps efficiently. In particular, they design algorithms for solving the ‘restricted’ and ‘augmented’ $k$-means problems (see Section 4.4) and implement the Robustify subroutine (see Section 5.3). While these subroutines are fast enough for general metric spaces (where we can only hope to get an update time of $\tilde{O}(k)$), there are too slow for our setting (where we can obtain an update time of $\tilde{O}(k^{\epsilon})$). For example, the time taken to implement Step 1 and determine the length of an epoch in [BCF25a] is $\tilde{O}(k\cdot(\ell+1))$. Since this can be amortized over an epoch of length $\ell$, this leads to $\tilde{O}(k)$ amortized update time. In the Euclidean setting, however, we would need to perform this task in $\tilde{O}(k^{\epsilon}\cdot(\ell+1))$ time to get our desired bounds. Thus, we must develop new geometric data structures that allow us to implement versions of these subroutines much more efficiently in Euclidean spaces. In Section 4, we present the new geometric data structures that we introduce for our purpose.
4 Data Structures in High-Dimensional Euclidean Spaces
In this section, we present new fundamental data structures required to implement the framework of [BCF25a] in high-dimensional Euclidean spaces. There are some parameters we use throughout this section, whose final values in our algorithm are being set as follows.
$\Gamma=\epsilon^{-3/2},\quad\Lambda=2^{O(\epsilon d)}\cdot\mathrm{poly}(d\log\Delta)=\tilde{O}(n^{O(\epsilon)}).$ (6)
These data structures are designed to be robust against adaptive adversaries, which is necessary in our application since we may generate queries based on the previous output from the structure.
Our data structures rely on a fundamental geometric tool called consistent hashing [JLN+05, CJK+22, Fil24], specifically $(\Gamma,\Lambda)$-hashes as in Definition 4.1. We provide the first time-efficient consistent hashing with competitive parameters, strengthening previous efficient constructions that only achieve a relaxed guarantee (see Section 4.1). Then in Sections 4.2, 4.3 and 4.4, we leverage our new consistent hashing scheme to design several dynamic data structures against adaptive adversaries, especially to aid the two key sub-tasks called restricted $k$-means and augmented $k$-means problems.
Although our final application is clustering, the structures we devise may be useful for more general tasks. Specifically, we devise a reduction-style approximate range query structure, that turns any structure whose query concerns only global dataset, into one that supports range query, and a typical application is to return the coreset [HM04] with respect to a ball range given as query. We also devise a structure that approximately maintains the approximate nearest-neighbor assignment between a dynamic dataset and a dynamic center set, which may be used to approximately draw $D^{2}$-samples [AV07].
In this section, when we say “with high probability” (w.h.p.), we mean the algorithm satisfies all stated guarantee simultaneously with probability $1-1/\mathrm{poly}(\Delta^{d})$. We assume the access to a $(\Gamma,\Lambda)$-hash family in Sections 4.2, 4.3 and 4.4, where we use the $\Gamma$ and $\Lambda$ parameters without further quantification.
4.1 Efficient Consistent Hashing
We give in Definition 4.1 the notion of consistent hashing, which was introduced in [CJK+22] and was also studied under a closely related notion called sparse partitions [JLN+05, Fil24]. Roughly speaking, a consistent hashing is a data-oblivious map on $[\Delta]^{d}$ that partitions the space $[\Delta]^{d}$ into hash buckets, such that each bucket has some bounded diameter $\rho$, and every small ball of diameter $\frac{\rho}{\Gamma}$ can intersect only $\Lambda$ parts.
Definition 4.1 (Consistent Hashing).
A map $\varphi$ on $[\Delta]^{d}$ is a $(\Gamma,\Lambda,\rho)$-hash if
  1. 1.
    (diameter) $\forall z\in\varphi([\Delta]^{d}),\mathrm{diam}(\varphi^{-1}(z))\leq\rho$.
  2. 2.
    (consistency) $\forall x\in[\Delta]^{d},|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\Lambda$.
We also need the hash function to evaluate efficiently. In particular, given a point $x\in[\Delta]^{d}$, we need to evaluate $\varphi(x)$ and $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ (allowing approximation) in $O(\Lambda)$ time (noticing that $|\varphi(\textsc{Ball}(x,\rho/\Gamma))|\leq O(\Lambda)$ by the consistency guarantee), as follows.
Definition 4.2 (Efficient consistent hashing).
We call a $(\Gamma,\Lambda,\rho)$-hash $\varphi$ efficient, if for every $x\in[\Delta]^{d}$, a) $\varphi(x)$ can be evaluated in $\tilde{O}(\Lambda)$ time and b) a set $\Phi$ can be computed in $\tilde{O}(\Lambda)$ time such that $\varphi(\textsc{Ball}(x,\rho/\Gamma))\subseteq\Phi\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$.
We devise efficient consistent hashing with parameter $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\mathrm{poly}(d)$. This tradeoff is close to the optimal existential bound of $\Lambda=2^{\Omega(d/\Gamma)}$ from [Fil24] (which runs in $\exp(d)$ time). Previously, efficient construction of consistent hashing is only known from [FJL+25], but their consistency guarantee is in expectation which is too weak to satisfy Definition 4.1. While such a weaker guarantee may be useful in many applications, it is nonetheless insufficient for our application especially for devising structures robust to adaptive adversary.
Lemma 4.3.
For any $\Gamma\geq 2\sqrt{2\pi}$, $\rho>0$ and $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-1/\mathrm{poly}(\Delta^{d})$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
Construction. We sketch the proof of Lemma 4.3. Our starting point is the above-mentioned weaker construction from [FJL+25] which only achieves an expected consistency guarantee. We aim to boost this expected guarantee to a high-probability one with only $\tilde{O}(1)$ loss on the consistency bound. Specifically, let $\tilde{\varphi}$ be a random hashing on $\mathbb{R}^{d}$, the expected consistency guarantee of $\tilde{\varphi}$ is
$\forall x\in[\Delta]^{d},\quad\mathbb{E}\big{[}|\tilde{\varphi}(\textsc{Ball}(x,{\rho}/{\Gamma}))|\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Lambda.$
We describe our construction. We first draw $m=O(d\log\Delta)$ independent weaker hashing $\{\tilde{\varphi}_{i}\}_{i\in[m]}$ [FJL+25] with expected consistency guarantee. Then, with high probability, for all points $x\in[\Delta]^{d}$, there exists $i\in[m]$ such that the strict consistency condition $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ holds. We pick for each point $x$ the smallest such $i\in[m]$ and call this the color of $x$, denoted as $c_{x}$. Our hash $\varphi$ is defined as $\varphi(x):=(c_{x},\tilde{\varphi}_{c_{x}}(x))$, $\forall x\in[\Delta]^{d}$.
Consistency. The diameter bound $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$ carries over from $\tilde{\varphi}$’s. We next show the consistency $|\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq O(m\Gamma)$ for $x\in[\Delta]^{d}$. Let $Y:=\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))$, and we need $|Y|\leq O(\Lambda)$. For $i\in[m]$ let $Y_{i}:=\{(c_{x},\tilde{\varphi}_{c_{x}}(x))\in Y:c_{x}=i\}$ be the subset of hash value coming from color $i$. Then $|Y|\leq\sum_{i}|Y_{i}|$. We show that $|Y_{i}|\leq O(\Gamma)$, and this leads to $|Y|\leq O(m\Gamma)$ as desired.
Fix some $i$. By our construction, an immediate bound is $|Y_{i}|\leq|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|$, so we are done if $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Gamma$. However, it could be that $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|>2\Gamma$. We show that if this happens, then no point in $\textsc{Ball}(x,\rho/(2\Gamma))$ can have color $i$, implying $|Y_{i}|=0$. This follows from the contrapositive: suppose some $y\in\textsc{Ball}(x,\rho/(2\Gamma))$ has color $c_{y}=i$. Then $\textsc{Ball}(x,\rho/(2\Gamma))\subseteq\textsc{Ball}(y,\rho/\Gamma)$, so the consistency guarantee at $y$ implies $|\phi_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
Efficiency. Our new construction makes the hashing more complicated to evaluate, particularly to determine the color $c_{x}$ for $x\in[\Delta]^{d}$. By the definition of $c_{x}$, this reduces to checking whether $N_{i}:=|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ for each color $i$. To do this efficiently, we leverage the special structure of $\tilde{\varphi}_{i}$ in [FJL+25]: each $\tilde{\varphi}_{i}$ is defined using a randomly shifted grid over $[\Delta]^{d}$, where each point is hashed based on the grid cell that it belongs to. Under this grid structure, $\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))$ corresponds to a connected set of grid cells, where each cell is adjacent to $O(d)$ others (i.e., differs in only one coordinate). We can explore these cells using a depth-first search, and once the depth-first search takes more than $\tilde{O}(\Lambda)$ time, we can terminate the search and conclude that $N_{i}>2\Lambda$, and no point around $x$ has color $i$. This entire process takes total time $N_{i}\cdot\mathrm{poly}(d)$. The computation of $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ follows similar ideas.
4.2 Approximate Range Query Structure
An immediate application of consistent hashing yields a range query structure. This works as a general reduction, that turns a data structure $\mathcal{A}$ which answers query regarding only the entire dataset, to one that can answer queries on a metric ball $\textsc{Ball}(x,r)$. Specifically, we consider algorithm $\mathcal{A}$ that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, and that it reports some element $e_{X}$ on some domain $U$. For example, elements in $U$ may be real numbers, or a subset of points. We say an element $e_{T}\in U$ admissible to $\mathcal{A}$ with respect to some dataset $T$, if it satisfies the specification of $\mathcal{A}$, such as having some approximation ratio $\beta$ on $T$.
Lemma 4.4.
Suppose an algorithm $\mathcal{A}$ maintains dynamic $X\subseteq[\Delta]^{d}$ and it reports $e_{X}\in U$ as query. There exists a dynamic algorithm that takes $r>0$ and dynamic $X\subseteq[\Delta]^{d}$ as input, and for every query $x\in[\Delta]^{d}$ reports $\{e_{1},\ldots,e_{s}\}$ for $s\leq O(\Lambda)$ such that there exists an $s$-partition $\{T_{1},\ldots,T_{s}\}$ of $\textsc{Ball}^{\prime}(x,r)\cap X$ that satisfies $\textsc{Ball}(x,r)\subseteq\textsc{Ball}^{\prime}(x,r)\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$, and $e_{i}$ is admissible with respect to $T_{i}$ ($\forall i\in[s]$). The update and query run in $\tilde{O}(\Lambda)$ times that of $\mathcal{A}$, respectively. This algorithm works against adaptive adversary if $\mathcal{A}$ does, and succeeds w.h.p.
Application: Range Query Coreset. This reduction is useful when the query output of $\mathcal{A}$ can be merged, i.e., $e_{1},\ldots,e_{s}$ can be used to derive $e_{\textsc{Ball}^{\prime}(x,r)}$ for dataset $\textsc{Ball}^{\prime}(x,r)\cap X$. In particular, we devise a structure that reports a coreset of size $O(k)$ for $k$-means with respect to $T$ for every query $x\in[\Delta]^{d}$. Here, a coreset with respect to some $T\subseteq[\Delta]^{d}$ is a (weighted) subset $D$ such that $\forall S\subseteq[\Delta]^{d},|S|\leq k$ one has $\textsc{cost}(D,S)=\Theta(1)\cdot\textsc{cost}(T,S)$ [HM04]. We can implement this coreset with range query by plugging in Lemma 4.4 a known dynamic coreset algorithm [HK20, lTHS24] as $\mathcal{A}$. Then we can obtain a coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ via a standard merge-and-reduce framework [HM04]. In our application, we need this coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ with $k=1$, so the size becomes $O(1)$, which enables us to compute approximate solutions and cost estimations over the approximate ball $\textsc{Ball}^{\prime}(x,r)$ very efficiently, which is stated in the following lemma.
Lemma 4.5 ($1$-Means Estimation).
There is an algorithm that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, such that given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, we can compute in time $\tilde{O}(\Lambda)$ a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimations $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$ such that there exists some point subset $B_{x}\subseteq X$ with $\textsc{Ball}(x,r)\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$ satisfying the following.
  1. 1.
    $b_{x}=w(B_{x})$.
  2. 2.
    $\textsc{cost}(B_{x},c)\leq\widehat{\textsc{cost}}_{c}\leq O(1)\cdot\textsc{cost}(B_{x},c)$, for either $c\in\{c^{\star},\ x\}$.
  3. 3.
    $\textsc{cost}(B_{x},c^{\star})\leq O(1)\cdot\mathrm{OPT}_{1}(B_{x})$.
The algorithm has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
Application: ANN. In addition, Lemma 4.4 can also be applied to yield an $O(\Gamma)$-approximate nearest neighbor (ANN) structure with update and query time $\tilde{O}(\Lambda)$, against adaptive adversary. Similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing. We summarize this guarantee as the following lemma which is needed in our main algorithm.
Lemma 4.6 (ANN Distance).
The exists a data structure that explicitly maintains, for every $s\in S$, a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
$\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq O(\Gamma)\cdot\mathrm{dist}(s,S-s).$
The data structure has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
Proof Sketch of Lemma 4.4. The algorithm starts with imposing a $(\Gamma,\Lambda,\rho:=\Gamma r)$-consistent hashing $\varphi$ (Lemma 4.3), and run an independent instance of $\mathcal{A}$ for each bucket. To answer the query, one compute $\varphi(\textsc{Ball}(x,r))$, and return the query result of $\mathcal{A}$ for each bucket $z\in\varphi(\textsc{Ball}(x,r))$. The analysis is straightforward and follows easily from Definition 4.2. Previously, an MPC version of Lemma 4.4 was introduced in [CGJ+24], but their algorithm is not efficient due to the lack of efficient consistent hashing which we devise (Lemma 4.3).
4.3 Approximate Assignment Structure
Another novel use of consistent hashing is to devise a structure for maintaining approximate assignment. This task asks to take two dynamic sets $X,S\subseteq[\Delta]^{d}$ as input, and at any time maintains an explicit assignment $\sigma:X\to S$, such that each $x\in X$ is assigned to an approximately nearest neighbor in $S$.
However, there is no finite ratio with small recourse for this task, since one can keep on inserting/deleting a point of $S$ to change the assignment drastically. To bypass this barrier, instead of maintaining the full assignment $\sigma:S\to X$, our idea is to hash the dataset into buckets $\varphi(X)$ and explicitly maintain only an intermediate assignment $\sigma^{\prime}$ between bucket (ID) and $S$. The final assignment $S\to X$ is implicitly defined by assigning all points in the same bucket to the same bucket partner in the intermediate assignment.
The formal statement goes in Lemma 4.7 (which summarizes our investigations in Section 10), and the actual structure we maintain needs to be more sophisticated than simply hash buckets of $\varphi$. In this lemma, we take a two-level hashing approach, namely low-level and high-level hashes. Denote by $L$ the ID set of low-level buckets and $H$ for high-level. The set of data points in each low-level bucket is explicitly maintained, and we maintain a subset $H^{\prime}\subseteq H$ of high-level bucket IDs, each consists of a (dynamic) list of low-level bucket IDs (via map $f$). The intermediate assignment $\sigma^{\prime}$ is defined between $H^{\prime}$ and $S$. Importantly, one can only afford to maintain the list of low-level bucket ID in a high-level bucket $h\in H$, since it is possible that a small change to this ID list result in big changes in the point set in $h$.
Lemma 4.7.
There exists data-oblivious sets $L,H$, and an algorithm that takes dynamic sets $X,S\subseteq[\Delta]^{d}$ as input, and maintains [^5] The size of $L,H$ may be $\Delta^{d}$, so full description of $\operatorname{bk}(\ell)$ and $f:H\to 2^{L}$ is not efficient. Here, we assume both are described explicitly only on their support, i.e., $\operatorname{bk}(\ell)\neq\emptyset$ and $f(h)\neq\emptyset$. in $\tilde{O}(\Lambda^{2})$ time per update against adaptive adversary w.h.p.:
  • (low-level bucket) $\operatorname{bk}(\ell)\subseteq X$ for $\ell\in L$
  • (high-level bucket) map $f:H\to 2^{L}$, denoting $\operatorname{bk}(h):=\bigcup_{\ell\in f(h)}\operatorname{bk}(\ell)$, such that $\forall h\in H,\ell_{1}\neq\ell_{2}\in f(h)$, $\operatorname{bk}(\ell_{1})\cap\operatorname{bk}(\ell_{2})=\emptyset$ and $\forall h_{1}\neq h_{2}\in H$, $f(h_{1})\cap f(h_{2})=\emptyset$
  • (partition) $H^{\prime}\subseteq H$ such that $\forall x\in X$ there is unique $H^{\prime}(x)\in H^{\prime}$ where $x\in\operatorname{bk}(H^{\prime}(x))$
  • (assignment) $\sigma^{\prime}:H^{\prime}\to S$, and implicitly $\sigma:x\mapsto\sigma^{\prime}(H^{\prime}(x))$ ($x\in X$)
such that the following properties holds:
  1. 1.
    (ANN) $\forall x\in X$, $\mathrm{dist}(x,\sigma(x))\leq O(\Gamma^{2})\cdot\mathrm{dist}(x,S)$;
  2. 2.
    (equidistant) $\forall h\in H^{\prime}$, $\max_{x\in\operatorname{bk}(h)}\mathrm{dist}(x,S)\leq O(\Gamma^{2})\min_{x\in\operatorname{bk}(h)}\mathrm{dist}(x,S)$.
Lemma 4.7 implies key lemmas that we use for restricted $k$-Means and augmented $k$-Means (see Section 4.4 for the definitions of these sub-tasks). Similar to Lemma 4.7, these results maintain dynamic $X$ and $S$, work against an adaptive adversary and succeeds with high probability.
Application: Cluster Weight Estimation. An immediate application is to compute the injectivity $w_{s}$ of points $s\in S$ in the assignment $\sigma$ (which is $\mathrm{poly}(\Gamma)$-approximate by Property 1 of Lemma 4.7), via maintaining $|\operatorname{bk}(h)|$ (for every $h\in H^{\prime}$) and $\sum_{h\in H^{\prime}:\sigma^{\prime}(h)=s}|\operatorname{bk}(h)|$. It is well known that the weighted set $(S,w)$ is a good proxy for $X$, which we will use in our restricted $k$-Means subroutine.
Application: $D^{2}$-Sampling. Another more interesting application is to implement $D^{2}$-sampling, where the task is to sample an $x\in X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ [AV07]. To implement this, we compute for each high-level bucket $\operatorname{bk}(h)$ $(h\in H^{\prime})$ and the low-level buckets therein the sum of squared distance to the assigned $\sigma^{\prime}(h)$. Next, to do the sampling, first sample a high-level bucket ID $h\in H^{\prime}$ proportional to the sum-of-squared distance to $\sigma^{\prime}(h)$, then further sample a low-level bucket $\ell\in f(h)$ within $f(h)$ again proportional to the sum-of-squared distance, and finally do a uniform sample in the sampled (low-level) bucket $\operatorname{bk}(\ell)$. Here, uniform sampling works since Property 2 in Lemma 4.7 ensures points in a bucket have similar distance to $S$, and by this same property we can estimate for high-level bucket $h\in H$ the $\sum_{x\in\operatorname{bk}(h)}\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ by $|\operatorname{bk}(h)|\cdot\mathrm{dist}^{2}(x,\sigma^{\prime}(h))$ where $x$ is an arbitrary point in $\operatorname{bk}(h)$ (and note that both $|\operatorname{bk}(h)|$ and this point $x$ can be maintained), and similarly for a low-level bucket $\ell\in L$. This eventually enables us to sample $x$ proportional to $\mathrm{dist}^{2}(x,S)$, up to $O(\mathrm{poly}(\Gamma))$ factor. This approximate $D^{2}$-sampler is sufficient to be used in our another key subroutine, augmented $k$-Means.
Proof Sketch for Lemma 4.7. Our dynamic algorithm is a dynamization of the following static algorithm. For every distance scale $i$, we impose a ($\Gamma,\Lambda,\rho_{i}:=\Gamma 2^{i})$-hash $\varphi_{i}$. We define the high-level bucket IDs as $H:=\bigcup_{i}\varphi_{i}([\Delta]^{d})$. The low-level buckets are not needed at this point and will be discussed later. For $i=1,2\ldots$, for each $s\in S$, take all non-empty (i.e., with non-empty intersection with $X$) buckets (ID) in $\varphi_{i}(\textsc{Ball}(s,2^{i}))$ into $H^{\prime}$, assign them in $\sigma^{\prime}$ to $s$, remove the data points in all these buckets and continue to the next scale. This ensures an invariant that a $\varphi_{i}$-bucket, whose diameter is $\rho_{i}=O(\Gamma)2^{i}$, is picked only when its distance to the closest $s\in S$ is roughly $2^{i}$, and this can imply both Property 1 and Property 2.
Representing Partial High-level Buckets by Low-level Buckets. The caveat is that the level-$i$ buckets $\varphi_{i}(\textsc{Ball}(s,2^{i}))$ which we add to $H^{\prime}$ should actually exclude data points already handled/deleted from levels $\leq i-1$. In other words, only a part of a high-level bucket should be added. To describe these partial bucket, we introduce low-level buckets, which are “refinement” of high-level buckets, defined as follows: for each $i$, for each bucket ID $z\in\varphi_{i}([\Delta]^{d})$, partition the level-$i$ high-level bucket $\varphi^{-1}_{i}(z)$ with respect to $\varphi_{i-1}$ (i.e., the level-$(i-1)$ high-level buckets). The set of all these parts (i.e., the intersection with $(i-1)$-level buckets) compromise the low-level buckets $L$. Now, each high-level bucket with ID $h\in H$ should maintain a list of existing/non-empty intersecting low-level buckets, in order to determine the set of points belong $h$. This is precisely our $f:H\to 2^{L}$.
Locality of Low-level Buckets. One might wonder if we need also to consider partitioning a level-$i$ high-level bucket by buckets from $i-2,i-3,\ldots$ levels and include all these into $L$ and potentially also to $f(h)$. We show that this is not necessary and only level-$(i-1)$ is needed; specifically, when we remove a high-level bucket $h$ from some level $i$, then all points in this $h$ must all be removed (possibly from levels less than $i$), without leaving partially some low-level buckets behind. This is a crucial property to control the time complexity.
Dynamic Implementation. These properties lead to efficient dynamic implementation of the static algorithm. Take the insertion operation for example. Suppose $s\in S$ is inserted, we go level $i=1,2,\ldots$, examine the intersecting high-level buckets around $s$, and try to update the subset $H^{\prime}$ as well as the assignment. Now, before we proceed to insert $s$ to level-$(i+1)$, we populate the change of $H^{\prime}$ (caused by level $i$) to level $i+1$, by updating the list of associated low-level buckets in affected high-level buckets. We can then go to level-$(i+1)$ and repeat the process. The deletion works similarly and we omit the discussion.
Time complexity. The running time of the dynamic algorithm is governed by the consistency guarantee of consistent hashing. In particular, when we examine $\varphi_{i}(\textsc{Ball}(s,2^{i}))$, we only need to consider at most $O(\Lambda)$ intersecting high-level buckets. Moreover, the consistency also guarantees that each (high-level) bucket in level-($i-1$) may result in at most $O(\Lambda)$ low-level buckets in level $i$.
4.4 Restricted and Augmented $k$-Means
We discuss how to implement the two main subroutines in the algorithm of [BCF25a]: restricted $k$-means and augmented $k$-means. The restricted $k$-means algorithm is used in Step 1, 2, 4 of the framework in Section 3.2, and augmented $k$-means is used in Step 4. Both problems take a parameter $r>0$, as well as dataset $X$ and center set $S$ as dynamic input. The restricted $k$-Means aims to remove $O(r)$ points from $S$, and augmented $k$-Means aims to add $O(r)$ points to $S$, both to minimize the resultant clustering cost.
4.4.1 Restricted $k$-Means
Lemma 4.8 (Restricted $k$-Means).
There exists a data structure that handles two dynamic sets $X,S\subseteq[\Delta]^{d}$, where updates are generated by an adaptive adversary. For every query with an integer $1\leq r\leq|S|$, it runs in time $\tilde{O}(2^{\epsilon d}\cdot r+r^{1+\epsilon})$ to return a set $R\subseteq S$ of $r$ points such that the following holds:
$\textsc{cost}(X,S-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
where $\mathrm{OPT}_{|S|-r}^{S}(X):=\min_{S^{\prime}\subseteq S:|S^{\prime}|=|S|-r}\textsc{cost}(X,S^{\prime})$. Each update is handled in amortized time $\tilde{O}(2^{\epsilon d})$, and the algorithm succeeds w.h.p.
As a standard reduction, we may first move each point $x\in X$ to its nearest center in $S$, and then consider the problem on the induced weighted set $(S,w_{S})$ instead of the point set $X$. Therefore, in what follows, we assume $X$ is a weighted set supported on $S$.
A Sketch-Based Approach.
The high-level idea is to extract a tiny sketch of $S$, then solve the restricted $k$-Means problem on this sketch. Let $\widehat{\mathrm{dist}}$ be a (dynamic) approximate ANN distance array as in Lemma 4.6. The sketch is constructed as follows:
  1. 1.
    Pick the top $O(r)$ centers from $S$ whose $w_{S}(s)\cdot\widehat{\mathrm{dist}}^{2}(s,S-s)$ value is the smallest, denoted as $T_{1}$.
  2. 2.
    Let $T_{2}\subseteq S-T_{1}$ be a set of $O(r)$ centers such that $\forall s\in T_{1},\mathrm{dist}(s,T_{2})\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{dist}(s,S-T_{1})$, i.e. the set of approximate nearest neighbors in $S-T_{1}$ for all centers in $T_{1}$.
The sketch is defined as $T:=T_{1}\cup T_{2}$, and we show that $T$ approximately preserves the optimal value of the restricted $k$-means objective:
$\mathrm{OPT}^{T}_{|T|-r}(T)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
Finally, it suffices to run a vanilla $(|T|-r)$-means algorithm on $T$ and obtain a $\mathrm{poly}(\epsilon^{-1})$-approximate solution $R\subseteq T\subseteq S$:
$\textsc{cost}(R,T-R)=\textsc{cost}(T,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{T}_{|T|-r}(T)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}_{|S|-r}^{S}(S).$
This step takes $\tilde{O}(|T|^{1+\epsilon})=\tilde{O}(r^{1+\epsilon})$ time by applying static almost-linear time $k$-means algorithms [lTS24, JJLL25]. The correctness of $R$ as a solution on the full set $S$ follows from a straightforward derivation:
$\textsc{cost}(S,S-R)=\textsc{cost}(R,S-R)\leq\textsc{cost}(R,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
Dynamic Data Structure.
The key step in making the above process dynamic, is to maintain the center weights $w_{S}$. This can be solved using the cluster weight estimation structure which we introduced in Section 4.3. Once this is addressed, the construction of the sketch $T=T_{1}\cup T_{2}$ can be implemented via ANN data structures (using e.g., the one we mentioned in Section 4.2), in total time $O(r\cdot\Lambda)$.
4.4.2 Augmented $k$-Means
Lemma 4.9 (Augmented $k$-Means).
There exists a data structure that handles two dynamic sets $X,S\subseteq[\Delta]^{d}$, where updates are generated by an adaptive adversary. For every query with an integer $a\geq 1$, it runs in time $\tilde{O}(a\cdot 2^{\epsilon d})$ to return a set $A\subseteq[\Delta]^{d}$ of $O(a\cdot\mathrm{poly}(\epsilon^{-1})\cdot d\log\Delta)$ points such that the following holds w.h.p.:
$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|=a}\textsc{cost}(X,S+A^{\star}).$
Each update is handled in amortized time $\tilde{O}(2^{\epsilon d})$, and the algorithm succeeds w.h.p.
Our approach is to iteratively select and add points to $S$ until a good augmentation is achieved. We borrow the idea from the classical $k$-means++ algorithm [AV07], which employs $D^{2}$-sampling – i.e., sampling a point from $X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ – to iteratively select centers and achieve an $O(\log k)$-approximation for the vanilla $k$-means problem. A follow-up work [ADK09] considers sampling $O(k)$ centers to achieve a bi-criteria $O(1)$-approximation for vanilla $k$-means, and we adapt their analysis to show that repeated $D^{2}$-sampling similarly yields a good approximate solution to the augmented $k$-means problem in our setting. We employ an approximate $D^{2}$-sampling in Section 4.3 to implement this idea in dynamic setting.
5 Our Dynamic Algorithm: An Overview
In this section, we give an overview of our algorithm, which we obtain by implementing a significantly modified version of the algorithm from Section 3, using the data structures from Section 4. In this extended abstract, we provide our algorithm with $\tilde{O}(n^{\epsilon})$ update time. It is possible to reduce the update time to $\tilde{O}(k^{\epsilon})$ by sparsification techniques. The reader can find a thorough description of how to achieve $\tilde{O}(k^{\epsilon})$ update time in Section 19.
Roadmap.
In Section 5.1, we present some preliminaries for Section 5. In Section 5.2, we describe how to implement an epoch of our algorithm. The remaining subsections describe how to implement some specific tasks and subroutines.
5.1 Preliminaries
Preprocessing.
Throughout this section, we assume that the input space to our algorithm is $[\Delta]^{d}$, where the dimension $d$ is $O(\log n)$. This assumption can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume the input space is $\mathbb{R}^{m}$. We initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$. Whenever a point $x\in\mathbb{R}^{m}$ is inserted to the dynamic space, we first apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$. As a result, we can feed this new dynamic space into our algorithm. Since $d=O(\log n)$, we conclude that the running time performance of all the data structures in Section 4 becomes $\tilde{O}(n^{O(\epsilon)})$. Then, by scaling $\epsilon$, we assume that the running time of these data structures is $\tilde{O}(n^{\epsilon})$. For simplicity, we consider a parameter $p_{\epsilon}$ which is a polynomial of $(1/\epsilon)$ and is an upper bound on the approximation guarantees of all data structures in Section 4. So, we consider the following throughout the entire Section 5.
[…]
approximation guarantee at most $\text{approximation guarantee at most }p_{\epsilon}\text{ and running time at most }\tilde{O}(n^{\epsilon}).$
First, we extend the building blocks of Definition 3.1 to work with approximate balls, because of the following reason: While using the fast data structures in Section 4, for every $x\in[\Delta]^{d}$ and $r>0$, some of the points in the space that has a distance less than $r$ to $x$ might be missing in the approximate ball, as well as some point that has distance more than $r$ to $x$ might be mistakenly included in the ball. This motivates us to modify the definition of robust centers, as follows.
Definition 5.1.
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
  1. 1.
    For each $0\leq i\leq t$, we have $\textsc{Ball}(x_{i},p_{\epsilon}^{6i})\subseteq B_{i}\subseteq\textsc{Ball}(x_{i},p_{\epsilon}^{6i+2})$.
  2. 2.
    For each $1\leq i\leq t$, at least one of the following hold,
    • Either
      $\frac{\textsc{cost}(B_{i},x_{i})}{w(B_{i})}\geq p_{\epsilon}^{12i-8}\ \text{and}\ x_{i-1}=x_{i},$ (7)
    • Or
      $\frac{\textsc{cost}(B_{i},x_{i})}{w(B_{i})}\leq p_{\epsilon}^{12i-4}\ \text{and}\ \textsc{cost}(B_{i},x_{i-1})\leq\min\{(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{i}),\ \textsc{cost}(B_{i},x_{i})\}.$ (8)
Note that both of these cases can happen simultaneously, and in the second case, $x_{i-1}$ can be equal to $x_{i}$ as well. Moreover, we call a point $x$, $t$-robust if there exists a $t$-robust sequence $(x_{0},x_{1},\ldots,x_{t})$ such that $x=x_{0}$.
Definition 5.2.
We call a center set $S$, robust w.r.t. $X$ if and only if each center $u\in S$ satisfies the following
$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\mathrm{dist}(u,S-u)/p_{\epsilon}^{20}.$ (9)
These definitions are relaxations of Definition 3.1 and Definition 3.2, which enable us to work with approximate balls while keeping the main properties of robust center (with a small overhead on the approximation ratio).
5.2 Implementation
We provide the implementation of our algorithm for an epoch lasting for $(\ell+1)$ updates. At the beginning of the epoch, we satisfy the following invariant, which is a relaxed version of A.
Invariant B.
At the start of an epoch, the set $S$ is robust and $\textsc{cost}(S,X)\leq(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X)$.
We start by providing the implementation of Steps 1 to 4 of the framework of the algorithm by using data structures in Section 4. Then, we provide the Robustify subroutine in Section 5.3.
Implementing Step 1.
We find an estimation of $\ell^{\star}$ instead of the exact value. We define the exact value of $\ell^{\star}$ to be the largest integer satisfying the following
$\mathrm{OPT}_{k-\ell^{\star}}(X^{(0)})\leq p_{\epsilon}^{88}\cdot\mathrm{OPT}_{k}(X^{(0)}).$ (10)
Now, we proceed with the procedure of estimating $\ell^{\star}$. For each $i\in[0,\log_{2}k]$ define $s_{i}:=2^{i}$, and let $s_{-1}:=0$. We now run a for loop, as described in Algorithm 1.
Algorithm 1 Computing an estimate $\widehat{\ell}$ of the value of $\ell^{\star}$.
$i=0$to $\log_{2}k$
[…]
After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where ^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot p_{\epsilon}^{102}}\right\rfloor$.
Implementing Step 2.
Instead of finding the optimum set of $(k-\ell)$ centers within $S_{\textnormal{{init}}}$, we approximate it using restricted $k$-means (Lemma 4.8) by setting $S=S_{\textnormal{{init}}}$ and $r=\ell$. We compute a set of $(k-\ell)$ centers $S^{(0)}\subseteq S_{\textnormal{{init}}}$ such that $\textsc{cost}(X^{(0)},S^{(0)})\leq p_{\epsilon}\cdot\mathrm{OPT}_{k-\ell}^{S_{\textnormal{{init}}}}(X^{(0)})$.
Implementing Step 3.
Trivially, we can implement each of these updates in constant time.
Implementing Step 4.
We approximate $A^{\star}$ with augmented $k$-means (Lemma 4.9) by setting $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$ and $a=p_{\epsilon}^{104}\cdot(\ell+1)$ to get $A\subseteq[\Delta]^{d}$ of size $\tilde{O}(a)$ such that
$\textsc{cost}(X,S_{\textnormal{{init}}}+A)\leq p_{\epsilon}\cdot\min\limits_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|\leq a}\textsc{cost}(X,S_{\textnormal{{init}}}+A^{\star}),\quad a=p_{\epsilon}^{104}\cdot(\ell+1).$ (11)
Then, we set $S^{\prime}=S_{\textnormal{{init}}}+A$ and $T^{\prime}:=S^{\prime}+\left(X^{(\ell+1)}-X^{(0)}\right)$ (approximations of $S^{\star}$ and $T^{\star}$ respectively). Next, we compute $W^{\prime}$ (approximation of $W^{\star}$) using restricted $k$-means (Lemma 4.8) that satisfies
$\textsc{cost}(X^{(\ell+1)},W^{\prime})\leq p_{\epsilon}\cdot\mathrm{OPT}_{k}^{T^{\prime}}(X^{(\ell+1)}).$ (12)
Finally, we explain below how we implement the call to $\texttt{Robustify}(W^{\prime})$ (see Equation 5).
5.3 Implementing Robustify
The goal of this subroutine is to make the solution $W^{\prime}$ at the end of the epoch, robust (see Definition 5.2) w.r.t. the current dataset. Robustify can be divided into two tasks. First, identifying non-robust centers, i.e. the centers that violate Condition (9). Next, making them robust w.r.t the current data set.
Searching for Non-Robust Centers.
In the algorithm, we only need to make sure that we do not miss any center violating Condition (9). In order to do this, for every center $u\in S$ in the main solution, we maintain an integer $t[u]$ indicating that the center $u$ is $t[u]$-robust. We also maintain an approximation of $\mathrm{dist}(u,S-u)$ denoted by $\widehat{\mathrm{dist}}(u,S-u)$ using Lemma 4.6 (We use the notation $\widehat{\mathrm{dist}}$ extensively throughout this section). We consider the following condition.
$t[u]\geq t,\text{ where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$ (13)
Since $\widehat{\mathrm{dist}}(u,S-u)\geq\mathrm{dist}(u,S-u)$, it is obvious that if $u$ satisfies Condition (13), it definitely satisfies Condition (9). We will show how to maintain $t[u]$ and how to determine if $u$ violates Condition (13) in Section 5.4.
Making Centers Robust.
After identifying a center $u$ that violates Condition (13), we find the smallest integer $t$ satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$, swap $u$ with a close center $v$ which is $t$-robust through a call to subroutine $\texttt{MakeRobust}(u)$, and set $t[v]:=t$. Note that here the denominator is $p_{\epsilon}^{14}$ instead of $p_{\epsilon}^{20}$, which is intentional and swaps $u$ with a center $v$ that is more robust than needed in Condition (9). This helps us to make sure that as long as $\mathrm{dist}(v,S-v)$ is not change significantly, the center $v$ remains robust enough. We explain how $\texttt{MakeRobust}(u)$ works in Section 5.5.
5.4 Searching for Non-Robust Centers
We can split this task into two parts as follows.
  1. 1.
    Identify all centers $u$ that might not be $t[u]$-robust anymore.
  2. 2.
    Identify a center $u$ that violates Condition (13).
At the end of part 1, we make all the identified centers robust (via a call to MakeRobust), and update their $t[u]$ value. Only after that, we follow with part 2.
We show that every center $u\in S$ will be $t[u]$-robust, i.e. the detected centers in part 1 become robust via a call to MakeRobust, and a center $u$ that is not detected in part 1 is still $t[u]$-robust. This fact (Claim 5.3) helps us to implement part 2 later.
Claim 5.3.
Each center $u$ not detected in part 1, remains $t[u]$-robust w.r.t. the final space $X^{(\ell+1)}$.
The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.10).
Partitioning $W^{\prime}$.
We consider a partitioning $W^{\prime}=W_{1}\cup W_{2}\cup W_{3}$ as follows. $W_{1}$ consists of the new centers added to the solution, i.e. $W_{1}:=W^{\prime}-S_{\textnormal{{init}}}$. $W_{2}$ consists of all contaminated centers. A center $u$ is called contaminated, if throughout the epoch, the ball of radius $p_{\epsilon}^{6t[u]+3}$ around $u$ is changed, i.e.,
$\textsc{Ball}(u,p_{\epsilon}^{6t[u]+2})\cap\left(X^{(0)}\oplus X^{(\ell+1)}\right)\neq\emptyset.$
$W_{3}$ is then the rest of the centers $W_{3}:=W^{\prime}-(W_{1}\cup W_{2})$.
5.4.1 Implementing Part 1
According to Claim 5.3, the only centers which we do not have any guarantee about their robustness are centers in $W_{1}\cup W_{2}$. Hence, we find $W_{1}$ and $W_{2}$ and consider centers $W_{1}\cup W_{2}$ for part 1. It is obvious how to find $W_{1}$ explicitly since $W_{1}=W^{\prime}-S_{\textnormal{{init}}}$. We only need to keep track of the newly added centers during the epoch. In order to find $W_{2}$, we provide a property of contaminated centers as follows.
Claim 5.4.
For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$, there is at most one center $u\in S_{\textnormal{{init}}}$ such that $t[u]=i$ and $\mathrm{dist}(x,u)\leq p_{\epsilon}^{6i+4}$.
This claim shows that every update in the data set can contaminate at most one center in each scale $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$. The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.11).
Partitioning According to $t[u]$ Value.
Claim 5.4 provides us with a way to find contaminated centers efficiently. We maintain a partitioning of the main solution $S$ throughout the entire algorithm as follows. $S=\cup_{i=0}^{\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil}S[i]$, where $S[i]$ consists of centers $u\in S$ with $t[u]=i$. Whenever a new center $u$ is inserted into the solution, after computing $t[u]$, we insert $u$ into the appropriate part $S[t[u]]$. Now, for each set $S[i]$, we invoke the approximate nearest neighbor oracle (ANN) that given $x\in[\Delta]^{d}$, returns a $u\in S[i]$ such that $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$. In the following claim, we show that $S[i]$ has a special structure, such that the approximate nearest neighbor is the only center that can be contaminated by $x$ within $S[i]$ at the end of each epoch. The proof of this claim is deferred to Section 6.2.1.
Claim 5.5.
For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$, if $u\in S[i]$ satisfies $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$, then the only center in $S[i]$ that might be contaminated by $x$ is $u$.
This claim shows how to identify $W_{2}$. For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$, we call the nearest neighbor oracle on $S[i]$ to find $u\in S[i]$ satisfying $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$. Then we check if $\mathrm{dist}(x,u)\leq p_{\epsilon}^{6i+3}$. If this happens, we see that $x$ contaminates $u$, otherwise $x$ does not contaminate any center in $S[i]$ according to Claim 5.5. Hence, we can identify all centers in $W_{2}$.
5.4.2 Implementing Part 2
For this part, if we iterate over all centers in $S$ and check whether they satisfy Condition (13), the update time might be as large as $\Omega(k)$. Since we make centers more robust when they violate Condition (13), we conclude that as long as the value of $\mathrm{dist}(u,S-u)$ is not increased significantly, they remain robust. Hence, to identify if a center $u$ violates Condition (13), we only need to identify the centers $u$ such that the value of $\mathrm{dist}(u,S-u)$ is increased significantly.
Required Data Structure. In order to do this, for every $i\in[0,\lceil\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta)\rceil]$, we maintain the following data structure for parameter $\gamma=p_{\epsilon}^{2i}$.
Lemma 5.6.
Given any parameter $\gamma\geq 0$, there is a data structure $D_{\gamma}(S)$ that maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ for every center $s\in S$, and after every update (insertion/deletion) to $S$, it reports all the centers $s$ that $b_{\gamma}(s,S)$ changes. The following properties are satisfied, for every $s\in S$.
  1. 1.
    If $\mathrm{dist}(s,S-s)\leq\gamma$, then $b_{\gamma}(s,S)=1$.
  2. 2.
    If $\mathrm{dist}(s,S-s)>\mathrm{poly}(1/\epsilon)\cdot\gamma$, then $b_{\gamma}(s,S)=0$.
Each update in $S$ is handled in amortized $\tilde{O}(n^{\epsilon})$ time.
This lemma follows from combining the ANN and [BGJ+24, Theorem 3.1]. We have already implemented ANN in Section 4, and in [BGJ+24], it is explained how we can implement the above data structure by using an ANN.
Implementation Using Data Structure $D_{\gamma}(S)$. Now, assume we have data structures $D_{p_{\epsilon}^{2i}}(S)$ of Lemma 5.6 (for parameter $\gamma:=p_{\epsilon}^{2i}$) for every $i\in[0,\lceil\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta)\rceil]$. According to the guarantee of this data structure, if the bit $b_{p_{\epsilon}^{2i}}(u,S)$ has not changed for all values $i$, then $\mathrm{dist}(u,S-u)$ has not changed by more than a factor of $p_{\epsilon}$. This provides a way to implement part 2 as follows.
After each change in $S$, we first identify all centers $u\in S$ such that for at least one $i$, the bit $b_{p_{\epsilon}^{2i}}(u,S)$ has changed, and add all of these centers to a set $\mathcal{Y}$ that contains possible centers $u$ violating Condition (13). Now, for each center $u\in\mathcal{Y}$, we find the smallest integer $t$ satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$. If $t[u]\geq t$, then $u$ is already $t$-robust (according to Claim 5.3) and it satisfies Condition (13). Hence, there is no need to call a MakeRobust on $u$. Otherwise, (if $t[u]<t$), we call MakeRobust on $u$. Note that it is also possible in this case that $u$ is actually satisfying Condition (9), but we still call MakeRobust on $u$ since we should not miss any center violating the main Condition (9).
5.4.3 Termination of Robustify
After each change in the center set $S$, the distances of centers $u\in S$ to other centers might change. This means that new centers might be inserted into $\mathcal{Y}$ and they might need a subsequent call to MakeRobust. Similar to [BCF25a], it is possible to show that if a call MakeRobust is made on a center in Robustify, it will continue to satisfy Conditions (13) and (9) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
Lemma 5.7.
Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop. Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
The proof of this lemma follows from [BCF25a] (although we have new definitions), and we provide the proof in Part III for the sake of completeness (see Section 20.9).
We summarize our implementation of Robustify in Algorithm 2. Throughout this procedure, after each insertion or deletion on $S$, all of the data structures ($\widehat{\mathrm{dist}}(u,S-u)$, $D_{p_{\epsilon}^{2i}}(S)$, and nearest neighbor oracle on $S[i]$) maintained related to $S$ are updated.
Algorithm 2 Implementation of the call to $\texttt{Robustify}(W^{\prime})$ at the end of an epoch.
[…]
6:$u\leftarrow$pop$(\mathcal{Y})$
[…]
5.5 Make-Robust
In this section, we introduce the subroutine MakeRobust that given $u\in S$, swaps this center with a close center $v$ which is robust (i.e. satisfies Condition Equation 13 as well as Equation 9). We start by computing the smallest integer $t$ satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$. The goal is to follow the Definition 5.1. We define $x_{t}:=u$ and by iterating over $j=t$ down to 00, we find $x_{j-1}$ and $B_{j}$ satisfying the properties in Definition 5.1. We use Lemma 4.5, and make a query on this data structure for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$. It returns a center $c^{\star}\in[\Delta]^{d}$, together with numbers $\widehat{\textsc{cost}}_{c^{\star}}$, $\widehat{\textsc{cost}}_{x_{j}}$, and $b_{x_{j}}$.
Defining $B_{j}$.
We let $B_{j}:=B$, which is the existential ball $B$ in the guarantees of Lemma 4.5. Note that we do not need to know $B_{j}$ explicitly, and in Definition 5.1, the existence of this set is sufficient. According to the guarantees in Lemma 4.5, we have $b_{x_{j}}=w(B_{j})$ and $\textsc{Ball}(x_{j},p_{\epsilon}^{6j})\subseteq B_{j}\subseteq\textsc{Ball}(x_{j},p_{\epsilon}\cdot p_{\epsilon}^{6j})$. Since $p_{\epsilon}\cdot p_{\epsilon}^{6j}\leq p_{\epsilon}^{6j+2}$, we conclude that $B_{j}=B$ satisfies the first condition in Definition 5.1 for $i=j$. Now, we proceed with defining $x_{j-1}$ that satisfies the second condition.
Computing $x_{j-1}$.
We compare $\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}$ and $p_{\epsilon}^{6j-3}$. There are two cases as follows.
  1. 1.
    $\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq p_{\epsilon}^{6j-3}$. In this case, we let $x_{j-1}=x_{j}$, and show in Claim 6.4 that Equation 7 holds for $i=j$.
  2. 2.
    $\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq p_{\epsilon}^{6j-3}$. In this case, we compare $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}$ and $\widehat{\textsc{cost}}_{c^{\star}}$. If, $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{\star}}$, we let $x_{j-1}=x_{j}$, otherwise, we let $x_{j-1}=c^{\star}$. We show in Claim 6.5 that Equation 8 holds for $i=j$.
As a result, in all cases, the second condition in Definition 5.1 holds for $i=j$. Finally, it follows that the sequences $(x_{0},x_{1},\ldots,x_{t})$ and $(B_{0},B_{1},\ldots,B_{t})$ [^6] We define $B_{0}:=\textsc{Ball}(x_{0},1)$ at the end. satisfy the conditions in Definition 5.1. Hence, $x_{0}$ is a $t$-robust center. Below, we summarize the implementation of MakeRobust in Algorithm 3.
Algorithm 3 Implementation of a call to $\texttt{MakeRobust}(u)$ at the end of an epoch.
[…]
3:$x_{j-1}\leftarrow x_{j}$Else
[…]
5:$v\leftarrow x_{0}$
6:$S\leftarrow S-u+v$
7:Save $t[v]:=t$ together with $v$
6 Analysis of Our Algorithm: Proof (Sketch) of Theorem 1.1
In this section, we describe the analysis of our algorithm. We summarize the approximation, recourse, and update time guarantees of our algorithm with the following lemmas.
Lemma 6.1 (Approximation Ratio).
The approximation ratio of the algorithm is $\mathrm{poly}(1/\epsilon)$ w.h.p.
The analysis of the approximation ratio of the algorithm follows from the prior work [BCF25a] with deliberate changes in the parameters inside the proofs. There are only two important considerations as follows. First, our new definition of robust centers still satisfy the main properties of robust centers as in [BCF25a] (see Section 15.1.1). Second, the analysis of [BCF25a] depends on the fact that the integrality gap of the standard LP-relaxation for $k$-median is constant. Since the same fact holds for $k$-means (see Theorem 1.1 in [ANFSW20]), we can extend the analysis for our purpose. For the sake of completeness, we provide a thorough analysis in Part III.
Lemma 6.2 (Recourse).
The amortized recourse of the algorithm is at most $\tilde{O}(1)$.
The recourse analysis of the algorithm also follows from the previous work [BCF25a] with deliberate changes. There are only some considerations as follows. First, in augmented $k$-means (Lemma 4.9), instead of finding a set $A$ of size exactly $a$, we find a set of size $\tilde{O}(a)$. This increases the recourse of the algorithm by some $\log$ factors, but the same analysis works as in [BCF25a]. Second, we changed the Condition (9) to a more restricted Condition (13), and we might call MakeRobust more often. But, the same analysis as in [BCF25a] can be applied here to show that the total amortized recourse of Robustify is still $\tilde{O}(1)$. For the sake of completeness, we provide a thorough analysis in Part III.
Lemma 6.3 (Update Time).
The amortized update time of the algorithm is at most $\tilde{O}(n^{\epsilon})$.
We can reduce the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ by sparsification techniques. The reader can find a thorough description of how to achieve $\tilde{O}(k^{\epsilon})$ update time in Section 19. The rest of this section is devoted to proving Lemma 6.3.
6.1 Update Time Analysis
6.1.1 Maintaining the Background Data Structures
Each update in the main data set $X$ and in the main center set $S$ is considered an update for all the background data structures that are maintained in our algorithm. Consider a sequence of $T$ updates on $X$. The number of updates on $S$ is bounded by $\tilde{O}(T)$ according to Lemma 6.2. Hence, the total number of updates on every background data structure is bounded by $\tilde{O}(T)$.
Since every background data structures has $\tilde{O}(n^{\epsilon})$ update time (see Section 4), the total time spent to maintain all of these data structures is at most $\tilde{O}(T\cdot n^{\epsilon})$. Hence, the amortized update time for maintaining all of the background data-structures is $\tilde{O}(n^{\epsilon})$ (w.r.t. the original update sequence on the main dataset $X$).
6.1.2 Update Time of Each Step
Now, we proceed with the update time of the main algorithm.
Step 1.
The running time of Algorithm 1 is dominated by the last call (say at iteration $i=i^{\star}$) to restricted $k$-means (Lemma 4.8) that takes $\tilde{O}(n^{\epsilon}\cdot s_{i^{\star}})$ time. We have $s_{i^{\star}}=O(\widehat{\ell})=\tilde{O}(\ell+1)$. Hence, the running time of this step is $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$.
Step 2.
We spend $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time by calling restricted $k$-means (Lemma 4.8) to find the desired $S^{(0)}\subseteq S_{\textnormal{{init}}}$ of size $k-\ell$.
Step 3.
This step is trivial, and each update can be done in $\tilde{O}(1)$ time.
Step 4.
The call to augmented $k$-means (Lemma 4.9) for finding $A$ of size $\tilde{O}(a)$ (where $a=p_{\epsilon}^{104}\cdot(\ell+1)$) takes $\tilde{O}(n^{\epsilon}\cdot a)=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time. Computing $S^{\prime}=S_{\textnormal{{init}}}+A$ and $T^{\prime}=S^{\prime}+(X^{(\ell+1)}-X^{(0)})$ is trivial. Then, the call to restricted $k$-means (Lemma 4.8) on $T^{\prime}$ takes $\tilde{O}(n^{\epsilon}\cdot(|T^{\prime}|-|W^{\prime}|))=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time. Finally, we call Robustify on $W^{\prime}$.
Amortized Update Time.
Apart from Robustify, all of the other parts of the algorithm in this epoch, take at most $\tilde{O}(n^{\epsilon}(\ell+1))$ time. As a result, the amortized update time over the length of the epoch is bounded by $\tilde{O}(n^{\epsilon})$.
6.1.3 Running Time of Robustify
Recall that the Robustify is divided into two tasks. Searching for non-robust centers, and making them robust. The search for non-robust centers is divided into two parts. In part 1, we identify $W_{1}$ and $W_{2}$ (recall the partitioning $W^{\prime}=W_{1}\cup W_{2}\cup W_{3}$). Then, in part 2, we search for centers violating Condition (13).
Time Spent on Part 1.
It is trivial that we can identify $W_{1}=W^{\prime}\setminus S_{\textnormal{{init}}}$ (new centers) in $\tilde{O}(|W_{1}|)$ time, which is $\tilde{O}(\ell+1)$. According to the implementation in Algorithm 2, to find $W_{2}$, for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$, we will call the approximate nearest neighbor oracle (ANN) on $S[i]$. The oracle takes $\tilde{O}(n^{\epsilon})$ time to answer, which concludes the total time spend to find $W_{2}$ is at most $\tilde{O}(n^{\epsilon}\cdot|X^{(0)}\oplus X^{(\ell+1)}|\cdot\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil)=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$. Hence, the amortized time spent on part 1 for searching non-robust centers is $\tilde{O}(n^{\epsilon})$.
Time Spent on Part 2.
We need a global type of analysis for this part. It is possible that a single call to Robustify takes a huge amount of time. But, considering the entire sequence of $T$ updates, we show that the amortized update time of this part of the Robustify subroutine is at most $\tilde{O}(n^{\epsilon})$. During these $T$ updates, $S$ will be updated by at most $\tilde{O}(T)$ centers.
Elements are being inserted into $\mathcal{Y}$ by data structures $D_{p_{\epsilon}^{2i}}(S)$ (Lemma 5.6) for each $i\in[0,\lceil\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta)\rceil]$ after every update on the main data set $X$ or the main center set $S$. Since this data structure has an amortized update time of $\tilde{O}(n^{\epsilon})$, this implicitly means that the number of centers $u$ inserted into $\mathcal{Y}$ by data structure $D_{p_{\epsilon}^{2i}}(S)$ (i.e. whose bit $b_{p_{\epsilon}^{2i}}(u,S)$ has changed) is at most $\tilde{O}(T)\cdot\tilde{O}(n^{\epsilon})$. As a result, the total number of elements inserted into $\mathcal{Y}$ throughout these $T$ updates is at most $\tilde{O}((\lceil\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta)\rceil+1)\cdot T\cdot n^{\epsilon})$. In amortization, there are $\tilde{O}(n^{\epsilon})$ centers inserted into $\mathcal{Y}$.
For each of these centers $u\in\mathcal{Y}$, during 6 to 7, the algorithm decides in $\tilde{O}(1)$ time whether or not a call to MakeRobust will be made on $u$ [^7] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 4.6. As a result, the amortized time spend on part 2 is $\tilde{O}(n^{\epsilon})$.
Time Spent on Making Centers Robust.
According to the recourse analysis of the algorithm, the amortized number of calls to MakeRobust is bounded by $\tilde{O}(1)$. We show in Section 6.1.4 that each call to MakeRobust takes $\tilde{O}(n^{\epsilon})$ time. Hence, the amortized time spent to make centers robust is bounded by $\tilde{O}(n^{\epsilon})$.
6.1.4 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 4.5) used in 2 of Algorithm 3. It is obvious that all other parts of the implementation take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
6.2 Deferred Proofs
6.2.1 Proof of Claim 5.5
We consider two cases.
  1. 1.
    $\mathrm{dist}(x,u)\leq p_{\epsilon}^{6i+4}$. According to Claim 5.4, there does not exists $v\in S_{\textnormal{{init}}}-u$ satisfying $t[v]=i$ and $\mathrm{dist}(x,v)\leq p_{\epsilon}^{6i+4}$. Since $\textsc{Ball}(v,p_{\epsilon}^{6i+3})\subseteq\textsc{Ball}(v,p_{\epsilon}^{6i+4})$, we conclude that the distance between $x$ and $v$ is strictly greater than $p_{\epsilon}^{6i+3}$. So, $x$ does not contaminate $v$. Hence, there does not exist any $v\in S[i]-u$ contaminated by $x$, and the only center in $S[i]$ that might be contaminated by $x$ is $u$.
  2. 2.
    $\mathrm{dist}(x,u)>p_{\epsilon}^{6i+4}$. In this case, we conclude that $p_{\epsilon}^{6i+4}<\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$. Hence, the distance of any center $v$ in $S[i]$ to $x$ is strictly greater than $p_{\epsilon}^{6i+4}/p_{\epsilon}=p_{\epsilon}^{6i+3}$, which concludes there is no center $v\in S[i]$ contaminated by $x$.
6.2.2 Correctness of MakeRobust
Claim 6.4.
In Case 1, Equation 7 holds for $i=j$.
Proof.
We have
$p_{\epsilon}^{6j-3}\leq\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq p_{\epsilon}\cdot\textsc{cost}(B_{j},x_{j})/w(B_{j}),$
where the second inequality holds by the second guarantee in Lemma 4.5. It follows that
$\textsc{cost}(B_{j},x_{j})/w(B_{j})\geq p_{\epsilon}^{6j-4}.$
Claim 6.5.
In Case 2, Equation 8 holds for $i=j$.
Proof.
We have
$\textsc{cost}(B_{j},x_{j})/w(B_{j})\leq\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq p_{\epsilon}^{6j-2}/p_{\epsilon}\leq p_{\epsilon}^{6j-2},$ (14)
where the first inequality holds by the second guarantee in Lemma 4.5. Now, we consider two cases.
  1. 1.
    If $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{\star}}$, we have $x_{j-1}=x_{j}$ and
    $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},x_{j})$ $\leq\widehat{\textsc{cost}}_{x_{j}}\leq p_{\epsilon}\cdot\widehat{\textsc{cost}}_{c^{\star}}$
    $\leq(p_{\epsilon})^{2}\cdot\textsc{cost}(B_{j},c^{\star})\leq(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{j}).$
    The first, third, and fourth inequalities hold by the second and third guarantees in Lemma 4.5, respectively. Hence, $\textsc{cost}(B_{j},x_{j-1})\leq\min\{(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{j}),\textsc{cost}(B_{j},x_{j})\}$. Combining with Equation 14, we see that Equation 8 holds for $i=j$.
  2. 2.
    If $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\geq\widehat{\textsc{cost}}_{c^{\star}}$, we have $x_{j-1}=c^{\star}$ and
    $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},c^{\star})$ $\leq\widehat{\textsc{cost}}_{c^{\star}}\leq\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}$
    $\leq p_{\epsilon}\cdot\textsc{cost}(B_{j},x_{j})/p_{\epsilon}=\textsc{cost}(B_{j},x_{j}),$
    where the first and third inequalities hold by the second guarantee in Lemma 4.5. Moreover, $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},c^{\star})\leq p_{\epsilon}\cdot\mathrm{OPT}_{1}(B_{j})$, by the third guarantee of Lemma 4.5. Combining with Equation 14, we see that Equation 8 holds for $i=j$.
Part II Data Structures in High Dimensional Euclidean Spaces
Roadmap.
In this part of the paper, we provide several dynamic data structures that are useful in subsequent parts. We begin by summarizing the formal statement of these data structures in Section 7. As a crucial foundation of our data structures, we design a time-efficient consistent hashing scheme in Section 8. Then, we elaborate on the construction and analysis of each data structure in Sections 9, 10, 11 and 12.
Notation.
Throughout this part, we fix a small constant $10<\epsilon<$ 1, denote by $X\subseteq[\Delta]^{d}$ and $w:X\to\mathbb{R}_{\geq 0}$ a dynamic weighted point set and its weight function, respectively, and denote by $S\subseteq[\Delta]^{d}$ a dynamic (unweighted) center set. Without loss of generality, we assume that both $X$ and $S$ are initially empty. We use $\tilde{O}(\cdot)$ to hide the $\mathrm{poly}(d\epsilon^{-1}\log(\Delta))$ factors.
7 Main Data Structures
In this section, we summarize the main data structures that will be used later in Part III. All of them maintain both the dynamic weighted point set $X\subseteq[\Delta]^{d}$ and the dynamic unweighted center set $S\subseteq[\Delta]^{d}$; we thus omit these details from their statements for simplicity.
The following Lemmas 7.1, 7.2, 7.3 and 7.4 are established in Section 9, where we design a unified range query structure for handling a class of “mergeable” query problems. Specifically, Lemmas 7.1, 7.2 and 7.3 address a set of approximate nearest neighbor (ANN) problems, while Lemma 7.4 concerns the $1$-means problem over approximate balls. Here, an “approximate ball” refers to a subset that is sandwiched between two balls with the same center and (up to a constant) roughly the same radii.
Lemma 7.1 (ANN Oracle; see Lemma 9.2).
There exists a data structure that, for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update and query time $\tilde{O}(2^{\epsilon d})$.
Lemma 7.2 (ANN Indicators; see Lemma 9.3).
For any parameter $\gamma\geq 0$, there exists a data structure that maintains, for every $s\in S$, a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
  1. 1.
    If $\mathrm{dist}(s,S-s)\leq\gamma$, then $b_{\gamma}(s,S)=1$.
  2. 2.
    If $\mathrm{dist}(s,S-s)>\Omega(\epsilon^{-3/2})\cdot\gamma$, then $b_{\gamma}(s,S)=0$.
After every update to $S$, the data structure reports all the centers $s$ that $b_{\gamma}(s,S)$ changes. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Lemma 7.3 (ANN Distance; see Lemma 9.4).
The exists a data structure that maintains, for every $s\in S$, a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
$\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(s,S-s).$
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Lemma 7.4 ($1$-Means over Approximate Balls; see Lemma 9.5).
There exists a data structure that, for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
  1. 1.
    $\textsc{Ball}(x,r)\cap X\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\epsilon^{-3/2})\cdot r)\cap X$ and $b_{x}=w(B_{x})$.
  2. 2.
    $\textsc{cost}(B_{x},c)\leq\widehat{\textsc{cost}}_{c}\leq O(1)\cdot\textsc{cost}(B_{x},c)$, for either $c\in\{c^{\star},\ x\}$.
  3. 3.
    $\textsc{cost}(B_{x},c^{\star})\leq O(1)\cdot\mathrm{OPT}_{1}(B_{x})$.
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update and query time $\tilde{O}(2^{\epsilon d})$.
The following Lemmas 7.5 and 7.6 address two crucial subproblems – restricted $k$-means and augmented $k$-means – respectively, which play a central role in our dynamic algorithm described in Part III. We present the data structures of Lemmas 7.5 and 7.6 in Sections 11 and 12, respectively, both of which are built on a common data structure that (implicitly) maintains an approximate assignment from the point set $X$ to the center set $S$, as described in Section 10.
Lemma 7.5 (Restricted $k$-Means; see Lemma 11.1).
There exists a data structure that, on query of an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
where $\mathrm{OPT}_{|S|-r}^{S}(X):=\min_{S^{\prime}\subseteq S:|S^{\prime}|=|S|-r}\textsc{cost}(X,S^{\prime})$. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Lemma 7.6 (Augmented $k$-Means; see Lemma 12.1).
There exists a data structure that, on query of an integer $a\geq 1$, computes a size-$\Theta(a\cdot\epsilon^{-6}\cdot d\log\Delta)$ subset $A\subseteq[\Delta]^{d}$ in time $\tilde{O}(2^{\epsilon d}\cdot a)$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
8 Efficient Consistent Hashing
In this section, we describe our construction of an efficient consistent hashing scheme. We begin by introducing the definition of efficient consistent hashing and then state our main result.
Definition 8.1 (Efficient Consistent Hashing).
A mapping $\varphi:[\Delta]^{d}\to\Phi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash, for some codomain $\Phi$, if
  1. 1.
    (diameter) $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$, for every hash value $z\in\varphi([\Delta]^{d})$.
  2. 2.
    (consistency) $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\Lambda$, for every point $x\in[\Delta]^{d}$.
  3. 3.
    (efficiency) For every point $x\in[\Delta]^{d}$, both the hash value $\varphi(x)$ and a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ satisfying $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho)$ can be computed in time $\tilde{O}(\Lambda)$.
Lemma 8.2 (Efficient Consistent Hashing).
For any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
8.1 Proof of Lemma 8.2
Our construction of an efficient consistent hash builds on a previous construction from [FJL+25], which provides a weaker guarantee of consistency that holds in expectation. In addition, we leverage the special structure of their hash function – based on randomly shifted grids – and show that the set of hash values over any given ball can be computed efficiently (see Item 4 below). We summarize these results in the following lemma.
Lemma 8.3 (Hashing with Weaker Consistency Guarantees [FJL+25]).
Given any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\widetilde{\varphi}:[\Delta]^{d}\to\mathbb{Z}^{d}$, such that
  1. 1.
    $\mathrm{diam}(\widetilde{\varphi}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$.
  2. 2.
    $\mathbb{E}\big{[}|\widetilde{\varphi}(\textsc{Ball}_{\mathbb{R}^{d}}(x,\frac{2\rho}{\Gamma}))|\big{]}\leq 2^{O(d/\Gamma^{2/3})}\cdot\mathrm{poly}(d)$, for every point $x\in\mathbb{R}^{d}$.
  3. 3.
    Given a point $x\in[\Delta]^{d}$, the hash value $\widetilde{\varphi}(x)$ can be evaluated in time $\mathrm{poly}(d)$.
  4. 4.
    Given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, the set $\widetilde{\varphi}(\textsc{Ball}(x,r))$ can be computed in time $|\widetilde{\varphi}(\textsc{Ball}(x,r))|\cdot\mathrm{poly}(d)$.
Proof.
Items 1, 2 and 3 are explicitly stated in [FJL+25]. We now prove Item 4.
We review the construction from [FJL+25]: Let $v\sim\operatorname{Unif}([0,r/\sqrt{d}]^{d})$ be a uniformly random real vector. The hash function is defined as $\tilde{\varphi}:x\mapsto\lfloor\frac{x+v}{r/\sqrt{d}}\rfloor$ for $x\in[\Delta]^{d}$. Herein, the integer vector $\lfloor z\rfloor:=(\lfloor z_{1}\rfloor,\dots,\lfloor z_{d}\rfloor)\in\mathbb{Z}^{d}$ coordinatewise rounds down a real vector $z=(z_{1},\dots,z_{d})\in\mathbb{R}^{d}$.
Consider such an (infinite) grid graph $G$: its vertex set is $\mathbb{Z}^{d}$, and every vertex $z\in\mathbb{Z}^{d}$ is adjacent to the $2d$ other vertices $z^{\prime}\in\mathbb{Z}^{d}$ with $\|z-z^{\prime}\|=1$, i.e., exactly one coordinate differs by $1$.
The subgraph induced in $G$ by the subset $\widetilde{\varphi}(\textsc{Ball}(x,r))=\{z\in\mathbb{Z}^{d}:\mathrm{dist}(x,\widetilde{\varphi}^{-1}(z))\leq r\}$ is connected, shown as follows. For notational clarification, let $z^{x}=\widetilde{\varphi}(x)$ redenote the “sink” vertex. Start from a different “source” vertex $z\in\widetilde{\varphi}(\textsc{Ball}(x,r))\setminus\{z^{x}\}$ [^8] If otherwise $\widetilde{\varphi}(\textsc{Ball}(x,r))\setminus\{z^{x}\}=\emptyset$, the singleton $\widetilde{\varphi}(\textsc{Ball}(x,r))=\{z^{x}\}=\{\widetilde{\varphi}(x)\}$ is trivially connected. – without loss of generality, $z_{i}<z^{x}_{i}$ for some coordinate $i\in[d]$ – and consider the movement to one of its neighbors $z^{\prime}=(z_{i}+1,z_{-i})$. By construction, the hash preimages $\widetilde{\varphi}^{-1}(z)$ and $\widetilde{\varphi}^{-1}(z^{\prime})$ are hypercubes of the form
$\widetilde{\varphi}^{-1}(z)$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z_{j}\cdot r/\sqrt{d}-v_{j},(z_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)},$
$\widetilde{\varphi}^{-1}(z^{\prime})$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z^{\prime}_{j}\cdot r/\sqrt{d}-v_{j},(z^{\prime}_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)}.$
[…]
To find the entire subset $\widetilde{\varphi}(\textsc{Ball}(x,r))$, given the connectivity, we can start from $z^{x}=\varphi(x)$ and run a depth-first search; the running time is dominated by the product of two factors:
  • The number of vertices $z\in\mathbb{Z}^{d}$ visited by the depth-first search. Such a vertex either belongs to $\widetilde{\varphi}(\textsc{Ball}(x,r))$ or is a neighbor thereof, thus at most $|\widetilde{\varphi}\left(\textsc{Ball}(x,r)\right)|\cdot(1+2d)$ many ones.
  • The time to compute a distance $\mathrm{dist}(x,\widetilde{\varphi}^{-1}(z))$ and compare it with the radius $r$. This can be done in time $O(d)$, precisely because the hash preimage $\widetilde{\varphi}^{-1}(z)=\prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)$ is a hypercube and thus $\mathrm{dist}(x,\widetilde{\varphi}^{-1}(z))=(\sum_{j\in[d]}\mathrm{dist}^{2}(x_{j},\widetilde{\varphi}_{j}^{-1}(z)))^{1/2}$.
In combination, the total running time is $|\widetilde{\varphi}\left(\textsc{Ball}(x,r)\right)|\cdot\mathrm{poly}(d)$. This finishes the proof. ∎
We now elaborate on our construction for achieving Lemma 8.2.
The Construction
Our consistent hash function $\varphi$ builds on $C$ independent weakly consistent hash functions $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$ (Lemma 8.3), with a sufficiently large integer $C=\Theta(d\log\Delta)$; for notational clarification, we refer to the indexes of $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$ as colors $c\in[C]$.
Following Lemma 8.3 and using the sufficiently large integer $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$ chosen in the statement of Lemma 8.2, every $\widetilde{\varphi}_{c}$ satisfies that:
  • $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$. $\triangleright$ Item 1 of Lemma 8.3
  • $\mathbb{E}\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big{]}\leq\frac{\Lambda}{2C}$, for every point $x\in[\Delta]^{d}$. $\triangleright$ Item 2 of Lemma 8.3
The following Claim 8.4 measures the likelihood of a certain random event, which turns out to be the success probability of the construction of our consistent hash function $\varphi$.
Claim 8.4 (Success Probability).
Over the randomness of the weakly consistent hash functions $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$, the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
$\forall x\in[\Delta]^{d},\ \exists c\in[C]:$ $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\tfrac{2\rho}{\Gamma}))|\leq\tfrac{\Lambda}{C}.$ (15)
Proof.
Let $\mathcal{E}_{x,c}$ denote the event that Equation 15 holds for a specific pair $(x,c)\in[\Delta]^{d}\times[C]$. By Item 2 of Lemma 8.3 and Markov’s inequality, the event $\mathcal{E}_{x,c}$ occurs with probability
$\Pr[\mathcal{E}_{x,c}]=\Pr\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\leq\frac{\Lambda}{C}\big{]}\geq\Pr\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\leq 2\cdot\mathbb{E}\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big{]}\big{]}\geq\frac{1}{2}$.
Following the independence of $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$ and the union bound, when $C=\Theta(d\log\Delta)$ is large enough, Equation 15 holds with probability $\Pr[\wedge_{x\in[\Delta]^{d}}\vee_{c\in[C]}\mathcal{E}_{x,c}]\geq 1-\Delta^{d}\cdot\frac{1}{2^{C}}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
This finishes the proof. ∎
Since Equation 15 holds with high enough probability, we would safely assume so hereafter. Thus, every point $x\in[\Delta]^{d}$ has at least one color $c\in[C]$ satisfying Equation 15; we designate the color of this point $c_{x}\in[C]$ as, say, the smallest such $c\in[C]$. (We need not determine the colors $\{c_{x}\}_{x\in[\Delta]^{d}}$ at this moment – just their existence is sufficient.)
Based on the weakly consistent hash functions $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$, we define our consistent hash function $\varphi:[\Delta]^{d}\to\Phi$ as follows, where the output space $\Phi:=[C]\times\mathbb{Z}^{d}$; following Lemma 8.3, the entire construction takes time $C\cdot O(d)=\mathrm{poly}(d\log\Delta)$.
$\varphi(x)\leavevmode\nobreak\ :=\leavevmode\nobreak\ (c_{x},\widetilde{\varphi}_{c_{x}}(x)),$ $\forall x\in[\Delta]^{d}.$ (16)
The Performance Guarantees
Below, we show that the constructed $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash (Definition 8.1), conditioned on the high-probability event given in Equation 15.
First, the diameter bound and the consistency properties (Items 1 and 2 of Definition 8.1) follow from Claims 8.5 and 8.6, respectively.
Claim 8.5 (Diameter).
$\mathrm{diam}(\varphi^{-1}(c,z))\leq\rho$, for every hash value $(c,z)\in\varphi([\Delta]^{d})$.
Proof.
By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ satisfies $\mathrm{diam}(\varphi^{-1}(c,z))\leq\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, where the last step applies Item 1 of Lemma 8.3. This finishes the proof. ∎
Claim 8.6 (Consistency).
$|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\Lambda$, for every point $x\in[\Delta]^{d}$.
Proof.
By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ induces a natural partition of the input space $[\Delta]^{d}=\cup_{c\in[C]}G_{c}$, where $G_{c}:=\{x\in[\Delta]^{d}\mid c_{x}=c\}$, $\forall c\in[C]$. In this manner, we have $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\sum_{c\in[C]}|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|$, so it suffices to show that
$\forall c\in[C]:$ $|\varphi(\textsc{Ball}(x,\tfrac{\rho}{\Gamma})\cap G_{c})|\leq\tfrac{\Lambda}{C}.$
Consider a specific color $c\in[C]$ and a specific point $y\in\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c}$. [^9] If otherwise $\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c}=\emptyset$, we trivially have $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|=0\leq\frac{\Lambda}{C}$. Note that:
(i) $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})\subseteq\{(c,z):z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))\}$, by construction.
(ii) $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))\subseteq\widetilde{\varphi}_{c}(\textsc{Ball}(y,\frac{2\rho}{\Gamma}))$, since $y\in\textsc{Ball}(x,\frac{\rho}{\Gamma})$ and thus $\textsc{Ball}(x,\frac{\rho}{\Gamma})\subseteq\textsc{Ball}(y,\frac{2\rho}{\Gamma})$.
(iii) $|\widetilde{\varphi}_{c}(\textsc{Ball}(y,\frac{2\rho}{\Gamma}))|\leq\frac{\Lambda}{C}$, following Claim 8.4, Equation 16, and that $c_{y}=c\iff y\in G_{c}$.
In combination, we directly conclude with $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|\leq\frac{\Lambda}{C}$, as desired.
This finishes the proof. ∎
Second, the efficiency guarantee (Item 3 of Definition 8.1) directly follows from a combination of Claims 8.7 and 8.8.
Claim 8.7 (Hash Value Query).
Given a point $x\in[\Delta]^{d}$, the hash value $\varphi(x)$ can be computed in time $\tilde{O}(\Lambda)$.
Proof.
By definition (Equation 16), we have $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$.
First, we can find the color $c_{x}\in[C]$ – the smallest $c\in[C]$ satisfying Equation 15, and we have assumed its existence – as follows. Run such a for-loop $c=1,2,\dots,C$:
Try the depth-first search for Item 4 of Lemma 8.3 to compute $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))$ – it starts with the empty set $\emptyset$ and adds elements in $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))$ one by one – but stop it once we are aware of violation $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$ of Equation 15.
If it answers $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))$ before stop – namely satisfaction $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\leq\frac{\Lambda}{C}$ of Equation 15 – return $c_{x}\leftarrow c$ as the color (and terminate the for-loop $c=1,2,\dots,C$).
Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, the depth-first search for every single $c=1,2,\dots,C$ stops in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$.
Next, provided the above color $c_{x}\in[C]$, we can simply compute $\widetilde{\varphi}_{c_{x}}(x)\in\mathbb{Z}^{d}$ using Lemma 8.3 in time $O(d)=\tilde{O}(1)$. This finishes the proof. ∎
Claim 8.8 (Range Query).
Given a point $x\in[\Delta]^{d}$, a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ such that $\varphi(\textsc{Ball}(x,\tfrac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$ can be computed in time $\tilde{O}(\Lambda)$.
Proof.
The algorithm for constructing the subset $\Phi_{x}$ is very similar to the one for Claim 8.7. Start with the empty set $\Phi_{x}\leftarrow\emptyset$ and run such a for-loop $c=1,2,\dots,C$:
Try the depth-first search for Item 4 of Lemma 8.3 to compute $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ – it starts with the empty set $\emptyset$ and adds elements in $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ one by one – but stop it once we are aware of $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$.
If it answers $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ before stop – namely $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\frac{\Lambda}{C}$ – augment the subset $\Phi_{x}\leftarrow\Phi_{x}\cup\{(c,z)\}_{z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))}$ (and move on to the next iteration of the for-loop $c=1,2,\dots,C$).
Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, every iteration $c=1,2,\dots,C$ will stop in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$ and augment the subset $\Phi_{x}$ by at most $\frac{\Lambda}{C}$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$ and the total size $|\Phi_{x}|\leq\Lambda$.
It remains to establish that $\varphi(\textsc{Ball}(x,\tfrac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$.
First, consider a specific hash value $(c,z)\in\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ – there exists a point $y\in\textsc{Ball}(x,\frac{\rho}{\Gamma})$ such that $\varphi(y)=(c,z)\implies c_{y}=c$. Note that:
(i) $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\tfrac{\rho}{\Gamma}))|\leq|\widetilde{\varphi}_{c}(\textsc{Ball}(y,\tfrac{2\rho}{\Gamma}))|\leq\tfrac{\Lambda}{C}$, given $\textsc{Ball}(x,\tfrac{\rho}{\Gamma})\subseteq\textsc{Ball}(y,\tfrac{2\rho}{\Gamma})$ and Equation 15.
(ii) $z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$, given our construction of the hash function $\varphi$ (Equations 15 and 16).
Thus, our subset $\Phi_{x}$ by construction will include the considered hash value $\Phi_{x}\ni(c,z)$.
Then, the arbitrariness of $(c,z)\in\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ implies that $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})$.
Second, consider a specific hash value $(c,z)\in\Phi_{x}\cap\varphi([\Delta]^{d})$ – there exists a point $y\in[\Delta]^{d}$ such that $\varphi(y)=(c,z)\implies y\in\widetilde{\varphi}_{c}^{-1}(z)$ (Equation 16); we deduce from the triangle inequality that
[…]
Here, the second step applies $\mathrm{dist}(x,\widetilde{\varphi}_{c}^{-1}(z))\leq\frac{\rho}{\Gamma}\impliedby z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ (by our construction of the subset $\Phi_{x}$) and $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$ (Item 1 of Lemma 8.3).
Then, the arbitrariness of $(c,z)\in\Phi_{x}\cap\varphi([\Delta]^{d})$ implies that $\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$.
This finishes the proof. ∎
9 Approximate Range Query Structure
We present a range query data structure in Lemma 9.1. A query problem can be formalized as a pair $(U,f)$, where $U$ is the candidate solution space, and $f$ is a mapping from (input) a weighted point set $X\subseteq[\Delta]^{d}$ to (output) a set of valid solutions $f(X)\subseteq U$.
Lemma 9.1 (Range Query Structure).
For a query problem $(U,f)$ and a parameter $r>0$, assume access to an efficient $(\Gamma,\Lambda,r\cdot\Gamma)$-hash, and suppose there exists a data structure $\mathcal{A}$ that maintains a dynamic weighted set $X\subseteq[\Delta]^{d}$, and returns a solution $Y\in f(X)$ upon query.
Then, there exists a data structure that maintains a dynamic weighted set $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$, returns $Y_{1},\dots,Y_{t}\in U$ for some $1\leq t\leq\Lambda$, with the following guarantees: There exists a collection of disjoint subsets $P_{1},\dots,P_{t}\subseteq X$ such that
  • $\forall i\in[t]$, $Y_{i}\in f(P_{i})$, and
  • $\textsc{Ball}(x,r)\cap X\subseteq\bigcup_{i\in[t]}P_{i}\subseteq\textsc{Ball}(x,3\Gamma\cdot r)\cap X$.
The update time is that of $\mathcal{A}$ plus $\tilde{O}(\Lambda)$, and the query time is $\tilde{O}(\Lambda)$ times that of $\mathcal{A}$. The data structure works against an adaptive adversary if $\mathcal{A}$ does.
We provide the proof of Lemma 9.1 in Section 9.1.
Lemma 9.1 is particularly useful for mergeable query problems, which means that for any two disjoint weighted point sets $X_{1},X_{2}\subseteq[\Delta]^{d}$ and their respective valid solutions $Y_{1}\in f(X_{1})$ and $Y_{2}\in f(X_{2})$, one can compute from $\{Y_{1},Y_{2}\}$ a valid solution $Y\in f(X_{1}+X_{2})$ for the union. For such problems, the output $\{Y_{1},\dots,Y_{t}\}$ returned by Lemma 9.1 can be integrated into a single solution $Y\in f(\bigcup_{i\in[t]}P_{i})$, where $\bigcup_{i\in[t]}P_{i}$ is “sandwiched” between the ball $\textsc{Ball}(x,r)$ (the given query range) and a slightly larger ball $\textsc{Ball}(x,O(\Gamma)\cdot r)$. Specifically, in Section 9.2, as a warm-up, we show how Lemma 9.1 can be used to derive the fundamental approximate nearest neighbor data structures; in Section 9.3, we design a data structure for the $1$-means problem on any approximate ball, which serves as an important component of our dynamic $k$-means algorithm.
9.1 Proof of Lemma 9.1
Let $\varphi:[\Delta]^{d}\to\Phi$ denote the given efficient $(\Gamma,\Lambda,r\cdot\Gamma)$-hash. For every $z\in\Phi$, let $X_{z}:=X\cap\varphi^{-1}(z)$. The data structure maintains the partition $\{X_{z}\}_{z\in\varphi(X)}$ induced by the hash function $\varphi$ and, for every $z\in\varphi(X)$, an instance $\mathcal{A}_{z}$ that handles dynamic updates to $X_{z}$.
Upon each insertion (resp. deletion) of a point $x\in[\Delta]^{d}$, we compute its hash value $z_{x}:=\varphi(x)$ in $\tilde{O}(\Lambda)$ time (Definition 8.1). We then update the partition via $X_{z_{x}}\leftarrow X_{z_{x}}+x$ (resp. $X_{z_{x}}\leftarrow X_{z_{x}}-x$), and feed the update to $\mathcal{A}_{z_{x}}$ (creating it if it does not exist). Clearly, this procedure correctly maintains the data structure, and it involves one update operation of $\mathcal{A}$ with an additional $\tilde{O}(\Lambda)$ time.
Now, we explain how we answer queries. For a given $x\in[\Delta]^{d}$, we compute a set $\Phi_{x}$ such that
$\varphi(\textsc{Ball}(x,r))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}\cap\varphi([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi(\textsc{Ball}(x,2\Gamma\cdot r)).$ (17)
Since $\varphi$ is an efficient $(\Gamma,\Lambda,r\cdot\Gamma)$-hash (Definition 8.1), the computation of $\Phi_{x}$ takes $\tilde{O}(\Lambda)$ time and we have $|\Phi_{x}|\leq\Lambda$. Since we explicitly maintain the partition $\{X_{z}\}_{z\in\varphi(X)}$, we can remove all $z$ with $X_{z}=\emptyset$ from $\Phi_{x}$ in $\tilde{O}(\Lambda)$ time. Henceforth, we assume $X_{z}\neq\emptyset$ (and thus $\mathcal{A}_{z}$ exists) for every $z\in\Phi_{x}$. We query $\mathcal{A}_{z}$ to obtain $Y_{z}\in f(X_{z})$ for every $z\in\Phi_{x}$, and return the collection $\{Y_{z}\}_{z\in\Phi_{x}}$. The query time consists of $\tilde{O}(\Lambda)$ queries to $\mathcal{A}$ and $\tilde{O}(\Lambda)$ overhead, resulting in a total of $\tilde{O}(\Lambda)$ times the query time of $\mathcal{A}$.
Clearly, the sets $\{X_{z}\}_{z\in\Phi_{x}}$ are pairwise disjoint. It remains to show that $\bigcup_{z\in\Phi_{x}}X_{z}$ forms an approximate ball such that
$\textsc{Ball}(x,r)\cap X\subseteq\bigcup_{z\in\Phi_{x}}X_{z}\subseteq\textsc{Ball}(x,3\Gamma\cdot r)\cap X,$ (18)
thereby ensuring the correctness of $\{Y_{z}\}_{z\in\Phi_{x}}$.
To this end, consider an arbitrary point $y\in\textsc{Ball}(x,r)\cap X$. By Equation 17, we have $\varphi(y)\in\Phi_{x}$, which implies that $y\in\bigcup_{z\in\Phi_{x}}(X\cap\varphi^{-1}(z))=\bigcup_{z\in\Phi_{x}}X_{z}$. The arbitrariness of $y\in\textsc{Ball}(x,r)\cap X$ implies that $\textsc{Ball}(x,r)\cap X\subseteq\bigcup_{z\in\Phi_{x}}X_{z}$.
On the other hand, consider a specific $y\in X_{z}$ for some $z\in\Phi_{x}$. By Equation 17, there exists $y^{\prime}\in\textsc{Ball}(x,2\Gamma\cdot r)$ such that $\varphi(y)=\varphi(y^{\prime})$. Since $\varphi$ is $(\Gamma,\Lambda,r\cdot\Gamma)$-consistent, we have $\mathrm{dist}(y,y^{\prime})\leq r\cdot\Gamma$ (see Definition 8.1). By the triangle inequality,
$\mathrm{dist}(x,y)\leq\mathrm{dist}(x,y^{\prime})+\mathrm{dist}(y^{\prime},y)\leq 3\Gamma\cdot r\implies y\in\textsc{Ball}(x,3\Gamma\cdot r)\cap X.$
The arbitrariness of $y\in\bigcup_{z\in\Phi_{x}}X_{z}$ implies $\bigcup_{z\in\Phi_{x}}X_{z}\subseteq\textsc{Ball}(x,3\Gamma\cdot r)\cap X$.
Together, the above discussions imply Equation 18 and thus complete the proof.
9.2 Application: Approximate Nearest Neighbor Structure
As a warm-up application of Lemma 9.1, we describe how to design an approximate nearest neighbor (ANN) oracle – a data structure that, given a dynamic center set $S\subseteq[\Delta]^{d}$, returns for each query point $x$ a (distinct) approximate nearest center $s\in S-x$. We note that similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
Without loss of generality, we assume hereafter that $S$ contains at least two centers, which guarantees the existence of a distinct nearest neighbor for any query point. Also, the case $|S|\leq 1$ is trivial.
Lemma 9.2 (ANN Oracle).
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic center set $S\subseteq[\Delta]^{d}$, and for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update and query time $\tilde{O}(2^{\epsilon d})$.
Proof.
Let $L:=\lceil\log(\sqrt{d}\Delta)\rceil$. For each $0\leq i\leq L$, we sample a $(\Gamma,\Lambda,2^{i}\cdot\Gamma)$-hash $\varphi_{i}$ from Lemma 8.2, where $\Gamma=O(\epsilon^{-3/2})$ and $\Lambda=\tilde{O}(2^{\epsilon d})$. By a union bound, with probability at least $1-\frac{L+1}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, all hash functions $\varphi_{0},\dots,\varphi_{L}$ are successfully sampled.
Consider a simple query problem $(U,f)$, where $U=[\Delta]^{d}\cup\{\perp\}$, and for every $P\subseteq[\Delta]^{d}$, we define $f(P)=\{\{x\}:x\in P\}$ if $P\neq\emptyset$, and $f(P)=\{\perp\}$ if $P=\emptyset$; that is, we return an arbitrary point from the set $P$ if it is non-empty. One can easily design a data structure that maintains a dynamic set $S$ and reports $a\in f(S)$ with both update and query time $\tilde{O}(1)$, while working against an adaptive adversary. Let $\mathcal{A}$ denote this data structure. For every $i\in[0,L]$, by plugging the $(\Gamma,\Lambda,2^{i}\cdot\Gamma)$-hash $\varphi_{i}$ and $\mathcal{A}$ into Lemma 9.1, we obtain a data structure $\mathcal{D}_{i}$ with update time $\tilde{O}(\Lambda)$. Our ANN data structure for Lemma 9.2 simply maintains $\{\mathcal{D}_{i}\}_{i\in[0,L]}$, so the update time is $(L+1)\cdot\tilde{O}(\Lambda)=\tilde{O}(2^{\epsilon d})$.
We then explain how to use $\{\mathcal{D}_{i}\}_{i\in[0,L]}$ to answer a query. First, we consider a query for a non-center point $x\in[\Delta]^{d}-S$. In this case, for every $i\in[0,L]$, we query $\mathcal{D}_{i}$ and obtain $a_{i,1},\dots,a_{i,t}\in U$ for some $1\leq t\leq\Lambda$. Since this query problem $(U,f)$ is clearly mergeable, we can compute a solution $a_{i}$ from $\{a_{i,j}\}_{j\in[t]}$ in $O(t)=O(\Lambda)$ time, such that, by Lemma 9.1, $a_{i}\in f(B_{i})$ for some $B_{i}\subseteq S$ satisfying that $\textsc{Ball}(x,2^{i})\cap S\subseteq B_{i}\subseteq\textsc{Ball}(x,3\Gamma\cdot 2^{i})\cap S$. Let $i^{*}\in[0,L]$ denote the smallest index such that $a_{i^{*}}\neq\perp$, and return $a_{i^{*}}$ as the answer. Such an index $i^{*}$ must exist, since $2^{L}\geq\sqrt{d}\cdot\Delta$ implies $S\subseteq B_{L}$. Hence, $a_{L}\in f(S)\neq\{\perp\}$.
We then prove the correctness: Notice that $a_{i^{*}}=f(B_{i^{*}})\neq\perp$ is equivalent to $B_{i^{*}}\neq\emptyset$ and $a_{i^{*}}\in B_{i^{*}}$. Since $B_{i^{*}}\subseteq\textsc{Ball}(x,3\Gamma\cdot 2^{i^{*}})\cap S$, it follows that $\mathrm{dist}(x,a_{i^{*}})\leq 3\Gamma\cdot 2^{i^{*}}$. On the other hand, if $i^{*}=1$, then $\mathrm{dist}(x,S)\geq 1$ (since $x\notin S$), and therefore $\mathrm{dist}(x,a_{i^{*}})\leq 3\Gamma\cdot\mathrm{dist}(x,S)$. If $i^{*}>1$, then $a_{i^{*}-1}=\perp$, which implies that $\textsc{Ball}(x,2^{i^{*}-1})\cap S\subseteq B_{i^{*}-1}=\emptyset$. Hence, $\mathrm{dist}(x,S)\geq 2^{i^{*}-1}$, and therefore $\mathrm{dist}(x,a_{i^{*}})\leq 6\Gamma\cdot\mathrm{dist}(x,S)$. Since $\Gamma=O(\epsilon^{-3/2})$, it follows that $a_{i^{*}}$ is a correct answer for the query $x\in[\Delta]^{d}-S$.
The query time is primarily due to querying $\mathcal{D}_{i}$ and merging the output for all $i\in[0,L]$, totaling $(L+1)\cdot\tilde{O}(\Lambda)=\tilde{O}(2^{\epsilon d})$.
Finally, for a query where the point $x\in S$ is a center, we temporarily remove $x$ from $S$, perform the query as above, and then reinsert $x$ into $S$. The total query time remains $\tilde{O}(2^{\epsilon d})$. ∎
An ANN oracle can be used to derive many useful data structures. By plugging Lemma 9.2 into the framework of [BGJ+24], we obtain the following ANN indicator data structure.
Lemma 9.3 (ANN Indicators).
For any $\gamma\geq 0$ and $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
  1. 1.
    If $\mathrm{dist}(s,S-s)\leq\gamma$, then $b_{\gamma}(s,S)=1$.
  2. 2.
    If $\mathrm{dist}(s,S-s)>\epsilon^{-3/2}\cdot\gamma$, then $b_{\gamma}(s,S)=0$.
After every update to $S$, the data structure reports all the centers $s$ that $b_{\gamma}(s,S)$ changes. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Proof.
Follows from Lemma 9.2 and [BGJ+24, Theorem 3.1]
Furthermore, based on the ANN indicators maintained in Lemma 7.2, we derive the following result, which is a data structure that maintains the approximate distance from each center in $S$ to its nearest neighbor.
Lemma 9.4 (ANN Distance).
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
$\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(s,S-s).$
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Proof.
The data structure for Lemma 9.4 builds upon the one used to maintain the nearest neighbor indicator in Lemma 9.3.
Without ambiguity, we simply write $b_{\ell}^{s}=b_{2^{\ell}}(s,S)$ throughout this proof. Let $L:=\lceil\log(\sqrt{d}\cdot\Delta)\rceil$.
For every scale $\ell\in[L]$, we build a data structure $D_{\ell}(S)$ from Lemma 7.2, which maintains a bit $b_{\ell}^{s}$ for every center $s\in S$ such that:
  • If $\mathrm{dist}(s,S-s)\leq 2^{\ell}$, then $b_{\ell}^{s}=1$. (If $b_{\ell}^{s}=0$, then $\mathrm{dist}(s,S-s)>2^{\ell}$.)
  • If $\mathrm{dist}(s,S-s)>\epsilon^{-3/2}\cdot 2^{\ell}$, then $b_{\ell}^{s}=0$. (If $b_{\ell}^{s}=1$, then $\mathrm{dist}(s,S-s)\leq\epsilon^{-3/2}\cdot 2^{\ell}$.)
We define $\widehat{\mathrm{dist}}(s,S-s):=\epsilon^{-3/2}\cdot 2^{\ell_{s}}$ for every center $s\in S$, where $\ell_{s}:=\min\{\ell\in[L]:b_{\ell}^{s}=1\}$ is the “threshold scale”; note that $\ell_{s}\in[L]$ must be well-defined, since $2^{L}=2^{\lceil\log(\sqrt{d}\Delta)\rceil}\geq\sqrt{d}\Delta=\mathrm{diam}([\Delta]^{d})$ and thus $b_{L}^{s}=1$.
First, we show the correctness of $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$. Consider a specific center $s\in S$. By construction, we have $b_{\ell_{s}}^{s}=1\implies\mathrm{dist}(s,S-s)\leq\epsilon^{-3/2}\cdot 2^{\ell_{s}}=\widehat{\mathrm{dist}}(s,S-s)$. When $\ell_{s}=1$, we have $\mathrm{dist}(s,S-s)\geq 1=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$, since two distinct points in the discrete space $[\Delta]^{d}$ have distance $\geq 1$. When $\ell_{s}\neq 1$, we have $b_{\ell_{s}-1}^{s}=0\implies\mathrm{dist}(s,S-s)>2^{\ell_{s}-1}=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$. In either case, we always have $\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq 2\epsilon^{-3/2}\cdot\mathrm{dist}(s,S-s)$.
Second, we consider the update time. Clearly, the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ can be maintained simultaneously with $\{b_{\ell}^{s}\}_{s\in S}$ for all scales $\ell\in[L]$. Specifically, for every update to the center set $S$, Lemma 7.2 reports all changes of the indicators $\{b_{\ell}^{s}\}_{s\in S,\ell\in[L]}$. Regarding a reported change $b_{\ell}^{s}$ (say), we can simultaneously recompute the threshold scale $\ell_{s}=\min\{\ell\in[L]:b_{\ell}^{s}=1\}$ and update the value $\widehat{\mathrm{dist}}(s,S-s)=\epsilon^{-3/2}\cdot 2^{\ell_{s}}$. This step takes time $O(L)=\tilde{O}(1)$. Hence, the update time for updating $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ is asymptotically equal to updating all $\{b_{\ell}^{s}\}_{s\in S}$ (up to an $\tilde{O}(1)$ factor), which overall takes $O(L)\cdot\tilde{O}(\Lambda)=\tilde{O}(\Lambda)$.
This finishes the proof. ∎
9.3 Application: $1$-Means Range Query
Here, we use Lemma 9.1 to devise a data structure that supports estimating both the $1$-means objective and its solution on any given approximate ball. This data structure is obtained by plugging a dynamic coreset algorithm from [HK20] (for the entire set $X$) into Lemma 9.1; hence, for each query, we first obtain a small-sized coreset and then compute the objective and/or solution based on it.
Lemma 9.5 ($1$-Means over Approximate Balls).
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic weighted dataset $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{*}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
  1. 1.
    $\textsc{Ball}(x,r)\cap X\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\epsilon^{-3/2})\cdot r)\cap X$ and $b_{x}=w(B_{x})$.
  2. 2.
    $\textsc{cost}(B_{x},c)\leq\widehat{\textsc{cost}}_{c}\leq O(1)\cdot\textsc{cost}(B_{x},c)$, for either $c\in\{c^{*},\ x\}$.
  3. 3.
    $\textsc{cost}(B_{x},c^{*})\leq O(1)\cdot\mathrm{OPT}_{1}(B_{x})$.
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update and query time $\tilde{O}(2^{\epsilon d})$.
Proof.
We first consider two extreme cases of the query radius $r$. In the case of a small radius $10\leq r<$ 1, we have $\textsc{Ball}(x,r)=\{x\}$ (recall that any two distinct points in the discrete space $[\Delta]^{d}$ have distance at least $1$), and we simply consider the point subset $B_{x}=\{x\}\cap X$. Either subcase – $B_{x}=\{x\}$ or $B_{x}=\emptyset$ – is trivial. In the case of a large radius $r>\sqrt{d}\Delta=\mathrm{diam}([\Delta]^{d})$, we have $\textsc{Ball}(x,r)\cap X=X$, and we simply consider the entire point set $B_{x}=X$. Thus, this extreme case $r>\sqrt{d}\Delta$ reduces to the boundary case $r=\sqrt{d}\Delta$.
Hereafter, we assume a bounded query radius $1\leq r\leq\sqrt{d}\Delta$. Let $L:=\lceil\log(\sqrt{d}\Delta)\rceil$. For each $i\in[L+1]$, we sample a $(\Gamma,\Lambda,2^{i}\cdot\Gamma)$-hash $\varphi_{i}$ from Lemma 8.2, where $\Gamma=O(\epsilon^{-3/2})$ and $\Lambda=\tilde{O}(2^{\epsilon d})$. By a union bound, with probability at least $1-\frac{L+1}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, all hash functions $\varphi_{0},\dots,\varphi_{L}$ are successfully sampled.
Fix an integer $i\in[L+1]$. We design a sub-data structure to handle queries with $2^{i-1}\leq r<2^{i}$. Since there are at most $L+1$ such sub-data structures, the overall update time for Lemma 9.5 is $L+1=\tilde{O}(1)$ times that of a single sub-data structure, and hence asymptotically the same.
The following claim provides a data structure that (1) maintains the total weight of the dataset (a trivial task), and (2) constructs a coreset for the entire dataset, following the approach of [HK20].101010[HK20] provides a general $(1+\epsilon)$-coreset construction for $k$-means. In this paper, we only need a constant-factor coreset for $1$-means (e.g., $k=1$, $\epsilon=0.5$), so we state this special case.
Claim 9.6.
There exists a data structure that maintains a dynamic weighted set $X\subseteq[\Delta]^{d}$, and, upon query, returns the total weight of $X$ and a weighted set $Y\subseteq[\Delta]^{d}$ such that the following guarantee holds with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
$\forall c\in[\Delta]^{d},\qquad\textsc{cost}(Y,c)=\Theta(\textsc{cost}(X,c)).$
The data structure has update and query time $\mathrm{poly}(d\log\Delta)$, and it works against an adaptive adversary.
We plug the $(\Gamma,\Lambda,2^{i}\cdot\Gamma)$-hash $\varphi_{i}$ and Claim 9.6 into Lemma 9.1 and obtain a data structure $\mathcal{D}_{i}$, which has an update time of $\tilde{O}(\Lambda\cdot\mathrm{poly}(d\log\Delta))=\tilde{O}(2^{\epsilon d})$.
We now explain how to use $\mathcal{D}_{i}$ to answer a query for a point $x\in[\Delta]^{d}$ and radius $r\in[2^{i-1},2^{i})$. We query $\mathcal{D}_{i}$ with $x$, and it returns, in time $\tilde{O}(\Lambda\cdot\mathrm{poly}(d\log\Delta))=\tilde{O}(2^{\epsilon d})$, weights $w_{1},\dots,w_{t}$ and weighted sets $Y_{1},\dots,Y_{t}$, for some $1\leq t\leq\Lambda=\tilde{O}(2^{\epsilon d})$. Let $Y:=\bigcup_{j\in[t]}Y_{j}$. By construction, the size of $Y$ is bounded by the query time, which is $\tilde{O}(2^{\epsilon d})$. Finally, we perform the following computations:
  • $b_{x}\leftarrow\sum_{j\in[t]}w_{j}$. This step takes $\tilde{O}(t)=\tilde{O}(2^{\epsilon d})$ time.
  • $c^{*}\leftarrow$ the $1$-means center of $Y$. This step takes $\tilde{O}(|Y|)=\tilde{O}(2^{\epsilon d})$ time.
  • $\widehat{\textsc{cost}}_{c^{*}}\leftarrow\textsc{cost}(Y,c^{*})$ and $\widehat{\textsc{cost}}_{x}\leftarrow\textsc{cost}(Y,x)$. This step takes $\tilde{O}(|Y|)=\tilde{O}(2^{\epsilon d})$ time.
Overall, the query time is $\tilde{O}(2^{\epsilon d})$. To see the correctness, by Lemma 9.1, there exists disjoint sets $P_{1},\dots,P_{t}\subseteq X$ such that
  • $\textsc{Ball}(x,2^{i})\cap X\subseteq\bigcup_{j\in[t]}P_{j}\subseteq\textsc{Ball}(x,O(\epsilon^{-3/2})\cdot 2^{i})\cap X$, which, combined with the fact that $r\in[2^{i-1},2^{i})$, implies that $\textsc{Ball}(x,r)\cap X\subseteq\bigcup_{j\in[t]}P_{j}\subseteq\textsc{Ball}(x,O(\epsilon^{-3/2})\cdot r)\cap X$;
  • for every $j\in[t]$, $w_{j}=w(P_{j})$; and
  • for every $j\in[t]$, the following holds with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. [^11] While Lemma 9.1 does not explicitly consider success probability, its analysis can be easily adapted to handle data structures that succeed with high probability.
    $\forall c\in[\Delta]^{d},\qquad\textsc{cost}(Y_{j},c)=\Theta(\textsc{cost}(P_{j},c)).$ (19)
Let $B_{x}:=\bigcup_{j\in[t]}P_{j}$. Item 1 then follows immediately. Moreover, by a union bound, with probability at least $1-\frac{t}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, Equation 19 holds for all $j\in[t]$ simultaneously. Therefore, it is easy to verify that
$\forall c\in[\Delta]^{d},\qquad\textsc{cost}(Y,c)=\Theta(\textsc{cost}(B_{x},c)).$ (20)
Hence, the $1$-means center $c^{*}$ on $Y$ is guaranteed to be an $O(1)$-approximation to the $1$-means center on $B_{x}$ (i.e., Item 2 holds), and the values $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$ are also $O(1)$-approximations to $\textsc{cost}(B_{x},c^{*})$ and $\textsc{cost}(B_{x},x)$, respectively (i.e., Item 3 holds).
This finishes the proof. ∎
10 Approximate Assignment Structure
In this section, we design a data structure that implicitly maintains an approximate assignment from a point set $X\subseteq[\Delta]^{d}$ to a center set $S\subseteq[\Delta]^{d}$, which supports dynamic updates to both sets. We first describe the data structure in Section 10.1 and analyze its update time in Section 10.2. We then present two key applications – maintaining center weights and supporting $D^{2}$-sampling – in Section 10.3 and Section 10.4, respectively. These are critical building blocks for the restricted $k$-means subroutine in Section 11 and the augmented $k$-means subroutine in Section 12.
10.1 Data Structure Description
In this part, we describe our data structure for maintaining an approximate assignment. We first present static and dynamic objects that are explicitly maintained by our data structure (Definition 10.1). Then, we show how these explicit objects define an implicit partition of the non-center points $X-S$ (Claim 10.5). Built on top of this implicit partition, we define an implicit assignment $\sigma:X-S\to S$ (Claim 10.7).
Explicit Objects
We summarize all the explicitly maintained objects below.
Definition 10.1 (Data Structure).
Firstly, the data structure explicitly maintains the following static objects, which remain unchanged under updates to the point set $X$ and the center set $S$.
  1. 1.
    Parameters $\Gamma:=O(\epsilon^{-3/2})$, $\Lambda:=\tilde{O}(2^{\epsilon d})$, $m:=\lceil\log(\sqrt{d}\cdot\Delta\cdot\Gamma)\rceil$, and $\rho_{i}:=\frac{1}{2}\cdot(3\Gamma)^{i}$, $\forall i\in[0,m]$.
  2. 2.
    Hash functions $\{\varphi_{i}\}_{i\in[0,m]}$; we assume without loss of generality that $\varphi_{0}:[\Delta]^{d}\to\Phi$ is an injective function, which is trivially an efficient $(\Gamma,\Lambda,\rho_{0})$-hash, and every other $\varphi_{i}:[\Delta]^{d}\to\Phi$ is an efficient $(\Gamma,\Lambda,\rho_{i})$-hash (Definition 8.1)
Secondly, defining (low-level buckets) $L:=[m]\times\Phi\times\Phi$ and (high-level buckets) $H:=[m]\times\Phi$, the data structure maintains the following dynamic objects. [^12] The sizes of $L$ and $H$ are $\mathrm{poly}(\Delta^{d})$, so a full representation of the dynamic objects in Items 3, 4, 5 and 6 would be inefficient. Instead, we store only the “supported” objects, i.e., those non-empty sets.
  1. 3.
    $X(i,z):=X\cap\varphi_{i}^{-1}(z)$, for every high-level bucket $(i,z)\in H$.
  2. 4.
    $\operatorname{bk}(i,z,z^{\prime}):=X\cap\varphi_{i}^{-1}(z)\cap\varphi_{i-1}^{-1}(z^{\prime})$, for every low-level bucket $(i,z,z^{\prime})\in L$.
  3. 5.
    $S(i,z)$, a subset of close centers for every high-level bucket $(i,z)\in H$ that, if $z\in\varphi_{i}([\Delta]^{d})$, then [^13] For any hash value $z\notin\varphi_{i}([\Delta]^{d})$, we impose no requirement on the corresponding center subset $S(i,z)$.
    $\big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\tfrac{\rho_{i}}{\Gamma}\big{\}}\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ S(i,z)\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\big{\}}.$ (21)
  4. 6.
    $f(i,z):=\{z^{\prime}\in\Phi:S(i-1,z^{\prime})=\emptyset\land\operatorname{bk}(i,z,z^{\prime})\neq\emptyset\}$, for every high-level bucket $(i,z)\in H$.
  5. 7.
    $H^{\prime}:=\{(i,z)\in H:S(i,z)\neq\emptyset\}$.
  6. 8.
    $\sigma^{\prime}(i,z)\in S(i,z)$, an arbitrary close center for every high-level bucket $(i,z)\in H^{\prime}$.
Lemma 10.2 (Initialization).
The data structure of Definition 10.1 can be initialized in time $\tilde{O}(1)$ and with success probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
Proof.
Initially, both $X$ and $S$ are empty =absent=$\emptyset$= ∅, so the initialization only needs to compute the $m+1$ hash functions $\{\varphi_{i}\}_{i\in[0,m]}$. By Lemma 8.2, every $\varphi_{i}$ can be constructed in $\tilde{O}(1)$ time with a success probability of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. Thus, the total initialization time is $(m+1)\cdot\tilde{O}(1)=\tilde{O}(1)$. Moreover, taking a union bound, the overall success probability is $1-\frac{m+1}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. ∎
Implicit Partition and Assignment
Finally, we show how the explicit objects (Definition 10.1) implicitly define a partition of the non-center points $X-S$ (Claim 10.5). This partition exhibits an approximate equidistant structure, where points in each part are roughly equally distant from the center set (Claim 10.6). We further define an implicit approximate assignment based on this partition (Claim 10.7).
Prerequisites: Properties of Close Centers.
Before defining the implicit partition and the implicit assignment, we establish some useful properties of the close centers $\{S(i,z)\}_{(i,z)\in H}$ as prerequisites. We introduce the following notations for convenience.
$\forall i\in[0,m],\ \forall x\in X:$ $S(i,x):=S(i,\varphi_{i}(x))$ (22)
$\forall(i,z)\in H:$ $\operatorname{bk}(i,z):=\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})$ (23)
Claim 10.3.
For every point $x\in X$ and every high-level bucket $(i,z)\in H^{\prime}$, if $x\in\operatorname{bk}(i,z)$, then $z=\varphi_{i}(x)$, $S(i,x)\neq\emptyset$ and $S(i-1,x)=\emptyset$.
Proof.
Consider a specific pair of $x\in X$ and $(i,z)\in H^{\prime}$ such that $x\in\operatorname{bk}(i,z)$; we must have $x\in\operatorname{bk}(i,z,z^{\prime})$ for some hash value $z^{\prime}\in f(i,z)$ (Equation 23). Then,
  • $S(i,z)\neq\emptyset$, given the definition of $H^{\prime}$ (Item 7 of Definition 10.1) and that $(i,z)\in H^{\prime}\subseteq H$;
  • $S(i-1,z^{\prime})=\emptyset$, given the definition of $f(i,z)$ (Item 6 of Definition 10.1) and that $z^{\prime}\in f(i,z)$;
  • $\varphi_{i}(x)=z$ and $\varphi_{i-1}(x)=z^{\prime}$, given the definition of $\operatorname{bk}(i,z,z^{\prime})$ (Item 4 of Definition 10.1) and that $x\in\operatorname{bk}(i,z,z^{\prime})$.
Putting everything together, we conclude that $z=\varphi_{i}(x)$, $S(i,x)\neq\emptyset$ and $S(i-1,x)=\emptyset$. ∎
Claim 10.4.
For every $x\in X$ and every $i\in[m]$, it holds that $S(i-1,x)\subseteq S(i,x)$.
Proof.
Let $z_{i}:=\varphi_{i}(x)$ and $z_{i-1}:=\varphi_{i-1}(x)$. Consider a specific center $s\in S(i-1,x)$. By Equation 21, it holds that $\mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))\leq 2\rho_{i-1}$. Therefore,
$\mathrm{dist}(s,x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))+\mathrm{diam}(\varphi_{i-1}^{-1}(z_{i-1}))$ (Triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\rho_{i-1}+\rho_{i-1}$ (Item 2 of Definition 10.1)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \rho_{i}/\Gamma.$ (Item 1 of Definition 10.1)
Then, $\mathrm{dist}(s,\varphi_{i}^{-1}(z_{i}))\leq\mathrm{dist}(s,x)\leq\rho_{i}/\Gamma\implies s\in S(i,z_{i})=S(i,x)$ by Equations 21 and 22. The arbitrariness of $s\in S(i-1,x)$ implies that $S(i-1,x)\subseteq S(i,x)$. This finishes the proof. ∎
Implicit Partition.
Now we are ready to define the implicit partition and the implicit assignment, using the explicit objects maintained in Definition 10.1.
Claim 10.5 (Partition).
$\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$.
Proof.
Consider a specific point $x\in X$ and let $z_{i}:=\varphi_{i}(x)$ for every $i\in[0,m]$.
In case of a center point $x\in S$, we must have $\mathrm{dist}(x,\varphi_{0}^{-1}(z_{0}))=\mathrm{dist}(x,x)=0\leq\frac{\rho_{0}}{\Gamma}$, which implies $x\in S(0,z_{0})$ (Equation 21) and $S(i,z_{i})\neq\emptyset$ for every $i\in[0,m]$ (Claim 10.4). Then, it follows from Claim 10.3 that $x\notin\operatorname{bk}(i,z)$ for all $(i,z)\in H^{\prime}$.
In case of a non-center point $x\in X-S$, since $x\notin S$ and $\varphi_{0}$ is an injective function, we have $\mathrm{dist}(s,\varphi_{0}^{-1}(z_{0}))=\mathrm{dist}(s,x)\geq 1\geq 2\rho_{0}$ for every $s\in S$ and thus $S(0,z_{0})=\emptyset$ (Equation 21). Also, since $(3\Gamma)^{L}/\Gamma\geq\sqrt{d}\Delta=\mathrm{diam}([\Delta]^{d})$, we have $S(L,z_{L})=S\neq\emptyset$ (Equation 21). Then, by Claim 10.4, there exists a unique index $1\leq j<L$ such that $S(i,x)=\emptyset$ when $i<j$ and $S(i,x)\neq\emptyset$ when $i\geq j$. By the definition of $H^{\prime}$ (Item 7 of Definition 10.1), it follows that $(i,z_{i})\in H^{\prime}$ for every $i\geq j$. Then, by Claim 10.3, we conclude that $(j,z_{j})\in H^{\prime}$ is the unique pair such that $x\in\operatorname{bk}(j,z_{j})$.
Since $x\in X$ is chosen arbitrarily, the proof is complete. ∎
Claim 10.6 (Equidistant).
For every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$, it holds that
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
Proof.
Consider a specific high-level bucket $(i,z)\in H^{\prime}$ and a specific point $x\in\operatorname{bk}(i,z)$; we have $\varphi_{i}(x)=z$ and $S(i,z)\neq\emptyset$ (Claim 10.3) and $x\in\operatorname{bk}(i,z,z^{\prime})$ for some hash value $z^{\prime}\in f(i,z)\subseteq\Phi$ (Equation 23). Thus, $\varphi_{i-1}(x)=z^{\prime}$ (Item 4 of Definition 10.1) and $z^{\prime}\in f(i,z)\implies S(i-1,z^{\prime})=\emptyset$ (Item 6 of Definition 10.1).
Since $S(i,z)\neq\emptyset$ and $z=\varphi_{i}(x)$, by Equation 21 and the fact that $\sigma^{\prime}(i,z)\in S(i,z)$ (Item 8 of Definition 10.1), we know
[…]
Together with the fact $\mathrm{diam}(\varphi_{i}^{-1}(z))\leq\rho_{i}=\frac{1}{2}\cdot(3\Gamma)^{i}$ (Item 1 of Definition 8.1), the distance from $x$ to $S$ can be bounded as:
$\mathrm{dist}(x,S)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))+\mathrm{diam}(\varphi_{i}^{-1}(z))$ ($\varphi_{i}(x)=z$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (3\Gamma)^{i}+\frac{1}{2}\cdot(3\Gamma)^{i}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
On the other hand, since $S(i-1,z^{\prime})=\emptyset$ and $z^{\prime}=\varphi_{i-1}(x)$, by Equation 21 again, we have
$\min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\rho_{i}}{\Gamma}\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
Therefore,
$\mathrm{dist}(x,S)\leavevmode\nobreak\ =\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(x,s)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
Overall, we have
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
The arbitrariness of $(i,z)\in H^{\prime}$ and $x\in\operatorname{bk}(i,z)$ finishes the proof. ∎
Implicit Assignment.
The following Claim 10.7 provides an implicit scheme for maintaining an approximate assignment, which follows directly from Claims 10.5 and 10.6.
Claim 10.7 (Assignment).
For every $x\in X-S$, define $\sigma(x):=\sigma^{\prime}(H^{\prime}(x))$, where $H^{\prime}(x)=(i_{x},z_{x})\in H$ denotes the unique pair such that $x\in\operatorname{bk}(i_{x},z_{x})$. The assignment $\sigma:X-S\to S$ satisfies that
$\forall x\in X-S,$ $\mathrm{dist}(x,\sigma(x))\leq(3\Gamma)^{2}\cdot\mathrm{dist}(x,S).$
10.2 Update Time Analysis
In this section, we analyze the update time for the data structure of Definition 10.1.
Lemma 10.8 (Update Time).
For each update to either the point set $X$ or the center set $S$, the data structure of Definition 10.1 can be maintained deterministically in $\tilde{O}(2^{O(\epsilon d)})$ time.
Since the updates are handled deterministically, the data structure naturally works against adaptive adversaries.
In the rest of this section, we would establish Lemma 10.8 by addressing all considered dynamic explicit objects one by one. Firstly, the preimages $\{X(i,z)\}_{(i,z)\in H}$ (Item 3 of Definition 10.1) are affected only by updates to the point set $X$, and not by updates to the center set $S$, and can be maintained in a straightforward manner.
Claim 10.9 (Maintenance of Preimages).
For every update (insertion or deletion) to the point set $X$, the preimages $\{X(i,z)\}_{(i,z)\in H}$ can be maintained in time $\tilde{O}(\Lambda)$.
Proof.
Given an insertion $X\leftarrow X+x$ (resp. a deletion $X\leftarrow X-x$), we compute all hash values $z_{i}=\varphi_{i}(x)$ for $i\in[0,m]$ in total time $(m+1)\cdot\tilde{O}(\Lambda)$ (Definition 8.1) and augment $X(i,z_{i})\leftarrow X(i,z_{i})+x$ (resp. diminish $X(i,z_{i})\leftarrow X(i,z_{i})-x$) all preimages for $i\in[0,m]$ in total time $(m+1)\cdot\tilde{O}(1)$.
The total update time is $(m+1)\cdot\tilde{O}(\Lambda)+(m+1)\cdot\tilde{O}(1)=\tilde{O}(\Lambda)$. ∎
Secondly, the low-level buckets $\{\operatorname{bk}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ (Item 4 of Definition 10.1) can likewise be maintained straightforwardly.
Claim 10.10 (Maintenance of Low-Level Buckets).
For every update (insertion or deletion) to the point set $X$, the low-level buckets $\{\operatorname{bk}(\ell,z,z^{\prime})\}_{(\ell,z,z^{\prime})\in L}$ can be maintained in time $\tilde{O}(\Lambda)$.
Proof.
Given an insertion $X\leftarrow X+x$ (resp. a deletion $X\leftarrow X-x$), we compute all hash values $z_{i}=\varphi_{i}(x)$ for $i\in[0,m]$ in total time $(m+1)\cdot\tilde{O}(\Lambda)$ (Definition 8.1) and augment $\operatorname{bk}(i,z_{i},z_{i-1})\leftarrow\operatorname{bk}(i,z_{i},z_{i-1})+x$ (resp. diminish $\operatorname{bk}(i,z_{i},z_{i-1})\leftarrow\operatorname{bk}(i,z_{i},z_{i-1})-x$) all low-level buckets $(i,z_{i},z_{i-1})$ for $i\in[1,m]$ in total time $\tilde{O}(m)$. The total update time is $(m+1)\cdot\tilde{O}(\Lambda)+(m+1)\cdot\tilde{O}(1)=\tilde{O}(\Lambda)$. ∎
Thirdly, the maintenance of the center subsets $\{S(i,z)\}_{(i,z)\in H}$ (Item 5 of Definition 10.1) only responds to updates to the center set $S$ and does not depend on updates to the point set $X$.
Claim 10.11 (Maintenance of Close Centers).
For every update (insertion or deletion) to the center set $S$, the center subsets $\{S(i,z)\}_{(i,z)\in H}$ can be maintained in time $\tilde{O}(\Lambda)$.
Proof.
We consider an insertion or an deletion separately.
To insert a new center $s\in[\Delta]^{d}-S$, we perform the following for every $i\in[0,m]$: compute a subset $\Phi_{i}\subseteq\Phi$ (Item 3 of Definition 8.1) such that
$\varphi_{i}(\textsc{Ball}(s,\tfrac{\rho_{i}}{\Gamma}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{i}\cap\varphi_{i}([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(s,2\rho_{i})).$
Then, augment $S(i,z)\leftarrow S(i,z)+s$, for every $z\in\Phi_{i}$. Since every $\Phi_{i}$ can be computed in $\tilde{O}(\Lambda)$ time and has size at most $\Lambda$ (Item 3 of Definition 8.1), the entire insertion procedure runs in total time $(m+1)\cdot\tilde{O}(\Lambda)=\tilde{O}(\Lambda)$.
To verify correctness, consider a specific pair $(i,z)\in H$ such that $\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\frac{\rho_{i}}{\Gamma}\iff z\in\varphi_{i}(\textsc{Ball}(s,\frac{\rho_{i}}{\Gamma}))\subseteq\Phi_{i}$. The above procedure inserts $s$ into $S(i,z)$. The arbitrariness of $(i,z)\in H$ implies that $\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\frac{\rho_{i}}{\Gamma}\}\subseteq S(i,z)$, for every $(i,z)\in H$, after the insertion. Further, consider a specific pair $(i,z)\in H$ with $z\in\varphi_{i}([\Delta]^{d})$ such that $s$ is inserted into $S(i,z)$ by the above procedure. Then $z\in\Phi_{i}\cap\varphi_{i}([\Delta]^{d})\subseteq\varphi_{i}(\textsc{Ball}(s,2\rho_{i}))\implies\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\cdot\rho_{i}$. Therefore, the arbitrariness of $(i,z)\in H$ ensures that $S(i,z)\subseteq\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\}$, for every $(i,z)\in H$, after the insertion. Overall, the above procedure maintains Item 5 of Definition 10.1.
To delete a current center $s\in S$, we perform the following for every $i\in[0,m]$: recompute the size-$(\leq\Lambda)$ subset $\Phi_{i}\subseteq\Phi$ in $\tilde{O}(\Lambda)$ time (Item 3 of Definition 8.1) – the same as the one when $s$ was inserted into $S$ – and diminish $S(i,z)\leftarrow S(i,z)-s$, for every $z\in\Phi_{i}$. Clearly, after this procedure, $s$ is no longer contained in any $S(i,z)$, so Equation 21 is maintained, and this deletion procedure runs in total time $(m+1)\cdot\tilde{O}(\Lambda)=\tilde{O}(\Lambda)$. ∎
Fourthly, in alignment with maintenance of the low-level buckets $\{\operatorname{bk}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ and the close center sets $\{S(i,z)\}_{(i,z)\in H}$, we can synchronously maintain the high-level bucket descriptions $f(i,z)=\{(i,z,z^{\prime})\in L:S(i-1,z^{\prime})=\emptyset,\ \operatorname{bk}(i,z,z^{\prime})\neq\emptyset\}$ for all $(i,z)\in H$ (Item 6 of Definition 10.1), as summarized in the following claim.
Claim 10.12 (High-Level Buckets Description).
The following guarantees hold.
  1. 1.
    When a single low-level bucket $\operatorname{bk}(i,z,z^{\prime})$ for some $(i,z,z^{\prime})\in L$ is updated, the high-level buckets $\{f(i,z)\}_{(i,z)\in H}$ can be synchronously maintained in time $\tilde{O}(1)$.
  2. 2.
    When a single close center set $S(i,z)$ for some $(i,z)\in H$ is updated, the high-level buckets $\{f(i,z)\}_{(i,z)\in H}$ can be synchronously maintained in time $\tilde{O}(\Lambda)$.
Proof.
We prove the two items separately.
Item 1: Consider an update to $\operatorname{bk}(i,z,z^{\prime})$ for some $(i,z,z^{\prime})\in L$. By definition (Item 6 of Definition 10.1), this update only affects $f(i,z)$. Specifically, if $S(i-1,z^{\prime})\neq\emptyset$ and $\operatorname{bk}(i,z,z^{\prime})$ changes from non-empty to empty, then $z^{\prime}$ should be removed from $f(i,z)$; conversely, if $S(i-1,z^{\prime})\neq\emptyset$ and $\operatorname{bk}(i,z,z^{\prime})$ changes from empty to non-empty, then $z^{\prime}$ should be added to $f(i,z)$. Clearly, either update can be performed in $\tilde{O}(1)$ time. In all other cases, $f(i,z)$ remains unchanged.
Item 2: Consider an update to a single close center set. If it is $S(m,z^{\prime})$ for some $z^{\prime}\in\Phi$ that is updated, then all $f(i,z)$ remain unchanged (Item 6 of Definition 10.1).
Otherwise, it is $S(i-1,z^{\prime})$ for some $i\in[1,m]$ and some $z^{\prime}\in\Phi$ that is updated. Regarding a hash value $z\in\Phi$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$ (Item 6 of Definition 10.1):
If $S(i-1,z^{\prime})$ changes from empty to non-empty, then $z^{\prime}$ should be deleted from $f(i,z)$.
If $S(i-1,z^{\prime})$ changes from non-empty to empty, then $z^{\prime}$ should be added to $f(i,z)$.
In all other cases, $f(i,z)$ remains unchanged.
Hence, it remains to identify all $z\in\Phi$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, which we refer to as to-update hash values. If $X(i-1,z^{\prime})=X\cap\varphi_{i-1}^{-1}(z^{\prime})=\emptyset$, then $\operatorname{bk}(i,z,z^{\prime})=X\cap\varphi_{i}^{-1}(z)\cap\varphi_{i-1}^{-1}(z^{\prime})=\emptyset$ for every $z\in\Phi$, so there is no to-update hash values. Otherwise, we pick a specific point $x\in X(i-1,z^{\prime})$ and compute a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ with $\varphi_{i}(\textsc{Ball}(x,\frac{1}{\Gamma}\cdot(3\Gamma)^{i}))\subseteq\Phi_{x}$ in time $\tilde{O}(\Lambda)$ (Item 3 of Definition 8.1). Then, to identify a to-update hash value $z\in\Phi_{x}$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, we can consider an arbitrary point $y\in\operatorname{bk}(i,z,z^{\prime})$; given that $\varphi_{i-1}(y)=z^{\prime}=\varphi_{i-1}(x)$ (Item 4 of Definition 10.1), we can deduce $\mathrm{dist}(x,y)\leq(3\Gamma)^{i-1}$ (Item 1 of Definition 8.1) and thus
$z\leavevmode\nobreak\ =\leavevmode\nobreak\ \varphi_{i}(y)\leavevmode\nobreak\ \in\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,(3\Gamma)^{i-1}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,\tfrac{1}{\Gamma}\cdot(3\Gamma)^{i}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}.$
Namely, to identify all to-update hash values $z\in\Phi_{x}$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, it suffices to enumerate all hash values $z\in\Phi_{x}$ and check the conditions $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, which runs in time $\tilde{O}(\Lambda)$.
This completes the proof. ∎
Since each update to the point set $X$ requires $\tilde{O}(\Lambda)$ time to maintain the low-level buckets $\{\operatorname{bk}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ (Claim 10.10), and each update to the center set $S$ takes $\tilde{O}(\Lambda)$ time to maintain the center subsets $\{S(i,z)\}_{(i,z)\in H}$ (Claim 10.11), by Claim 10.12, maintaining $\{f(i,z)\}_{(i,z)\in H}$ under each update to either $X$ or $S$ takes at most $\tilde{O}(\Lambda^{2})$ time.
Fifthly, in alignment with maintenance of the close center sets $\{S(i,z)\}_{(i,z)\in H}$, we can synchronously maintain $H^{\prime}=\{(i,z)\in H:S(i,z)\neq\emptyset\}$ and $\{\sigma^{\prime}(i,z)\}_{(i,z)\in H^{\prime}}$ (Items 7 and 8 of Definition 10.1).
Claim 10.13 (Partition and Assignment).
When a single close center set $S(i,z)$ for some $(i,z)\in H$ is updated, both $H^{\prime}$ and $\{\sigma^{\prime}(i,z)\}_{(i,z)\in H^{\prime}}$ can be maintained in time $\tilde{O}(1)$.
Proof.
When a single close center set $S(i,z)$ for some $(i,z)\in H$ is updated, we proceed as follows:
  • If $S(i,z)$ changes from empty to non-empty, insert $(i,z)$ to $H^{\prime}$ and set $\sigma^{\prime}(i,z)$ to an arbitrary point in $S(i,z)$.
  • If $S(i,z)$ changes from non-empty to empty, delete $(i,z)$ from $H^{\prime}$ and discard $\sigma^{\prime}(i,z)$.
  • Otherwise, if the update to $S(i,z)$ deletes $\sigma^{\prime}(i,z)$, set $\sigma^{\prime}(i,z)$ to another point in $S(i,z)$.
Clearly, this procedure correctly maintains both $H^{\prime}$ and $\{\sigma^{\prime}(i,z)\}_{(i,z)\in H^{\prime}}$, and takes $\tilde{O}(1)$ time. ∎
Proof of Lemma 10.8.
The total update time for the data structure of Definition 10.1 follows by combining the results from Claims 10.9, 10.10, 10.11, 10.12 and 10.13. Among these, the most expensive one is maintaining the mappings $\{f(i,z)\}_{(i,z)\in H}$, which requires $\tilde{O}(\Lambda^{2})$ time. Therefore, the overall update time for the data structure is $\tilde{O}(\Lambda^{2})$. ∎
10.3 Application: Maintaining Cluster Weight
In this part, we consider the dynamic point set $X\subseteq[\Delta]^{d}$ to be weighted, with a weight function $w:X\to\mathbb{R}_{\geq 0}$. Thus, each update to $X$ corresponds to the insertion or deletion of a weighted point.
We extend the data structure of Definition 10.1 to support explicit maintenance of center weights for the dynamic center set $S\subseteq[\Delta]^{d}$ (Lemma 10.14). These center weights correspond to the cluster sizes in some an approximate clustering.
Lemma 10.14 (Size of Clusters).
There exists a data structure for handling a dynamic weighted point set $X\subseteq[\Delta]^{d}$ and a dynamic center set $S\subseteq[\Delta]^{d}$. It maintains, for every $s\in S$, a weight $w_{S}(s)\geq 0$ such that for some point-to-center assignment $\sigma:X\to S$,
  1. 1.
    $\mathrm{dist}(x,\sigma(x))\leq{O(\epsilon^{-3})}\cdot\mathrm{dist}(x,S)$, for every point $x\in X$;
  2. 2.
    $w_{S}(s)=w(\sigma^{-1}(s))$, for every center $s\in S$.
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update time $\tilde{O}(2^{\epsilon d})$.
Before presenting our construction for Lemma 10.14, we state the following corollary, which follows as a consequence of Lemmas 10.14 and 9.4, and will be useful in the analysis of restricted $k$-means (Section 11).
Corollary 10.15 (Importance of Centers).
There exists a data structure that maintains an ordering $c_{1},c_{2},\ldots,c_{|S|}$ of centers $S$ such that for each $11\leq i\leq|S|-$ 1, the following holds,
$w_{S}(c_{i})\cdot\widehat{\mathrm{dist}}^{2}(c_{i},S-c_{i})\leq w_{S}(c_{i+1})\cdot\widehat{\mathrm{dist}}^{2}(c_{i+1},S-c_{i+1}),$
where $w_{S}$ is the weight function maintained by Lemma 10.14, and for every $i\in[|S|]$, the value $\widehat{\mathrm{dist}}(c_{i},S-c_{i})$ satisfies that
$\mathrm{dist}(c_{i},S-c_{i})\leq\widehat{\mathrm{dist}}(c_{i},S-c_{i})\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(c_{i},S-c_{i}).$
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
Proof.
Since both the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ and the weights $\{w_{S}(s)\}_{s\in S}$ are explicitly maintained by Lemmas 9.4 and 10.14, respectively, the desirable ordering $c_{1},c_{2},\ldots,c_{|S|}$ of the centers $S$ of centers $S$ can be maintained synchronously, with only a $\tilde{O}(1)$-factor increase in update time. ∎
We now turn to the construction of the data structure for Lemma 10.14.
Data Structure for Lemma 10.14.
Built on top of the data structure of Definition 10.1, we additionally maintain the following explicit objects:
  • $w_{L}(i,z,z^{\prime}):=w(\operatorname{bk}(i,z,z^{\prime}))=\sum_{x\in\operatorname{bk}(i,z,z^{\prime})}w(x)$, for every $(i,z,z^{\prime})\in L$.
  • $w_{H}(i,z):=\sum_{z^{\prime}\in f(i,z)}w_{L}(i,z,z^{\prime})=\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$, for every $(i,z)\in H$.
  • $w_{S}(s):=\sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(i,z)+w(s)$,141414 If $s\notin X$, we define $w(s)=0$. for every center $s\in S$.
By Lemma 10.2, the data structure of Definition 10.1 can be initialized with success probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$; we henceforth assume that this initialization succeeds.
Update Time.
Firstly, the values $\{w_{L}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ can be maintained synchronously with $\{\operatorname{bk}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ with only $\tilde{O}(1)$ overhead: whenever a point $x$ is inserted into (resp. deleted from) $\operatorname{bk}(i,z,z^{\prime})$ for some $(i,z,z^{\prime})\in L$, we increase (resp. decrease) $w_{L}(i,z,z^{\prime})$ by $w(x)$.
Secondly, the values $\{w_{H}(i,z)\}_{(i,z)\in H}$ can be maintained synchronously with $\{f(i,z)\}_{(i,z)\in H}$ and $\{w_{L}(i,z,z^{\prime})\}_{(i,z,z^{\prime})\in L}$ with only $\tilde{O}(1)$ overhead: whenever a value $z^{\prime}$ is inserted into (resp. deleted from) $f(i,z)$ for some $(i,z)\in H$, we increase (resp. decrease) $w_{H}(i,z)$ by $w_{L}(i,z,z^{\prime})$; whenever $w_{L}(i,z,z^{\prime})$ is updated for some $(i,z,z^{\prime})\in L$, we also update $w_{H}(i,z)$ accordingly if $z^{\prime}\in f(i,z)$.
Finally, the weights $\{w_{S}(s)\}_{s\in S}$ can be maintained synchronously with $H^{\prime}$ and $\{\sigma^{\prime}(i,z)\}_{(i,z)\in H}$ with only $\tilde{O}(1)$ overhead: once a pair $(i,z)$ is inserted into (resp. deleted from) $H^{\prime}$, we increase (resp. decrease) $w_{S}(\sigma^{\prime}(i,z))$ by $w_{H}(i,z)$; once $\sigma^{\prime}(i,z)$ changes from, say, $s$ to $s^{\prime}$, we update $w_{S}(s)\leftarrow w_{S}(s)-w_{H}(i,z)$ and $w_{S}(s^{\prime})\leftarrow w_{S}(s^{\prime})+w_{H}(i,z)$.
Overall, the update time of the data structure for Lemma 10.14 remains asymptotically the same as that of the data structure in Definition 10.1, which is $\tilde{O}(\Lambda^{2})=\tilde{O}(2^{O(\epsilon d)})$ by Lemma 10.8. Rescaling $\epsilon$ yields the update time $\tilde{O}(2^{\epsilon d})$ as in Lemma 10.14.
Correctness.
We construct an assignment from $X$ to $S$ as follows: Let $\sigma:X-S\to S$ be the assignment as defined in Claim 10.7; we naturally extend this definition by setting $\sigma(x):=x$ for every $x\in X\cap S$. Clearly, this construction gives us a full assignment from $X$ to $S$. We then show that the weights $\{w_{S}(s)\}_{s\in S}$ and this assignment $\sigma\colon X\to S$ satisfy both Items 1 and 2 of Lemma 10.14.
Item 1: For a center point $x\in X\cap S$, we have $\sigma(x)=x\in S$ by definition, so Item 1 follows directly. For a non-center point $x\in X-S$, we have $\sigma(x)=\sigma^{\prime}(H^{\prime}(x))$ by definition, so Item 1 follows directly from Claim 10.7 and that $\Gamma=O(\epsilon^{-3/2})$.
Item 2: Observe that for every $(i,z)\in H$, the low-level buckets $\{\operatorname{bk}(i,z,z^{\prime})\}_{z^{\prime}\in\Phi}$ are disjoint (Item 4 of Definition 10.1). Therefore,
$w_{H}(i,z)=\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))=w\Big{(}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\Big{)}=w(\operatorname{bk}(i,z)).$
where the last step follows from Equation 23. Now, for a specific center $s\in S$, we can write
$w_{S}(s)$ $\leavevmode\nobreak\ \leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(s)+w(s)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w(\operatorname{bk}(i,z))+w(s)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}\sum_{x\in\operatorname{bk}(i,z)}w(s)+w(s).$
Since $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$ (Claim 10.5) and $\sigma(x)=\sigma^{\prime}(i,z)$ for every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$ (Claim 10.7). It then follows that
$w_{S}(s)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X-S:\sigma(x)=s}w(x)+w(s)=\sum_{x\in X:\sigma(x)=s}w(x).$
This finishes the proof of Item 2.
10.4 Application: $D^{2}$-Sampling
In this part, we extend the data structure of Definition 10.1 to support $D^{2}$-sampling over the dynamic weighted point set $X\subseteq[\Delta]^{d}$ with respect to the dynamic center set $S\subseteq[\Delta]^{d}$.
Corollary 10.16 ($D^{2}$-Sampling).
There exists a data structure for handling a dynamic weighted point set $X\subseteq[\Delta]^{d}$ and a dynamic center set $S\subseteq[\Delta]^{d}$. At any time, it can generate independent random points from $X$ such that each random point $y\in X$ is generated in time $\tilde{O}(1)$ and satisfies that
$\forall x\in X:$ $\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{-6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$ (24)
The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update time $\tilde{O}(2^{\epsilon d})$.
Proof.
By Lemma 10.2, the data structure of Definition 10.1 can be initialized with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$; we henceforth assume that this initialization succeeds.
Built on top of the data structure of Definition 10.1, the data structure for Corollary 10.16 additionally maintain the following explicit objects:
  • $w_{L}(i,z,z^{\prime}):=w(\operatorname{bk}(i,z,z^{\prime}))=\sum_{x\in\operatorname{bk}(i,z,z^{\prime})}w(x)$, for every $(i,z,z^{\prime})\in L$.
  • $w_{H}(i,z):=(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w_{L}(i,z,z^{\prime})=\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$, for every $(i,z)\in H$.
Similar to the analysis for Lemma 10.14, the data structure for Corollary 10.16 has an update time of $\tilde{O}(2^{\epsilon d})$. The remainder of this proof is devoted to implementing the sampling procedure, which proceeds in three steps.
  1. 1.
    First, we sample a high-level bucket $(i,z)$ from $H^{\prime}$ with probability proportional to $w_{H}(i,z)$.
    Since both the support $H^{\prime}$ and the weights $\{w_{H}(i,z)\}_{(i,z)\in H^{\prime}}$ are explicitly maintained, this can be implemented by storing the weighted set $H^{\prime}$ in, e.g., a binary search tree, allowing for $\tilde{O}(1)$-time sampling and incurring only $\tilde{O}(1)$ overhead in the update time.
  2. 2.
    Second, we sample a value $z^{\prime}$ from $f(i,z)$ with probability proportional to $w_{L}(i,z,z^{\prime})$. Again, since both the support $f(i,z)$ and the weights $\{w_{L}(i,z,z^{\prime})\}_{z^{\prime}\in f(i,z)}$ are explicitly maintained, this step also takes $\tilde{O}(1)$ time.
  3. 3.
    Third, we sample a point $y$ from $\operatorname{bk}(i,z,z^{\prime})$ with probability proportional to $w(y)$. Again, since the weighted set $\operatorname{bk}(i,z,z^{\prime})$ is explicitly maintained (Item 4 of Definition 10.1), this step can also be implemented in $\tilde{O}(1)$ time.
Hence, the overall sampling time is $\tilde{O}(1)$.
It remains to show the correctness of the procedure. Clearly, the final sampled point $y$ satisfies
$y\leavevmode\nobreak\ \in\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\operatorname{bk}(i,z)\leavevmode\nobreak\ =\leavevmode\nobreak\ X-S.$ (25)
Consider a specific point $x\in X$. If $x\in S$ is a center, then by Equation 25, we have $\Pr[y=x]=0$. Since $\mathrm{dist}(x,S)=0$, Equation 24 trivially holds in this case. If $x\notin S$ is not a center, then there exists a unique pair $(i_{x},z_{x})\in H^{\prime}$ such that $x\in\operatorname{bk}(i_{x},z_{x})=\bigcup_{z^{\prime}\in f(i_{x},z_{x})}\operatorname{bk}(i,z,z^{\prime})$, since the collection $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ forms a partition of $X-S$ (Claim 10.5). Moreover, since the low-level buckets $\{\operatorname{bk}(i_{x},z_{x},z^{\prime})\}_{z^{\prime}\in f(i_{x},z_{x})}$ are disjoint (Item 4 of Definition 10.1), there exists a unique value $z_{x}^{\prime}\in f(i_{x},z_{x})$ such that $x\in\operatorname{bk}(i_{x},z_{x},z_{x}^{\prime})$. Thus, the point $x$ is sampled with probability
$\Pr[y=x]$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{w_{H}(i_{x},z_{x})}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}\cdot\frac{w_{L}(i_{x},z_{x},z^{\prime}_{x})}{\sum_{z^{\prime}\in f(i_{x},z_{x})}w_{L}(i_{x},z_{x},z^{\prime})}\cdot\frac{w(x)}{w(\operatorname{bk}(i_{x},z_{x},z^{\prime}_{x}))}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{(3\Gamma)^{2i_{x}}\cdot w(x)}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}.$
By Claim 10.6, the numerator $(3\Gamma)^{2i_{x}}\geq\Omega(1)\cdot\mathrm{dist}^{2}(x,S)$, and the denominator
$\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}w(\operatorname{bk}(i,z))\cdot(3\Gamma)^{2i}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}w(x)\cdot(3\Gamma)^{2i}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}O(\Gamma^{4})\cdot w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.6)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\sum_{x\in X-S}w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.5)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\textsc{cost}(X,S).$
Therefore, we conclude that
$\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\Gamma^{4})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$
This finishes the proof. ∎
11 Restricted $k$-Means
In this section, we consider the restricted $k$-means problem: Given a weighted point set $X\subseteq[\Delta]^{d}$, a center set $S\subseteq[\Delta]^{d}$, and an integer $r\geq 1$ as input, the goal is to select a size-$r$ subset $R\subseteq S$ that minimizes the $k$-means objective $\textsc{cost}(X,S-R)$.
Assuming that the weighted point set $X$ and the center set $S$ are maintained by the data structures previously discussed (specifically Corollaries 10.15, 9.2 and 10.14), we design a subroutine that computes an approximate restricted $k$-means solution in sublinear time.
Lemma 11.1 (Subroutine for Restricted $k$-Means).
Given access to the data structures from Corollaries 10.15, 9.2 and 10.14, there exists a subroutine that, on input an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
where $\mathrm{OPT}_{|S|-r}^{S}(X):=\min_{S^{\prime}\subseteq S:|S^{\prime}|=|S|-r}\textsc{cost}(X,S^{\prime})$.
Recall that all data structures from Corollaries 10.15, 9.2 and 10.14 support dynamic updates to both $X$ and $S$ against an adaptive adversary, each with update time $\tilde{O}(2^{\epsilon d})$. By packing them into a single data structure together with the subroutine in Lemma 11.1, we obtain Lemma 7.5 stated in Section 7.
Proof of Lemma 11.1.
We describe the subroutine in Algorithm 4. We then analyze its running time and correctness, which together establish Lemma 11.1.
Algorithm 4 Approximating restricted $k$-means.
1:Let $T_{1}\leftarrow\{c_{i}\}_{i\in[6r]}\subseteq S$ be the first $6r$ centers in the ordering maintained by Corollary 10.15.
2:Let $T_{2}\leftarrow\{s_{i}\}_{i\in[6r]}\subseteq S-T_{1}$ be such that $\forall i\in[6r]$, $\mathrm{dist}(c_{i},s_{i})\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(c_{i},S-T_{1})$.
3:Compute a size-$r$ subset $\widehat{R}\subseteq T:=T_{1}+T_{2}$ such that $\triangleright$ $T\subseteq S$.
$\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-\widehat{R})\leq O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-R),$
where $w_{S}:S\to\mathbb{R}_{\geq 0}$ is the weight function maintained by Lemma 10.14.
4:\$\widehat{R}$.
Running Time.
We analyze the running time of Algorithm 4 line by line.
1 takes trivially $\tilde{O}(r)$ time, since the entire ordering $c_{1},\dots,c_{|S|}$ is explicitly maintained by Corollary 10.15.
2 can be implemented in time $\tilde{O}(r\cdot 2^{\epsilon d})$: We temporarily delete all centers in $T_{1}$ from $S$ and update the data structure from Lemma 9.2. By querying this data structure for every center $c_{i}\in T_{1}$, we can obtain $s_{i}$ in time $\tilde{O}(2^{\epsilon d})$, thus constructing $T_{2}=\{s_{i}\}_{i\in[6r]}$ in time $\tilde{O}(r\cdot 2^{\epsilon d})$. Finally, we insert all centers from $T_{1}$ back into $S$ and update the data structure from Lemma 9.2.
3 can be implemented in time $\tilde{O}(r^{1+\epsilon})$: Since the weight function $w_{S}:S\to\mathbb{R}_{\geq 0}$ is explicitly maintained by Lemma 10.14, we can construct the weighted set $T=T_{1}\cup T_{2}$, where every point $s\in T$ has weight $w_{S}(s)$, in time $\tilde{O}(r)$. Then, to compute $\widehat{R}$, it suffices to run an $(|T|-r)$-means algorithm on the weighted set $T$, with centers selected from $T$, and obtain a solution $\widehat{S}\subseteq T$ with $|\widehat{S}|=|T|-r$. Then, we return $\widehat{R}\leftarrow T-\widehat{S}$ as the desired output; for this, we can apply the static algorithm from [JJLL25], which takes $\tilde{O}(r^{1+\epsilon})$ time.
Notice that each step has a success probability (either from the data structures or the static algorithm) of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$; therefore, by the union bound, all steps succeed simultaneously with high probability, resulting in a total running time of $\tilde{O}(2^{\epsilon d}\cdot r+r^{1+\epsilon})$.
Correctness Analysis.
It remains to establish the correctness of Algorithm 4. Specifically, we aim to show that the subset $\widehat{R}\subseteq S$ computed in 3 satisfies
$\textsc{cost}(X,S-\widehat{R})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$ (26)
We first have the following claim, which establishes a sufficient condition for Equation 26. Intuitively, this condition states that the optimal value of the restricted $k$-means objective on the weighted set $T$ is at most a $\mathrm{poly}(1/\epsilon)$ factor larger than that on the weighted set $S\supseteq T$.
Claim 11.2.
Equation 26 holds if
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$
We defer the proof of Claim 11.2 to Section 11.1. Claim 11.2 allows us to focus on the restricted problem for the weighted sets $S$ and $T$. From now on, for any center set $C\subseteq[\Delta]^{d}$, we denote the clustering objective of $C$ on $S$ and $T$ by $\textsc{cost}(S,C)=\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,C)$ and $\textsc{cost}(T,C)=\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,C)$, respectively. Hence, the sufficient condition stated in Claim 11.2 becomes
$\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R).$ (27)
Let $R^{*}:=\mathrm{argmin}_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ denote the optimal solution to the restricted $k$-means on $S$. The following Claim 11.3 shows the existence of an approximate solution $\widehat{R}^{*}$ by selecting points solely from $T_{1}\subseteq T$. Moreover, Claim 11.4 guarantees that this solution yields a comparable cost when evaluated on both $T$ and $S$.
Claim 11.3.
There exists a size-$r$ subset $\widehat{R}^{*}\subseteq T_{1}$ such that
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
Claim 11.4.
For any size-$r$ set $R\subseteq T_{1}$, it holds that
$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
We defer the proof of Claims 11.3 and 11.4 to Sections 11.2 and 11.4, respectively.
We now show how these two claims together imply Equation 27. Let $\widehat{R}^{*}\subseteq T_{1}$ be a size-$r$ subset that satisfies $\textsc{cost}(S,S-\widehat{R}^{*})\leq O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 11.3). Then, Equation 27 follows from the derivation below.
LHS of Equation 27 $=\leavevmode\nobreak\ \min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
$\leq\leavevmode\nobreak\ \textsc{cost}(T,T-\widehat{R}^{*})$ ($\widehat{R}^{*}\subseteq T_{1}\subseteq T$, $|\widehat{R}^{*}|=r$)
$\leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-\widehat{R}^{*})$ (Claim 11.4)
$\leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 11.3)
$=\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Definition of $R^{*}$)
[…]
By Claim 11.2, this completes the correctness analysis of Algorithm 4. ∎
11.1 Proof of Claim 11.2
Claim 11.2 (Restated).
Equation 26 (i.e., $\textsc{cost}(X,S-\widehat{R})\leq O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)$) holds if
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$ (28)
Proof.
Recall the property of the weight function $w_{S}$ from Lemma 10.14: There exists an assignment $\sigma:X\to S$ such that (i) $\forall x\in X$, $\mathrm{dist}(x,\sigma(x))\leq O(\epsilon^{-3})\cdot\mathrm{dist}(x,S)$; and (ii) $\forall s\in S$, $w_{S}(s)=w(\sigma^{-1}(s))$. Therefore,
$\textsc{cost}(X,S-\widehat{R})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot\mathrm{dist}^{2}(x,S-\widehat{R})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot(\mathrm{dist}(x,\sigma(x))+\mathrm{dist}(\sigma(x),S-\widehat{R}))^{2}$ (triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(\sigma(x),S-\widehat{R}))$ (AM-GM inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(X,S)+2\cdot\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$ (Properties (i) and (ii))
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+2\cdot\textsc{cost}(S,S-\widehat{R}).$
The last step applies $\textsc{cost}(X,S)\leq\mathrm{OPT}_{|S|-r}^{S}(X)$ and $\textsc{cost}(S,S-\widehat{R})=\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$. Since every center $s\in S-\widehat{R}$ has a zero distance to $S-\widehat{R}$, we have
$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($T\subseteq S$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($\widehat{R}\subseteq T$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$ (3 of Algorithm 4)
Let $R^{*}_{X}\subseteq S$ be the optimal solution to the restricted $k$-means on $X$, i.e., $|R^{*}_{X}|=r$ and $\textsc{cost}(X,S-R^{*}_{X})=\mathrm{OPT}_{|S|-r}^{S}(X)$. It follows that
$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Equation 28)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\textsc{cost}(S,S-R^{*}_{X})$ ($R^{*}_{X}\subseteq S$, $|R^{*}_{X}|=r$)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,s)+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\textsc{cost}(X,S)+O(\epsilon^{-7})\cdot\textsc{cost}(X,S-R^{*}_{X})$ (Property (i))
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
Putting everything together, we obtain Equation 26:
$\textsc{cost}(X,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
This finishes the proof. ∎
11.2 Proof of Claim 11.3
Claim 11.3 (Restated).
There exists a size-$r$ subset $\widehat{R}^{*}\subseteq T_{1}$ such that
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
The Construction of $\widehat{R}^{*}$.
Recall the definition of $R^{*}=\mathrm{argmin}_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$. To construct the desired approximate solution $\widehat{R}^{*}\in T_{1}$, we need the following technical claim, whose proof is deferred to Section 11.3.
Claim 11.5.
There exists a size-$r$ subset $T^{\prime}_{1}\subseteq T_{1}$ such that the following guarantees hold:
  1. 1.
    $T^{\prime}_{1}\cap R^{*}=\emptyset$.
  2. 2.
    For every $s\in R^{*}$, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})=\mathrm{dist}(s,S-R^{*})$.
  3. 3.
    For every $s\in T^{\prime}_{1}$, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})\leq 2\cdot\mathrm{dist}(s,S-s)$.
Our construction of $\widehat{R}^{*}$ based on the subset $T^{\prime}_{1}\subseteq T_{1}$ from Claim 11.5 proceeds in the following three steps.
  1. 1.
    Let $R^{*}_{1}:=R^{*}\cap T_{1}$ and $R^{*}_{2}:=R^{*}-T_{1}$.
  2. 2.
    Since $|R^{*}_{2}|\leq|R^{*}|=r$, we can pick an arbitrary size-$|R^{*}_{2}|$ subset of $T^{\prime}_{1}$, denoted by $\widehat{R}^{*}_{2}$.
  3. 3.
    Define $\widehat{R}^{*}:=R^{*}_{1}+\widehat{R}^{*}_{2}$.
Correctness of $\widehat{R}^{*}$.
Since $R^{*}_{1}\subseteq T_{1}$ and $\widehat{R}^{*}_{2}\subseteq T^{\prime}_{1}\subseteq T_{1}$, we have $\widehat{R}^{*}=R^{*}_{1}+\widehat{R}^{*}_{2}\subseteq T_{1}$. By Item 1 of Claim 11.5, we have $\widehat{R}^{*}_{2}\cap R^{*}_{1}=\emptyset$. Since $|\widehat{R}^{*}_{2}|=|R^{*}_{2}|$, it then follows that $|\widehat{R}^{*}|=|R^{*}_{1}|+|\widehat{R}^{*}_{2}|=|R^{*}_{1}|+|R^{*}_{2}|=r$. Hence, it remains to show that
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$ (29)
To this end, observe that
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$
$=\sum_{s\in\widehat{R}^{*}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$
$=\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})+\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*}).$
Similarly, we can obtain
$\textsc{cost}(S,S-R^{*})\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*})+\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$
As a result, we can further divide the proof of Equation 29 into the following two parts:
$\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}),$ (30)
$\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$ (31)
Equation 30: Consider a specific center $s\in R^{*}_{1}$. By Item 2 of Claim 11.5, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})=\mathrm{dist}(s,S-R^{*})$. Since $\widehat{R}^{*}=R^{*}_{1}+\widehat{R}^{*}_{2}\subseteq R^{*}+T^{\prime}_{1}$, it follows that $s^{\prime}\in S-\widehat{R}^{*}$. Therefore,
$\mathrm{dist}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,s^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathrm{dist}(s,S-R^{*}).$
The arbitrariness of $s\in R^{*}_{1}$ implies Equation 30.
Equation 31: Let $t:=|R^{*}_{2}|=|\widehat{R}^{*}_{2}|$, and write $R^{*}_{2}=\{r^{*}_{1},\dots,r^{*}_{t}\}$, $\widehat{R}^{*}_{2}=\{{\widehat{r}}^{*}_{1},\dots,{\widehat{r}}^{*}_{t}\}$. Recall that $T_{1}=\{c_{1},\dots,c_{6r}\}\subseteq S$ (1 of Algorithm 4) is the first $6r$ centers in the ordering maintained by Corollary 10.15. Since $\widehat{R}^{*}_{2}\subseteq T^{\prime}_{1}\subseteq T_{1}$ and $R^{*}_{2}=R^{*}-T_{1}\implies R^{*}_{2}\cap T_{1}=\emptyset$, it follows that
$\forall i\in[t]:$ $w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i}).$
Now consider a fixed $i\in[t]$. We can deduce from Corollary 10.15 that
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-r^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
Since ${\widehat{r}}^{*}_{i}\in\widehat{R}^{*}_{2}\subseteq T^{\prime}_{1}$, by Item 3 of Claim 11.5, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}({\widehat{r}}^{*}_{i},s^{\prime})\leq 2\cdot\mathrm{dist}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$. Recall that $\widehat{R}^{*}=R^{*}_{1}+\widehat{R}^{*}_{2}\subseteq R^{*}+T^{\prime}_{1}$. It follows that $s^{\prime}\in S-\widehat{R}^{*}$ and thus
$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-\widehat{R}^{*})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},s^{\prime})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 4\cdot w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
The arbitrariness of $i\in[t]$ implies Equation 31.
This finishes the proof of Claim 11.3.
11.3 Proof of Claim 11.5
Claim 11.5 (Restated).
There exists a size-$r$ subset $T^{\prime}_{1}\subseteq T_{1}$ such that the following guarantees hold:
  1. 1.
    $T^{\prime}_{1}\cap R^{*}=\emptyset$.
  2. 2.
    For every $s\in R^{*}$, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})=\mathrm{dist}(s,S-R^{*})$.
  3. 3.
    For every $s\in T^{\prime}_{1}$, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})\leq 2\cdot\mathrm{dist}(s,S-s)$.
Items 1 and 2.
By removing the following centers from $T_{1}$, we can ensure that Items 1 and 2 hold for every remaining subset $T^{\prime}_{1}\subseteq T_{1}$:
  • Let $T_{1}\leftarrow T_{1}-R^{*}$. It is then clear that Item 1, i.e., $T^{\prime}_{1}\cap R^{*}=\emptyset$, holds for any subset $T^{\prime}_{1}\subseteq T_{1}$.
  • For every $s\in R^{*}$, select an arbitrary $s^{\prime}\in S-R^{*}$ such that $\mathrm{dist}(s,s^{\prime})=\mathrm{dist}(s,S-R^{*})$, and let $T_{1}\leftarrow T_{1}-s^{\prime}$. This ensures that $s^{\prime}\in S-(R^{*}+T_{1})$, which implies Item 2, i.e., $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ for any subset $T^{\prime}_{1}\subseteq T_{1}$.
Clearly, the above process will remove at most $2\cdot|R^{*}|=2r$ centers from $T_{1}$, thus leaving at least $6r-2r=4r$ centers still in $T_{1}$ (1 of Algorithm 4).
Item 3.
Based on the above discussion, we can assume that $|T_{1}|\geq 4r$ and that Items 1 and 2 hold for every subset $T^{\prime}_{1}\subseteq T_{1}$. We now show the existence of a size-$r$ subset $T^{\prime}_{1}\subseteq T_{1}$ that also satisfies Item 3. To this end, let us consider the exact nearest neighbor assignment $g:T_{1}\to S$, i.e., $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$ for every $s\in T_{1}$. We then proceed with a case study.
Case 1: $|g(T_{1})|\geq 3r$.
We construct a desired subset $T^{\prime}_{1}\subseteq T_{1}$ as follows. First, consider the set $T^{\prime}_{1}=\bigcup_{y\in g(T_{1})-R^{*}}g^{-1}(y)$. This set has size
[…]
Further, for every $s\in T^{\prime}_{1}$, it holds that $g(s)\in g(T_{1})-R^{*}=S-R^{*}$ and $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$. We then apply the following claim from [Tho04].
Claim 11.6 ([Tho04, Lemma 21]).
Given a set $A$, suppose that every element $a\in A$ has at most one successor $s(a)\in A$. Then, we can mark at least $|A|/2$ elements such that, if $a$ is marked, then $s(a)$ is unmarked.
Applying Claim 11.6, we can mark at least $|T^{\prime}_{1}|/2\geq r$ centers in $T^{\prime}_{1}$ such that, for every marked $s\in T^{\prime}_{1}$, $g(s)$ is unmarked; we then remove all unmarked centers from $T^{\prime}_{1}$. Afterward, we still have $|T^{\prime}_{1}|\geq r$, and for every $s\in T^{\prime}_{1}$, $g(s)\in S-(R^{*}+T^{\prime}_{1})$; namely, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}(s,s^{\prime})=\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)\leq 2\cdot\mathrm{dist}(s,S-s)$. Therefore, any size-$r$ subset of $T^{\prime}_{1}$ satisfies Item 3.
Case 2: $|g(T_{1})|<3r$.
Let $t:=|g(T_{1})|<3r$ and write $g(T_{1})=\{g_{1},\dots,g_{t}\}$. We construct a desired subset $T^{\prime}_{1}\subseteq T_{1}$ as follows: First, we partition the subset $T_{1}$ into groups $\{g^{-1}(g_{i})\}_{i\in[t]}$. For every $i\in[t]$, we define $s^{\prime}_{i}:=\mathrm{argmin}_{s\in g^{-1}(g_{i})}\mathrm{dist}(s,g_{i})$. Then, the subset $T^{\prime}_{1}$ is defined as
$T^{\prime}_{1}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \bigcup_{i\in[t]}(g^{-1}(g_{i})-s^{\prime}_{i}).$ (32)
Clearly, we have $|T^{\prime}_{1}|\geq|T_{1}|-t>4r-3r=r$. Now consider a specific center $s\in T^{\prime}_{1}$; we have $g(s)=g_{i}$ for some $i\in[t]$. We will show that $s^{\prime}_{i}\in S-(R^{*}+T^{\prime}_{1})$ and $\mathrm{dist}(s,s^{\prime}_{i})\leq 2\cdot\mathrm{dist}(s,S-s)$, thus proving Item 3 for this $s\in T^{\prime}_{1}$.
Since $g(s)=g_{i}$, it follows that $\mathrm{dist}(s,g_{i})=\mathrm{dist}(s,S-s)$. Additionally, by the definition of $s^{\prime}_{i}$, we have $\mathrm{dist}(s^{\prime}_{i},g_{i})\leq\mathrm{dist}(s,g_{i})$. Hence, by applying the triangle inequality, we have
[…]
Since $s^{\prime}_{i}\in T_{1}\subseteq S$, it remains to show that $s^{\prime}_{i}\notin R^{*}+T^{\prime}_{1}$: Recall that $T_{1}$ already satisfies (Item 1) $T_{1}\cap R^{*}=\emptyset$, and thus $s^{\prime}_{i}\in T_{1}\implies s^{\prime}_{i}\notin R^{*}$. Moreover, by definition (Equation 32), $s^{\prime}_{i}\notin T^{\prime}_{1}$. In conclusion, Item 3 holds for this $s\in T^{\prime}_{1}$. The arbitrariness of $s\in T^{\prime}_{1}$ implies Item 3.
This completes the proof of Claim 11.5
11.4 Proof of Claim 11.4
Claim 11.4 (Restated).
For any size-$r$ set $R\subseteq T_{1}$, it holds that
$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
Proof.
Consider a specific size-$r$ subset $R\subseteq T_{1}$. Since $T_{1}\cap T_{2}=\emptyset$ (1 and 2 of Algorithm 4), we have $T-R=(T_{1}-R)+T_{2}$ and $S-R=(S-T_{1})+(T_{1}-R)$. Therefore, we have
$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\ \mathrm{dist}^{2}(s,T_{2})\}.$
Now, for every center $s\in R\subseteq T_{1}$, by 2 of Algorithm 4, there exists $s^{\prime}\in T_{2}$ such that $\mathrm{dist}(s,T_{2})\leq\mathrm{dist}(s,s^{\prime})\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(s,S-T_{1})$. Hence, it follows that
$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\mathrm{dist}^{2}(s,T_{2})\}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),O(\epsilon^{-3})\cdot\mathrm{dist}^{2}(s,S-T_{1})\}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
This finishes the proof. ∎
12 Augmented $k$-Means
In this section, we consider the augmented $k$-means problem: Given a weighted point set $X\subseteq[\Delta]^{d}$, a center set $S\subseteq[\Delta]^{d}$ and a parameter $a\geq 1$ as input, the goal is to compute a size-$a$ subset $A\subseteq[\Delta]^{d}$ that minimizes the $k$-means objective $\textsc{cost}(X,S+A)$.
Lemma 12.1 (Augmented $k$-Means).
Given access to the data structure of Corollary 10.16, there exists a subroutine that, on input integer $a\geq 1$, computes a size-$\Theta(a\cdot\epsilon^{-6}\cdot d\log\Delta)$ subset $A\subseteq[\Delta]^{d}$ in time $\tilde{O}(2^{\epsilon d}\cdot a)$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S+A)\leq O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d},|A|=a}\textsc{cost}(X,S+A^{\prime}).$
Recall that the data structure of Corollary 10.16 support dynamic updates to both $X$ and $S$ against an adaptive adversary, with update time $\tilde{O}(2^{\epsilon d})$. Hence, Lemma 12.1 directly implies Lemma 7.6, as stated in Section 7.
Without loss of generality, we assume throughout this section that $|X|>\Omega(a\cdot\epsilon^{-6}\cdot\log n)$ is sufficiently large, as otherwise the algorithm could simply return the entire dataset $X$ as the output.
We present the subroutine for Lemma 12.1 in Algorithm 5, and establish its running time and correctness in Claim 12.2 and Claim 12.3, respectively. Together, these results complete the proof of Lemma 12.1.
Algorithm 5 Approximating augmented $k$-Means.
[…]
2:Sample $t:=\Theta(\epsilon^{-6}\cdot d\log\Delta)$ i.i.d. points $\{x_{i}\}_{i\in[t]}$ from $X$, such that $\Pr[x_{i}=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}$ for every $y\in X$.
[…]
Claim 12.2 (Fast Implementation of Algorithm 5).
Given access to the data structure of Corollary 10.16, we can implement Algorithm 5 in time $\tilde{O}(n^{\epsilon}\cdot a)$.
Proof.
We modify Algorithm 5 as follows: in each of the $a+1$ iterations, instead of only adding the points $\{x_{i}\}_{i\in[t]}$ (sampled in 2) to the set of augmented centers $A$, we also add them to the center set $S$ and update the data structure of Corollary 10.16. This update takes a total of $\tilde{O}(2^{\epsilon d})$ time per iteration. With this modification, 2 becomes equivalent to sampling points $x\in X$ with probability at least $\Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}(x,S)}{\textsc{cost}(X,S)}$. By querying the data structure of Corollary 10.16 $t$ times, we obtain $t$ independent samples in total time $\tilde{O}(2^{\epsilon d})\cdot t=\tilde{O}(2^{\epsilon d})$. Hence, the total running time for the entire for loop is $\tilde{O}(2^{\epsilon d}\cdot a)$. After the for loop completes, we perform an additional step: we remove the augmented centers $A$ from $S$ and update the data structure from Corollary 10.16 accordingly. This step takes $|A|\cdot\tilde{O}(2^{\epsilon d})=\tilde{O}(2^{\epsilon d}\cdot a)$ time.
Therefore, the overall running time of this implementation is $\tilde{O}(2^{\epsilon d})$. ∎
Clearly, Algorithm 5 returns a set $A$ of size $\tilde{O}(a\cdot\epsilon^{-6}\cdot d\log\Delta)$. We now state the following claim regarding the correctness of $A$.
Claim 12.3 (Correctness of Algorithm 5).
Let $A$ denote the (random) output of Algorithm 5. Then, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$,
$\textsc{cost}(X,S+A)\leq 16\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
The proof of Claim 12.3 follows almost directly from [ADK09, Theorem 1], where a similar sampling procedure is analyzed to obtain a bi-criteria approximation to the vanilla clustering problem. For completeness, we provide a proof in Section 12.1.
12.1 Proof of Claim 12.3
Effect of $D^{2}$-sampling on reducing bad clusters
A crucial step in the analysis is to understand the behavior of the (distorted) $D^{2}$-sampling procedure, as described in 2 of Algorithm 5 – namely, sampling a random point $x\in X$ such that
$\forall y\in X,\quad\Pr[x=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}.$ (33)
To this end, we define a potential function for any center set $A$ that augments $S$: the number of bad clusters. This quantity will play a key role in the proof.
Let $A^{*}=\{c_{1},\dots,c_{a}\}$ denote the optimal set of $a$ centers that minimizes $\textsc{cost}(X,S+A^{*})$. Let
$X_{0}:=\{x\in X:\mathrm{dist}(x,S)\leq\mathrm{dist}(x,A^{*})\},$ (34)
and let $\{X_{1},\dots,X_{a}\}$ be the clustering of $X-X_{0}$ with respect to $A^{*}=\{c_{1},\dots,c_{a}\}$. Then,
$\forall j\in[a],\forall x\in X_{j},\quad\mathrm{dist}(x,c_{j})=\mathrm{dist}(x,S+A^{*}).$ (35)
The bad clusters with respect to $A$, denoted by $\mathcal{B}(A)$, are those whose clustering objective under $S+A$ remains large. Formally:
$\mathcal{B}(A):=\left\{X_{j}:\textsc{cost}(X_{j},S+A)>8\cdot\textsc{cost}(X_{j},S+A^{*})\right\}.$ (36)
We also define the good clusters as $\mathcal{G}(A):=\{X_{1},\dots,X_{a}\}\setminus\mathcal{B}(A)$. Throughout the proof, we slightly abuse notation by writing $j\in\mathcal{B}(A)$ to refer to the index $j$ such that $X_{j}\in\mathcal{B}(A)$.
We establish the following claim regarding the bad clusters.
Claim 12.4.
For any center sets $A,A^{\prime}$ with $A\subseteq A^{\prime}$, it holds that $|\mathcal{B}(A)|\leq a$ and $\mathcal{B}(A)\supseteq\mathcal{B}(A^{\prime})$.
Proof.
The bound $|\mathcal{B}(A)|\leq a$ follows directly from the definition of bad clusters (see Equation 36). To see that $\mathcal{B}(A)\supseteq\mathcal{B}(A^{\prime})$, consider any $X_{j}\in\mathcal{B}(A^{\prime})$. Since $A\subseteq A^{\prime}$, we have $\textsc{cost}(X_{j},S+A)\geq\textsc{cost}(X_{j},S+A^{\prime})>8\cdot\textsc{cost}(X_{j},S+A^{*})$, and hence $X_{j}\in\mathcal{B}(A)$ as well. ∎
The following claim shows that if all clusters are good, then the augmented center set $S+A$ achieves a good approximation:
Claim 12.5.
Let $A$ be a center set. If $\mathcal{B}(A)=\emptyset$, then $\textsc{cost}(X,S+A)\leq 8\cdot\textsc{cost}(X,S+A^{*})$.
Proof.
By Equation 34, we have $\textsc{cost}(X_{0},S+A)\leq\textsc{cost}(X_{0},S)=\textsc{cost}(X_{0},S+A^{*})$. Since all clusters are good, by Equation 36, we have $\textsc{cost}(X_{j},S+A)\leq 8\cdot\textsc{cost}(X_{j},S+A^{*})$ for all $j\in[a]$. Therefore,
$\textsc{cost}(X,S+A)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A)+\sum_{j=1}^{a}\textsc{cost}(X_{j},S+A)$
$\leq\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A^{*})+\sum_{j=1}^{a}8\cdot\textsc{cost}(X_{j},S+A^{*})$
$\leq\leavevmode\nobreak\ 8\cdot\textsc{cost}(X,S+A^{*}).$
The claim follows. ∎
Next, we investigate the behavior of $D^{2}$-sampling, and show its effectiveness in reducing the number of bad clusters, as stated below.
Claim 12.6 (Effect of $D^{2}$-Sampling on Reducing Bad Clusters).
Suppose that a center set $A$ satisfies $\textsc{cost}(X,S+A)>16\cdot\textsc{cost}(X,S+A^{*})$, then a random point $x\in X$ whose distribution satisfies Equation 33 leads to $|\mathcal{B}(A+x)|\leq|\mathcal{B}(A)|-1$ with probability at least $\Omega(\epsilon^{6})$.
Proof.
For every bad cluster $X_{j}\in\mathcal{B}(A)$ – its center $c_{j}$ satisfies Equation 35 – we define
$\widehat{X}_{j}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \left\{y\in X_{j}:\ \mathrm{dist}^{2}(y,c_{j})\leq\frac{2}{w(X_{j})}\cdot\textsc{cost}(X_{j},c_{j})\right\}.$ (37)
A random point $x\in X$ sampled via $D^{2}$-sampling (2 of Algorithm 5) turns out to satisfy:
  1. (a)
    Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]\geq\Omega(\epsilon^{6})$.
  2. (b)
    If $x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$, then $\textsc{cost}(X_{j},x)\leq 8\cdot\textsc{cost}(X_{j},S+A^{*})$.
We first leverage Items (a) and (b) to establish Claim 12.6: Suppose that the random point $x\in X$ satisfies $x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$ – this occurs with probability at least $\Omega(\epsilon^{6})$ (Item (a)) – then $\textsc{cost}(X_{j},S+A+x)\leq\textsc{cost}(X_{j},x)\leq 8\cdot\textsc{cost}(X_{j},S+A^{*})$ (Item (b)), so every such bad cluster $j\in\mathcal{B}(A)$ regarding $A$ becomes a good cluster $j\notin\mathcal{B}(A+x)$ regarding $A+x$ (Equation 36). And since $\mathcal{B}(A+x)\subseteq\mathcal{B}(A)$ (Claim 12.4), we conclude with $|\mathcal{B}(A+x)|\leq|\mathcal{B}(A)|-1$.
Item (a). According to the $D^{2}$-sampling procedure (2 of Algorithm 5), we have
Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]$ $\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}\Bigg{.}}_{(\dagger)}\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\Bigg{.}}_{(\ddagger)}.$
It remains to show that $(\dagger)=\Omega(1)$ and $(\ddagger)=\Omega(1)$.
For the term ()($\dagger$)( † ), we deduce that
(($\dagger$)( † ) = $\leavevmode\nobreak\ =\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A)+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}$
$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A^{*})+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}8\cdot\textsc{cost}(X_{j},S+A^{*})}{16\cdot\textsc{cost}(X,S+A^{*})}$
$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}.$
Here the second step applies a combination of $\textsc{cost}(X_{0},S+A^{*})=\textsc{cost}(X_{0},S)\geq\textsc{cost}(X_{0},S+A)$ (Equation 34), $\textsc{cost}(X_{j},S+A)\leq 8\cdot\textsc{cost}(X_{j},S+A^{*})$ when $j\neq 0\land j\notin\mathcal{B}(A)$ (Equation 36), and $\textsc{cost}(X,S+A)>16\cdot\textsc{cost}(X,S+A^{*})$ (the premise of Claim 12.6).
For the term ()($\ddagger$)( ‡ ), let us consider a specific bad cluster $j\in\mathcal{B}(A)$. Let $\alpha:=\sqrt{w(X_{j})}\cdot\mathrm{dist}(c_{j},S+A)$ and $\beta:=\sqrt{\textsc{cost}(X_{j},c_{j})}$ for brevity. We can upper-bound $\textsc{cost}(X_{j},S+A)$ using these parameters.
$\textsc{cost}(X_{j},S+A)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}\mathrm{dist}(y,c_{j})+\mathrm{dist}(c_{j},S+A)\big{)}^{2}$ (triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(c_{j},S+A)\big{)}$ (AM-GM inequality)
= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 2\beta^{2}+2\alpha^{2}.$
This together with $\textsc{cost}(X_{j},S+A)\geq 8\cdot\textsc{cost}(X_{j},c_{j})=8\beta^{2}$ (Equation 36) also gives $\alpha\geq\sqrt{3}\beta$.
Also, we can lower-bound $\textsc{cost}(\widehat{X}_{j},S+A)$ analogously. First, following the triangle inequality and Equation 37, we have $\mathrm{dist}(y,S+A)\geq\mathrm{dist}(c_{j},S+A)-\mathrm{dist}(y,c_{j})\geq(\alpha-\sqrt{2}\beta)/\sqrt{w(X_{j})}$, for every point $y\in\widehat{X}_{j}$. Second, following Equation 37 and a simple average argument, we have $w(X_{j}-\widehat{X}_{j})\leq\frac{w(X_{j})}{2}\leq w(\widehat{X}_{j})$. Altogether, we deduce that
$\textsc{cost}(\widehat{X}_{j},S+A)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{y\in\widehat{X}_{j}}w(y)\cdot\mathrm{dist}^{2}(y,S+A)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}\cdot(\alpha-\sqrt{2}\beta)^{2}.$
These two observations in combination with arbitrariness of $j\in\mathcal{B}(A)$ implies that
$(\ddagger)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\left(\alpha-\sqrt{2}\beta\right)^{2}}{4\beta^{2}+4\alpha^{2}}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{(\sqrt{3}-\sqrt{2})^{2}}{4+4\cdot(\sqrt{3})^{2}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{5-2\sqrt{6}}{16}.$
Here the last step follows from elementary algebra; specifically, the equality holds when $\alpha=\sqrt{3}\beta$.
Item (b). If $x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$, then the triangle and AM-GM inequalities together give
$\textsc{cost}(X_{j},x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(x,c_{j})\big{)}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\cdot\textsc{cost}(X_{j},c_{j})+4\cdot\textsc{cost}(X_{j},c_{j})$ (Equation 37)
= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 6\cdot\textsc{cost}(X_{j},S+A^{*}).$ (Equation 35)
This finishes the proof of Claim 12.6. ∎
Proof of Claim 12.3.
For every $i\in[a+1]$, let $A_{i}$ denote the center set $A$ after the $i$-th iteration of the for loop (1, 2 and 3) in Algorithm 5. Also, let $A_{0}:=\emptyset$. We define the following two events for every $i\in[a+1]$:
  • Event $\mathcal{E}_{i}$: $\textsc{cost}(X,S+A_{i-1})\leq 16\cdot\textsc{cost}(X,S+A^{*})$.
  • Event $\mathcal{F}_{i}$: $|\mathcal{B}(A_{i})|\leq|\mathcal{B}(A_{i-1})|-1$.
We prove Claim 12.3 in two steps.
Step 1: $\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})\implies$ Claim 12.3.
Conditional on $\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})$, at least one event $\mathcal{E}_{i}$ some $i\in[a+1]$ must occur; otherwise, we immediately see a contradiction $|\mathcal{B}(A_{a+1})|\leq|\mathcal{B}(A_{0})|-(a+1)\leq a-(a+1)=-1$ to the nonnegativity of $|\mathcal{B}(A_{a+1})|$.
Consider an arbitrary index $i\in[a+1]$ such that event $\mathcal{E}_{i}$ occurs. Algorithm 5 guarantees that $A_{i-1}\subseteq A_{a+1}$, and hence
$\textsc{cost}(X,S+A_{a+1})\leq\textsc{cost}(X,S+A_{i-1})\leq 16\cdot\textsc{cost}(X,S+A^{*}).$
Since $A_{a+1}$ is the output of Algorithm 5, this completes the proof of Claim 12.3.
Step 2: $\Pr[\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})]\geq 1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
It remains to lower-bound the probability of this desirable case:
$\Pr\big{[}\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})\big{]}=1-\Pr\big{[}\lor_{i\in[a+1]}(\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i})\big{]}\geq 1-\sum_{i\in[a+1]}\Pr\big{[}\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big{]},$ (38)
where the last step applies the union bound. Then, consider a fixed $i\in[a+1]$. We have
$\Pr\big{[}\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Pr\big{[}\neg\mathcal{F}_{i}\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Pr\big{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}.$
Recall that $A_{i}=A_{i-1}+\{x_{1},\dots,x_{t}\}$, where $x_{1},\dots,x_{t}$ are the independent points sampled in the $i$-th iteration by 2. According to Claim 12.6, we have
$\forall j\in[t],\quad\Pr\big{[}|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq 1-\Omega(\epsilon^{6}).$
Let $t=O(\epsilon^{-6}\cdot d\log(\Delta))$ be sufficiently large, we have
$\Pr\big{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq\left(1-\Omega(\epsilon^{6})\right)^{t}\leq\frac{1}{\mathrm{poly}(\Delta^{d})}.$
For every $j\in[t]$, by Claim 12.4, we have $\mathcal{B}(A_{i-1}+x_{j})\supseteq\mathcal{B}(A_{i})$. Hence, if $|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1$, then it must be that $\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1$. As a result,
$\Pr[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}]\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \Pr\bigg{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
$\leq\leavevmode\nobreak\ \Pr\bigg{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
$\leq\leavevmode\nobreak\ \frac{1}{\mathrm{poly}(\Delta^{d})}.$
Finally, we apply the union bound over all $i\in[a+1]$, plug the result into Equation 38, and obtain $\Pr[\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})]\geq 1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
This finishes the proof. ∎
Part III Our Dynamic Algorithm
Roadmap.
In this part of the paper, we provide our algorithm, its implementation, and its analysis. We start by explaining the framework of [BCF25a] in Section 13, with some major modifications that are necessary to ensure that we can implement it using our geometric data structures (see Part II). Next, in Section 14, we show how to implement this modified framework in Euclidean spaces with an update time of $\tilde{O}(n^{\epsilon})$. Finally, in Sections 15, 16, 17 and 18, we provide the analysis of the algorithm.
Theorem 12.7.
For every sufficiently small $\epsilon>0$, there is a randomized dynamic algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio, $\tilde{O}(n^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse. If the length of the update sequence is polynomially bounded by $n$ (the maximum number of input points at any time), then the approximation, update time and recourse guarantees of the algorithm hold with probability $1-1/\text{poly}(n)$ across the entire sequence of updates.
Later, in Section 19, we improve the update time of the algorithm to $\tilde{O}(k^{\epsilon})$ by considering Theorem 12.7 as a black box. We do this in a way that the probability of the guarantees of the algorithm remains $1-1/\mathrm{poly}(n)$. Hence, our final result is as follows.
Theorem 12.8.
For every sufficiently small $\epsilon>0$, there is a randomized dynamic algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio, $\tilde{O}(k^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse. If the length of the update sequence is polynomially bounded by $n$ (the maximum number of input points at any time), then the approximation, update time and recourse guarantees of the algorithm hold with probability $1-1/\text{poly}(n)$ across the entire sequence of updates.
Remark 12.9.
If we use the Johnson-Lindenstrauss transform [JL84] (see the paragraph before Section 13 for more details) to reduce the dimension $d$ to $O(\log n)$, the algorithm works against an oblivious adversary. Otherwise, if we are not using the Johnson-Lindenstrauss transform, then our dynamic algorithm works against an adaptive adversary, at the cost of incurring an exponential in $d$ factor in the update time.
Notation.
From this point forward, we assume that $p_{\epsilon}$ is a polynomial on $1/\epsilon$ which is an upper bound for all of the polynomials in Section 7. As a result, the guarantee of every one of these data structures and subroutines holds for the fixed polynomial $p_{\epsilon}$. If we fix the value of $\epsilon$, we can first compute the value of $p_{\epsilon}$ and then the parameters of the algorithm are tuned w.r.t. this value. We have a widely used parameter $\lambda$ in the algorithm that is set to $\lambda=(p_{\epsilon})^{2}$.
$p_{\epsilon}$ : A polynomial of $1/\epsilon$ as an upper bound on all the approximation guarantees
[…]
$\lambda$ $=(p_{\epsilon})^{2}.$ (39)
Throughout the rest of the paper, we use notation $\tilde{O}$ to hide $\mathrm{poly}(d\epsilon^{-1}\log(n\Delta))$. $d$ is the dimension of the space. $n$ is the maximum size of the space, and $[0,\Delta]$ is the range of the coordinates of the points. Note that the aspect ratio of the space would be $\sqrt{d}\Delta$ since we only consider points in the grid $[\Delta]^{d}$.
Preprocessing.
Throughout Sections 14, 15, 16, 17 and 18, we assume that the dimension of the input space $d$ is $O(\log n)$. Hence, the running time performance of all the data structures in Part II becomes $\tilde{O}(n^{\epsilon})$. The assumption $d=O(\log n)$ can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84]. Assume that the input space is $\mathbb{R}^{m}$. We can initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$, and whenever we have an insertion of a point $x\in\mathbb{R}^{m}$, we apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$. Then, we consider feeding this new dynamic space into our algorithm. Note that according to this transformation, we will have an additive $\tilde{O}(m)$ in the update time of the algorithm, but we omit to write this down from now on since it only appears because of this preprocessing step, and the rest of the algorithm will not depend on that.
13 The Algorithmic Framework of [BCF25a]
Roadmap.
In this section, we explain how the algorithmic framework of [BCF25a] works, with some significant modifications that are necessary for our purpose. We start by defining the modified version of the main tool used in the algorithm in Section 13.1, which is called the robust solution. Then, we continue with the description of the algorithm in Section 13.2.
Remark 13.1.
Although the algorithm of [BCF25a] is primarily designed for $k$-median, it also works for $k$-means. The only consideration is that the analysis of [BCF25a] depends on the fact that the integrality gap of the LP-relaxation for $k$-median is constant. Since this fact is also correct for $k$-means (see Theorem 1.3 in [ANFSW20]), we can extend the analysis of [BCF25a] for $k$-means. The constant factors in the analysis might change, and here we provide a thorough analysis of the algorithm specifically for Euclidean $k$-means.
13.1 Robust Solution
Before we explain the algorithm of [BCF25a], we start with the necessary definitions for describing the algorithm. Since our goal is to achieve an update time of $\tilde{O}(n^{\epsilon})$, we slightly changed some of the structural definitions of [BCF25a] that enable us to work with approximate balls around points instead of exact balls.
Definition 13.2.
Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$. We call this sequence $t$-robust w.r.t. the current data set $X$, if there exists a sequence $(B_{0},B_{1},\cdots,B_{t})$ of subsets of $X$ such that the following hold,
  1. 1.
    For each $0\leq i\leq t$, we have $\textsc{Ball}(x_{i},\lambda^{3i})\subseteq B_{i}\subseteq\textsc{Ball}(x_{i},\lambda^{3i+1})$.
  2. 2.
    For each $1\leq i\leq t$, at least one of the following hold,
    • Either
      $\frac{\textsc{cost}(B_{i},x_{i})}{w(B_{i})}\geq\lambda^{6i-4}\ \text{and}\ x_{i-1}=x_{i},$ (40)
    • Or
      $\frac{\textsc{cost}(B_{i},x_{i})}{w(B_{i})}\leq\lambda^{6i-2}\ \text{and}\ \textsc{cost}(B_{i},x_{i-1})\leq\min\{(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{i}),\ \textsc{cost}(B_{i},x_{i})\}.$ (41)
Note that both of these cases can happen simultaneously, and in the second case, $x_{i-1}$ can be equal to $x_{i}$ as well. Moreover, we call a point $x$, $t$-robust if there exists a $t$-robust sequence $(x_{0},x_{1},\ldots,x_{t})$ such that $x=x_{0}$.
Definition 13.3.
We call a center set $S$, robust w.r.t. $X$ if and only if each center $u\in S$ satisfies the following
$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}.$ (42)
13.2 Description of The Algorithm
In this section, we describe the algorithm framework of [BCF25a]. For simplicity. Later in Section 14, we provide our algorithm using this framework in Euclidean spaces that has a small update time of $\tilde{O}(n^{\epsilon})$. Note that since we are considering new Definition 13.2 and we will implement subroutines of the algorithm approximately, the parameters of the algorithm here change significantly. The constant values in different parts of the algorithm are extremely tied together. For simplicity, in this section, we use big $O$ notation to describe their algorithm, which makes it easier to read. We will provide the exact parameters in Section 14 and the proper proofs in Section 15.
The algorithm works in epochs; each epoch lasts for some consecutive updates in $X$. Let $S\subseteq X$ denote the maintained solution (set of $k$ centers). We satisfy the following invariant.
Invariant C.
At the start of an epoch, the set $S$ is robust and $\textsc{cost}(S,X)\leq O(1)\cdot\mathrm{OPT}_{k}(X)$.
We now describe how the dynamic algorithm works in a given epoch, in four steps.
Step 1: Determining the length of the epoch. At the start of an epoch, the algorithm computes the maximum $\ell^{\star}\geq 0$ such that $\mathrm{OPT}_{k-\ell^{\star}}(X)\leq c\cdot\mathrm{OPT}_{k}(X)$, and sets $\ell\leftarrow\lfloor\ell^{\star}/(\Theta(c))\rfloor$, where $c=O(1)$ is a large constant. The epoch will last for the next $\ell+1$ updates. [^15] Note that it is very well possible that $\ell=0$. From now on, consider the superscript $t\in[0,\ell+1]$ to denote the status of some object after the algorithm has finished processing the $t^{th}$ update in the epoch. For example, at the start of the epoch, $X=X^{(0)}$.
Step 2: Preprocessing at the start of the epoch. Let $S_{\textnormal{{init}}}\leftarrow S$ be the solution maintained by the algorithm after it finished processing the last update in the previous epoch. Before handling the very first update in the current epoch, the algorithm initializes the maintained solution by setting
$S^{(0)}\leftarrow\arg\min_{S^{\prime}\subseteq S_{\textnormal{{init}}}\,:\,|S^{\prime}|=k-\ell}\textsc{cost}(X^{(0)},S^{\prime}).$ (43)
Step 3: Handling the updates within the epoch. Consider the $t^{th}$ update in the epoch, for $t\in[1,\ell+1]$. The algorithm handles this update in a lazy manner, as follows. If the update involves the deletion of a point from $X$, then it does not change the maintained solution, and sets $S^{(t)}\leftarrow S^{(t-1)}$. In contrast, if the update involves the insertion of a point $p$ into $X$, then it sets $S^{(t)}\leftarrow S^{(t-1)}+p$.
Step 4: Post-processing at the end of the epoch. After the very last update in the epoch, the algorithm does some post-processing, and computes another set $S_{\textnormal{{final}}}\subseteq X$ of at most $k$ centers (i.e., $|S_{\textnormal{{final}}}|\leq k$) that satisfies Invariant C. Then the next epoch can be initiated with $S\leftarrow S_{\textnormal{{final}}}$ being the current solution. The post-processing is done as follows.
The algorithm adds $O(\ell+1)$ extra centers to the set $S_{\textnormal{{init}}}$, while minimizing the cost of the resulting solution w.r.t. $X^{(0)}$.
$A^{\star}\leftarrow\arg\min\limits_{\begin{subarray}{c}A\subseteq[\Delta]^{d}:|A|\leq O(\ell+1)\end{subarray}}\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+A),\text{ and }S^{\star}\leftarrow S_{\textnormal{{init}}}+A^{\star}.$ (44)
The algorithm then adds the newly inserted points within the epoch to the set of centers, so as to obtain the set $T^{\star}$. Next, the algorithm identifies the subset $W^{\star}\subseteq T^{\star}$ of $k$ centers that minimizes the $k$-means objective w.r.t. $X^{(\ell+1)}$.
$T^{\star}\leftarrow S^{\star}+\left(X^{(\ell+1)}-X^{(0)}\right),\text{ and }W^{\star}\leftarrow\arg\min\limits_{\begin{subarray}{c}W\subseteq T^{\star}\,:\,|W|=k\end{subarray}}\textsc{cost}(X^{(\ell+1)},W).$ (45)
Finally, it calls a subroutine referred to as Robustify on $W^{\star}$ and lets $S_{\textnormal{{final}}}$ be the set of $k$ centers returned by this subroutine. The goal of calling Robustifyis to make the final solution $S_{\textnormal{{final}}}$, robust w.r.t. the current space $X=X^{(\ell+1)}$. We will provide our implementation of this subroutine in Section 14.2. Finally, before starting the next epoch, the algorithm sets $S\leftarrow S_{\textnormal{{final}}}$.
$S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}(W^{\star}),\ \text{and}\ S\leftarrow S_{\textnormal{{final}}}.$ (46)
It is shown that the maintained set $S\subseteq[\Delta]^{d}$ is always of size at most $k$, and at the final step, the solution $S=S_{\textnormal{{final}}}$ satisfies Invariant C at the end of Step 4, that validates the analysis of the next epoch.
14 Our Dynamic Algorithm
Roadmap.
In this section, we provide a complete description of our dynamic algorithm. In Section 14.1, we implement different steps of the framework from Section 13. Then, we proceed with the implementation of the very last subroutine of Step 4 of the algorithm in Section 14.2, which is called Robustify.
14.1 Implementing Steps of an Epoch
In this section, we provide the implementation of different steps of the algorithmic framework (see Section 13) except the very last subroutine Robustify. We provide the implementation for a fixed epoch that lasts for $(\ell+1)$ updates.
Implementing Step 1.
We find an estimation of $\ell^{\star}$ instead of the exact value. We define the exact value of $\ell^{\star}$ to be the largest integer satisfying the following
$\mathrm{OPT}_{k-\ell^{\star}}(X^{(0)})\leq\lambda^{44}\cdot\mathrm{OPT}_{k}(X^{(0)}).$ (47)
Now, we proceed with the procedure of estimating $\ell^{\star}$. For each $i\in[0,\log_{2}k]$ define $s_{i}:=2^{i}$, and let $s_{-1}:=0$. We now run a for loop, as described below.
Algorithm 6 Computing an estimate $\widehat{\ell}$ of the value of $\ell^{\star}$.
[…]
After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where
^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot\lambda^{51}}\right\rfloor.$
With some extra calculations, with the same argument as in [BCF25a], we can show that $\ell+1=\tilde{\Omega}(\widehat{\ell}+1)=\tilde{\Omega}(\ell^{\star}+1)$ and
$\frac{\mathrm{OPT}_{k-{\ell}}(X^{(0)})}{\lambda^{51}}\leq\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+{\ell}}(X^{(0)}).$
The running time of this for loop is $\tilde{O}\left(n^{\epsilon}\cdot\sum_{j=0}^{i^{\star}}s_{j}\right)=\tilde{O}(n^{\epsilon}\cdot s_{i^{\star}})$, where $i^{\star}$ is the index s.t. $s_{i^{\star}-1}=\widehat{\ell}$. Thus, we have $s_{i^{\star}}=O(\widehat{\ell})=O(\ell+1)$, and hence we can implement Step 1 in $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time.
Implementing Step 2.
Instead of finding the optimum set of $(k-\ell)$ centers within $S_{\textnormal{{init}}}$, we approximate it using restricted $k$-means (Lemma 7.5). We compute a set of $(k-\ell)$ centers $S^{(0)}\subseteq S_{\textnormal{{init}}}$ such that
$\textsc{cost}(X^{(0)},S^{(0)})\leq p_{\epsilon}\cdot\mathrm{OPT}_{k-\ell}^{S_{\textnormal{{init}}}}(X^{(0)}).$ (48)
Note that the running time for this step is also $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$.
Implementing Step 3.
Trivially, we can implement each of these updates in constant time.
Implementing Step 4.
We first need to add $O(\ell+1)$ centers to $S_{\textnormal{{init}}}$, while minimizing the cost of the solution w.r.t. $X^{(0)}$. Instead of finding $A^{\star}$ (see Equation 44), we approximate it by augmented $k$-means (Lemma 7.6). By setting $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$ and $s=\lambda^{52}\cdot(\ell+1)$ in augmented $k$-means (Lemma 7.6), we can compute $A$ of size at most $\tilde{O}(s)$, in $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time such that
$\textsc{cost}(X,S_{\textnormal{{init}}}+A)\leq p_{\epsilon}\cdot\min\limits_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|\leq s}\textsc{cost}(X,S_{\textnormal{{init}}}+A^{\star}),\quad s=\lambda^{52}\cdot(\ell+1)$ (49)
At this stage, we compute $S^{\prime}=S_{\textnormal{{init}}}+A$ and $T^{\prime}:=S^{\prime}+\left(X^{(\ell+1)}-X^{(0)}\right)$ which is an approximation of $T^{\star}$ as in Equation 45, in only $\tilde{O}(\ell+1)$ time. Next, we compute $W^{\prime}$ which is an approximation of $W^{\star}$ (see Equation 45) using restricted $k$-means (Lemma 7.5), which again takes $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time, and satisfies
$\textsc{cost}(X^{(\ell+1)},W^{\prime})\leq p_{\epsilon}\cdot\mathrm{OPT}_{k}^{T^{\prime}}(X^{(\ell+1)}).$ (50)
Finally, we explain below how we implement the call to $\texttt{Robustify}(W^{\prime})$ (see Equation 46).
14.2 Implementing Robustify
In this section, we provide the implementation of the subroutine called Robustify, which is the very last subroutine called in an epoch. The main purpose of this subroutine is to make the solution $W^{\prime}$ at the end of the epoch, robust (see Definition 13.3) w.r.t. the current dataset.
Notation.
Throughout this section, we refer to $S$ as the main center set maintained in our algorithm and refer to $W^{\prime}$ as the center set computed in Step 4 of the epoch. So, if we say we maintain a data structure for $S$, we mean that we maintain it in the entire algorithm not a specific epoch.
Outline.
Robustify can be divided into two parts. First, identifying the centers that might violate Condition (42). Next, making those centers robust w.r.t the current space.
In our implementation, for each center $u\in S$ in our main solution, we always maintain an integer $t[u]$ that indicates the center $u$ is $t[u]$-robust, i.e. we want to maintain the variant that $t[u]$ is always greater than or equal to the smallest integer $t$ satisfying $\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}$ (see Condition (42)). Hence, to check whether a center $u$ satisfies Condition (42), it suffices to check whether $t[u]\geq t$ [^16] Note that there might be a center $u$ such that $t[u]<t$, but $u$ is still $t$-robust. More precisely, we maintain an approximation of $\mathrm{dist}(u,S-u)$ using Lemma 7.3 and find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$.
After identifying a center $u$, where $t[u]<t$, we use the $p_{\epsilon}$ approximation of $\mathrm{dist}(u,S-u)$ (using Lemma 7.3), and find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{7}$ and swap $u$ with a near center $v$ which is $t$-robust through a call to subroutine $\texttt{MakeRobust}(u)$ introduced in Section 14.2.2. Note that here the denominator is $\lambda^{7}$ instead of $\lambda^{10}$, which is intentional and make $u$ to be more robust that what we need in Condition (42).
In Section 14.2.1, we explain how to identify the centers that might violate Condition (42). Then in Section 14.2.2, we explain how to swap a center $u$ with a near center $v$ which is $t$-robust for a given $t\geq 0$.
14.2.1 Identifying Non-Robust Centers
At the end of each epoch, we partition the center set $W^{\prime}$ into three parts $W^{\prime}=W_{1}\cup W_{2}\cup W_{3}$. $W_{1}$ consists of the new centers added to the solution, i.e. $W_{1}:=W^{\prime}-S_{\textnormal{{init}}}$. We call a center $u$ contaminated, if throughout the epoch, the ball of radius $\lambda^{3t[u]+1.5}$ around $u$ is changed, i.e.
$\textsc{Ball}(u,\lambda^{3t[u]+1.5})\cap\left(X^{(0)}\oplus X^{(\ell+1)}\right)\neq\emptyset.$
$W_{2}$ consists of all contaminated centers. $W_{3}$ then is the rest of the centers $W_{3}:=W^{\prime}-(W_{1}\cup W_{2})$. We will show the following claim. The proof of this claim is deferred to Section 20.10.
Claim 14.1.
Each center $u\in W_{3}$ remains $t[u]$-robust w.r.t. the final space $X^{(\ell+1)}$.
The above claim shows that if a center $u$ is in $W_{3}$, we have the guarantee that $u$ is still $t[u]$-robust. The only centers that we do not know about their robustness are centers in $W_{1}\cup W_{2}$.
As a result, we do the following. First, we call MakeRobust on all centers $u\in W_{1}\cup W_{2}$ and set their $t[u]$ value. After that, every center $u$ in our current solution is $t[u]$-robust, and in order to figure out if they violate Condition (42), as described before, it is sufficient to check if $t[u]\geq t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,W^{\prime}-u)/\lambda^{10}$.
Bottleneck.
Since we need an amortized update time of $\tilde{O}(n^{\epsilon})$, it is not efficient to iterate over all centers $u$, find $t$ and check if $t[u]\geq t$. Instead, we need a new way to figure out if a center $u$ satisfies $t[u]\geq t$. It is also non-trivial how to find the contaminated centers very fast.
We will use Lemma 7.3 and Lemma 7.2 to identify a superset of centers that contains all centers that might violate Condition (42). The size of this superset is small enough such that the amortized recourse of our algorithm remains $\tilde{O}(1)$.
Identifying $W_{1}$.
It is obvious how to find $W_{1}$ explicitly since $W_{1}=W^{\prime}-S_{\textnormal{{init}}}$. We only need to keep track of the newly added centers during the epoch which can be done in $\tilde{O}(W^{\prime}-S_{\textnormal{{init}}})$ time. According to the implementation, we have $W^{\prime}-S_{\textnormal{{init}}}=\tilde{O}(\ell+1)$, which concludes we can find $W_{1}$ in a total of $\tilde{O}(1)$ time in amortization.
Identifying $W_{2}$.
We will show that each update within an epoch can contaminate at most $O(\log(\sqrt{d}\Delta))$ many centers. More precisely, we show something stronger as follows. We defer the proof of this claim to Section 20.11.
Claim 14.2.
For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$, there is at most one center $u\in S_{\textnormal{{init}}}$ such that $t[u]=i$ and $\mathrm{dist}(x,u)\leq\lambda^{3i+2}$.
This claim provides us with a way to find contaminated centers efficiently. We split our center set $S$ into $\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil+1$ parts. For each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$, we keep track of centers $u\in S$ with $t[u]=i$ and maintain a subset $S[i]\subseteq S$ that contains all centers $u\in S$ satisfying $t[u]=i$. For each set $S[i]$, we invoke the nearest neighbor oracle (Lemma 7.1) that given $x\in[\Delta]^{d}$, returns a $u\in S[i]$ such that $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$. We will show that $S[i]$ has a special structure, and this approximate nearest neighbor is the only center that can be contaminated by $x$ within $S[i]$ at the end of each epoch. More precisely, if $x$ contaminates a center $v$, then $v$ must be an exact nearest neighbor and there does not exist any other $p_{\epsilon}$-approximate nearest neighbor within $S[i]$ (all other centers in $S[i]$ have distance more than $p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$ to $x$). So, we show the following claim. The proof of this claim is deferred to Section 20.12.
Claim 14.3.
For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$, if $u\in S[i]$ satisfies $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$, then the only center in $S[i]$ that might be contaminated by $x$ is $u$.
This claim shows how to identify $W_{2}$. For each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$, we call the nearest neighbor oracle on $S[i]$ to find $u\in S[i]$ satisfying $\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$. Then we check if $\mathrm{dist}(x,u)\leq\lambda^{3i+1.5}$. If this happens, we see that $x$ contaminates $u$, otherwise $x$ does not contaminate any center in $S[i]$ according to Claim 14.3. Hence, we can identify all centers in $W_{2}$.
Make Centers in $W_{1}$ and $W_{2}$ Robust.
Until now, we found $W_{1}$ and $W_{2}$ explicitly. We made a call to MakeRobust on each center in $W_{1}$ and $W_{2}$ to make them robust. As a result, at this moment in time, we have a set of centers and each $u$ has an integer $t[u]$ that indicates $u$ is $t[u]$-robust w.r.t. the current dataset. This is because we called a MakeRobust explicitly on centers in $W_{1}$ and $W_{2}$, and for the rest of the center we use Claim 14.1. Now, using these integers, we proceed with an efficient way to identify centers violating Condition (42).
Identifying Other Centers Violating Condition (42).
We introduce a way to identify the centers that might violate Condition (42). The goal is to not miss any single center violating Condition (42). So, it is possible that our procedure reports a center $u$ that does not violate Condition (42), but if a center $u$ violates Condition (42), our procedure definitely reports that center. Note that after each change in the center set $S$, the distances of centers $u\in S$ to other centers might change. As a result, a call to MakeRobust on $u$ might make another center $v\in S$ to violate Condition (42). We show that although the distances between centers will change after each change in $S$, in the current call to Robustify, we will not call MakeRobust on a center twice, which means if we call MakeRobust on a center, it will continue to satisfy (42) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
Lemma 14.4.
Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop. Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
We provide the proof of this lemma in Section 20.9.
Now, we proceed with the explanation of the subroutine. For every $i\in[0,\lceil\log_{\lambda}(\sqrt{d}\Delta)\rceil]$, we maintain the data structure $D_{\lambda^{i}}(S)$ of Lemma 7.2 for parameter $\gamma:=\lambda^{i}$. Now, after each change in $S$, we do the following. First, we identify all centers $u\in S$ such that for at least one $i$, the bit $b_{\lambda^{i}}(u,S)$ has changed, and add all of these centers to a maintained set $\mathcal{Y}$ referred to as yellow set. This set will contain centers that might not satisfy Condition (42) and there is a possibility that we call MakeRobust on these centers.
Now, for each center $u$ in yellow set, we do the following. We find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$. Then, check whether $t[u]\geq t$. If $t[u]\geq t$, we are sure that $u$ is already $t$-robust and it satisfies Condition (42) since
$\lambda^{3t[u]}\geq\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}\geq\mathrm{dist}(u,S-u)/\lambda^{10}.$
Hence, there is no need to call a MakeRobust on $u$. Otherwise (if $t[u]<t$), we call MakeRobust on $u$ that find the smallest integer $t^{\prime}$ satisfying $\lambda^{3t^{\prime}}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{7}$  [^17] Note the discrepancy between $\lambda^{7}$ here and $\lambda^{10}$ in Condition (42) and swaps $u$ with a close center which is $t^{\prime}$-robust. Note that it is also possible in this case that $u$ is satisfying Condition (42), but we call MakeRobust on $u$. We will show that the number of these calls is $\tilde{O}(n)$ throughout the entire algorithm. The main reason is that if this case happens, then the distance between $u$ and other centers must have increased by a multiplicative factor, and this can happen at most $O(\log(\sqrt{d}\Delta))$ many times in a row before $u$ is contaminated or become removed from the center set.
We summarize our implementation of Robustify in Algorithm 7. Throughout this procedure, after each insertion or deletion on $S$, all of the data structures maintained related to $S$ are updated. For instance, $D_{\lambda^{i}}(S)$ might add some centers to $\mathcal{Y}$ whose bits have changed in this data structure. Note that we do not need to update these data structures in the previous steps of our algorithm. We can stay until the end of the epoch and update them all together.
Algorithm 7 Implementation of the call to $\texttt{Robustify}(W^{\prime})$ at the end of an epoch.
[…]
3:$S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$Foreach $u\in W_{1}\cup W_{2}$
[…]
5:$u\leftarrow$pop$(\mathcal{Y})$
[…]
14.2.2 Make-Robust
In this section, we introduce the subroutine that swaps a center with a near center which is robust. Assume we are given $u\in S$. First, we compute the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{7}$. Now, we want to find a $t$-robust sequence $(x_{0},x_{1},\ldots,x_{t})$ such that $x_{t}=u$. Then swap $u$ in the solution with $x_{0}$ which is $t$-robust. To do this, we follows the Definition 13.2. Starting from $j=t$, at each iteration, we find $x_{j}$ such that it satisfies the properties in Definition 13.2. Initially, $x_{t}=u$. Now assume that we have found $x_{t},x_{t-1},\ldots,x_{j}$ and we want to find $x_{j-1}$ such that the second condition in Definition 13.2 holds for all $i=j$.
We use the data structure of Lemma 7.4. We make a query on this data structure for $x=x_{j}$ and $r=\lambda^{3j}$. Hence, we get $c^{*}\in[\Delta]^{d}$, $\widehat{\textsc{cost}}_{x_{j}}$, $\widehat{\textsc{cost}}_{c^{*}}$, and $b_{x_{j}}$ explicitly such that there exists a ball $B_{x_{j}}\subseteq X$ with the specified guarantees. We let $B_{j}=B_{x_{j}}$ in Definition 13.2. Note that we do not need to know $B_{j}$ explicitly, and in Definition 13.2, the existence of this set is sufficient. According to the first guarantee in Lemma 7.4, we have $\textsc{Ball}(x_{j},\lambda^{3j})\leq B_{j}\leq\textsc{Ball}(x_{j},p_{\epsilon}\cdot\lambda^{3j})$. Since $\lambda=(p_{\epsilon})^{2}$, we have $p_{\epsilon}\cdot\lambda^{3j}\leq\lambda^{3j+1}$, and we conclude that $B_{x_{j}}=B_{j}$ satisfies the first condition in Definition 13.2 for $i=j$. Now, we proceed with the second condition.
We compare $\widehat{\textsc{cost}}_{x_{j}}/{b_{x_{j}}}$ and $\lambda^{3j-1}/p_{\epsilon}$. There are two cases as follows.
Case 1.
$\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq\lambda^{3j-1}/p_{\epsilon}$. In this case, we let $x_{j-1}=x_{j}$, and show that Equation 40 holds for $i=j$. We have
$\lambda^{3j-1}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq p_{\epsilon}\cdot\textsc{cost}(B_{j},x_{j})/w(B_{j}),$
where the second inequality holds by the second and the fifth guarantees in Lemma 7.4. Since $\lambda=(p_{\epsilon})^{2}$, it follows that $\textsc{cost}(B_{j},x_{j})/w(B_{j})\geq\lambda^{3j-2}$. Equation 40 then holds since we set $x_{j-1}=x_{j}$.
Case 2.
$\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq\lambda^{3j-1}/p_{\epsilon}$. In this case, we conclude
$\textsc{cost}(B_{j},x_{j})/w(B_{j})\leq\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\leq\lambda^{3j-1}/p_{\epsilon}\leq\lambda^{3j-1},$ (51)
where the first inequality holds by the second and the fifth guarantees in Lemma 7.4. Now, we compare $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}$ and $\widehat{\textsc{cost}}_{c^{*}}$. We have two cases as follows.
  • $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{*}}$. In this case, we let $x_{j-1}=x_{j}$. We have
    $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},x_{j})$ $\leq\widehat{\textsc{cost}}_{x_{j}}\leq p_{\epsilon}\cdot\widehat{\textsc{cost}}_{c^{*}}$
    $\leq(p_{\epsilon})^{2}\cdot\textsc{cost}(B_{j},c^{*})\leq(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{j}).$
    The first, third, and fourth inequalities are followed by the second, third, and fourth guarantees in Lemma 7.4 respectively. Since $x_{j-1}=x_{j}$, we obviously have $\textsc{cost}(B_{j},x_{j-1})\leq\textsc{cost}(B_{j},x_{j})$. Together with Equation 51, we see that Equation 41 holds in this case.
  • $\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\geq\widehat{\textsc{cost}}_{c^{*}}$. In this case, we let $x_{j-1}=y$. Similarly,
    $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},c^{*})$ $\leq\widehat{\textsc{cost}}_{c^{*}}\leq\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}$
    $\leq p_{\epsilon}\cdot\textsc{cost}(B_{j},x_{j})/p_{\epsilon}=\textsc{cost}(B_{j},x_{j}),$
    where the first, and the third inequalities are followed by the third and the second guarantees in Lemma 7.4. We also have
    $\textsc{cost}(B_{j},x_{j-1})=\textsc{cost}(B_{j},c^{*})\leq p_{\epsilon}\cdot\mathrm{OPT}_{1}(B_{j}).$
    Together with Equation 51, we see that Equation 41 holds in this case.
As a result, in all cases, the second condition in Definition 13.2 holds for $j=i$. Finally, it follows that the sequences $(x_{0},x_{1},\ldots,x_{t})$ and $(B_{0},B_{1},\ldots,B_{t})$181818we define $B_{0}:=\textsc{Ball}(x_{0},1)$ at the end. satisfy the conditions in Definition 13.2. Hence, $x_{0}$ is a $t$-robust center. Below, we summarize the implementation of MakeRobust in Algorithm 8.
Algorithm 8 Implementation of a call to $\texttt{MakeRobust}(u)$ at the end of an epoch.
1:$t\leftarrow$ Smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{7}$
[…]
4:$x_{j-1}\leftarrow x_{j}$Else
[…]
6:$u_{0}\leftarrow x_{0}$
7:$S\leftarrow S-u+u_{0}$
8:Save $t[u_{0}]\leftarrow t$ together with $u_{0}$
15 Analysis of Our Dynamic Algorithm: Proof of Theorem 12.7
In this section, we provide the analysis of our algorithm. We start by providing some useful lemmas in Section 15.1 that are used to analyze our algorithm. We defer the proof of these main lemmas in Section 20. We provide the approximation ratio analysis in Section 16. Then, we proceed with the recourse analysis in Section 17. Finally, we provide the update time analysis in Section 18.
15.1 Key Lemmas
In this section, we provide the statements of the main lemmas that are used to analyze the algorithm.
15.1.1 Structure of Robust Centers
Lemma 15.1.
If $(x_{0},x_{1},\ldots,x_{t})$ is a $t$-robust sequence, then for each $1\leq j\leq t$, the following hold,
  1. 1.
    $\mathrm{dist}(x_{j-1},x_{j})\leq 2\cdot\lambda^{3j-1}$,
  2. 2.
    $B_{j-1}\subseteq B_{j}$,
  3. 3.
    $\mathrm{dist}(x_{0},x_{j})\leq 4\cdot\lambda^{3j-1}$.
Lemma 15.2.
Let $(x_{0},x_{1},\ldots,x_{t})$ be a $t$-robust sequence w.r.t. $X\subseteq[\Delta]^{d}$. Then, for every $0\leq i\leq t$ and every $B_{i}\subseteq U\subseteq X$, we have $\textsc{cost}(U,x_{0})\leq 4\cdot\textsc{cost}(U,x_{i})$.
Lemma 15.3.
If $S$ is the output of $\texttt{Robustify}(W)$, then $\textsc{cost}(X,S)\leq 4\cdot\textsc{cost}(X,W)$.
15.1.2 Well-Separated Pairs
Definition 15.4.
Assume $S,T\subseteq[\Delta]^{d}$ are two center sets. We call $(u,v)\in S\times T$ a well-separated pair if and only if
$\mathrm{dist}(u,v)\leq\frac{1}{\lambda^{20}}\cdot\min\{\mathrm{dist}(u,S-u),\mathrm{dist}(v,T-v)\}.$
Lemma 15.5.
Consider any two sets of centers $S,V\subseteq[\Delta]^{d}$ such that $S$ is robust. Then, for every well-separated pair $(u,v)\in S\times V$, we have
$\textsc{cost}(C_{v},u)\leq 4\cdot(p_{\epsilon})^{3}\cdot\textsc{cost}(C_{v},v),$
where $C_{v}$ is the cluster of $v$ w.r.t. the center set $V$, i.e. $C_{v}=\{x\in X\mid\mathrm{dist}(x,v)=\mathrm{dist}(x,V)\}$.
Lemma 15.6 ($k$-Means Version of Lemma 4.2 [BCF25a]).
Let $r\geq 0$ and $m\in[0,k]$. Consider any two sets of centers $S,V\subseteq X$ such that $|S|=k$ and $|V|=k+r$. If the number of well-separated pairs w.r.t. $(S,V)$ is $k-m$, then there exists a subset $\tilde{S}\subseteq S$ of size at most $k-\lfloor(m-r)/4\rfloor$ such that
$\textsc{cost}(X,\tilde{S})\leq 81\lambda^{40}\cdot\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right).$
15.1.3 Other Key Lemmas
Lemma 15.7 (Projection Lemma).
Consider any set of centers $C\subseteq[\Delta]^{d}$ of size $|C|\geq k$, where $k$ is a positive integer. Then we have $\mathrm{OPT}_{k}^{C}(X)\leq 2\cdot\textsc{cost}(X,C)+8\cdot\mathrm{OPT}_{k}(X)$.
Lemma 15.8 (Lazy-Updates Lemma).
Consider any two sets of input points $X,X^{\prime}\subseteq[\Delta]^{d}$ such that $|X\oplus X^{\prime}|\leq s$. Then for every $k\geq 1$, we have $\mathrm{OPT}_{k+s}(X^{\prime})\leq\mathrm{OPT}_{k}(X)$.
Lemma 15.9 ($k$-Means Version of Double-Sided Stability (Lemma 4.1 [BCF25a])).
Consider any $r\in[0,k-1]$ and any $\eta\geq 1$. If $\mathrm{OPT}_{k-r}(X)\leq\eta\cdot\mathrm{OPT}_{k}(X)$, then we must have $\mathrm{OPT}_{k}(X)\leq 12\cdot\mathrm{OPT}_{k+\lfloor r/(36\eta)\rfloor}(X)$.
16 Approximation Ratio Analysis of Our Algorithm
In this section, we provide the approximation ratio analysis of our algorithm as follows.
Lemma 16.1.
The approximation ratio of our algorithm is $\mathrm{poly}(1/\epsilon)$ w.h.p.
Roadmap.
In Section 16.1, we show an upper bound on the cost of the solution during the epoch, given that the cost of the initial solution in the epoch is bounded. Then in Section 16.2, we show that the cost of the initial solution at any epoch is bounded.
16.1 Approximation Ratio During the Epoch
In this section, we show that the cost of the solution during any epoch is bounded by $\mathrm{poly}(1/\epsilon)$.
Assumption.
Consider an epoch of length $\ell+1$. We assume that the initial solution in the epoch satisfies
$\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})\leq(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X^{(0)}).$ (52)
We will show that throughout the epoch the solution is always a $\lambda^{53}$ approximation w.r.t. the current dataset. According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Step 1 of our algorithm (see Algorithm 6, Algorithm 6), we infer that
$\mathrm{OPT}_{k-\widehat{\ell}}(X^{(0)})\leq\textsc{cost}(X^{(0)},\widehat{S}_{i-1})\leq(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}).$
Hence, we conclude by Equation 52 that
$\mathrm{OPT}_{k-\widehat{\ell}}(X^{(0)})\leq(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})\leq(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X^{(0)})\leq\lambda^{51}\cdot\mathrm{OPT}_{k}(X^{(0)}),$
where the last inequality follows since $\lambda=(p_{\epsilon})^{2}$. According to Double-Sided Stability Lemma 15.9 for $X=X^{(0)}$, $r=\widehat{\ell}$, and $\eta=\lambda^{51}$, we conclude that
$\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor}(X^{(0)}).$
Finally, since =^$\ell=\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor\leq\widehat{\ell}$, we have
$\frac{\mathrm{OPT}_{k-\ell}(X^{(0)})}{\lambda^{51}}\leq\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+\ell}(X^{(0)}).$ (53)
According to the definition of $S^{(0)}$ in Step 2 of our algorithm, we have
$\textsc{cost}(X^{(0)},S^{(0)})$ $\leq p_{\epsilon}\cdot\mathrm{OPT}_{k-\ell}^{S_{\textnormal{{init}}}}(X^{(0)})$
$\leq p_{\epsilon}\cdot(2\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})+8\cdot\mathrm{OPT}_{k-\ell}(X^{(0)}))$
$\leq p_{\epsilon}\cdot(2\cdot(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X^{(0)})+8\cdot\lambda^{51}\cdot\mathrm{OPT}_{k}(X^{(0)}))$
$\leq\lambda^{52}\cdot\mathrm{OPT}_{k}(X^{(0)})$
$\leq 12\cdot\lambda^{52}\cdot\mathrm{OPT}_{k+\ell}(X^{(0)}).$ (54)
The first inequality follows from Equation 48, the second inequality follows from Projection Lemma 15.7, the third inequality follows from Equation 52 and Equation 53, the fourth inequality follows from the fact that $\lambda=(p_{\epsilon})^{2}$ is a large number, and the last inequality follows from Equation 53.
For each $t\in[0,\ell]$, we have $\left|X^{(t)}\oplus X^{(0)}\right|\leq t\leq\ell$, which concludes
$\mathrm{OPT}_{k+\ell}\left(X^{(0)}\right)\leq\mathrm{OPT}_{k}\left(X^{(t)}\right),$ (55)
As a result, for each $t\in[0,\ell]$, it follows that
$\textsc{cost}(X^{(t)},S^{(t)})$ $\leq\textsc{cost}(X^{(0)},S^{(0)})\leq 12\cdot\lambda^{52}\cdot\mathrm{OPT}_{k+\ell}(X^{(0)})$
$\leq 12\cdot\lambda^{52}\cdot\mathrm{OPT}_{k}(X^{(t)})\leq\lambda^{53}\cdot\mathrm{OPT}_{k}(X^{(t)}).$
The first inequality follows from $S^{(t)}=S^{(0)}+(X^{(t)}-X^{(0)})$, the second inequality follows from Equation 54, the third inequality follows from Equation 55, and the last one follows since $\lambda$ is a large number. This shows that at all times within an epoch, the set $S^{(t)}$ maintained by our algorithm remains a valid $\lambda^{53}$-approximate solution to the $k$-median problem on the current dataset $X^{(t)}$. It remains to show that our algorithm successfully restores Equation 52 when the epoch terminates after the $(\ell+1)^{th}$ update. We do the analysis of this part in the following section.
16.2 Approximation Ratio At the End of the Epoch
In this section, we show that the cost of the final solution at the end of any epoch (which is the initial solution for the next epoch) is bounded by $\mathrm{poly}(1/\epsilon)$ as follows.
Lemma 16.2.
At the end of the epoch, the set $S=S_{\textnormal{{final}}}$ satisfies
$\textsc{cost}(X^{(\ell+1)},S)\leq(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X^{(\ell+1)}).$
The claim below bounds the cost of the solution $S^{\prime}$ w.r.t. the dataset $X^{(0)}$.
Claim 16.3.
In step 4 of our algorithm, we have $\textsc{cost}(X^{(0)},S^{\prime})\leq 4\cdot(p_{\epsilon})^{4}\cdot\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$.
Before proving Claim 16.3, we explain how it implies Lemma 16.2. Towards this end, note that:
$\textsc{cost}(X^{(\ell+1)},S_{\textnormal{{final}}})$$\leq$ $4\cdot\textsc{cost}(X^{(\ell+1)},W^{{}^{\prime}})$ (56)
$\leq$ $4p_{\epsilon}\cdot\mathrm{OPT}_{k}^{T^{\prime}}(X^{(\ell+1)})$ (57)
$\leq$ $4p_{\epsilon}\cdot\left(2\cdot\textsc{cost}(X^{(\ell+1)},T^{\prime})+8\cdot\mathrm{OPT}_{k}(X^{(\ell+1)})\right)$ (58)
$\leq$ $4p_{\epsilon}\cdot\left(2\cdot\textsc{cost}(X^{(0)},S^{\prime})+8\cdot\mathrm{OPT}_{k}(X^{(\ell+1)})\right)$ (59)
$\leq$ $4p_{\epsilon}\cdot\left(8(p_{\epsilon})^{4}\cdot\mathrm{OPT}_{k+\ell+1}(X^{(0)})+8\cdot\mathrm{OPT}_{k}(X^{(\ell+1)})\right)$ (60)
$\leq$ $4p_{\epsilon}\cdot\left(8(p_{\epsilon})^{4}\cdot\mathrm{OPT}_{k}(X^{(\ell+1)})+8\cdot\mathrm{OPT}_{k}(X^{(\ell+1)})\right)$ (61)
$\leq$ $(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}\left(X^{(\ell+1)}\right).$ (62)
In the above derivation, the first step (56) follows from Lemma 15.3. The second step (57) follows from Equation 50. The third step (58) follows from Projection Lemma 15.7. The fourth step (59) follows from the fact that $T^{\prime}=S^{\prime}+(X^{(\ell+1)}-X^{(0)})$. The fifth step (60) follows from Claim 16.3. The fourth step (61) follows from the Lazy-Updates Lemma 15.8 and the observation that $\left|X^{(\ell+1)}\oplus X^{(0)}\right|\leq\ell+1$. The last step (62) follows from the fact that $p_{\epsilon}$ is a large number.
16.2.1 Proof of Claim 16.3
Let $V\subseteq[\Delta]^{d}$ be an optimal $(k+\ell+1)$-median solution for the dataset $X^{(0)}$, i.e., $\left|V\right|=k+\ell+1$ and $\textsc{cost}(X^{(0)},V)=\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$. Let $m\in[0,k]$ be the unique integer such that there are $(k-m)$ well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},T\right)$. Let $\{(u_{1},v_{1}),(u_{2},v_{2}),\ldots,(u_{k-m},v_{k-m})\}\subseteq S_{\textnormal{{init}}}\times V$ denote the collection of $(k-m)$ well-separated pairs w.r.t. $\left(S_{\textnormal{{init}}},V\right)$. Define the set $\tilde{A}:=V\setminus\{v_{1},\ldots,v_{k-m}\}$. It is easy to verify that:
$\left|\tilde{A}\right|=\left|V\right|-(k-m)=m+\ell+1.$ (63)
Claim 16.4.
We have $m\leq 300\cdot\lambda^{51}\cdot(\ell+1)$.
Claim 16.5.
We have $\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})\leq 4\cdot(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$.
By (63), Claim 16.4 and Claim 16.5, there exists a set $\tilde{A}\subseteq[\Delta]^{d}$ of $m+\ell+1\leq(300\cdot\lambda^{51}+1)\cdot(\ell+1)\leq\lambda^{52}\cdot(\ell+1)=s$ centers such that $\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})\leq 4(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$. Accordingly, from (49), we get
$\textsc{cost}(X^{(0)},S^{\prime})$ $\leq p_{\epsilon}\cdot\min\limits_{A\subseteq[\Delta]^{d}:|A|\leq s}\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+A)$
$\leq p_{\epsilon}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})$
$\leq 4\cdot(p_{\epsilon})^{4}\cdot\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right),$
which implies Claim 16.3.
It now remains to prove Claim 16.4 and Claim 16.5.
Proof of Claim 16.4
We apply Lemma 15.6, by setting $r=\ell+1$, $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$. This implies the existence of a set $\tilde{S}\subseteq S_{\textnormal{{init}}}$ of at most $(k-b)$ centers, with $b=\lfloor(m-\ell-1)/4\rfloor$, such that
$\textsc{cost}(X^{(0)},\tilde{S})$$\leq$ $81\lambda^{40}\cdot\left(\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})+\textsc{cost}(X^{(0)},V)\right)$
$\leq$ $81\lambda^{40}\cdot\left((p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right)+\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)\right)\leq\lambda^{44}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right).$
In the above derivation, the second inequality follows from the initial assumption at the beginning of the epoch (52), and the last inequality holds because $\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)\leq\mathrm{OPT}_{k}(X^{(0)})$ and $\lambda=(p_{\epsilon})^{2}$ is a large number (see Part III). Since $\mathrm{OPT}_{k-b}\left(X^{(0)}\right)\leq\textsc{cost}(X^{(0)},\tilde{S})$, we get
$\mathrm{OPT}_{k-b}\left(X^{(0)}\right)\leq\lambda^{44}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right).$ (64)
Recall the way $\ell,\widehat{\ell}$ and $\ell^{\star}$ are defined in Step 1 of our algorithm and our implementation (Algorithm 6). From (64), it follows that $b\leq\ell^{\star}$. According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Algorithm 6, for $s_{i}=2s_{i-1}=2\widehat{\ell}$, from (Algorithm 6), we have
$(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})<\textsc{cost}(X^{(0)},\widehat{S}_{i})\leq p_{\epsilon}\cdot\mathrm{OPT}^{S_{\textnormal{{init}}}}_{k-2\widehat{\ell}}(X^{(0)})$
We also have,
$\mathrm{OPT}^{S_{\textnormal{{init}}}}_{k-2\widehat{\ell}}(X^{(0)})$ $\leq 2\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})+8\cdot\mathrm{OPT}_{k-2\widehat{\ell}}(X^{(0)})$
$\leq 2\cdot(p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}(X^{(0)})+8\cdot\mathrm{OPT}_{k-2\widehat{\ell}}(X^{(0)})$
$\leq(p_{\epsilon})^{7}\cdot\mathrm{OPT}_{k-2\widehat{\ell}}(X^{(0)})$
The first inequality follows from Lemma 15.7, the second inequality follows from Equation 52, and the last inequality follows since $p_{\epsilon}$ is a large number. Hence,
$(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\mathrm{OPT}_{k}(X^{(0)})<p_{\epsilon}\cdot\mathrm{OPT}^{S_{\textnormal{{init}}}}_{k-2\widehat{\ell}}(X^{(0)})\leq(p_{\epsilon})^{8}\cdot\mathrm{OPT}_{k-2\widehat{\ell}}(X^{(0)}),$
which concludes
$\lambda^{44}\cdot\mathrm{OPT}_{k}(X^{(0)})<\mathrm{OPT}_{k-2\widehat{\ell}}(X^{(0)}).$
This shows that $2\widehat{\ell}\geq\ell^{\star}$ by the definition of $\ell^{\star}$.
Now, we have $b\geq\frac{m-\ell-1}{4}-1=\frac{m-\ell-5}{4}$ and ^$\ell\geq\frac{\widehat{\ell}}{36\cdot\lambda^{51}}-1$, we get
$\frac{m-\ell-5}{4}\leq b\leq\ell^{\star}\leq 2\widehat{\ell}\leq 2\cdot 36\cdot\lambda^{51}\cdot(\ell+1)$
and hence
$m\leq 288\cdot\lambda^{51}\cdot(\ell+1)+\ell+5\leq 300\cdot\lambda^{51}\cdot(\ell+1).$
This concludes the proof of the claim.
Proof of Claim 16.5
For each $v_{i}\in V$, denote the cluster of $v_{i}\in V$ w.r.t. the center set $V$ by $C_{v_{i}}$, i.e. $C_{v_{i}}=\{x\in X^{(0)}\mid\mathrm{dist}(x,v_{i})=\mathrm{dist}(x,V)\}$. We define assignment $\sigma:X^{(0)}\rightarrow S_{\textnormal{{init}}}+\tilde{A}$, as follows.
  • If $p\in C_{v_{i}}$ for some $i\in[1,k-m]$, then $\sigma(p):=u_{i}$.
  • Otherwise, $\sigma(p):=\pi_{V}(p)$.
In words, for every well-separated pair $(u_{i},v_{i})\in S_{\textnormal{{init}}}\times V$ all the points in the cluster of $v_{i}$ get reassigned to the center $u_{i}$, and the assignment of all other points remain unchanged (note that their assigned centers are present in $S_{\textnormal{{init}}}+\tilde{A}$ as well as $V$). Now, recall that at the beginning of the epoch, the set of centers $S_{\textnormal{{init}}}$ is robust (see Invariant C). Hence, by applying Lemma 15.5, we infer that
$\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})\leq\sum_{x\in X^{(0)}}\mathrm{dist}(x,\sigma(x))\leq 4\cdot(p_{\epsilon})^{3}\cdot\textsc{cost}(X^{(0)},V).$
The claim then follows since $\textsc{cost}(X^{(0)},V)=\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$ by definition.
17 Recourse Analysis of Our Algorithm
In this section, we provide the recourse analysis of our algorithm as follows.
Lemma 17.1.
The amortized recourse of the algorithm is at most $\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$, where $d$ is the dimension of the space, $n$ is the upper bound on the number of the points in the data-set, and $[0,\Delta]$ is the range of the coordinates of the points.
Roadmap.
We start by showing the bound on the recourse throughout an epoch except the call to Robustify in Section 17.1. Then, we proceed by analysis of the recourse overhead of Robustify subroutine in Section 17.2.
17.1 Recourse Before Calling Robustify
Fix an epoch (say) $\mathcal{E}$ of length $(\ell+1)$. The total recourse incurred by our algorithm during this epoch is
$R_{\mathcal{E}}\leq\left|S_{\textnormal{{init}}}\oplus S^{(0)}\right|+\left(\sum_{t=1}^{\ell+1}\left|S^{(t)}\oplus S^{(t-1)}\right|\right)+\left|S^{(\ell+1)}\oplus S_{\textnormal{{final}}}\right|.$ (65)
$S^{(0)}$ is obtained by deleting $\ell$ centers from $S_{\textnormal{{init}}}$ in Step 2, and hence we have $\left|S_{\textnormal{{init}}}\oplus S^{(0)}\right|=\ell$. During Step 3, we incur a worst-case recourse of at most one per update. Specifically, we have $\left|S^{(t)}\oplus S^{(t-1)}\right|\leq 1$ for all $t\in[1,\ell+1]$, and hence $\left|S^{(\ell+1)}\oplus S_{\textnormal{{init}}}\right|\leq\left|S^{(\ell+1)}\oplus S^{(0)}\right|+\left|S^{(0)}\oplus S_{\textnormal{{init}}}\right|\leq(\ell+1)+\ell=2\ell+1$. Thus, from (65) we get:
$R_{\mathcal{E}}$$\leq$ $\ell+(\ell+1)+\left|S^{(\ell+1)}\oplus S_{\textnormal{{final}}}\right|$ (66)
$\leq$ $(2\ell+1)+\left|S^{(\ell+1)}\oplus S_{\textnormal{{init}}}\right|+\left|S_{\textnormal{{init}}}\oplus S_{\textnormal{{final}}}\right|$
$\leq$ $(4\ell+2)+\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|+\left|W^{\prime}\oplus S_{\textnormal{{final}}}\right|$
Next, we bound $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$. According to Step 4 of our algorithm, we infer that
$\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ $\leq\left|S_{\textnormal{{init}}}\oplus S^{\prime}\right|+\left|S^{\prime}\oplus T^{\prime}\right|+\left|T^{\prime}\oplus W^{\prime}\right|$
According to Equation 49, $S^{\prime}\supseteq S_{\textnormal{{init}}}$ satisfies $\left|S_{\textnormal{{init}}}\oplus S^{\prime}\right|\leq\tilde{O}(\ell+1)$. The $\left|S^{\prime}\oplus T^{\prime}\right|$ term is bounded by $\ell+1$ since $T^{\prime}=S^{\prime}+(X^{(\ell+1)}-X^{(0)})$. Since $\left|S^{\prime}\right|=k+\tilde{O}(\ell+1)$, and $W^{\prime}$ is a subset of $T^{\prime}$ of size $k$, we get $\left|T^{\prime}\oplus W^{\prime}\right|\leq\tilde{O}(\ell+1)$. This concludes
$\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|\leq\tilde{O}(\ell+1).$ (67)
Since the epoch lasts for $\ell+1$ updates, the $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ and $24\ell+$ 2 terms in the right hand side of (66) contributes an amortized recourse of $\tilde{O}(1)$. Moreover, the term $\left|W^{\prime}\oplus S_{\textnormal{{final}}}\right|$ is proportional to the number of MakeRobust calls made while computing $S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}\left(W^{\prime}\right)$. In Lemma 17.2, we show that the amortized number of calls made to the MakeRobust subroutine throughout the entire sequence of updates is at most $\tilde{O}(1)$, which proves Lemma 17.1.
Lemma 17.2.
The algorithm makes at most $\mathrm{poly}(1/\epsilon)\cdot\log n\cdot\log^{2}(\sqrt{d}\Delta)$ many calls to MakeRobust, amortized over the entire sequence of updates (spanning multiple epochs).
17.2 Proof of Lemma 17.2
At the end of the epoch, we set $S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}\left(W^{\prime}\right)$ Recall the subroutine Robustify, and specifically, the partition $W^{\prime}=W_{1}\cup W_{2}\cup W_{3}$.
We can partition the calls to MakeRobust that are made by $\texttt{Robustify}\left(W^{\prime}\right)$ into the following three types.
Type I. A call to $\texttt{MakeRobust}(w)$ for some $w\in W_{1}$. The set $W_{1}=W^{\prime}\setminus S_{\textnormal{{init}}}$, has size at most $\tilde{O}(\ell+1)$ by Equation 67. Since the epoch lasts for $(\ell+1)$ updates, the amortized number of Type I calls to MakeRobust, per update, is $\tilde{O}(1)$.
Type II. A call to $\texttt{MakeRobust}(w)$ for some $w\in W_{2}$ that is contaminated at the end of the epoch. We can bound the number of these calls by $(\ell+1)\cdot\log(\sqrt{d}\Delta)$. This bound comes from Claim 14.3 which concludes each $x\in X^{(\ell+1)}\oplus X^{(0)}$ can contaminate at most $\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil+1\leq\log(\sqrt{d}\Delta)$ centers. The amortized number of such Type II MakeRobust calls, per update, is then at most $\log(\sqrt{d}\Delta)$.
Type III. A call to $\texttt{MakeRobust}(w)$ for some $w\in W_{3}$. Note that according to Lemma 14.4, after a call $u^{\text{new}}\leftarrow\texttt{MakeRobust}(u)$ for some $u\in W_{1}\cup W_{2}$, we will not make a call to MakeRobust on $u^{\text{new}}$ in the current Robustify. Hence, our classification of calls to MakeRobust is comprehensive.
To bound the amortized number of Type III calls, we need to invoke a more “global” argument, that spans across multiple epochs. Consider a maximal “chain” of $j$ many Type III calls (possibly spanning across multiple different epochs), in increasing order of time:
$w_{1}\leftarrow\texttt{MakeRobust}(w_{0}),w_{2}\leftarrow\texttt{MakeRobust}(w_{1}),$
$\ldots,w_{j}\leftarrow\texttt{MakeRobust}(w_{j-1}).$
Note that the calls in the above chain can be interspersed with other Type I, Type II, or Type III calls that are not part of the chain.
Claim 17.3.
The length of this chain is at most $\log(\sqrt{d}\Delta)$.
Proof.
Consider one Type III call $w_{i}\leftarrow\texttt{MakeRobust}(w_{i-1})$. This call is made on 8 in Algorithm 7, which means that $t[w_{i-1}]$ satisfies $t[w_{i-1}]<t$ where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}$. Hence,
$\lambda^{3t[w_{i-1}]}<\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}\leq p_{\epsilon}\cdot\mathrm{dist}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}\leq\mathrm{dist}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{9}.$
The second inequality follows by the guarantee of $\widehat{\mathrm{dist}}$ in Lemma 7.3 and the third inequality follows since $\lambda=(p_{\epsilon})^{2}$ is a large number. Now, in the call $\texttt{MakeRobust}(w_{i-1})$, we set $t[w_{i}]$ to be the smallest integer satisfying $\lambda^{3t[w_{i}]}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{7}$. It follows that,
$\lambda^{3t[w_{i}]}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{7}$ $\geq\mathrm{dist}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{7}$
$>\mathrm{dist}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{9}>\lambda^{3t[w_{i-1}]},$
which concludes $t[w_{i}]>t[w_{i-1}]$.
As a result, $t[w_{0}]<t[w_{1}]<\ldots<t[w_{j}]$. Since the aspect ratio of the space is $(\sqrt{d}\Delta)$, all of the integers $t[w_{i}]$ are between 00 and $\log_{\lambda^{3}}(\sqrt{d}\Delta)<\log(\sqrt{d}\Delta)$ which concludes $j\leq\log(\sqrt{d}\Delta)$. ∎
For the chain to start in the first place, we must have had a Type I or Type II call to MakeRobust which returned the center $w_{0}$. We can thus “charge” the length (total number of Type III calls) in this chain to the Type I or Type II call that triggered it (by returning the center $w_{0}$). In other words, the total number of Type III calls ever made is at most $\log(\sqrt{d}\Delta)$ times the total number of Type I plus Type II calls. Since the amortized number of Type I and Type II calls per update is at most $\tilde{O}(1)$, the amortized number of Type III calls per update is also at most $\tilde{O}(1)$. Recall that $(\sqrt{d}\Delta)=\sqrt{d}\Delta$ is the aspect ratio of the space, and we hide $\mathrm{poly}(1/\epsilon),\mathrm{poly}\log(n\Delta)$ and $\mathrm{poly}(d)$ in the $\tilde{O}$ notation. Hence, the proof of Lemma 17.2 is complete.
18 Update Time Analysis
In this section, we analyze the update time of our algorithm as follows.
Lemma 18.1.
The amortized update time of the algorithm is at most $n^{\epsilon}\cdot\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$.
As discussed in different Steps of our algorithm in Section 14, all of the subroutines, except Robustify, will spend at most $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time.
In Step 1, the running time is dominated by the last call to restricted $k$-means (Lemma 7.5) that takes $\tilde{O}(n^{\epsilon}\cdot s_{i^{\star}})$ time, where $s_{i^{\star}}=O(\widehat{\ell})=O(\ell+1)$. Step 2 is similar to Step 1, where we spend $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time by calling restricted $k$-means to find the desired $S^{(0)}\subseteq S_{\textnormal{{init}}}$ of size $k-\ell$. Step 3 is trivial and each update can be done in $\tilde{O}(1)$ time (apart from updating the underlying data structures). In Step 4, we call the augmented $k$-means (Lemma 12.1) to find the desire $S^{\prime}=S_{\textnormal{{init}}}+A$ of size $k+\lambda^{52}\cdot(\ell+1)$. This subroutine spends $\tilde{O}(n^{\epsilon}\cdot\lambda^{52}\cdot(\ell+1))=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time. Computing $T^{\prime}=S^{\prime}+(X^{(\ell+1)}-X^{(0)})$ is trivial. We will call restricted $k$-means on $T^{\prime}$ to get $W^{\prime}\subseteq T^{\prime}$ of size $k$. This takes $\tilde{O}(|T^{\prime}|-|W^{\prime}|)=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time as well. Finally, we call Robustify on $W^{\prime}$.
Apart from Robustify and updating the background data structures, all of the other parts of the algorithm in this epoch, take at most $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$ time, which can be amortized on the length of the epoch.
18.1 Running Time of Robustify
Recall that at the end of each epoch, $W^{\prime}$ is partitioned into three sets $W^{\prime}=W_{1}\cup W_{2}\cup W_{3}$.
Time Spent on $W_{1}\cup W_{2}$.
At the end of an epoch of length $\ell+1$, Since $W_{1}=W^{\prime}\setminus S_{\textnormal{{init}}}$, it is obvious that we can find $W_{1}$ in a total of $|S^{\prime}\setminus S_{\textnormal{{init}}}|=\tilde{O}(\ell+1)$. According to the subroutine in Algorithm 7, to find $W_{2}$, for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ and each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$, we will call the nearest neighbor oracle $\textsc{NN}(x,S[i])$ (Lemma 7.1). The oracle takes $\tilde{O}(n^{\epsilon})$ time, which concludes the total time spend to find $W_{2}$ is at most $\tilde{O}(n^{\epsilon}\cdot|X^{(0)}\oplus X^{(\ell+1)}|\cdot\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil)=\tilde{O}(n^{\epsilon}\cdot(\ell+1))$. Next, for each $u\in W_{1}\cup W_{2}$ a call to $\texttt{MakeRobust}(u)$ is made that takes $\tilde{O}(n^{\epsilon})$ according to Section 18.3. In total, the time spend to handle centers in $W_{1}$ and $W_{2}$ (Lines 1 to 9 in Algorithm 7) is at most $\tilde{O}(n^{\epsilon}\cdot(\ell+1))$, which can be amortized over the length of the epoch.
Time Spent on $W_{3}$.
We need a global type of analysis for this part. It is possible that Robustify takes a huge amount of time for handling points in $W_{3}$. But, considering the entire sequence of updates, we show that the amortized update time of this part of the Robustify subroutine is at most $\tilde{O}(n^{\epsilon})$.
According to our algorithm, for each center $u$ in the yellow set $\mathcal{Y}$, in $\tilde{O}(1)$ time (5 to 6) [^20] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 7.3 ., our algorithm decides whether or not a call to MakeRobust will be made on $u$. Each call to MakeRobust also takes $\tilde{O}(n^{\epsilon})$ time according to Section 18.3. As a result, if the number of elements added to $\mathcal{Y}$ is $N$ and the number of calls to MakeRobust is $M$ throughout the entire algorithm, we conclude that the running time of handling $W_{3}$ is at most $N\cdot\tilde{O}(1)+M\cdot\tilde{O}(n^{\epsilon})$. According to the recourse analysis of our algorithm, the amortized number of calls to MakeRobust is bounded by $M\leq\tilde{O}(1)$.
Elements are added to $\mathcal{Y}$ by data structures $D_{\lambda^{i}}(S)$ for each $i\in[0,\lceil\log_{\lambda}(\sqrt{d}\Delta)\rceil]$. Each update on the main center set $S$ (insertion or deletion) is considered an update for this data structure. This data structure has an amortized update time of $\tilde{O}(n^{\epsilon})$, which implicitly means that the number of centers $u$ whose bit $b_{\lambda^{i}}(u,S)$ has changed is at most $\tilde{O}(n^{\epsilon})$ in amortization (w.r.t. the sequence of updates on $S$). Hence, for each $i$, the number of elements added to $\mathcal{Y}$ by the data structure $D_{\lambda^{i}}(S)$ is at most $\tilde{O}(M\cdot n^{\epsilon})$. As a result, the total number of elements added to $\mathcal{Y}$ throughout the entire algorithm is at most $N\leq\tilde{O}((\lceil\log_{\lambda}(\sqrt{d}\Delta)\rceil+1)\cdot M\cdot n^{\epsilon})=\tilde{O}(n^{\epsilon})$
Finally, the amortized update time of Robustify spent on handling $W_{3}$ throughout the entire algorithm is $\tilde{O}(n^{\epsilon})$.
18.2 Update Time of Background Data Structures
Each update in the main data set $X$ and in the main center set $S$ is considered an update for the background data structures that are maintained in our algorithm. Throughout a sequence of $T$ updates on $X$, the number of updates on $S$ is bounded by $\tilde{O}(T)$ according to the recourse analysis of our algorithm (see Lemma 17.1). Since each of the data structures maintained in the background has $\tilde{O}(n^{\epsilon})$ update time (see Section 7), the total time spent to update these data structures during these $T+\tilde{O}(T)$ updates on either $X$ or $S$, is at most $\tilde{O}(T\cdot n^{\epsilon})$. Hence, the amortized update time of all the background data structures is $\tilde{O}(n^{\epsilon})$ in amortization (w.r.t. the original update sequence on the main dataset $X$).
18.3 Running Time of MakeRobust
The main for loop in the implementation consists of at most $O(\log_{\lambda}(\sqrt{d}\Delta))$ many iterations. Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 7.4) used in 3 of the implementation. It is obvious that all other parts of the algorithm take at most $\tilde{O}(1)$ time. Hence, the total time spent to perform one call to MakeRobust is at most $\tilde{O}(n^{\epsilon})$.
19 Improving the Update Time to $\tilde{O}(k^{\epsilon})$
Roadmap.
In this section, we show how to improve the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$. We do this by using dynamic sparsification as a black box, in a similar manner to previous dynamic clustering algorithms [BCG+24, BCF25a]. In [BCF25a], the authors give an algorithm that maintains a $O(1)$-approximation with $\tilde{O}(1)$ recourse and $\tilde{O}(n)$ update time. Using the dynamic $k$-means algorithm of [BCLP23] as a dynamic sparsifier in a black-box manner, they can then improve the update time of their algorithm to $\tilde{O}(k)$, while only incurring $O(1)$ overhead in the approximation and $\tilde{O}(1)$ overhead in the recourse. We speed up the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ using a completely analogous approach, except that we use the sparsifier of [DHS24], which is more well-suited to Euclidean spaces.
Our Sparsified Dynamic Algorithm.
The following theorem summarises the relevant properties of the dynamic algorithm of [DHS24], which we call Sparsifier. [^21] The dynamic algorithm of [DHS24] dynamically maintains a coreset of the input space. For the sake of simplicity, we do not define the notion of a coreset and instead focus on the fact that it can be used as a sparsifier. We use this algorithm as a black-box.
Theorem 19.1 ([DHS24]).
There is an algorithm Sparsifier that, given a dynamic weighted $d$-dimensional Euclidean space $X$, maintains a weighted subspace $U\subseteq X$ of size $\tilde{O}(k)$ with $\tilde{O}(1)$ amortized update time and $\tilde{O}(1)$ amortized recourse. Furthermore, after each update, the following holds with probability at least $1-1/\mathrm{poly}(n)$,
  • Any $\beta$-approximation to the $k$-means problem on the weighted subspace $U$ is also a $O(\beta)$-approximation to the $k$-means problem on the space $X$.
The algorithm Theorem 19.1 shows that we can efficiently maintain a dynamic sparsifier of the space $X$ with high probability. The following theorem summarizes our main algorithm, which we call ALG.
Theorem 19.2.
There is an algorithm ALG that, given a dynamic weighted $d$-dimensional Euclidean space of size $n$ and a sufficiently small parameter $\epsilon>0$, maintains a $\mathrm{poly}(1/\epsilon)$-approximate solution to the $k$-means problem with $\tilde{O}(n^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse. The approximation guarantee holds with probability at least $1-1/\mathrm{poly}(n)$ after each update.
We now show how to combine these algorithms in order to improve the running time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$. [^22] We note that the main challenge here is doing this while ensuring that the success probability of our algorithm remains $1-1/\mathrm{poly}(n)$, since the naive approach leads to a success probability that is $1-1/\mathrm{poly}(k)$, which can be significantly smaller.
19.1 Algorithm Description
We initialize a copy of the algorithm Sparsifier and feed it the dynamic space $X$ as input. In response, the algorithm Sparsifier maintains a weighted subspace $U\subseteq X$, with the properties outlined in Theorem 19.1. We also initialize one ‘primary’ copy of the algorithm ALG, which we denote by $\mathrm{\textsc{ALG}}^{\star}$, along with $L=O(\log n)$ ‘verification’ copies of ALG, which we denote by $\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$. Each of these algorithms is fed the dynamic space $U$ as input. In respose, the algorithms $\mathrm{\textsc{ALG}}^{\star},\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ maintain solutions $S^{\star},S_{1}^{\prime},\dots,S_{L}^{\prime}$ to the $k$-means problem on the space $U$. Thus, whenever the space $X$ is updated by some point insertion or deletion, the algorithm Sparsifier updates the subspace $U$. In turn, the algorithms $\mathrm{\textsc{ALG}}^{\star},\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ then update the solutions $S^{\star},S_{1}^{\prime},\dots,S_{L}^{\prime}$.
At any point in time, the output of our algorithm is the solution $S^{\star}$. In order to make sure that the solution $S^{\star}$ has a good approximation guarantee, we make the following modification to this algorithm.
Detecting When $S^{\star}$ Expires.
Our algorithm maintains an estimate $\mathcal{E}:=\min_{i\in[L]}\textsc{cost}(U,S_{i}^{\prime})$ of $\mathrm{OPT}_{k}(U)$. Let $\alpha\leq\mathrm{poly}(1/\epsilon)$ denote an upper bound on the approximation ratio of $\mathrm{\textsc{ALG}}^{\star}$. After handling each update, we check if $\textsc{cost}(U,S^{\star})>\alpha\cdot\mathcal{E}$. If this is the case, we completely recompute the primary instance $\mathrm{\textsc{ALG}}^{\star}$ from scratch using fresh randomness, which reconstructs the solution $S^{\star}$. We repeat this process until $\textsc{cost}(U,S^{\star})\leq\alpha\cdot\mathcal{E}$.
19.2 Analysis
We now show that this modified version of the algorithm, which uses sparsification, has an update time of $\tilde{O}(k^{\epsilon})$ without sacrificing any other guarantees. In particular, we prove the following theorem.
Theorem 19.3.
There is an algorithm that, given a dynamic weighted $d$-dimensional Euclidean space of size at most $n$, a polynomially bounded sequence of updates to the space $X$, and a sufficiently small parameter $\epsilon>0$, maintains a $\mathrm{poly}(1/\epsilon)$-approximate solution to the $k$-means problem with $\tilde{O}(k^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse. The approximation, update time, and recourse guarantees hold with probability at least $1-1/\mathrm{poly}(n)$ across the entire sequence of updates.
19.2.1 Approximation Ratio Analysis
We first show that, at any point in time, we have that $\textsc{cost}(X,S^{\star})\leq O(\alpha^{2})\cdot\mathrm{OPT}_{k}(X)$ with probability at least $1-1/\mathrm{poly}(n)$. We begin with the following lemma, which shows that $\mathcal{E}$ is an $\alpha$-approximation to $\mathrm{OPT}_{k}(U)$ with high probability.
Lemma 19.4.
After each update, we have $\mathcal{E}\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $1-1/\mathrm{poly}(n)$.
Proof.
By Theorem 19.2, at each point in time, we have that $\textsc{cost}(U,S_{i}^{\prime})\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $21-1/\mathrm{poly}(k)\geq 1/$ 2 for each $i\in[L]$. Since each of the algorithms $\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ uses independent randomness, the probability that $\textsc{cost}(U,S_{i}^{\prime})>\alpha\cdot\mathrm{OPT}_{k}(U)$ for all $i\in[L]$ is at most $(1/2)^{L}=1/\mathrm{poly}(n)$. ∎
Since our algorithm maintains that $\textsc{cost}(U,S^{\star})\leq\alpha\cdot\mathcal{E}$ after each update, it follows from Lemma 19.4 that $\textsc{cost}(U,S^{\star})\leq\alpha^{2}\cdot\mathrm{OPT}_{k}(U)$ with probability at least $1-1/\mathrm{poly}(n)$ after each update. By Theorem 19.1, we know that if $\textsc{cost}(U,S^{\star})\leq\alpha^{2}\cdot\mathrm{OPT}_{k}(U)$, then $\textsc{cost}(X,S^{\star})\leq O(\alpha^{2})\cdot\mathrm{OPT}_{k}(X)$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, our approximation guarantee follows from union bounding over these two events.
19.2.2 Recourse and Update Time Analysis
Consider a sequence of $T=O(\mathrm{poly}(n))$ updates $\sigma_{1},\dots,\sigma_{T}$ defining the dynamic space $X$. For each $i\in[T]$, let $R_{i}$ denote the recourse caused by the update $\sigma_{i}$, i.e. the total number of points that are either inserted or deleted from $S^{\star}$ by our algorithm while handling the update, and let $R:=\sum_{i=1}^{T}R_{i}$. In other words, $R$ is the total recourse of the solution $S^{\star}$ caused by handling the updates $\sigma_{1},\dots,\sigma_{T}$. Thus, $R/T$ is the amortized recourse. Similarly, for each $i\in[T]$, let $r_{i}$ denote the indicator for the event that the algorithm restarts the algorithm $\mathrm{\textsc{ALG}}^{\star}$ using fresh randomness after handling the update $\sigma_{i}$, and let $r:=\sum_{i=1}^{T}r_{i}$.
Lemma 19.5.
We have that $R=\tilde{O}(T)+O(kr)$.
Proof.
Since the algorithms Sparsifier and ALG both have an amortized recourse of $\tilde{O}(1)$, it follows that the algorithm obtained by composing them also has an amortized recourse of $\tilde{O}(1)$. Thus, ignoring the recourse caused by resetting the algorithm $\mathrm{\textsc{ALG}}^{\star}$, the algorithm has an amortized recourse of $\tilde{O}(1)$, and thus the total recourse across $T$ updates is at most $\tilde{O}(T)$. If the algorithm resets the algorithm $\mathrm{\textsc{ALG}}^{\star}$, it incurs an additional recourse of $O(k)$ since the solution might change completely. Thus, we get an additive $O(kr)$ term in the recourse of our algorithm to account for the resets. ∎
Now, let $p(n)=\mathrm{poly}(n)=\Omega(n)$ be the polynomial such that the approximation guarantee of ALG holds with probability at least $1-1/p(n)$ after each update, as described in Theorem 19.2. Consider some subsequence of updates $\sigma_{\ell+1},\dots,\sigma_{\ell+q}$, where $q\leq p(k)/2$. We denote $\sigma_{\ell+j}$ and $r_{\ell+j}$ by $\sigma^{\prime}_{j}$ and $r^{\prime}_{j}$ respectively, and let $r^{\prime}:=\sum_{j=1}^{q}r^{\prime}_{j}$.
Lemma 19.6.
We have that $\Pr[r^{\prime}\geq\Omega(\log n)]\leq 1/\mathrm{poly}(n)$.
Proof.
We begin with the following simple claim, which bounds the probability of our algorithm resetting $\mathrm{\textsc{ALG}}^{\star}$ after an update.
Claim 19.7.
For each $j\in[q]$, we have that $\Pr[r^{\prime}_{j}=1]\leq 1/p(k)$.
Proof.
If the space $U$ has size less than $k$ after handling the update $\sigma^{\prime}_{j}$, then $S^{\star}$ is the optimum solution, and we will not restart $\mathrm{\textsc{ALG}}^{\star}$. Otherwise, by Theorem 19.2, we know that $S^{\star}$ is an $\alpha$-approximation to $\mathrm{OPT}_{k}(U)$ with probability at least $1-1/p(|U|)\geq 1-1/p(k)$, which implies that $\textsc{cost}(U,S^{\star})\leq\alpha\cdot\mathrm{OPT}_{k}(U)\leq\alpha\cdot\mathcal{E}$. Thus, we have that $\textsc{cost}(U,S^{\star})>\alpha\cdot\mathcal{E}$, and hence $r^{\prime}_{j}=1$, with probability at most $1/p(k)$. ∎
Claim 19.8.
For each $i\in[k]$, we have that $\Pr[r^{\prime}\geq i\mid r^{\prime}\geq i-1]\leq 1/2$.
Proof.
We refer to an update $\sigma_{j}^{\prime}$ as bad if $r_{j}^{\prime}=1$, i.e. if we restart the algorithm $\mathrm{\textsc{ALG}}^{\star}$ with fresh randomness at the end of the update. Suppose that $r^{\prime}\geq i-1$ and let $\sigma_{j_{1}}^{\prime},\dots,\sigma_{j_{i-1}}^{\prime}$ be the first $i-1$ bad updates in $\sigma^{\prime}_{1},\dots,\sigma^{\prime}_{q}$. Since we restart $\mathrm{\textsc{ALG}}^{\star}$ after update $\sigma^{\prime}_{j_{i-1}}$, the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ that effect whether any of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ are bad are independent of the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ to handle the previous updates. Thus, conditioned on this event, each of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ is bad with probability at most $1/p(k)$ by Claim 19.7. Taking a union bound, we get that none of them is bad with probability at least $21-q/p(k)\geq 1/$ 2. ∎
By applying Claim 19.8, we get that
$\Pr[r^{\prime}\geq\Omega(\log n)]=\Pr[r^{\prime}\geq 0]\cdot\prod_{i=1}^{\Omega(\log n)}\Pr[r^{\prime}\geq i\mid r^{\prime}\geq i-1]\leq\prod_{i=1}^{\Omega(\log n)}\frac{1}{2}=2^{-\Omega(\log n)}=\frac{1}{\mathrm{poly}(n)}.$
Corollary 19.9.
We have that $r=O(T\log n/k)$ with probability at least $1-1/\mathrm{poly}(n)$.
Proof.
By splitting up the sequence of updates $\sigma_{1},\dots,\sigma_{T}$ into $O(T/p(k))$ many subsequences of length $p(k)/2$, we can see that our algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ at most $O(\log n)$ many times during while handling each of these subsequences with probability at least $1-1/\mathrm{poly}(n)$. Applying a union bound, it follows that $r=O(\log n)\cdot O(T/p(k))\leq O(T\log n/k)$ with probability at least $1-1/\mathrm{poly}(n)$. ∎
It follows from Lemmas 19.5 and 19.9 that $R=\tilde{O}(T)+k\cdot\tilde{O}(T/k)=\tilde{O}(T)$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized recourse of our algorithm is $\tilde{O}(1)$ with high probability.
Update Time Analysis.
Let $r^{\star}$ denote the total number of times that the algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ across the sequence of updates $\sigma_{1},\dots,\sigma_{T}$. Note that we might have $r^{\star}>r$ if the algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ multiple times between two updates. The following lemma shows that this can not happen too many times.
Claim 19.10.
We have that $r^{\star}=O(r\log n)$ with probability at least $1-1/\mathrm{poly}(n)$.
Proof.
For each $i\in[T]$, let $r^{\star}_{i}$ denote the total number of times that our algorithm resets $\mathrm{\textsc{ALG}}^{\star}$ after handling the update $\sigma_{i}$. Note that $r_{i}\geq r_{i}$ and $r^{\star}=\sum_{i=1}^{T}r_{i}^{\star}$. Now, suppose that we have restarted $\mathrm{\textsc{ALG}}^{\star}$ $j$ times so far after handling the update $\sigma_{i}$, i.e. that $r^{\star}_{i}\geq j$. Since the algorithm uses fresh randomness, the probability that we restart it again (i.e. that $r_{i}^{\star}\geq j+1$) is at most $21/p(k)\leq 1/$ 2. Thus, it follows that
$\Pr[r^{\star}_{i}\geq\Omega(\log n)\mid r^{\star}_{i}\geq 1]=\prod_{j=1}^{\Omega(\log n)}\Pr[r^{\star}_{i}\geq j+1\mid r^{\star}_{i}\geq j]\leq 2^{-\Omega(\log n)}=\frac{1}{\mathrm{poly}(n)}.$
Since $r_{i}^{\star}=0$ if $r_{i}=0$, it follows from a union bound that $r^{\star}=O(r\log n)$ with probability at least $1-1/\mathrm{poly}(n)$. ∎
Lemma 19.11.
The total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot(T+kr^{\star})$.
Proof.
The total time taken to update our algorithm is the sum of (i) the time taken to maintain Sparsifier, (ii) the time taken to maintain our $O(\log n)$ copies of ALG, and (iii) the time taken to reset $\mathrm{\textsc{ALG}}^{\star}$. Since Sparsifier has an amortized update time of $\tilde{O}(1)$, this can be maintained in time $\tilde{O}(1)\cdot T$. Since Sparsifier has an amortized recourse of $\tilde{O}(1)$ and $\mathrm{\textsc{ALG}}^{\star}$ has an amortized update time of $\tilde{O}(k^{\epsilon})$, the total time taken to update the $O(\log n)$ copies of ALG maintained by our algorithm is $O(\log n)\cdot\tilde{O}(T)\cdot\tilde{O}(k^{\epsilon})=\tilde{O}(k^{\epsilon})\cdot T$. Finally, each time we reset $\mathrm{\textsc{ALG}}^{\star}$, it takes us $\tilde{O}(k^{1+\epsilon})$ time to reconstruct the data structure using fresh randomness. Thus, the total time spent resetting $\mathrm{\textsc{ALG}}^{\star}$ is $\tilde{O}(k^{1+\epsilon})\cdot r^{\star}$. ∎
It follows from Lemmas 19.11, 19.10 and 19.9 that the total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot T$ with probability at least $1-1/\mathrm{poly}(n)$. Thus, the amortized update time of our algorithm is $\tilde{O}(k^{\epsilon})$ with high probability.
20 Deferred Proofs
20.1 Proof of Lemma 15.1
According to the second property of a $t$-robust sequence, we have two cases. If Equation 40 holds, then $x_{j-1}=x_{j}$ and it is trivial that $\mathrm{dist}(x_{j-1},x_{j})\leq 2\cdot\lambda^{3j-1}$. Now, assume Equation 41 holds. We have
$\mathrm{dist}(x_{j},x_{j-1})^{2}$ $=\frac{1}{w(B_{j})}\cdot\sum_{x\in B_{j}}w(x)\cdot\mathrm{dist}(x_{j},x_{j-1})^{2}$
$\leq\frac{1}{w(B_{j})}\cdot\sum_{x\in B_{j}}w(x)\cdot(2\cdot\mathrm{dist}(x_{j},x)^{2}+2\cdot\mathrm{dist}(x_{j-1},x)^{2})$
$=\frac{1}{w(B_{j})}\cdot\left(2\cdot\textsc{cost}(B_{j},x_{j})+2\cdot\textsc{cost}(B_{j},x_{j-1})\right)$
$\leq\frac{4}{w(B_{j})}\cdot\textsc{cost}(B_{j},x_{j})\leq 4\cdot\lambda^{6j-2}.$
This concludes that $\mathrm{dist}(x_{j},x_{j-1})\leq 2\cdot\lambda^{3j-1}$ which proves the first property.
Now, assume $x\in B_{j-1}$ is arbitrary. We show that $x\in B_{j}$, which concludes $B_{j-1}\subseteq B_{j}$. Since $B_{j-1}\subseteq\textsc{Ball}(x_{j},\lambda^{3(j-1)+1})$, we have $\mathrm{dist}(x,x_{j})\leq\lambda^{3j-2}$. Hence,
$\mathrm{dist}(x,x_{j})\leq\mathrm{dist}(x,x_{j-1})+\mathrm{dist}(x_{j-1},x_{j})\leq\lambda^{3j-2}+2\cdot\lambda^{3j-1}\leq\lambda^{3j},$
where the last inequality follows from the fact that we assumed $\lambda$ is a large constant. Finally, since $\textsc{Ball}(x_{j},\lambda^{3j})\subseteq B_{j}$, we conclude that $x\in B_{j}$.
Property 3 follows from
$\mathrm{dist}(x_{0},x_{j})\leq\sum_{i=1}^{j}\mathrm{dist}(x_{i-1},x_{i})\leq 2\cdot\sum_{i=1}^{j}\lambda^{3i-1}\leq 2\cdot\frac{\lambda^{3(j+1)}-1}{\lambda(\lambda^{3}-1)}\leq 4\cdot\lambda^{3j-1},$
where the last inequality follows from $\lambda$ being a large constant.
20.2 Proof of Lemma 15.2
Since $(x_{0},x_{1},\ldots,x_{t})$ is $t$-robust, for every $1\leq j\leq t$, we know
$\textsc{cost}(B_{j},x_{j})\geq\textsc{cost}(B_{j},x_{j-1}).$ (68)
Assume $y\in X\setminus B_{j}$. We have
$\mathrm{dist}(y,x_{j})\geq\lambda^{3j}\geq(\lambda/4)\cdot\mathrm{dist}(x_{0},x_{j})\geq\mathrm{dist}(x_{0},x_{j}).$
The first inequality follows from $\textsc{Ball}(x_{j},\lambda^{3j})\subseteq B_{j}$, the second inequality follows from Lemma 15.1, and the last one follows from $\lambda$ being a large constant. Hence,
$2\cdot\mathrm{dist}(y,x_{j})=\mathrm{dist}(y,x_{j})+\mathrm{dist}(y,x_{j})\geq\mathrm{dist}(y,x_{j})+\mathrm{dist}(x_{0},x_{j})\geq\mathrm{dist}(y,x_{0}),$
which means
$\mathrm{dist}(y,x_{j})^{2}\geq\frac{1}{4}\cdot\mathrm{dist}(y,x_{0})^{2}.$ (69)
Finally, Since $B_{1}\subseteq\cdots\subseteq B_{i}\subseteq U$ (by Lemma 15.1) we can apply Equation 68 and Equation 69 repeatedly to get
$\textsc{cost}(U,x_{i})$ $=\textsc{cost}(B_{i},x_{i})+\textsc{cost}(U\setminus B_{i},x_{i})$
$\geq\textsc{cost}(B_{i},x_{i-1})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{i},x_{0})$
$=\textsc{cost}(B_{i-1},x_{i-1})+\textsc{cost}(B_{i}\setminus B_{i-1},x_{i-1})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{i},x_{0})$
$\geq\textsc{cost}(B_{i-1},x_{i-2})+\frac{1}{4}\cdot\textsc{cost}(B_{i}\setminus B_{i-1},x_{0})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{i},x_{0})$
$=\textsc{cost}(B_{i-1},x_{i-2})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{i-1},x_{0})$
[…]
$\geq\textsc{cost}(B_{1},x_{0})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{1},x_{0})\geq\frac{1}{4}\cdot\textsc{cost}(U,x_{0}).$
20.3 Proof of Lemma 15.3
Assume $S\leftarrow\texttt{Robustify}(W)$ is the center set returned by calling Robustify on $W$. According to Lemma 14.4, Robustify calls MakeRobust on each center at most once. Let $r$ be the number of centers in $W$ for which a call to MakeRobust has happened. Assume $W=\{w_{1},\ldots,w_{k}\}$ is the ordering by the time that MakeRobust is called on the centers $w_{1},w_{2},\ldots,w_{r}$ and the last elements are ordered arbitrarily. Assume also that $S=\{w^{\prime}_{1},\ldots,w^{\prime}_{k}\}$ where $w_{i}^{\prime}$ is obtained by the call MakeRobust on $w_{i}$ for each $1\leq i\leq r$ and $w_{i}^{\prime}=w_{i}$ for each $r+1\leq i\leq k$.
For every $1\leq j\leq r$, integer $t[w_{j}^{\prime}]$ is the smallest integer satisfying
$\lambda^{3t[w_{j}^{\prime}]}\geq\widehat{\mathrm{dist}}(w_{j},\{w_{1}^{\prime},\ldots,w_{j-1}^{\prime},w_{j+1},\ldots,w_{k}\})/\lambda^{7}.$
According to the guarantee of $\widehat{\mathrm{dist}}$ in Lemma 7.3, we conclude
$\lambda^{3t[w_{j}^{\prime}]}<\widehat{\mathrm{dist}}(w_{j},\{w_{1}^{\prime},\ldots,w_{j-1}^{\prime},w_{j+1},\ldots,w_{k}\})/\lambda^{4}\leq\mathrm{dist}(w_{j},\{w_{1}^{\prime},\ldots,w_{j-1}^{\prime},w_{j+1},\ldots,w_{k}\})/\lambda^{4}.$
Hence,
$\lambda^{3t[w_{j}^{\prime}]}\leq\mathrm{dist}(w_{j},w_{i}^{\prime})/\lambda^{4}\quad\forall i<j$ (70)
and
$\lambda^{3t[w_{j}^{\prime}]}\leq\mathrm{dist}(w_{j},w_{i})/\lambda^{4}\quad\forall i>j.$ (71)
Now, for every $1\leq i<j\leq r$, we conclude
$\lambda^{4}\cdot\lambda^{3t[w_{j}^{\prime}]}$$\leq$ $\mathrm{dist}(w_{i}^{\prime},w_{j})\leq\mathrm{dist}(w_{i},w_{j})+\mathrm{dist}(w_{i},w_{i}^{\prime})$
$\leq$ $\mathrm{dist}(w_{i},w_{j})+4\cdot\lambda^{3t[w_{i}^{\prime}]-1}\leq(1+4/\lambda^{5})\cdot\mathrm{dist}(w_{i},w_{j}).$
The first inequality holds by Equation 70, the third one holds by Lemma 15.1 (note that $w_{i}^{\prime}=$ MakeRobust$(w_{i})$) and the last one holds by Equation 71 (by exchanging $i$ and $j$ in this equation since $i<j$). Since $\lambda$ is a large number, we conclude that for each $i<j$, the following holds
$\mathrm{dist}(w_{i},w_{j})>\lambda^{4}/(1+4/\lambda^{5})\cdot\lambda^{3t[w_{j}^{\prime}]}>2\cdot\lambda^{3t[w_{j}^{\prime}]}.$
The same inequality holds for each $i>j$ by Equation 71 and $\lambda$ being large. As a result, $\mathrm{dist}(w_{i},w_{j})>2\cdot\lambda^{3t[w_{j}^{\prime}]}$ for all $i\neq j$. This concludes that all the points around $w_{j}$ of distance at most $\lambda^{3t[w_{j}^{\prime}]}$ belong to the cluster of center $w_{j}$ in solution $W$. Denote by $C_{w_{j}}$ the cluster of $w_{j}$ in solution $W$, i.e., $C_{w_{j}}:=\{x\in X\mid\mathrm{dist}(x,w_{j})=\mathrm{dist}(x,W)\}$ (breaking the ties arbitrarily). We conclude
$\textsc{Ball}(w_{j},\lambda^{3t[w_{j}^{\prime}]})\subseteq C_{w_{j}},$
for every $1\leq j\leq r$. It follows from Lemma 15.2 that
$\textsc{cost}(C_{w_{j}},w_{j}^{\prime})\leq 4\cdot\textsc{cost}(C_{w_{j}},w_{j})\quad\forall 1\leq j\leq r.$
Finally, since $C_{w_{j}}$ for $1\leq j\leq k$ form a partition of $X$, we conclude
$\textsc{cost}(X,S)$ $=$ $\sum_{j=1}^{k}\textsc{cost}(C_{w_{j}},S)\leq\sum_{j=1}^{k}\textsc{cost}(C_{w_{j}},w_{j}^{\prime})$
$=$ $\sum_{j=1}^{r}\textsc{cost}(C_{w_{j}},w_{j}^{\prime})+\sum_{j=r+1}^{k}\textsc{cost}(C_{w_{j}},w_{j}^{\prime})$
$\leq$ $4\cdot\sum_{j=1}^{r}\textsc{cost}(C_{w_{j}},w_{j})+\sum_{j=r+1}^{k}\textsc{cost}(C_{w_{j}},w_{j})$
$\leq$ $4\cdot\sum_{j=1}^{k}\textsc{cost}(C_{w_{j}},w_{j})=4\cdot\textsc{cost}(X,W).$
20.4 Proof of Lemma 15.5
Denote the set of points in the cluster of $v\in V$ in the center set $V$ by $C_{v}$, i.e. $C_{v}=\{x\in X\mid\mathrm{dist}(x,v)=\mathrm{dist}(x,V)\}$. If $\mathrm{dist}(u,v)=0$, the statement is trivial. Now assume $\mathrm{dist}(u,v)\geq 1$. Since $S$ is robust, then $u$ is $t$-robust where $t$ is the smallest integer such that $\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}\geq(\lambda^{20}/\lambda^{10})\cdot\mathrm{dist}(u,v)=\lambda^{10}\cdot\mathrm{dist}(u,v)$. The second inequality is because $(u,v)$ is well-separated pair. Assume $t^{*}$ is the integer satisfying
$\lambda^{10}\cdot\mathrm{dist}(u,v)\leq\lambda^{3t^{*}}<\lambda^{13}\cdot\mathrm{dist}(u,v).$ (72)
Note that $t^{*}\geq 1$ (since we assumed $\mathrm{dist}(u,v)\geq 1$). We have $t^{*}\leq t$ and $u$ is $t$-robust, then it is also $t^{*}$-robust which means there is a $t^{*}$-robust sequence $(x_{0},x_{1},\ldots,x_{t^{*}})$ such that $x_{0}=u$. Consider approximate balls $B_{i}$ around $x_{i}$ in the definition of the robust sequence.
Claim 20.1.
For each $0\leq i\leq t^{*}$, we have $B_{i}\subseteq C_{v}$.
Proof.
Assume $y\in B_{i}$. Then,
$\mathrm{dist}(y,v)$$\leq$ $\mathrm{dist}(y,x_{i})+\mathrm{dist}(x_{i},u)+\mathrm{dist}(u,v)=\mathrm{dist}(y,x_{i})+\mathrm{dist}(x_{i},x_{0})+\mathrm{dist}(u,v)$
$\leq$ $\lambda^{3i+1}+4\cdot\lambda^{3i-1}+\lambda^{3t^{*}-10}\leq(2\lambda)\cdot\lambda^{3t^{*}}.$
Second inequality follows from $y\in B_{i}\subseteq\textsc{Ball}(x_{i},\lambda^{3i+1})$, Lemma 15.1 and Equation 72. The last inequality follows from $\lambda$ being a large constant. Now, for every $v^{\prime}\in V-v$ we have
$\mathrm{dist}(v,v^{\prime})\geq\mathrm{dist}(v,V-v)\geq\lambda^{20}\cdot\mathrm{dist}(u,v)>\lambda^{3t^{*}+7}\geq(4\lambda)\cdot\lambda^{3t^{*}}\geq 2\cdot\mathrm{dist}(y,v),$
where the second inequality is by $(u,v)$ being well-separated, the third one is by Equation 72, and the fourth one is since $\lambda$ is a large constant. This concludes $\mathrm{dist}(y,v^{\prime})\geq\mathrm{dist}(v,v^{\prime})-\mathrm{dist}(y,v)>\mathrm{dist}(y,v)$ which means $v$ is the closest center to $y$ in $T$. So, $y\in C_{v}$. Since $y\in B_{i}$ was arbitrary, we conclude that $B_{i}\subseteq C_{v}$. ∎
Now, according to the definition of a $t$-robust sequence, we have two cases as follows.
Case 1.
Equation 40 holds for $i=t^{*}$. So, $\textsc{cost}(B_{t^{*}},x_{t^{*}})/w(B_{t^{*}})\geq\lambda^{6t^{*}-4}$ and $x_{t^{*}-1}=x_{t^{*}}$. We conclude
$\mathrm{dist}(v,x_{t^{*}})$ $=$ $\mathrm{dist}(v,x_{t^{*}-1})\leq\mathrm{dist}(v,x_{0})+\mathrm{dist}(x_{0},x_{t^{*}-1})=\mathrm{dist}(v,u)+\mathrm{dist}(x_{0},x_{t^{*}-1})$
$\leq$ $\lambda^{3t^{*}-10}+4\cdot\lambda^{3(t^{*}-1)-1}\leq\lambda^{3t^{*}-3}.$
The second inequality holds by Lemma 15.1 and Equation 72, and the last one holds since $\lambda$ is a large constant. So,
$\mathrm{dist}(v,x_{t^{*}})\leq\lambda^{3t^{*}-3}.$ (73)
For every $y\in C_{v}-B_{t^{*}}$, we have $\mathrm{dist}(y,v)\geq\mathrm{dist}(x_{t^{*}},y)-\mathrm{dist}(v,x_{t^{*}})\geq\lambda^{3t^{*}}-\mathrm{dist}(v,x_{t^{*}})\geq\mathrm{dist}(v,x_{t^{*}})$. The second inequality follows by $y\in C_{v}-B_{t^{*}}\subseteq C_{v}-\textsc{Ball}(x_{t^{*}},\lambda^{3t^{*}})$, and the last inequality follows by Equation 73 and $\lambda$ being a large constant. So, $\mathrm{dist}(y,v)\geq(\mathrm{dist}(y,v)+\mathrm{dist}(v,x_{t^{*}}))/2\geq\mathrm{dist}(y,x_{t^{*}})/2$, which implies
$\textsc{cost}(C_{v}-B_{t^{*}},v)\geq\frac{\textsc{cost}(C_{v}-B_{t^{*}},x_{t^{*}})}{4}.$ (74)
We also have
$\textsc{cost}(B_{t^{*}},x_{t^{*}})/w(B_{t^{*}})\geq\lambda^{6t^{*}-4}\geq 4\cdot\mathrm{dist}(v,x_{t^{*}})^{2},$ (75)
by the assumption of this case, Equation 73, and $\lambda$ begin a large constant. Now, we conclude
$\textsc{cost}(C_{v},v)$ $=$ $\textsc{cost}(B_{t^{*}},v)+\textsc{cost}(C_{v}-B_{t^{*}},v)$
$=$ $w(B_{t^{*}})\cdot\textsc{cost}(B_{t^{*}},v)/w(B_{t^{*}})+\textsc{cost}(C_{v}-B_{t^{*}},v)$
$\geq$ $w(B_{t^{*}})\cdot(\textsc{cost}(B_{t^{*}},x_{t^{*}})/(2w(B_{t^{*}}))-\mathrm{dist}(v,x_{t^{*}})^{2})+\textsc{cost}(C_{v}-B_{t^{*}},v)$
$\geq$ $w(B_{t^{*}})\cdot\frac{\textsc{cost}(B_{t^{*}},x_{t^{*}})}{4\cdot w(B_{t^{*}})}+\textsc{cost}(C_{v}-B_{t^{*}},v)$
$=$ $\frac{\textsc{cost}(B_{t^{*}},x_{t^{*}})}{4}+\textsc{cost}(C_{v}-B_{t^{*}},v)$
$\geq$ $\frac{\textsc{cost}(B_{t^{*}},x_{t^{*}})}{4}+\frac{\textsc{cost}(C_{v}-B_{t^{*}},x_{t^{*}})}{4}=\frac{\textsc{cost}(C_{v},x_{t^{*}})}{4}.$
The first equality holds by $B_{t^{*}}\subseteq C_{v}$ (Claim 20.1). The first inequality holds by $\mathrm{dist}(y,v)^{2}\geq\frac{1}{2}\cdot\mathrm{dist}(y,x_{t^{*}})^{2}-\mathrm{dist}(v,x_{t^{*}})^{2}$ for all $y\in B_{t^{*}}$, the second inequality holds by Equation 75 and the last inequality holds by $B_{t^{*}}\subseteq C_{v}$ (Claim 20.1) and Equation 74. Finally, since $B_{t^{*}}\subseteq C_{v}$ (by Claim 20.1), we can apply Lemma 15.2 to get
$\textsc{cost}(C_{v},u)=\textsc{cost}(C_{v},x_{0})\leq 4\cdot\textsc{cost}(C_{v},x_{t^{*}})\leq 16\cdot\textsc{cost}(C_{v},v)\leq 4\cdot(p_{\epsilon})^{3}\cdot\textsc{cost}(C_{v},v).$
Case 2.
Equation 41 holds for $i=t^{*}$. So,
[…]
In this case, for every $y\in C_{v}-B_{t^{*}}$,
$\mathrm{dist}(y,x_{t^{*}-1})\geq\mathrm{dist}(y,x_{t^{*}})-\mathrm{dist}(x_{t^{*}},x_{t^{*}-1})\geq\lambda^{3t^{*}}-\lambda^{3t^{*}-1}\geq\lambda^{3t^{*}-1}.$
The second inequality holds by $q\in C_{v}-B_{t^{*}}\subseteq C_{v}-\textsc{Ball}(x_{t^{*}},\lambda^{3t^{*}})$ and Lemma 15.1, and the last one holds by $\lambda$ being a large constant. So,
$\mathrm{dist}(y,x_{t^{*}-1})\geq\lambda^{3t^{*}-1}.$ (76)
We also have
$\mathrm{dist}(v,x_{t^{*}-1})$$\leq$ $\mathrm{dist}(v,u)+\mathrm{dist}(u,x_{t^{*}-1})$
$=$ $\mathrm{dist}(v,u)+\mathrm{dist}(x_{0},x_{t^{*}-1})\leq\lambda^{3t^{*}-10}+4\cdot\lambda^{3(t^{*}-1)-1}\leq\lambda^{3t^{*}-3}.$
The second inequality holds by Equation 72 and Lemma 15.1, and the last one holds by $\lambda$ being a large constant. So,
$\mathrm{dist}(v,x_{t^{*}-1})\leq\lambda^{3t^{*}-3}.$ (77)
Combining Equation 76, Equation 77 and $\lambda$ begin a large constant, we have
$\mathrm{dist}(v,x_{t^{*}-1})\leq\lambda^{3t^{*}-3}\leq\frac{1}{2}\cdot\lambda^{3t^{*}-1}\leq\frac{1}{2}\cdot\mathrm{dist}(y,x_{t^{*}-1}),$
which concludes
$\frac{\mathrm{dist}(y,x_{t^{*}-1})}{\mathrm{dist}(y,v)}\leq\frac{\mathrm{dist}(y,x_{t^{*}-1})}{\mathrm{dist}(y,x_{t^{*}-1})-\mathrm{dist}(x_{t^{*}-1},v)}\leq\frac{\mathrm{dist}(y,x_{t^{*}-1})}{\mathrm{dist}(y,x_{t^{*}-1})-\frac{1}{2}\cdot\mathrm{dist}(y,x_{t^{*}-1})}=2.$
So, we have $\mathrm{dist}(y,x_{t^{*}-1})\leq 2\cdot\mathrm{dist}(y,v)$. Using this inequality for all $y\in C_{v}-B_{t^{*}}$, we conclude
$\textsc{cost}(C_{v}-B_{t^{*}},x_{t^{*}-1})\leq 4\cdot\textsc{cost}(C_{v}-B_{t^{*}},v).$ (78)
We also have
$\textsc{cost}(B_{t^{*}},x_{t^{*}-1})\leq(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{t^{*}})\leq(p_{\epsilon})^{3}\cdot\textsc{cost}(B_{t^{*}},v).$ (79)
The first inequality holds by the assumption of this case. Hence,
$\textsc{cost}(C_{v},x_{t^{*}-1})$ $=$ $\textsc{cost}(B_{t^{*}},x_{t^{*}-1})+\textsc{cost}(C_{v}-B_{t^{*}},x_{t^{*}-1})$
$\leq$ $(p_{\epsilon})^{3}\cdot\textsc{cost}(B_{t^{*}},v)+4\cdot\textsc{cost}(C_{v}-B_{t^{*}},v)\leq(p_{\epsilon})^{3}\cdot\textsc{cost}(C_{v},v).$
The equality holds since $B_{t^{*}}\subseteq C_{v}$ (by Claim 20.1) and the first inequality holds by Equation 78 and Equation 79. Finally, applying Lemma 15.2 implies
$\textsc{cost}(C_{v},u)=\textsc{cost}(C_{v},x_{0})\leq 4\cdot\textsc{cost}(C_{v},x_{t^{*}-1})\leq 4\cdot(p_{\epsilon})^{3}\cdot\textsc{cost}(C_{v},v).$
In both cases, we showed that $\textsc{cost}(C_{v},u)\leq 4\cdot(p_{\epsilon})^{3}\cdot\textsc{cost}(C_{v},v)$, which completes the proof.
20.5 Proof of Lemma 15.6
Consider the standard LP relaxation for the weighted $k$-means problem. We consider the set of potential centers to open as $S$, and the goal is to open at most $k-(m-r)/4$ many centers, i.e., find $\tilde{S}\subseteq S$.
$\min$ $\sum_{p\in X}\sum_{u\in S}w(p)\cdot\mathrm{dist}(u,p)^{2}\cdot x_{up}$
s.t. $\quad x_{up}\leq y_{u}$ $\forall u\in S,p\in X$
$\sum_{u\in S}x_{up}\geq 1$ $\forall p\in X$
$\sum_{u\in S}y_{u}\leq k-(m-r)/4$
$x_{up},y_{u}\geq 0$ $\forall u\in S,p\in X$
Now, we explain how to construct a fractional solution for this LP.
Fractional Opening of Centers. Consider the projection $\pi_{S}:V\rightarrow S$ function. Assume $S=S_{I}+S_{F}$ is a partition of $S$ where $S_{I}$ contains those centers $u\in S$ satisfying at least one of the following conditions:
  • $u$ forms a well-separated pair with one center in $V$.
  • $|\pi^{-1}_{S}(u)|\geq 2$.
For every $u\in S_{I}$, set $y_{u}=1$ and for every $u\in S_{F}$ set $y_{u}=1/2$. First, we show that
$\sum_{u\in S}y_{u}\leq k-(m-r)/4.$
Each center $u\in S$ that forms a well-separated pair with a center $v\in V$ has $|\pi^{-1}_{S}(u)|\geq 1$ since $u$ must be the closest center to $v$ in $S$. Since the number of well-separated pairs is $k-m$, we have
$k+r\geq|V|=\sum_{u\in S}|\pi^{-1}(u)|\geq\sum_{u\in S_{I}}|\pi^{-1}(u)|\geq k-m+2\cdot\left(|S_{I}|-(k-m)\right).$
Hence,
$|S_{I}|\leq\frac{k+r+(k-m)}{2}=k-\frac{m-r}{2}.$
Finally, we conclude
$\sum_{u\in S}y_{u}$ $=$ $\sum_{u\in S_{I}}y_{u}+\sum_{u\in S_{F}}y_{u}$
$\leq$ $k-\frac{m-r}{2}+\frac{1}{2}\cdot\frac{m-r}{2}$
$=$ $k-\frac{m-r}{4}.$
Fractional Assignment of Points. For every $p\in X$, assume $v_{p}=\pi_{V}(p)$ is the closest center to $p$ in $V$ and $u_{p}=\pi_{S}(v_{p})$ is the closest center in $S$ to $v_{p}$. We have three cases:
  • If $y_{u_{p}}=1$, then set $x_{u_{p}p}=1$. The cost of this assignment would be $w(p)\cdot\mathrm{dist}(u_{p},p)^{2}$.
  • If $y_{u_{p}}=1/2$ and there is a center $u^{\prime}_{p}\in S-u_{p}$ such that $\mathrm{dist}(u_{p},u^{\prime}_{p})\leq\lambda^{20}\cdot\mathrm{dist}(u_{p},v_{p})$, then set $x_{u_{p}p}=x_{u_{p}^{\prime}p}=1/2$. Note that $u_{p}^{\prime}\neq u_{p}$ which means point $p$ is assigned one unit to centers. The cost of this assignment would be
    $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+\mathrm{dist}(p,u_{p}^{\prime})^{2}\right)$
    $\leq$ $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+2\cdot\mathrm{dist}(p,u_{p})^{2}+2\cdot\mathrm{dist}(u^{\prime}_{p},u_{p})^{2}\right)$
    $\leq$ $w(p)\left(\frac{3}{2}\cdot\mathrm{dist}(p,u_{p})^{2}+\lambda^{40}\cdot\mathrm{dist}(u_{p},v_{p})^{2}\right).$
    In the first inequality, we have used the fact that $\mathrm{dist}(x,y)^{2}\leq 2\cdot\mathrm{dist}(x,z)^{2}+2\cdot\mathrm{dist}(z,y)^{2}$ for all $x,y,z\in\mathbb{R}^{d}$.
  • If $y_{u_{p}}=1/2$ and the previous case does not hold, then since $(u_{p},v_{p})$ is not a well-separated pair, there is a center $v^{\prime}_{p}\in V-v_{p}$ such that $\mathrm{dist}(v_{p},v^{\prime}_{p})\leq\lambda^{20}\cdot\mathrm{dist}(u_{p},v_{p})$. Let $u_{p}^{\prime}=\pi_{S}(v^{\prime}_{p})$ and set $x_{u_{p}p}=x_{u^{\prime}_{p}p}=1/2$. First, we show that $u^{\prime}_{p}\neq u_{p}$. Since $y_{u_{p}}=1/2$, we have $u_{p}\in S_{F}$ which concludes $|\pi^{-1}_{S}(u_{p})|\leq 1$. We also know that $\pi_{S}(v_{p})=u_{p}$. So, $v_{p}$ is the only center in $V$ mapped to $u_{p}$ which implies $\pi_{S}(v^{\prime}_{p})\neq u_{p}$ or $u_{p}^{\prime}\neq u_{p}$ (note that $v_{p}^{\prime}\neq v_{p}$). So, point $p$ is assigned one unit to centers. The cost of this assignment would be
    $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+\mathrm{dist}(p,u^{\prime}_{p})^{2}\right)$
    $\leq$ $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+(\mathrm{dist}(p,u_{p})+\mathrm{dist}(u_{p},v^{\prime}_{p})+\mathrm{dist}(v^{\prime}_{p},u^{\prime}_{p}))^{2}\right)$
    $\leq$ $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+(\mathrm{dist}(p,u_{p})+\mathrm{dist}(u_{p},v^{\prime}_{p})+\mathrm{dist}(v^{\prime}_{p},u_{p}))^{2}\right)$
    $\leq$ $\frac{1}{2}w(p)\left(\mathrm{dist}(p,u_{p})^{2}+2\cdot\mathrm{dist}(p,u_{p})^{2}+8\cdot\mathrm{dist}(u_{p},v^{\prime}_{p})^{2}\right)$
    $=$ $w(p)\left(\frac{3}{2}\cdot\mathrm{dist}(p,u_{p})^{2}+4\cdot\mathrm{dist}(u_{p},v_{p}^{\prime})^{2}\right)$
    $\leq$ $w(p)\left(\frac{3}{2}\cdot\mathrm{dist}(p,u_{p})^{2}+4\cdot(\mathrm{dist}(u_{p},v_{p})+\mathrm{dist}(v_{p},v_{p}^{\prime}))^{2}\right)$
    $\leq$ $w(p)\left(\frac{3}{2}\cdot\mathrm{dist}(p,u_{p})^{2}+4(\lambda^{20}+1)^{2}\ \mathrm{dist}(u_{p},v_{p})^{2}\right).$
    The second inequality holds because of the choice of $u^{\prime}_{p}=\pi_{S}(v_{p}^{\prime})$ and the last inequality holds since $\mathrm{dist}(v_{p},v_{p}^{\prime})\leq\lambda^{20}\cdot\mathrm{dist}(u_{p},v_{p})$.
Bounding the Cost. Assume $u_{p}^{*}=\pi_{S}(p)$ for each $p\in X$. We have
$\mathrm{dist}(u_{p},v_{p})\leq\mathrm{dist}(u_{p}^{*},v_{p})\leq\mathrm{dist}(u_{p}^{*},p)+\mathrm{dist}(p,v_{p}).$
The first inequality is by the choice of $u_{p}=\pi_{S}(v_{p})$. As a result,
$\sum_{p\in X}w(p)\ \mathrm{dist}(u_{p},v_{p})^{2}\leq\sum_{p\in X}w(p)\left(2\cdot\mathrm{dist}(u^{*}_{p},p)^{2}+2\cdot\mathrm{dist}(p,v_{p})^{2}\right)=2\cdot\textsc{cost}(X,S)+2\cdot\textsc{cost}(X,V).$ (80)
In each of the cases of fractional assignments of points to centers, the cost of assigning a point $p\in C_{v}(V,X)$ is at most $w(p)\left(\frac{3}{2}\cdot\mathrm{dist}(p,u_{p})^{2}+4(\lambda^{20}+1)^{2}\cdot\mathrm{dist}(u_{p},v_{p})^{2}\right)$. As a result, the total cost of this assignment is upper bounded by
$\sum_{p\in P}w(p)\left(\frac{3}{2}\mathrm{dist}(p,u_{p})^{2}+4(\lambda^{20}+1)^{2}\cdot\mathrm{dist}(u_{p},v_{p})^{2}\right)$
$\leq$ $\sum_{p\in X}w(p)\left(\frac{3}{2}\cdot(\mathrm{dist}(p,v_{p})+\mathrm{dist}(u_{p},v_{p}))^{2}+4(\lambda^{20}+1)^{2}\ \mathrm{dist}(u_{p},v_{p})^{2}\right)$
$\leq$ $\sum_{p\in X}w(p)\left(3\cdot\mathrm{dist}(p,v_{p})^{2}+4(\lambda^{20}+2)^{2}\ \mathrm{dist}(u_{p},v_{p})^{2}\right)$
$\leq$ $3\cdot\textsc{cost}(X,V)+8(\lambda^{20}+2)^{2}\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right)$
$\leq$ $9\lambda^{40}\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right).$
The third inequality follows by Equation 80.
Finally, since the integrality gap of the LP relaxation is known to be at most $9$ [ANFSW20]. [^23] Note that although it is not explicitly mentioned, the result of [ANFSW20] works for the weighted case. One way to show this, which only uses the unweighted case as a black box is as follows. First, we can round all the real weights to some rational weights with an additive error of at most $\epsilon$. This will not change the feasible solutions and change the optimal objective function only by a factor of $1\pm\epsilon$. Then, we scale up all the weights to have integer weights. Now, we replace every point $p$ of weight $w(p)$ with $w(p)$ many unweighted points $p_{1},p_{2},\cdots,p_{w(p)}$. It is easy to find a correspondence between any optimal solution (either fractional or integral) of the original weighted LP and any optimal solution of the new unweighted LP. Hence, we conclude that the integrality gap of the weighted LP is also at most $9+O(\epsilon)$. Since this is correct for every $\epsilon>0$, we conclude that the integrality gap is at most $9$. As a result, there exist an integral solution whose cost is at most $81\lambda^{40}\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right)$, and this solution opens at most $k-\lfloor\frac{m-r}{4}\rfloor$ centers which completes the proof.
20.6 Proof of Lemma 15.7
Assume $S^{\star}$ is the optimal $k$-means solution on $X$. Define $S^{\prime}:=\{\pi_{C}(s)\mid s\in S^{\star}\}$ be the projection of $S^{\star}$ on $C$. Let $x\in X$ be arbitrary. Let $s^{\star}=\pi_{S^{\star}}(x)$, $s^{\prime}=\pi_{C}(s^{\star})\in S^{\prime}$, and $s=\pi_{C}(x)$. We have
$\mathrm{dist}(x,S^{\prime})^{2}$ $\leq\mathrm{dist}(x,s^{\prime})^{2}\leq(\mathrm{dist}(x,s^{\star})+\mathrm{dist}(s^{\prime},s^{\star}))^{2}\leq(\mathrm{dist}(x,s^{\star})+\mathrm{dist}(s,s^{\star}))^{2}$
$\leq(2\cdot\mathrm{dist}(x,s^{\star})+\mathrm{dist}(s,x))^{2}\leq 8\cdot\mathrm{dist}(x,s^{\star})^{2}+2\cdot\mathrm{dist}(s,x)^{2}.$
As a result, $\mathrm{OPT}_{k}^{C}(X)\leq\textsc{cost}(X,S^{\prime})\leq 2\cdot\textsc{cost}(X,C)+8\cdot\mathrm{OPT}_{k}(X)$.
20.7 Proof of Lemma 15.8
Assume $S\subseteq[\Delta]^{d}$ of size at most $k$ is such that $\textsc{cost}(X,S)=\mathrm{OPT}_{k}(X)$. Define $S^{\prime}=S+(X\oplus X^{\prime})$. Obviously, $S^{\prime}$ is a feasible solution for the $(k+s)$-means problem on $X^{\prime}$ and since $S^{\prime}$ contains $X\oplus X^{\prime}$, we conclude
$\mathrm{OPT}_{k+s}(X^{\prime})\leq\textsc{cost}(X^{\prime},S^{\prime})=\textsc{cost}(X^{\prime},S+(X\oplus X^{\prime}))\leq\textsc{cost}(X,S)=\mathrm{OPT}_{k}(X).$
20.8 Proof of Lemma 15.9
Consider the LP relaxation for the $k$-means problem on $X$ for each $k$ as follows.
$\min$ $\sum_{p\in X}\sum_{c\in[\Delta]^{d}}w(p)\cdot\mathrm{dist}(c,p)^{2}\cdot x_{cp}$
s.t. $\quad x_{cp}\leq y_{c}$ $\forall c\in[\Delta]^{d},p\in X$
$\sum_{c\in[\Delta]^{d}}x_{cp}\geq 1$ $\forall p\in X$
$\sum_{c\in[\Delta]^{d}}y_{c}\leq k$
$x_{cp},y_{c}\geq 0$ $\forall c\in[\Delta]^{d},p\in X$
Denote the cost of the optimal fractional solution for this LP by $\text{FOPT}_{k}$. Since the space $X$ is fixed here, we denote $\mathrm{OPT}_{k}(X)$ by $\mathrm{OPT}_{k}$. It is known that the integrality gap of this relaxation is at most $9$ [ANFSW20]. So, for every $k$ we have
$\text{FOPT}_{k}\leq\mathrm{OPT}_{k}\leq 9\ \text{FOPT}_{k}.$ (81)
Claim 20.2.
For every $k_{1}$, $k_{2}$ and $0\leq\alpha,\beta\leq 1$ such that $\alpha+\beta=1$, we have
$\text{FOPT}_{\alpha k_{1}+\beta k_{2}}\leq\alpha\ \text{FOPT}_{k_{1}}+\beta\ \text{FOPT}_{k_{2}}.$
Proof.
Assume optimal fractional solutions $(x^{*},y^{*})$ and $(z^{*},t^{*})$ for above LP relaxation of $k_{1}$ and $k_{2}$-means problems respectively. It is easy to verify that $(\alpha x^{*}+\beta z^{*},\alpha y^{*}+\beta t^{*})$ is a feasible solution for fractional $(\alpha k_{1}+\beta k_{2})$-means problem, whose cost is $\alpha\ \text{FOPT}_{k_{1}}+\beta\ \text{FOPT}_{k_{2}}$, which concludes the claim. ∎
Now, plug $k_{1}=k-r$, $k_{2}=k+r/(36\eta)$, $\alpha=1/(36\eta)$ and $\beta=1-\alpha$ in the claim. We have
$\alpha k_{1}+\beta k_{2}=\frac{1}{36\eta}(k-r)+\left(1-\frac{1}{36\eta}\right)\left(k+\frac{r}{36\eta}\right)=k-\frac{r}{(36\eta)^{2}}\leq k.$
As a result,
$\text{FOPT}_{k}\leq\text{FOPT}_{\alpha k_{1}+\beta k_{2}}\leq\alpha\ \text{FOPT}_{k_{1}}+\beta\ \text{FOPT}_{k_{2}}.$
Together with Equation 81, we have
$\mathrm{OPT}_{k}\leq 9\alpha\ \mathrm{OPT}_{k_{1}}+9\beta\ \mathrm{OPT}_{k_{2}}.$
We also have the assumption that $\mathrm{OPT}_{k_{1}}=\mathrm{OPT}_{k-r}\leq\eta\ \mathrm{OPT}_{k}$, which implies
$\mathrm{OPT}_{k}\leq 9\alpha\eta\ \mathrm{OPT}_{k}+9\beta\ \mathrm{OPT}_{k_{2}}.$
Finally
$\mathrm{OPT}_{k}\leq\frac{9\beta}{1-9\alpha\eta}\cdot\mathrm{OPT}_{k_{2}}\leq 12\ \mathrm{OPT}_{k_{2}}\leq 12\ \mathrm{OPT}_{k+\lfloor r/(36\eta)\rfloor}.$
The second inequality follows from
$\frac{9\beta}{1-9\alpha\eta}\leq\frac{9}{1-9\alpha\eta}=\frac{9}{1-1/4}=12.$
20.9 Proof of Lemma 14.4
Assume a MakeRobust call to a center $w\in W^{\prime}$ is made and it is replaced by $w_{0}$. Let $W^{\text{old}}$ be the set of centers just before this call to MakeRobust is made on $w^{\prime}$. At this point of time, $t[w_{0}]$ is the smallest integer satisfying $\lambda^{3t[w_{0}]}\geq\widehat{\mathrm{dist}}(w,W^{\text{old}}-w)/\lambda^{7}$. For the sake of contradiction, assume that another call to MakeRobust is made on $w_{0}$. We assume that this pair $w$, $w_{0}$ is the first pair for which this occurs. Let $w^{\prime}\in W^{\text{old}}-w$ be such that $\mathrm{dist}(w,\mathcal{W}^{\text{old}}-w)=\mathrm{dist}(w,w^{\prime})$. Then,

### v2

@@
 Along the way, we design several novel geometric data structures that are of independent interest.
-Specifically, one of our main contributions is designing the first consistent hashing scheme [Czumaj, Jiang, Krauthgamer, Veselý, Yang; FOCS’22] that achieves $\text{poly}(d)$ running time per point evaluation with competitive parameters.
+Specifically, one of our main contributions is designing the first consistent hashing scheme [Czumaj, Jiang, Krauthgamer, Veselý, Yang; FOCS’22] that achieves $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters.
 Part I Extended Abstract
@@
 Section 4 presents an overview of our geometric data structures.
-One of our key contributions here is the first consistent hashing scheme [CJK+22] that achieves $\mathrm{poly}(d)$ running time per point evaluation with competitive parameters.
+One of our key contributions here is the first consistent hashing scheme [CJK+22] that achieves $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters.
 This new consistent hashing is crucial to ensure that our geometric data structures work against an adaptive adversary.
@@
 Then,
+$\lambda^{3t[w_{0}]}\geq\widehat{\mathrm{dist}}(w,W^{\text{old}}-w)/\lambda^{7}\geq\mathrm{dist}(w,w^{\prime})/\lambda^{7},$ (82)
+and
+$\lambda^{3t[w_{0}]}<\widehat{\mathrm{dist}}(w,W^{\text{old}}-w)/\lambda^{4}\leq\mathrm{dist}(w,w^{\prime})\cdot p_{\epsilon}/\lambda^{4},$ (83)
+by definition of $t[w_{0}]$ and the guarantee about $\widehat{\mathrm{dist}}(w,W^{\text{old}}-w)$ in Lemma 7.3. So, by Lemma 15.1, we have
+$\mathrm{dist}(w_{0},w)\leq 4\cdot\lambda^{3t[w_{0}]-1}<\mathrm{dist}(w,w^{\prime})\cdot 4p_{\epsilon}/\lambda^{5}.$ (84)
+Assume $W^{\text{new}}$ is the set of centers just after the call to MakeRobust is made on $w_{0}$.
+Since we assumed $w,w_{0}$ is the first pair that another call to MakeRobust on $w_{0}$ is made, we have one of the two following cases.
+Case 1.
+$w^{\prime}\in W^{\text{new}}$. In this case, we have
+$\mathrm{dist}(w_{0},W^{\text{new}}-w_{0})$ $\leq\mathrm{dist}(w_{0},w^{\prime})\leq\mathrm{dist}(w_{0},w)+\mathrm{dist}(w,w^{\prime})\leq(4p_{\epsilon}/\lambda^{5}+1)\cdot\mathrm{dist}(w,w^{\prime})$
+$<2\cdot\mathrm{dist}(w,w^{\prime})$ (85)
+The first inequality holds since $w^{\prime}$ is present in $W^{\text{new}}$, the second inequality is triangle inequality, the third inequality holds by Equation 84 ,and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number.
+According to 6 in Algorithm 7, we must have $t[w_{0}]<t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}$, as otherwise we should not made a call to MakeRobust on $w_{0}$.
+Hence,
+$\lambda^{3t[w_{0}]}$ $<\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}\leq\mathrm{dist}(w_{0},W^{\text{new}}-w_{0})\cdot p_{\epsilon}/\lambda^{10}$
+$\leq\mathrm{dist}(w,w^{\prime})\cdot 2\cdot p_{\epsilon}/\lambda^{10}<\mathrm{dist}(w,w^{\prime})/\lambda^{7},$
+where the second inequality holds by the guarantee of $\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})$ in Lemma 7.3, the third inequality holds by Section 20.9, and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number.
+This is in contradiction with Equation 82.
+Case 2.
+There is $w^{\prime}_{0}\in W^{\text{new}}$ that replaced $w^{\prime}$ by a single call to MakeRobust.
+When MakeRobust is called on $w^{\prime}$, the algorithm selects the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w^{\prime},W^{\text{mid}}-w^{\prime})/\lambda^{7}$ and sets $t[w_{0}^{\prime}]\leftarrow t$, where $W^{\text{mid}}$ is the status of the center set when a call to $\texttt{MakeRobust}(w^{\prime})$ is made.
+At that point in time, $w_{0}$ was present in $W^{\text{mid}}$, which concludes,
+$\mathrm{dist}(w^{\prime},w_{0}^{\prime})$ $\leq 4\cdot\lambda^{3t[w_{0}^{\prime}]-1}<\widehat{\mathrm{dist}}(w^{\prime},W^{\text{mid}}-w^{\prime})\cdot 4/\lambda^{5}\leq\mathrm{dist}(w^{\prime},W^{\text{mid}}-w^{\prime})\cdot 4p_{\epsilon}/\lambda^{5}$
+$\leq\mathrm{dist}(w^{\prime},w_{0})\cdot 4p_{\epsilon}/\lambda^{5}\leq(\mathrm{dist}(w^{\prime},w)+\mathrm{dist}(w,w_{0}))\cdot 4p_{\epsilon}/\lambda^{5}$
+$\leq\mathrm{dist}(w,w^{\prime})\cdot(1+4p_{\epsilon}/\lambda^{5})\cdot 4p_{\epsilon}/\lambda^{5}$ (86)
+The first inequality holds by Lemma 15.1, the second inequality holds by the definition of $t[w_{0}^{\prime}]$ at the time of call to $\texttt{MakeRobust}(w^{\prime})$, the third inequality holds by the guarantee of $\widehat{\mathrm{dist}}(w^{\prime},W^{\text{mid}}-w^{\prime})$ in Lemma 7.3, the fourth inequality holds since $w_{0}$ was present in $W^{\text{mid}}$ in the time of call to $\texttt{MakeRobust}(w^{\prime})$, the fifth inequality is triangle inequality and the last inequality holds by Equation 84.
+As a result,
+$\mathrm{dist}(w_{0},W^{\text{new}}-w_{0})$ $\leq\mathrm{dist}(w_{0},w_{0}^{\prime})\leq\mathrm{dist}(w_{0},w)+\mathrm{dist}(w,w^{\prime})+\mathrm{dist}(w^{\prime},w_{0}^{\prime})$
+$\leq\left((4p_{\epsilon}/\lambda^{5})+(1)+(1+4p_{\epsilon}/\lambda^{5})\cdot 4p_{\epsilon}/\lambda^{5}\right)\cdot\mathrm{dist}(w,w^{\prime})<2\cdot\mathrm{dist}(w,w^{\prime}).$
+The first inequality holds since $w^{\prime}_{0}$ is present in $W^{\text{new}}$, the second inequality is triangle inequality, the third inequality follows by Equation 83 and Section 20.9, and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number.
+So, we derived the same inequality as Section 20.9, and the contradiction follows exactly like the previous case.
+20.10 Proof of Claim 14.1
+Since $u$ is $t[u]$-robust at the beginning of the epoch (w.r.t.
+$X^{(0)}$), according to Definition 13.2, we conclude that there exists a sequence of points $(x_{0},x_{1},\ldots,x_{t[u]})$ together with a sequence of sets $(B_{0},B_{1},\cdots,B_{t[u]})$ such that $x_{0}=u$ and the guarantees in Definition 13.2 holds for these sequences w.r.t.
+$X^{(0)}$.
+Since $u\in W_{3}$ is not contaminated, we infer that the ball around $u=x_{0}$ of radius $\lambda^{3t[u]+1.5}$ is not changes.
+This shows that non of the sets $B_{i}$ has changes.
+This is because if there exists an $0\leq i\leq t[u]$ such that $B_{i}\cap(X^{(0)}\oplus X^{(\ell+1)})\neq\emptyset$, by taking an arbitrary $y\in B_{i}\cap(X^{(0)}\oplus X^{(\ell+1)})$, we conclude that
+$\mathrm{dist}(y,x_{0})\leq\mathrm{dist}(y,x_{i})+\mathrm{dist}(x_{i},x_{0})\leq\lambda^{3i+1}+4\cdot\lambda^{3i-1}\leq\lambda^{3t[u]+1.5},$
+where the second inequality follows by $y\in B_{i}\subseteq\textsc{Ball}(x_{i},\lambda^{3i+1})$ and Lemma 15.1, and the third inequality follows since $i\leq t[u]$ and $\lambda$ is a large number.
+As a result, $y\in\textsc{Ball}(x_{0},\lambda^{3t[u]+1.5})\cap(X^{(0)}\oplus X^{(\ell+1)})$, which is in contradiction with $u=x_{0}$ not being contaminated.
+Hence, all of the $B_{i}$s remain unchanged during the epoch and it is easy to observe that the same sequences $(x_{0},x_{1},\ldots,x_{t[u]})$ and $(B_{0},B_{1},\cdots,B_{t[u]})$ also have the guarantees in Definition 13.2 w.r.t.
+the new dataset $X^{(\ell+1)}$, which means $u=x_{0}$ is $t[u]$-robust w.r.t.
+$X^{(\ell+1)}$.
+20.11 Proof of Claim 14.2
+For the sake of contradiction assume $u,v\in S_{\textnormal{{init}}}$ such that $t[u]=t[v]=i$ and $\mathrm{dist}(x,u)\leq\lambda^{3i+2}$ and $\mathrm{dist}(x,v)\leq\lambda^{3i+2}$.
+We conclude $\mathrm{dist}(u,v)\leq\mathrm{dist}(x,u)+\mathrm{dist}(x,v)\leq 2\cdot\lambda^{3i+2}$.
+By symmetry between $u$ and $v$ assume that $v$ is added to the center set after $u$ through a call $\texttt{MakeRobust}(v^{\text{old}})$ for some $v^{\text{old}}\in S^{\text{old}}$ that is replaced by $v$ in the solution and set $t[v]=t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(v^{\text{old}},S^{\text{old}}-v^{\text{old}})/\lambda^{7}$ at that time, where $\widehat{\mathrm{dist}}(v^{\text{old}},S^{\text{old}}-v^{\text{old}})$ is a $p_{\epsilon}$ approximation of the distance of $v^{\text{old}}$ to the center set.
+Hence, $\lambda^{3(t-1)}<\widehat{\mathrm{dist}}(v^{\text{old}},S^{\text{old}}-v^{\text{old}})/\lambda^{7}$ by the definition of $t$ and $\widehat{\mathrm{dist}}(v^{\text{old}},S^{\text{old}}-v^{\text{old}})\leq p_{\epsilon}\cdot\mathrm{dist}(v^{\text{old}},u)$ since $u$ was there in the center set at the time of the call $\texttt{MakeRobust}(v^{\text{old}})$.
+As a result,
+$\lambda^{3(t-1)}$ $<\widehat{\mathrm{dist}}(v^{\text{old}},S^{\text{old}}-v^{\text{old}})/\lambda^{7}$
+$\leq(p_{\epsilon}/\lambda^{7})\cdot\mathrm{dist}(v^{\text{old}},u)$
+$\leq(p_{\epsilon}/\lambda^{7})\cdot(\mathrm{dist}(v^{\text{old}},v)+\mathrm{dist}(v,u))$
+$\leq(p_{\epsilon}/\lambda^{7})\cdot(4\cdot\lambda^{3t-1}+2\cdot\lambda^{3i+2})$
+$\leq 4\cdot\lambda^{3t-7}+2\cdot\lambda^{3i-4}.$
+The fourth inequality follows by Lemma 15.1 and the fifth inequality follows since $\lambda=(p_{\epsilon})^{2}$ (see Part III). As a result,
+$\lambda^{3t-3}\leq 2\cdot(\lambda^{3t-3}-4\cdot\lambda^{3t-7})\leq 4\cdot\lambda^{3i-4}<\lambda^{3i-3}.$
+Here we used the fact that $\lambda$ is a large number.
+Thus, $t<i=t[v]$.
+This is in contradiction with the definition of $t[v]$ for $v$ since when we made the call to $\texttt{MakeRobust}(v^{\text{old}})$, we set $t[v]=t$ which indicated that $v$ is $t[v]$-robust.
+To explain more, the only time that the value of $t[v]$ can be increased is when we have another call on $\texttt{MakeRobust}(v)$ that replaces $v$ with $v^{\text{new}}$ which is $t[v^{\text{new}}]=t^{\prime}$-robust for a larger integer $t^{\prime}>t$.
+But, we assumed the last call to MakeRobust and the value of $t[v]=t$ is set in this call.
+20.12 Proof of Claim 14.3
+Note that $t[u]=i$ since $u\in S[i]$. We consider two cases.
+Case 1.
+$\mathrm{dist}(x,u)\leq\lambda^{3i+2}$.
+According to Claim 14.2, there does not exists $v\in S_{\textnormal{{init}}}-u$ satisfying $t[v]=i$ and $\mathrm{dist}(x,v)\leq\lambda^{3i+2}$.
+Since $\textsc{Ball}(v,\lambda^{3i+1.5})\subseteq\textsc{Ball}(v,\lambda^{3i+2})$, we conclude that the distance between $x$ and $v$ is strictly greater than $\lambda^{3i+1.5}$.
+So, $x$ does not contaminate $v$.
+Hence, there does not exist any $v\in S[i]-u$ contaminated by $x$ and the only center in $S[i]$ that might be contaminated by $x$ is $u$.
+Case 2.
+$\mathrm{dist}(x,u)>\lambda^{3i+2}$.
+In this case, we conclude that $\lambda^{3i+2}<\mathrm{dist}(x,u)\leq p_{\epsilon}\cdot\mathrm{dist}(x,S[i])$.
+Hence, the distance of any center $v$ in $S[i]$ to $x$ is strictly greater than $\lambda^{3i+2}/p_{\epsilon}=\lambda^{3i+1.5}$ (since $\lambda=(p_{\epsilon})^{2}$ by Part III), which concludes there is no center $v\in S[i]$ contaminated by $x$.
+Appendix A Barriers Towards Adapting the Algorithm of [BCG+24]
+In [BCG+24], the authors give a dynamic algorithm for $k$-means on general metric spaces, with $O(1/\epsilon)$-approximation ratio and $\tilde{O}(k^{\epsilon})$ update time.
+With the current state-of-the-art tools for designing geometric algorithms in Euclidean spaces, we cannot avoid a tradeoff between a $\mathrm{poly}(1/\epsilon)$ term in the approximation and a $k^{\epsilon}$ term in the update time for this problem.
+[^24] Such tradeoffs are standard in the geometric setting (e.g.
+from using LSH).
+Bypassing them, even in the static setting, is a major open problem.
+Thus, it might appear that building on top of the work of [BCG+24] might be a simpler and more natural approach than building on [BCF25a], since we will eventually incur such a tradeoff in our final algorithm.
+Unfortunately, the algorithm of [BCG+24] is not as well suited to implementation in high-dimensional Euclidean spaces, and it is not clear how to use this algorithm without avoiding an exponential blowup in the approximation ratio, leading to a $\exp(1/\epsilon)$-approximation.
+The Algorithm of [BCG+24].
+The high level approach taken by [BCG+24] is to maintain a hierarchy of nested solutions $X\supseteq S_{0}\supseteq\dots\supseteq S_{\ell}$, where $S_{\ell}$ has size $k$ and is the output of the algorithm.
+Using a new randomized variant of local search, they can maintain this hierarchy by periodically reconstructing these solutions and maintaining them lazily between reconstructions, leading to an update time of $\tilde{O}(k^{1+1/\ell})$ and a recourse of $\tilde{O}(k^{1/\ell})$.
+To bound their approximation ratio, they show that their randomized local search almost matches the guarantees provided by the projection lemma, which says that there must exist some $S_{i+1}\subseteq S_{i}$ such that $\textsc{cost}(X,S_{i+1})\leq\textsc{cost}(X,S_{i})+O(\mathrm{OPT}_{k}(X))$.
+Applying this argument $O(\ell)$ times, they show that $\textsc{cost}(X,S_{\ell})\leq O(\ell)\cdot\mathrm{OPT}_{k}(X)$.
+The Barrier to Using [BCG+24].
+If we want to implement [BCG+24] efficiently in high-dimensional Euclidean spaces, the most natural approach is to modify some components within the algorithm to use approximate nearest neighbor data structures, which can be implemented efficiently in this setting.
+However, the hierarchical nature of this algorithm requires us to have very strong guarantees on the quality of the solutions in order for the analysis to go through.
+For example, a guarantee that $\textsc{cost}(X,S_{i+1})\leq O(\textsc{cost}(X,S_{i})+\mathrm{OPT}_{k}(X))$ for the solutions in the hierarchy is not enough to obtain an approximation ratio of $O(\ell)$, and instead would lead to an approximation ratio of $2^{O(\ell)}$ if analyzed in the naive manner.
+Unfortunately, if we use approximate nearest neighbor data structures to speed up the randomized local search, it is not clear how we can get the sufficiently strong guarantees required to obtain an approximation ratio of $\mathrm{poly}(\ell)$ with this algorithm.
+To summarize, the lack of robustness to the constants in the guarantees makes the algorithm less well-suited to implementation in high-dimensional Euclidean spaces.

### v3

@@
 In Euclidean $k$-means clustering, the input consists of a set of $n$ points $X\subseteq\mathbb{R}^{d}$ in a $d$-dimensional Euclidean space, where each point $x\in X$ has a nonnegative weight $w(x)\geq 0$, and a positive integer $k\in[n]$.
-We have to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$, so as to minimize the objective $\sum_{x\in X}w(x)\cdot\left(\mathrm{dist}(x,S)\right)^{2}$, where $\mathrm{dist}(x,S):=\min_{y\in S}\{\mathrm{dist}(x,y)\}$ and $d(x,y)$ denotes the Euclidean distance between the points $x$ and $y$.
+We have to compute a set of $k$ centers $S\subseteq\mathbb{R}^{d}$, so as to minimize the objective $\sum_{x\in X}w(x)\cdot\left(\mathrm{dist}(x,S)\right)^{2}$, where $\mathrm{dist}(x,S):=\min_{y\in S}\{\mathrm{dist}(x,y)\}$ and $\mathrm{dist}(x,y)$ denotes the Euclidean distance between the points $x$ and $y$.
 Our focus in this paper is on a general setting, where the parameters $k,d$ can not be treated as absolute constants.
 Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84].
-The study of Euclidean $k$-means dates back to 1950s, starting with the celebrated Lloyd’s algorithm [Llo82].
-It is regarded as one of the most fundamental clustering problems, and serves as a basic primitive in unsupervised learning.
-The problem is APX-hard in Euclidean spaces of dimension $d=\Omega(\log n)$ [ACKS15].
-In contrast, for every sufficiently small $\epsilon>0$, it admits a $\text{poly}(1/\epsilon)$-approximation algorithm that runs in $\tilde{O}(n^{1+\epsilon})$ time [lTS24, JJLL25].
-[^1] We use $\tilde{O}(\cdot)$ notation to hide polylogarithmic factors in $n$ and the aspect ratio, and polynomial factors in $1/\epsilon$.
-Using standard coreset based sparsification [DSS24], the runtime can be further reduced to $\tilde{O}(n+k^{1+\epsilon})$.
-This tradeoff between $\mathrm{poly}(1/\epsilon)$-approximation and $\tilde{O}(n+k^{1+\epsilon})$ runtime is believed to be essentially the best possible.222 When $n=k-1$, $k$-means reduces to closest pair (CP) problem on $n$ points, whose current tradeoff is $\mathrm{poly}(1/\epsilon)$-approximation and $n^{1+\epsilon}$ time via locality sensitive hashing [AI06, AR15].
-Whether this tradeoff for CP can be improved, i.e., achieving $\tilde{O}(n)$ time for $O(1)$-approximation, is a long-standing open question and is believed that a lower bound exists [Rub18].
+[…]
 Dynamic $k$-Means Clustering.
@@
   • (iii) Recourse, which denotes the number of changes (i.e., point insertions/deletions) in the maintained solution $S$ per update, and indicates how stable the solution is over time.
-A naive way to solve this problem would be to run the best possible static algorithm from scratch after each update [lTS24, DSS24, JJLL25]; leading to $\text{poly}(1/\epsilon)$-approximation, $\tilde{O}(n+k^{1+\epsilon})$ update time and $O(k)$ recourse.
+A naive way to solve this problem would be to run the best possible static algorithm from scratch after each update [laTourS24, DraganovSS24, abs-2504-03513]; leading to $\text{poly}(1/\epsilon)$-approximation, $\tilde{O}(n+k^{1+\epsilon})$ update time and $O(k)$ recourse.
 But can we do substantially better?
@@
 Theorem 1.1.
-For every sufficiently small $\epsilon>0$, there is a randomized dynamic algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio, $\tilde{O}(k^{\epsilon})$ amortized update time and $\tilde{O}(1)$ amortized recourse.
-If the length of the update sequence is polynomially bounded by $n$ (the maximum number of input points at any time), then the approximation, update time and recourse guarantees of the algorithm hold with probability $1-1/\text{poly}(n)$ across the entire sequence of updates.
+[…]
 Remarks.
@@
   3. 3.
-    In recent years, significant attention has been devoted towards designing dynamic algorithms for $k$-means and $k$-median objectives [LV17, CHP+19, HK20, FLNS21, BCLP23, BCG+24, BCF25a].
-All these papers, however, focus on solving the problem in general metric spaces, where it is known that any dynamic algorithm has to incur an update time of $\Omega(k)$ [CS04].
+    In recent years, significant attention has been devoted towards designing dynamic algorithms for $k$-means and $k$-median objectives [LattanziV17, Cohen-AddadHPSS19, HenzingerK20, FichtenbergerLN21, nips/BhattacharyaCLP23, BhattacharyaCGLP24, BCF24].
+All these papers, however, focus on solving the problem in general metric spaces, where it is known that any dynamic algorithm has to incur an update time of $\Omega(k)$ [CzumajS04].
 In contrast, ours is the first work that designs an algorithm for this problem specifically tailored to the Euclidean setting, and accordingly bypasses the $\Omega(k)$ update time barrier.
   4. 4.
-    Using our dynamic algorithm as a black-box subroutine, we can obtain a static algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(n+k^{1+\epsilon})$ runtime: Indeed, we can first reduce the number of input points to $\tilde{O}(k)$ by computing a coreset [DSS24] in $\tilde{O}(n)$ time, and then insert each of these (weighted) input points into our dynamic algorithm.
+    Using our dynamic algorithm as a black-box subroutine, we can obtain a static algorithm for Euclidean $k$-means with $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(n+k^{1+\epsilon})$ runtime: Indeed, we can first reduce the number of input points to $\tilde{O}(k)$ by computing a coreset [DraganovSS24] in $\tilde{O}(n)$ time, and then insert each of these (weighted) input points into our dynamic algorithm.
 Our algorithm would handle these $\tilde{O}(k)$ updates in $\tilde{O}(k^{1+\epsilon})$ total time, and in the end it will return a $\text{poly}(1/\epsilon)$-approximate solution to the original set of input points.
-This matches the current best tradeoff between approximation ratio and runtime in the static setting [lTS24, DSS24, JJLL25], which is also believed to be (almost) optimal (see Footnote 2).
+This matches the current best tradeoff between approximation ratio and runtime in the static setting [laTourS24, DraganovSS24, abs-2504-03513], which is also believed to be (almost) optimal (see Footnote 2).
 Furthermore, it is easy to show that any nontrivial dynamic $k$-means algorithm must necessarily incur $\Omega(1)$ recourse, whereas the recourse guarantee achieved by our algorithm is only $\tilde{O}(1)$.
 In summary, the approximation ratio, update time and recourse bounds of our algorithm are almost optimal.
-Our Techniques.
+Other Related Work.
+The Euclidean $k$-means problem has also been extensively studied in the special setting of low-dimensional spaces, where the dimension $d$ is assumed to be a constant.
+This problem in NP-hard even when $d=2$ [MahajanNV12].
+In contrast, it is known to admit a static $(1+\epsilon)$-approximation algorithm that runs in $\tilde{O}\left(2^{(1/\epsilon)^{O(d^{2})}}\cdot n\right)$ time [Cohen-AddadFS21].
+Very recently, for the related Euclidean $k$-median problem, a dynamic $O(\mathrm{poly}(\epsilon^{-1})\log n)$ approximation algorithm with $O(\mathrm{poly}(d)\cdot n^{\epsilon})$ update time has been obtained [GJK+26].
+However, because of a fundamental distortion lower bound of the tree embedding technique which this result relies on, it does not generalize to $k$-means and it cannot yield $o(\log n)$ approximation ratio.
+Hence, this result is largely incomparable to ours.
+Finally, we note that an influential line of recent work has focused on two other canonical clustering problems in the dynamic setting: facility location [GoranciHL18, Cohen-AddadHPSS19, BhattacharyaLP22, BhattacharyaGJQ24] and $k$-center [ChanGS18, GoranciHLSS21, BateniEFHJMW23, CrucianiFGNS24, LackiHGJR24, icml25, ForsterS25].
+2 Our Techniques
 At a high level, our approach towards proving Theorem 1.1 consists of two stages.
@@
 (ii) Speed up the update time of the concerned algorithm in Euclidean spaces, without significantly degrading its approximation ratio and recourse, by designing appropriate geometric data structures.
-At present, there are only two such dynamic algorithms with close to optimal guarantees (in terms of approximation ratio, update time and recourse) on general metrics [BCG+24, BCF25a].
-Among them, the dynamic algorithm developed by [BCG+24] incurs $\Theta(k^{\epsilon})$ recourse.
+At present, there are only two such dynamic algorithms with close to optimal guarantees (in terms of approximation ratio, update time and recourse) on general metrics [BhattacharyaCGLP24, BCF24].
+Among them, the dynamic algorithm developed by [BhattacharyaCGLP24] incurs $\Theta(k^{\epsilon})$ recourse.
 Furthermore, it seems that implementing this framework in Euclidean spaces would inherently incur an exponential in $1/\epsilon$ factor loss in the approximation ratio (see Appendix A for a more detailed discussion).
-Since we are aiming for $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(1)$ recourse, we cannot use the algorithm of [BCG+24] as the starting point of our approach.
-Instead, we build on top of the dynamic algorithm of [BCF25a], which achieves $O(1)$ approximation ratio, $\tilde{O}(k)$ update time and $\tilde{O}(1)$ recourse on general metrics.
-To achieve our $k^{\epsilon}$ update time, we must break the $\tilde{O}(k)$ update time barrier of general metrics, and improve the framework of [BCF25a] to utilize the Euclidean structure.
-In Section 3, we review the basic framework of this algorithm, and outline the major challenges we need to overcome to implement a (modified) version of it in the geometric setting.
-This particularly requires us to implement several geometric data structures that work against adaptive adversary.
-Section 4 presents an overview of our geometric data structures.
-One of our key contributions here is the first consistent hashing scheme [CJK+22] that achieves $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters.
-This new consistent hashing is crucial to ensure that our geometric data structures work against an adaptive adversary.
-Beyond clustering, our data structures generally address more fundamental tasks including dynamic range queries and maintenance of approximately the nearest assignment, which may be of independent interest.
-Finally, Section 5 outlines how we modify the framework of [BCF25a] and implement the resulting algorithm using our geometric data structures, leading us to Theorem 1.1.
-Parts II and III of the paper contains the full version of our algorithm and analysis.
-Other Related Work.
-The Euclidean $k$-means problem has also been extensively studied in the special setting of low-dimensional spaces, where the dimension $d$ is assumed to be a constant.
-This problem in NP-hard even when $d=2$ [MNV12].
-In contrast, it is known to admit a static $(1+\epsilon)$-approximation algorithm that runs in $\tilde{O}\left(2^{(1/\epsilon)^{O(d^{2})}}\cdot n\right)$ time [CFS21].
-Finally, we note that an influential line of recent work has focused on two other canonical clustering problems in the dynamic setting: facility location [GHL18, CHP+19, BLP22, BGJ+24] and $k$-center [CGS18, GHL+21, BEF+23, CFG+24, LHG+24, BCF+25b, FS25].
+Since we are aiming for $\text{poly}(1/\epsilon)$-approximation ratio and $\tilde{O}(1)$ recourse, we cannot use the algorithm of [BhattacharyaCGLP24] as the starting point of our approach.
+Instead, we build on top of the dynamic algorithm of [BCF24], which achieves $O(1)$ approximation ratio, $\tilde{O}(k)$ update time and $\tilde{O}(1)$ recourse on general metrics.
+To achieve our $k^{\epsilon}$ update time, we break the $\tilde{O}(k)$ update time barrier of general metrics, and adapt the framework of [BCF24] to utilize the Euclidean structure.
+Although our algorithm is built on the skeleton of the [BCF24] framework, we face intricate challenges that we highlight below, as well as our key technical contributions that show how to get around these challenges.
+2.1 Challenge I: Approximate Neighborhood
+All of the main objects used in the algorithm of [BCF24] work with exact balls around points in general metric spaces.
+On the other hand, in Euclidean spaces, the fast geometric data structures can only guarantee an approximate neighborhood of some radius $r$ around a point $x$ if we use any fast geometric data structure, i.e., there might be some points of distance $(1+\epsilon)r$ to $x$ that are mistakenly included in the neighborhood.
+As a result, the first challenge here is to extend the framework of [BCF24] and come up with new objects that work even with an approximate neighborhood.
+Unfortunately, the very first main crucial object in [BCF24] called ‘robust center’ must be reconsidered, and does not work with an approximate neighborhood.
+This means that after our modification, we need to rebuild the framework of [BCF24] from scratch.
+We modify the definition deliberately and provide a more complicated definition of ‘robust center’ such that: 1) it can be handled with the approximate neighborhood oracles, and 2) the rest of the framework of [BCF24] can be extended to work w.r.t.
+the new definition, without introducing a conceptual barrier inside the analysis.
+Intuitively, a robust center $x$ must satisfy specific properties w.r.t.
+the balls around $x$ of geometrically increasing radius $10^{0},10^{1},10^{2},\ldots$ (see Definition 4.1).
+The new definition of ‘robust center’ that we consider generalizes the previous definition in an abstract way, such that no matter what the approximate neighborhood around a point $x$ is (a subset $B$ of the data-set that contains all points of distance at most $r$ to $x$ and some of the points of distance at most $(1+\epsilon)r$ to $x$), it always works with this neighborhood (as well as any other arbitrary neighborhood).
+This is a non-trivial modification which is too technical, and we refer the reader to Definition 4.1 and Definition 6.1 for a direct comparison.
+2.2 Challenge II: Geometric Data Structures Against Adaptive Adversaries
+To break the $O(k)$ time barrier of [BCF24], we need to design efficient data structures that leverage the Euclidean structure and support several fundamental geometric operations required in the algorithm of [BCF24].
+In our setting, the data structures must handle both a dynamic dataset $X$, which is updated by an adversary, and a dynamic solution (center set) $S$, which is explicitly maintained and updated by the [BCF24] algorithm.
+Therefore, another major challenge arises: even if we assume that $X$ is updated by an oblivious adversary (so that the whole dynamic $k$-means algorithm faces an oblivious adversary), the updates to $S$ cannot be considered as oblivious, since the [BCF24] algorithm is randomized and makes subsequent updates to $S$ randomly according to the current outputs/states of the data structures.
+As a result, it is necessary to provide data structures that work against an adaptive adversary.
+This challenge arises from adapting any randomized algorithm not specifically [BCF24].
+To build such data structures, one of our key contributions is a new construction of a consistent hashing scheme [CJKVY22] that, for the first time, achieves sublinar $\tilde{O}(n^{\epsilon})$ running time per point evaluation with competitive parameters.
+Based on this new consistent hashing, we design robust data structures that can handle two fundamental tasks, answering range queries and maintaining approximate assignments, with $\tilde{O}(n^{\epsilon})$ update and query time; see Section 5 for a more detailed exposition of our consistent hashing and data structure design.
+These data structures play an important role in speeding up the [BCF24] algorithm, especially for the key subroutines discussed in the below Section 2.3, and we believe they may be of independent interest, as they address fundamental tasks beyond clustering.
+2.3 Challenge III: New Subroutines
+All of the critical subroutines in the [BCF24] algorithm can not be directly implemented with fast geometric data structures in the Euclidean spaces, in particular, the three subroutines: ‘Restricted $k$-Means’, ‘Augmented $k$-Means’, and ‘Robustify’.
+In the following, we explain what the conceptual barriers are towards directly implementing these subroutines as stated in [BCF24], and provide our brand new algorithms for these subroutines that are implementable in Euclidean spaces.
+Except for ‘Robustify’, we believe that the other two subroutines are handy, and our new algorithms for these subroutines might be of independent interest to the clustering community.
+Restricted $k$-Means.
+In the restricted $k$-means subroutine, the goal is to remove a given number of centers $r\leq k$ from the solution set $S$ so as to minimize the cost among all possible deletions of $r$ centers out of $S$.
+The [BCF24] algorithm uses randomized local search for this subroutine, which is designed for general metric spaces and thus the inherent $\Omega(kr)$ time lower bound prevents it from achieving sublinear $o(k)$ update time.
+Here we take a completely different coreset-based approach: we show that there exists a small $O(r)$-sized center subset of $S$ such that solving (approximately) the restricted $k$-means on this subset already suffices.
+This coreset-based algorithm is more promising in leveraging the Euclidean structure.
+Indeed, we reduce the construction of such a coreset to approximate nearest neighbor search (which is supported by our data structures), and by applying an efficient algorithm for Euclidean $k$-means [laTourS24, abs-2504-03513] to this coreset, we achieve a nearly tight running time of $\tilde{O}(r^{1+\epsilon})$ for the restricted $k$-means subroutine.
+Please see Section 5.4.1 for a more complete explanation of this part.
+Augmented $k$-Means.
+In the augmented $k$-means subroutine, the goal is to add $r$ more centers to the solution set $S$ so as to minimize the cost among all possible additions of $r$ centers to $S$, for a given $r>0$.
+The algorithm proposed in [BCF24] contracts the centers in $S$ and introduces a new metric space such that any $(r+1)$-means solution in the new space corresponds to a solution for the augmented $k$-means problem.
+It is obvious that the new metric space is far from being Euclidean, and we can not use our Euclidean data structures in order to solve the $r$-means problem in the new metric space.
+Hence, we need a new algorithm to solve augmented $k$-means.
+Our key observation here is that the $k$-means++ algorithm [ArthurV07], which iteratively performs $D^{2}$-sampling to add centers, can in fact be used to yields a good approximation for augmented $k$-means.
+Notice that our data structures can maintain (implicitly and approximately) the nearest-center assignments efficiently.
+With this support, this new $D^{2}$-sampling-based algorithm can be readily implemented in sublinear time $\tilde{O}(r\cdot n^{\epsilon})$.
+Please see Section 5.4.2 for more details.
+Robustify.
+In the robustify subroutine, the goal is to detect all the centers in $S$ that might violate the condition of being a robust center, and then swap these centers with a robust center in their neighborhood.
+The algorithm that [BCF24] proposes here is simply iterating over all elements of $S$, checking the robustness condition, and swap non-robust centers with robust ones.
+Obviously, this takes $\Omega(k)$ time, which is not affordable for our purpose.
+According to our modification of the main objects of [BCF24], in order to have a fast ‘Robustify’ subroutine, the main operation that we have to carry out deep inside the subroutine is to detect the centers $u$ in $S$ whose distances to their closest center in the SuS-u has been changed significantly, i.e., multiplicatively by a constant factor.
+Due to our modification, for each geometrically increasing radius $10^{i}$, we combine our approximate nearest neighbor oracle with one of the data structures in [BhattacharyaGJQ24] in order to detect such centers.
+Please see Section 6.4 for a more complete explanation of this part.
 3 Notations and Preliminaries
 Parameters and Assumptions.
-We fix the following parameters in the paper: $\epsilon$ is a sufficiently small constant, $n$ is the number of points in the data-set, $d$ is the dimension of the space, and $\Delta$ is the aspect ratio, i.e., the max distance divided by the smallest distance between distinct points.
+We fix the following parameters in the paper: $\epsilon$ is a sufficiently small constant, $n$ is the number of points in the data-set, $d$ is the dimension of the space, and $\Delta$ is the aspect ratio, i.e., the maximum distance divided by the smallest distance between distinct points.
 As a typical setup, we assume $\Delta\leq\mathrm{poly}(n)$ is an integer.
-Without loss of generality, the input points that we consider are from the discrete grid $[\Delta]^{d}:=\{(x_{1},\ldots,x_{d})\mid 1\leq x_{i}\leq\Delta\ \forall 1\leq i\leq d\}$, by a standard rounding and rescaling.
+Without loss of generality, the input points that we consider are from the discrete grid $\Delta]^{d}:=\{(x_{1},\ldots,x_{d})\mid 1\leq x_{i}\leq\Delta\ \forall 1\leq i\leq d\}$, by a standard rounding and rescaling.
 Also without loss of generality, we assume $d=O(\log n)$ by a standard Johnson-Lindenstrauss transform [JL84] (see the first paragraph in Section 6.1 for a detailed discussion).
@@
 We now introduce some key notations that will be used in subsequent sections.
-For integer $n\geq 1$, denote $[n]:=\{1,\ldots,n\}$.
+For integer $n\geq 1$, denote $n]:=\{1,\ldots,n\}$.
 Consider a generic function $f:U\to V$; for $T\subseteq U$ we write $f(T):=\{f(u):u\in T\}$, and for $v\in V$ we write $f^{-1}(v):=\{u\in U\mid f(u)=v\}$.
-For every two sets $A$ and $B$, we use $A-B$ instead of $A\setminus B$, and whenever $B=\{b\}$ is a singleton, we write $A-b$ instead of $A-\{b\}$ for simplicity.
+For every two sets $A$ and $B$, we use ABA-B instead of $A\setminus B$, and whenever $B=\{b\}$ is a singleton, we write AbA-b instead of $A-\{b\}$ for simplicity.
 For $x\in[\Delta]^{d},r\geq 0$, let $\textsc{Ball}(x,r):=\{y\in[\Delta]^{d}:\mathrm{dist}(x,y)\leq r\}$ be the metric ball.
@@
 $\mathrm{OPT}_{k}^{W}(X):=\min_{S\subseteq W,|S|=k}\textsc{cost}(X,S)$.
-3 Overview of the Dynamic Algorithm of [BCF25a]
-In this section, we explain the algorithm of [BCF25a] and its main ingredients that works for general metric spaces.
+Organization.
+We provide an overview of the algorithm of [BCF24] in Section 4.
+We provide our geometric data structures in Section 5.
+In Section 6, we provide our modification of the algorithm of [BCF24].
+Finally, in Section 7, we provide the sketch of the analysis of our algorithm.
+We refer the reader to Part II for a thorough description and analysis of our geometric data structures, and to Part III for a complete description and analysis of all parts of our algorithm.
+4 Overview of the Dynamic Algorithm of [BCF24]
+In this section, we explain the algorithm of [BCF24] and its main ingredients that works for general metric spaces.
 Although their algorithm is primarily designed for $k$-median, it can be seamlessly extended to work for $k$-means as well.
@@
 4.1 Robust Centers
-Definition 3.1 (Robust Center, Definition 3.2 in the arxiv version of [BCF25a]).
+Definition 4.1 (Robust Center, Definition 3.2 in the arxiv version of [BCF24]).
 Assume $X\subseteq[\Delta]^{d}$.
-Let $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points in $[\Delta]^{d}$, and let $B_{i}=\textsc{Ball}_{X}(x_{i},10^{i})$ for each $i\in[0,t]$.
+Let $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points in $\Delta]^{d}$, and let $B_{i}=\textsc{Ball}_{X}(x_{i},10^{i})$ for each $i\in[0,t]$.
 We refer to $(x_{0},x_{1},\ldots,x_{t})$ as a $t$-robust sequence w.r.t.
@@
 We say that a point $x\in[\Delta]^{d}$ is $t$-robust w.r.t. $X$ iff there exists a $t$-robust sequence $(x_{0},x_{1},\ldots,x_{t})$ w.r.t. $X$ such that $x_{0}=x$.
-Definition 3.2 (Robust Solution, Definition 3.5 in the arxiv version of [BCF25a]).
+Definition 4.2 (Robust Solution, Definition 3.5 in the arxiv version of [BCF24]).
 We call a center set $S\subseteq[\Delta]^{d}$, robust w.r.t. $X$ iff each center $u\in S$ satisfies the following
-$u\ \text{is}\ t\text{-robust w.r.t.\ $X$, where}\ t\ \text{is the smallest integer satisfying}\ 10^{t}\geq\mathrm{dist}(u,S-u)/200.$ (1)
+[…]
 4.2 Description of The Algorithm
@@
 Invariant A.
-At the start of an epoch, the set $S$ is robust and $\textsc{cost}(S,X)\leq O(1)\cdot\mathrm{OPT}_{k}(X)$.
+At the start of an epoch, the set $S$ is robust and $\textsc{cost}(X,S)\leq O(1)\cdot\mathrm{OPT}_{k}(X)$.
 We now describe how the dynamic algorithm works in a given epoch, in four steps.
@@
 Barriers to Implementation in Euclidean Spaces.
-To implement this algorithmic framework in general metric spaces, [BCF25a] give various different subroutines for each of these steps efficiently.
+To implement this algorithmic framework in general metric spaces, [BCF24] gives various different subroutines for each of these steps efficiently.
 In particular, they design algorithms for solving the ‘restricted’ and ‘augmented’ $k$-means problems (see Section 5.4) and implement the Robustify subroutine (see Section 6.3).
-While these subroutines are fast enough for general metric spaces (where we can only hope to get an update time of $\tilde{O}(k)$), there are too slow for our setting (where we can obtain an update time of $\tilde{O}(k^{\epsilon})$).
-For example, the time taken to implement Step 1 and determine the length of an epoch in [BCF25a] is $\tilde{O}(k\cdot(\ell+1))$.
+While these subroutines are fast enough for general metric spaces (where we can only hope to get an update time of $\tilde{O}(k)$), they are too slow for our setting (where we can obtain an update time of $\tilde{O}(k^{\epsilon})$).
+For example, the time taken to implement Step 1 and determine the length of an epoch in [BCF24] is $\tilde{O}(k\cdot(\ell+1))$.
 Since this can be amortized over an epoch of length $\ell$, this leads to $\tilde{O}(k)$ amortized update time.
@@
 5 Data Structures in High-Dimensional Euclidean Spaces
-In this section, we present new fundamental data structures required to implement the framework of [BCF25a] in high-dimensional Euclidean spaces.
+In this section, we present new fundamental data structures required to implement the framework of [BCF24] in high-dimensional Euclidean spaces.
 There are some parameters we use throughout this section, whose final values in our algorithm are being set as follows.
@@
 These data structures are designed to be robust against adaptive adversaries, which is necessary in our application since we may generate queries based on the previous output from the structure.
-Our data structures rely on a fundamental geometric tool called consistent hashing [JLN+05, CJK+22, Fil24], specifically $(\Gamma,\Lambda)$-hashes as in Definition 4.1.
+Our data structures rely on a fundamental geometric tool called consistent hashing [JiaLNRS05, CJKVY22, Filtser24], specifically $(\Gamma,\Lambda)$-hashes as in Definition 5.1.
 We provide the first time-efficient consistent hashing with competitive parameters, strengthening previous efficient constructions that only achieve a relaxed guarantee (see Section 5.1).
-Then in Sections 4.2, 4.3 and 4.4, we leverage our new consistent hashing scheme to design several dynamic data structures against adaptive adversaries, especially to aid the two key sub-tasks called restricted $k$-means and augmented $k$-means problems.
+Then in Sections 5.2, 5.3 and 5.4, we leverage our new consistent hashing scheme to design several dynamic data structures against adaptive adversaries, especially to aid the two key sub-tasks called restricted $k$-means and augmented $k$-means problems.
 Although our final application is clustering, the structures we devise may be useful for more general tasks.
-Specifically, we devise a reduction-style approximate range query structure, that turns any structure whose query concerns only global dataset, into one that supports range query, and a typical application is to return the coreset [HM04] with respect to a ball range given as query.
-We also devise a structure that approximately maintains the approximate nearest-neighbor assignment between a dynamic dataset and a dynamic center set, which may be used to approximately draw $D^{2}$-samples [AV07].
+Specifically, we devise a reduction-style approximate range query structure, that turns any structure whose query concerns only global dataset, into one that supports range query, and a typical application is to return the coreset [DBLP:conf/stoc/Har-PeledM04] with respect to a ball range given as query.
+We also devise a structure that approximately maintains the approximate nearest-neighbor assignment between a dynamic dataset and a dynamic center set, which may be used to approximately draw $D^{2}$-samples [ArthurV07].
 In this section, when we say “with high probability” (w.h.p.), we mean the algorithm satisfies all stated guarantee simultaneously with probability $1-1/\mathrm{poly}(\Delta^{d})$.
-We assume the access to a $(\Gamma,\Lambda)$-hash family in Sections 4.2, 4.3 and 4.4, where we use the $\Gamma$ and $\Lambda$ parameters without further quantification.
+We assume the access to a $(\Gamma,\Lambda)$-hash family in Sections 5.2, 5.3 and 5.4, where we use the $\Gamma$ and $\Lambda$ parameters without further quantification.
 5.1 Efficient Consistent Hashing
-We give in Definition 4.1 the notion of consistent hashing, which was introduced in [CJK+22] and was also studied under a closely related notion called sparse partitions [JLN+05, Fil24].
-Roughly speaking, a consistent hashing is a data-oblivious map on $[\Delta]^{d}$ that partitions the space $[\Delta]^{d}$ into hash buckets, such that each bucket has some bounded diameter $\rho$, and every small ball of diameter $\frac{\rho}{\Gamma}$ can intersect only $\Lambda$ parts.
+We give in Definition 5.1 the notion of consistent hashing, which was introduced in [CJKVY22] and was also studied under a closely related notion called sparse partitions [JiaLNRS05, Filtser24].
+Roughly speaking, a consistent hashing is a data-oblivious map on $\Delta]^{d}$ that partitions the space $\Delta]^{d}$ into hash buckets, such that each bucket has some bounded diameter $\rho$, and every small ball of diameter $\frac{\rho}{\Gamma}$ can intersect only $\Lambda$ parts.
 Definition 5.1 (Consistent Hashing).
-A map $\varphi$ on $[\Delta]^{d}$ is a $(\Gamma,\Lambda,\rho)$-hash if
+A map $\varphi$ on $\Delta]^{d}$ is a $(\Gamma,\Lambda,\rho)$-hash if
   1. 1.
@@
 We also need the hash function to evaluate efficiently.
-In particular, given a point $x\in[\Delta]^{d}$, we need to evaluate $\varphi(x)$ and $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ (allowing approximation) in $O(\Lambda)$ time (noticing that $|\varphi(\textsc{Ball}(x,\rho/\Gamma))|\leq O(\Lambda)$ by the consistency guarantee), as follows.
+In particular, given a point $x\in[\Delta]^{d}$, we need to evaluate $\varphi(x)$ and $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ (allowing approximation) in $\tilde{O}(\Lambda)$ time (noticing that $|\varphi(\textsc{Ball}(x,\rho/\Gamma))|\leq\Lambda$ by the consistency guarantee), as follows.
 Definition 5.2 (Efficient consistent hashing).
@@
 We devise efficient consistent hashing with parameter $\Lambda=2^{O(d/\Gamma^{2/3})}\cdot\mathrm{poly}(d)$.
-This tradeoff is close to the optimal existential bound of $\Lambda=2^{\Omega(d/\Gamma)}$ from [Fil24] (which runs in $\exp(d)$ time).
-Previously, efficient construction of consistent hashing is only known from [FJL+25], but their consistency guarantee is in expectation which is too weak to satisfy Definition 4.1.
+This tradeoff is close to the optimal existential bound of $\Lambda=2^{\Omega(d/\Gamma)}$ from [Filtser24] (which runs in $\exp(d)$ time).
+Previously, efficient construction of consistent hashing is only known from [random-shift], but their consistency guarantee is in expectation which is too weak to satisfy Definition 5.1.
 While such a weaker guarantee may be useful in many applications, it is nonetheless insufficient for our application especially for devising structures robust to adaptive adversary.
 Lemma 5.3.
-For any $\Gamma\geq 2\sqrt{2\pi}$, $\rho>0$ and $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-1/\mathrm{poly}(\Delta^{d})$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
+For any $\Gamma\geq 2\sqrt{2\pi}$, $\rho>0$ and $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, there exists a random hash $\varphi$ on $\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-1/\mathrm{poly}(\Delta^{d})$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
 Construction.
 We sketch the proof of Lemma 5.3.
-Our starting point is the above-mentioned weaker construction from [FJL+25] which only achieves an expected consistency guarantee.
+Our starting point is the above-mentioned weaker construction from [random-shift] which only achieves an expected consistency guarantee.
 We aim to boost this expected guarantee to a high-probability one with only $\tilde{O}(1)$ loss on the consistency bound.
 Specifically, let $\tilde{\varphi}$ be a random hashing on $\mathbb{R}^{d}$, the expected consistency guarantee of $\tilde{\varphi}$ is
-$\forall x\in[\Delta]^{d},\quad\mathbb{E}\big{[}|\tilde{\varphi}(\textsc{Ball}(x,{\rho}/{\Gamma}))|\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Lambda.$
+$\forall x\in[\Delta]^{d},\quad\mathbb{E}\big[|\tilde{\varphi}(\textsc{Ball}(x,{\rho}/{\Gamma}))|\big]\penalty 10000\ \leq\penalty 10000\ \Lambda.$
 We describe our construction.
-We first draw $m=O(d\log\Delta)$ independent weaker hashing $\{\tilde{\varphi}_{i}\}_{i\in[m]}$ [FJL+25] with expected consistency guarantee.
+We first draw $m=O(d\log\Delta)$ independent weaker hashing $\{\tilde{\varphi}_{i}\}_{i\in[m]}$ [random-shift] with expected consistency guarantee.
 Then, with high probability, for all points $x\in[\Delta]^{d}$, there exists $i\in[m]$ such that the strict consistency condition $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ holds.
@@
 Consistency.
-The diameter bound $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$ carries over from $\tilde{\varphi}$’s.
-We next show the consistency $|\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq O(m\Gamma)$ for $x\in[\Delta]^{d}$.
-Let $Y:=\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))$, and we need $|Y|\leq O(\Lambda)$.
-For $i\in[m]$ let $Y_{i}:=\{(c_{x},\tilde{\varphi}_{c_{x}}(x))\in Y:c_{x}=i\}$ be the subset of hash value coming from color $i$.
+The diameter bound $\mathrm{diam}(\varphi^{-1}(z))\leq\rho$ carries over from $\tilde{\varphi}_{i}$’s.
+We next show the consistency $|\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq O(m\Lambda)$ for $x\in[\Delta]^{d}$.
+Let $Y:=\varphi(\textsc{Ball}(x,\rho/(2\Gamma)))$, and we need $|Y|\leq O(m\Lambda)$.
+For $i\in[m]$, let $Y_{i}:=\{(c_{x},\tilde{\varphi}_{c_{x}}(x))\in Y:c_{x}=i\}$ be the subset of hash values coming from color $i$.
 Then $|Y|\leq\sum_{i}|Y_{i}|$.
-We show that $|Y_{i}|\leq O(\Gamma)$, and this leads to $|Y|\leq O(m\Gamma)$ as desired.
+We show that $|Y_{i}|\leq O(\Lambda)$, and this leads to $|Y|\leq O(m\Lambda)$ as desired.
 Fix some $i$.
-By our construction, an immediate bound is $|Y_{i}|\leq|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|$, so we are done if $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Gamma$.
-However, it could be that $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|>2\Gamma$.
+By our construction, an immediate bound is $|Y_{i}|\leq|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|$, so we are done if $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
+However, it could be that $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|>2\Lambda$.
 We show that if this happens, then no point in $\textsc{Ball}(x,\rho/(2\Gamma))$ can have color $i$, implying $|Y_{i}|=0$.
 This follows from the contrapositive: suppose some $y\in\textsc{Ball}(x,\rho/(2\Gamma))$ has color $c_{y}=i$.
-Then $\textsc{Ball}(x,\rho/(2\Gamma))\subseteq\textsc{Ball}(y,\rho/\Gamma)$, so the consistency guarantee at $y$ implies $|\phi_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
+Then $\textsc{Ball}(x,\rho/(2\Gamma))\subseteq\textsc{Ball}(y,\rho/\Gamma)$, so the consistency guarantee at $y$ implies $|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/(2\Gamma)))|\leq 2\Lambda$.
 Efficiency.
@@
 By the definition of $c_{x}$, this reduces to checking whether $N_{i}:=|\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))|\leq 2\Lambda$ for each color $i$.
-To do this efficiently, we leverage the special structure of $\tilde{\varphi}_{i}$ in [FJL+25]: each $\tilde{\varphi}_{i}$ is defined using a randomly shifted grid over $[\Delta]^{d}$, where each point is hashed based on the grid cell that it belongs to.
+To do this efficiently, we leverage the special structure of $\tilde{\varphi}_{i}$ in [random-shift]: each $\tilde{\varphi}_{i}$ is defined using a randomly shifted grid over $\Delta]^{d}$, where each point is hashed based on the grid cell that it belongs to.
 Under this grid structure, $\tilde{\varphi}_{i}(\textsc{Ball}(x,\rho/\Gamma))$ corresponds to a connected set of grid cells, where each cell is adjacent to $O(d)$ others (i.e., differs in only one coordinate).
 We can explore these cells using a depth-first search, and once the depth-first search takes more than $\tilde{O}(\Lambda)$ time, we can terminate the search and conclude that $N_{i}>2\Lambda$, and no point around $x$ has color $i$.
-This entire process takes total time $N_{i}\cdot\mathrm{poly}(d)$.
-The computation of $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ follows similar ideas.
+This entire process takes total time $\tilde{O}(\Lambda)\cdot\mathrm{poly}(d)$.
+The computation of $\varphi(\textsc{Ball}(x,\rho/\Gamma))$ follows from similar ideas.
 5.2 Approximate Range Query Structure
@@
 This works as a general reduction, that turns a data structure $\mathcal{A}$ which answers query regarding only the entire dataset, to one that can answer queries on a metric ball $\textsc{Ball}(x,r)$.
-Specifically, we consider algorithm $\mathcal{A}$ that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, and that it reports some element $e_{X}$ on some domain $U$.
-For example, elements in $U$ may be real numbers, or a subset of points.
-We say an element $e_{T}\in U$ admissible to $\mathcal{A}$ with respect to some dataset $T$, if it satisfies the specification of $\mathcal{A}$, such as having some approximation ratio $\beta$ on $T$.
+To model this, suppose $U$ is the domain for the output, then for each input dataset $T$, we identify a subset $U_{T}\subseteq U$ in the output domain as admissible output to the dataset $T$.
+For example, elements in $U$ may be real numbers, or a subset of points, and admissible output $U_{T}$ may consist of the solutions that have certain approximation ratio to the input $T$.
+Under this modeling, we assume that the algorithm $\mathcal{A}$ generates admissible output for the entire input dataset, and our structure, stated as follows in Lemma 5.4, uses this algorithm in a black-box way to generate admissible outputs for a ball range query $\textsc{Ball}(x,r)$.
+Specifically, we would use consistent hashing technique to define some approximate $\textsc{Ball}^{\prime}(x,r)$ that is sandwiched between $\textsc{Ball}(x,r)$ (which is the desired ball) and $\textsc{Ball}(x,O(\Gamma)r)$ (which is the ball with slightly enlarged radius).
+By the consistent hashing guarantee, the approximate ball $\textsc{Ball}^{\prime}(x,r)$ may be represented as a disjoint union of $s:=O(\Lambda)$ buckets/subsets $T_{1},\ldots,T_{s}$.
+We then run $\mathcal{A}$ on $T_{i}$’s to generate admissible output $e_{1},\ldots,e_{s}$.
+We note that the following Lemma 5.4 stops at generating these admissible $e_{i}$’s, but to eventually obtain an admissible output for the (approximate) query $\textsc{Ball}^{\prime}(x,r)$, we still need a downstream algorithm that aggregates $e_{1},\ldots,e_{s}$.
+We choose not to include this step in Lemma 5.4 since this step may vary significantly for different applications, and we will discuss this aggregation step in concrete applications.
 Lemma 5.4.
-Suppose an algorithm $\mathcal{A}$ maintains dynamic $X\subseteq[\Delta]^{d}$ and it reports $e_{X}\in U$ as query.
-There exists a dynamic algorithm that takes $r>0$ and dynamic $X\subseteq[\Delta]^{d}$ as input, and for every query $x\in[\Delta]^{d}$ reports $\{e_{1},\ldots,e_{s}\}$ for $s\leq O(\Lambda)$ such that there exists an $s$-partition $\{T_{1},\ldots,T_{s}\}$ of $\textsc{Ball}^{\prime}(x,r)\cap X$ that satisfies $\textsc{Ball}(x,r)\subseteq\textsc{Ball}^{\prime}(x,r)\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$, and $e_{i}$ is admissible with respect to $T_{i}$ ($\forall i\in[s]$).
+Suppose we have an algorithm $\mathcal{A}$ on a dynamic dataset $X\subseteq[\Delta]^{d}$ and it reports some $e_{X}\in U$ that is admissible to $X$.
+Then there exists a dynamic algorithm that takes $r>0$ and dynamic $X\subseteq[\Delta]^{d}$ as input, and for each query point $x\in[\Delta]^{d}$ reports $e_{1},\ldots,e_{s}\in U$ that are admissible to some point sets $T_{1},\ldots,T_{s}$ respectively, where $s\leq O(\Lambda)$, such that $\{T_{1},\ldots,T_{s}\}$ is a an $s$-partition of some approximate ball $\textsc{Ball}^{\prime}(x,r)\cap X$ that satisfies $\textsc{Ball}(x,r)\subseteq\textsc{Ball}^{\prime}(x,r)\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$.
 The update and query run in $\tilde{O}(\Lambda)$ times that of $\mathcal{A}$, respectively.
@@
 This reduction is useful when the query output of $\mathcal{A}$ can be merged, i.e., $e_{1},\ldots,e_{s}$ can be used to derive $e_{\textsc{Ball}^{\prime}(x,r)}$ for dataset $\textsc{Ball}^{\prime}(x,r)\cap X$.
-In particular, we devise a structure that reports a coreset of size $O(k)$ for $k$-means with respect to $T$ for every query $x\in[\Delta]^{d}$.
-Here, a coreset with respect to some $T\subseteq[\Delta]^{d}$ is a (weighted) subset $D$ such that $\forall S\subseteq[\Delta]^{d},|S|\leq k$ one has $\textsc{cost}(D,S)=\Theta(1)\cdot\textsc{cost}(T,S)$ [HM04].
-We can implement this coreset with range query by plugging in Lemma 4.4 a known dynamic coreset algorithm [HK20, lTHS24] as $\mathcal{A}$.
-Then we can obtain a coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ via a standard merge-and-reduce framework [HM04].
+In particular, we devise a structure that reports a coreset of size $O(k)$ for $k$-means with respect to $\textsc{Ball}^{\prime}(x,r)\cap X$ for every query $x\in[\Delta]^{d}$.
+Here, a coreset with respect to some $T\subseteq[\Delta]^{d}$ is a (weighted) subset $D$ such that $\forall S\subseteq[\Delta]^{d},|S|\leq k$ one has $\textsc{cost}(D,S)=\Theta(1)\cdot\textsc{cost}(T,S)$ [DBLP:conf/stoc/Har-PeledM04].
+We can implement this coreset with range query by plugging in Lemma 5.4 a known dynamic coreset algorithm [HenzingerK20, DBLP:conf/esa/TourHS24] as $\mathcal{A}$.
+Then we can obtain a coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ via a standard merge-and-reduce framework [DBLP:conf/stoc/Har-PeledM04].
 In our application, we need this coreset for $\textsc{Ball}^{\prime}(x,r)\cap X$ with $k=1$, so the size becomes $O(1)$, which enables us to compute approximate solutions and cost estimations over the approximate ball $\textsc{Ball}^{\prime}(x,r)$ very efficiently, which is stated in the following lemma.
 Lemma 5.5 ($1$-Means Estimation).
-There is an algorithm that maintains dynamic dataset $X\subseteq[\Delta]^{d}$, such that given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, we can compute in time $\tilde{O}(\Lambda)$ a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimations $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$ such that there exists some point subset $B_{x}\subseteq X$ with $\textsc{Ball}(x,r)\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)$ satisfying the following.
+There is an algorithm on a dynamic dataset $X\subseteq[\Delta]^{d}$, such that given a point $x\in[\Delta]^{d}$ and a radius $r\geq 0$, it can compute in time $\tilde{O}(\Lambda)$ a real number $b_{x}\geq 0$, a center $c^{\star}\in[\Delta]^{d}$, and two estimations $\widehat{\textsc{cost}}_{c^{\star}}$ and $\widehat{\textsc{cost}}_{x}$ such that there exists some point subset $B_{x}\subseteq X$ with $\textsc{Ball}(x,r)\cap X\subseteq B_{x}\subseteq\textsc{Ball}(x,O(\Gamma)\cdot r)\cap X$ satisfying the following.
   1. 1.
@@
     $\textsc{cost}(B_{x},c^{\star})\leq O(1)\cdot\mathrm{OPT}_{1}(B_{x})$.
-The algorithm has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
+The algorithm has amortized update time $\tilde{O}(\Lambda)$ (w.r.t. updates on $X$), works against an adaptive adversary, and succeeds w.h.p.
 Application: ANN.
 In addition, Lemma 5.4 can also be applied to yield an $O(\Gamma)$-approximate nearest neighbor (ANN) structure with update and query time $\tilde{O}(\Lambda)$, against adaptive adversary.
-Similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
-We summarize this guarantee as the following lemma which is needed in our main algorithm.
+Similar guarantee of ANN has also been obtained in e.g., [DBLP:conf/nips/0001DFGHJL24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
+By plugging this ANN data structure in the algorithm of [BhattacharyaGJQ24], we obtain a new data structure that maintains, for each point, its distance to the approximate nearest neighbor.
+We summarize this result in the following lemma, which is needed in our main algorithm.
 Lemma 5.6 (ANN Distance).
@@
 $\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq O(\Gamma)\cdot\mathrm{dist}(s,S-s).$
-The data structure has amortized update time $\tilde{O}(\Lambda)$, works against an adaptive adversary, and succeeds w.h.p.
+The data structure has amortized update time $\tilde{O}(\Lambda)$ (w.r.t. updates on $S$), works against an adaptive adversary, and succeeds w.h.p.
 Proof Sketch of Lemma 5.4.
@@
 The analysis is straightforward and follows easily from Definition 5.2.
-Previously, an MPC version of Lemma 4.4 was introduced in [CGJ+24], but their algorithm is not efficient due to the lack of efficient consistent hashing which we devise (Lemma 4.3).
+Previously, an MPC version of Lemma 5.4 was introduced in [DBLP:conf/icalp/CzumajGJK024], but their algorithm is not efficient due to the lack of efficient consistent hashing which we devise (Lemma 5.3).
 5.3 Approximate Assignment Structure
@@
 However, there is no finite ratio with small recourse for this task, since one can keep on inserting/deleting a point of $S$ to change the assignment drastically.
-To bypass this barrier, instead of maintaining the full assignment $\sigma:S\to X$, our idea is to hash the dataset into buckets $\varphi(X)$ and explicitly maintain only an intermediate assignment $\sigma^{\prime}$ between bucket (ID) and $S$.
-The final assignment $S\to X$ is implicitly defined by assigning all points in the same bucket to the same bucket partner in the intermediate assignment.
+To bypass this barrier, instead of maintaining the full assignment $\sigma:X\to S$, our idea is to hash the dataset into buckets $\varphi(X)$ and explicitly maintain only an intermediate assignment $\sigma^{\prime}$ between bucket (ID) and $S$.
+The final assignment $X\to S$ is implicitly defined by assigning all points in the same bucket to the same bucket partner in the intermediate assignment.
 The formal statement goes in Lemma 5.7 (which summarizes our investigations in Section 11), and the actual structure we maintain needs to be more sophisticated than simply hash buckets of $\varphi$.
@@
 Lemma 5.7 implies key lemmas that we use for restricted $k$-means and augmented $k$-means (see Section 5.4 for the definitions of these sub-tasks).
-Similar to Lemma 4.7, these results maintain dynamic $X$ and $S$, work against an adaptive adversary and succeeds with high probability.
+Similar to Lemma 5.7, these results maintain dynamic $X$ and $S$, work against an adaptive adversary and succeed with high probability.
 Application: Cluster Weight Estimation.
-An immediate application is to compute the injectivity $w_{s}$ of points $s\in S$ in the assignment $\sigma$ (which is $\mathrm{poly}(\Gamma)$-approximate by Property 1 of Lemma 4.7), via maintaining $|\operatorname{bk}(h)|$ (for every $h\in H^{\prime}$) and $\sum_{h\in H^{\prime}:\sigma^{\prime}(h)=s}|\operatorname{bk}(h)|$.
-It is well known that the weighted set $(S,w)$ is a good proxy for $X$, which we will use in our restricted $k$-Means subroutine.
+An immediate application is to compute the injectivity $w_{S}(s)$ of points $s\in S$ in the assignment $\sigma$ (which is $\mathrm{poly}(\Gamma)$-approximate by Property 1 of Lemma 5.7), via maintaining $|\operatorname{bk}(h)|$ (for every $h\in H^{\prime}$) and $w_{S}(s):=\sum_{h\in H^{\prime}:\sigma^{\prime}(h)=s}|\operatorname{bk}(h)|$.
+It is well known that the weighted set $(S,w_{S})$ is a good proxy for $X$, which we will use in our restricted $k$-means subroutine.
 Application: $D^{2}$-Sampling.
-Another more interesting application is to implement $D^{2}$-sampling, where the task is to sample an $x\in X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ [AV07].
+Another more interesting application is to implement $D^{2}$-sampling, where the task is to sample an $x\in X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ [ArthurV07].
 To implement this, we compute for each high-level bucket $\operatorname{bk}(h)$ $(h\in H^{\prime})$ and the low-level buckets therein the sum of squared distance to the assigned $\sigma^{\prime}(h)$.
@@
 In other words, only a part of a high-level bucket should be added.
-To describe these partial bucket, we introduce low-level buckets, which are “refinement” of high-level buckets, defined as follows: for each $i$, for each bucket ID $z\in\varphi_{i}([\Delta]^{d})$, partition the level-$i$ high-level bucket $\varphi^{-1}_{i}(z)$ with respect to $\varphi_{i-1}$ (i.e., the level-$(i-1)$ high-level buckets).
-The set of all these parts (i.e., the intersection with $(i-1)$-level buckets) compromise the low-level buckets $L$.
+To describe these partial bucket, we introduce low-level buckets, which are “refinement” of high-level buckets, defined as follows: for each $i$, for each bucket ID $z\in\varphi_{i}([\Delta]^{d})$, partition the level-$i$ high-level bucket $\varphi^{-1}_{i}(z)$ with respect to $\varphi_{i-1}$ (i.e., the level-(i1)(i-1) high-level buckets).
+The set of all these parts (i.e., the intersection with (i1)(i-1)-level buckets) compromise the low-level buckets $L$.
 Now, each high-level bucket with ID $h\in H$ should maintain a list of existing/non-empty intersecting low-level buckets, in order to determine the set of points belong $h$.
@@
 One might wonder if we need also to consider partitioning a level-$i$ high-level bucket by buckets from $i-2,i-3,\ldots$ levels and include all these into $L$ and potentially also to $f(h)$.
-We show that this is not necessary and only level-$(i-1)$ is needed; specifically, when we remove a high-level bucket $h$ from some level $i$, then all points in this $h$ must all be removed (possibly from levels less than $i$), without leaving partially some low-level buckets behind.
+We show that this is not necessary and only level-(i1)(i-1) is needed; specifically, when we remove a high-level bucket $h$ from some level $i$, then all points in this $h$ must all be removed (possibly from levels less than $i$), without leaving partially some low-level buckets behind.
 This is a crucial property to control the time complexity.
@@
 In particular, when we examine $\varphi_{i}(\textsc{Ball}(s,2^{i}))$, we only need to consider at most $O(\Lambda)$ intersecting high-level buckets.
-Moreover, the consistency also guarantees that each (high-level) bucket in level-($i-1$) may result in at most $O(\Lambda)$ low-level buckets in level $i$.
+Moreover, the consistency also guarantees that each (high-level) bucket in level-(i1i-1) may result in at most $O(\Lambda)$ low-level buckets in level $i$.
 5.4 Restricted and Augmented $k$-Means
-We discuss how to implement the two main subroutines in the algorithm of [BCF25a]: restricted $k$-means and augmented $k$-means.
+We discuss how to implement the two main subroutines in the algorithm of [BCF24]: restricted $k$-means and augmented $k$-means.
 The restricted $k$-means algorithm is used in Step 1, 2, 4 of the framework in Section 4.2, and augmented $k$-means is used in Step 4.
 Both problems take a parameter $r>0$, as well as dataset $X$ and center set $S$ as dynamic input.
-The restricted $k$-Means aims to remove $O(r)$ points from $S$, and augmented $k$-Means aims to add $O(r)$ points to $S$, both to minimize the resultant clustering cost.
+The restricted $k$-means aims to remove $r$ points from $S$, and augmented $k$-means aims to add $r$ points to $S$, both to minimize the resultant clustering cost.
 5.4.1 Restricted $k$-Means
@@
 Each update is handled in amortized time $\tilde{O}(2^{\epsilon d})$, and the algorithm succeeds w.h.p.
-As a standard reduction, we may first move each point $x\in X$ to its nearest center in $S$, and then consider the problem on the induced weighted set $(S,w_{S})$ instead of the point set $X$.
-Therefore, in what follows, we assume $X$ is a weighted set supported on $S$.
-A Sketch-Based Approach.
-The high-level idea is to extract a tiny sketch of $S$, then solve the restricted $k$-Means problem on this sketch.
-Let $\widehat{\mathrm{dist}}$ be a (dynamic) approximate ANN distance array as in Lemma 4.6.
-The sketch is constructed as follows:
-  1. 1.
-    Pick the top $O(r)$ centers from $S$ whose $w_{S}(s)\cdot\widehat{\mathrm{dist}}^{2}(s,S-s)$ value is the smallest, denoted as $T_{1}$.
-  2. 2.
-    Let $T_{2}\subseteq S-T_{1}$ be a set of $O(r)$ centers such that $\forall s\in T_{1},\mathrm{dist}(s,T_{2})\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{dist}(s,S-T_{1})$, i.e.
-the set of approximate nearest neighbors in $S-T_{1}$ for all centers in $T_{1}$.
-The sketch is defined as $T:=T_{1}\cup T_{2}$, and we show that $T$ approximately preserves the optimal value of the restricted $k$-means objective:
-$\mathrm{OPT}^{T}_{|T|-r}(T)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
-Finally, it suffices to run a vanilla $(|T|-r)$-means algorithm on $T$ and obtain a $\mathrm{poly}(\epsilon^{-1})$-approximate solution $R\subseteq T\subseteq S$:
-$\textsc{cost}(R,T-R)=\textsc{cost}(T,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{T}_{|T|-r}(T)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}_{|S|-r}^{S}(S).$
-This step takes $\tilde{O}(|T|^{1+\epsilon})=\tilde{O}(r^{1+\epsilon})$ time by applying static almost-linear time $k$-means algorithms [lTS24, JJLL25].
-The correctness of $R$ as a solution on the full set $S$ follows from a straightforward derivation:
-$\textsc{cost}(S,S-R)=\textsc{cost}(R,S-R)\leq\textsc{cost}(R,T-R)\leq\mathrm{poly}(\epsilon^{-1})\cdot\mathrm{OPT}^{S}_{|S|-r}(S).$
-Dynamic Data Structure.
-The key step in making the above process dynamic, is to maintain the center weights $w_{S}$.
-This can be solved using the cluster weight estimation structure which we introduced in Section 4.3.
-Once this is addressed, the construction of the sketch $T=T_{1}\cup T_{2}$ can be implemented via ANN data structures (using e.g., the one we mentioned in Section 4.2), in total time $O(r\cdot\Lambda)$.
+For ease of illustration, we may assume that $X=S$; this is natural, since each point $x\in X$ can be moved to its (approximate) nearest center in $S$ with only a constant-factor increase in the clustering objective.
+Under this assumption, the goal of the restricted $k$-means problem becomes finding a set $R$ of $r$ points in $S$ that minimizes
+$\textsc{cost}(S,S-R)\penalty 10000\ =\penalty 10000\ \textsc{cost}(R,S-R)\penalty 10000\ =\penalty 10000\ \sum_{s\in R}\mathrm{dist}^{2}(s,S-R).$
+The first observation here is that the restricted $k$-means problem with $r$ centers to delete is equivalent to the $k^{\prime}$-means problem with $k^{\prime}=k-r$, where the solution is restricted to be selected from $S$.
+Hence, a naïve approach would be to directly apply an efficient algorithm for Euclidean $k$-means (e.g.
+[laTourS24, abs-2504-03513]); however, this would only achieve a superlinear $|S|^{1+\epsilon}$ running time.
+Our Coreset-based Algorithm.
+Our algorithm, at a high level, is to speed up this naïve approach by designing a coreset for the restricted $k$-means problem.
+Roughly speaking, this coreset guarantees that running an $O(1)$-approximation algorithm for the restricted $k$-means on it yields a solution whose ratio is larger but still constant for the full set $S$.
+We give a coreset construction with size $O(r)$, and this immediately helps to speed up the naïve approach to a nearly tight $\tilde{O}(r^{1+\epsilon})$ time while obtaining an $O(1)$-approximation for the restricted $k$-means on $S$.
+This notion of coreset is new and is specifically devised for our restricted $k$-means problem.
+This notion is particularly different than the standard coreset for clustering in the literature (see e.g.
+[DBLP:conf/stoc/Har-PeledM04]), and a main conceptual difference is that ours is a coreset with respect to the points to be deleted (i.e., $R$), whereas a standard coreset preserves costs for the center set (i.e., SRS-R).
+Moreover, the techniques for constructing our new coreset is also very different than commonly used sampling techniques [DBLP:journals/siamcomp/Chen09, DBLP:conf/stoc/FeldmanL11, DBLP:conf/stoc/Cohen-AddadSS21, DBLP:conf/stoc/Cohen-AddadLSS22, DBLP:conf/focs/0001CPSS24, Huang0024, DBLP:conf/soda/Cohen-AddadD0SS25] as in standard coresets, and we instead employ a greedy algorithm to build the coreset.
+Next, we give an overview of our coreset construction, which consists of two major steps.
+To start, in the first step, we greedily select $6r$ points $s\in S$ with the smallest values of $\mathrm{dist}(s,S-{s})$ as the candidate centers.
+Denoting the resulting set by $T_{1}$, we show (in Claim 5.9) that the candidate set $T_{1}$ contains, as a subset, a good solution to the restricted $k$-means problem.
+Claim 5.9.
+There exists a subset of $r$ points from $T_{1}$ that is an $O(1)$-approximation to the restricted $k$-means on the full set $S$.
+However, the existence of such a good solution does not imply that $T_{1}$ is a coreset, since the clustering objective may be significantly distorted on $T_{1}$, and thus an approximation algorithm run on $T_{1}$ may not yield the good solution guaranteed by Claim 5.9.
+Therefore, the second step is to address this issue, and we proceed in the following way.
+For each selected point $s\in T_{1}$, we find its (approximate) nearest neighbors among the unselected points $S-T_{1}$.
+Denoting these (approximate) nearest neighbors by $T_{2}$ and letting $T:=T_{1}\cup T_{2}$, we prove the following claim.
+Claim 5.10.
+For any subset $R\subseteq T_{1}$, it holds that
+$\textsc{cost}(S,S-R)\penalty 10000\ \leq\penalty 10000\ \textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\textsc{cost}(S,S-R).$
+Claim 5.10 implies that the objectives of all candidate solutions are preserved on $T$ within a constant factor.
+Therefore, by combining Claims 5.9 and 5.10, it can be shown that $T=T_{1}\cup T_{2}$ is the desired coreset.
+In what follows, we explain the intuition behind Claims 5.9 and 5.10 separately, and then discuss the dynamic implementation of this construction.
+Candidate Centers via Greedy Selection (Claim 5.9).
+Recall that $T_{1}$ consists of the $6r$ points $s\in S$ with the smallest values of $\mathrm{dist}(s,S-{s})$.
+Here, we consider the case where the optimal solution $R^{*}$ to the restricted $k$-means problem on $S$ is disjoint from $T_{1}$, for ease of illustration.
+(The case where $R^{*}$ and $T_{1}$ overlap is handled similarly.)
+The key to the proof is that, among the points in $T_{1}$ we show the existence of a subset of $r$ points, denoted by $R^{\prime}$, such that
+$\forall s\in R^{\prime},\quad\mathrm{dist}(s,S-s)\penalty 10000\ \geq\penalty 10000\ \Omega(1)\cdot\mathrm{dist}(s,S-R^{\prime}).$
+Since $R^{\prime}\subseteq T_{1}$ and $R^{*}\cap T_{1}=\emptyset$, and given that $T_{1}$ consists of the points $s\in S$ with the smallest $\mathrm{dist}(s,S-s)$ values, it follows that
+$\forall s_{1}\in R^{\prime},\forall s_{2}\in R^{*},\quad\mathrm{dist}(s_{1},S-s_{1})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s_{2},S-s_{2}).$ (7)
+As a result,
+$\sum_{s_{2}\in R^{*}}\mathrm{dist}^{2}(s_{2},S-R^{*})\geq\sum_{s_{2}\in R^{*}}\mathrm{dist}^{2}(s_{2},S-s_{2})\geq\sum_{s_{1}\in R^{\prime}}\mathrm{dist}^{2}(s_{1},S-s_{1})\geq\Omega(1)\cdot\sum_{s_{1}\in R^{\prime}}\mathrm{dist}^{2}(s_{1},S-R^{\prime}).$
+This concludes that $R^{\prime}\subseteq T_{1}$ is an $O(1)$-approximation to the restricted $k$-means problem, thereby proving Claim 5.9.
+Nearest Unselected Neighbors for Candidates (Claim 5.10).
+The inequality $\textsc{cost}(S,S-R)\leq\textsc{cost}(T,T-R)$ (equivalently, $\textsc{cost}(R,S-R)\leq\textsc{cost}(R,T-R)$) follows directly from the fact that $T$ is a subset of $S$.
+We then focus on the inequality $\textsc{cost}(T,T-R)\leq O(1)\cdot\textsc{cost}(S,S-R)$.
+Consider a fixed candidate solution $R^{\prime}\subseteq T_{1}$ and suppose that its restricted $k$-means objective $\textsc{cost}(T_{1},T_{1}-R^{\prime})=\textsc{cost}(R^{\prime},T_{1}-R^{\prime})$ on the candidate set $T_{1}\subseteq S$ is severely distorted, i.e., significantly larger than its objective $\textsc{cost}(S,S-R^{\prime})=\textsc{cost}(R^{\prime},S-R^{\prime})$ on the full set $S$.
+This distortion occurs only if, for some point $s\in R^{\prime}$, the distance $\mathrm{dist}(s,T_{1}-R^{\prime})$ is significantly larger than $\mathrm{dist}(s,S-R^{\prime})$.
+This further means that the nearest neighbor of $s$ in $S-R^{\prime}$ is not contained in $T_{1}-R^{\prime}$; such a neighbor must lie in $S-T_{1}$ since $R^{\prime}\subseteq T_{1}\subseteq S$.
+Therefore, this issue can be directly fixed by adding, for every candidate point $s\in T_{1}$, its (approximate) nearest neighbor in $S-T_{1}$ to the coreset, which is precisely our second step.
+This outlines the main idea behind the proof of Claim 5.10.
+Dynamic Implementations.
+This coreset construction is well suited for the dynamic setting, since it only relies on maintaining/querying approximate nearest neighbors, which is supported by our data structures introduced in Section 5.2.
+Moreover, we note that all the above arguments hold for the case where $X=S$.
+Therefore, we actually need an additional step that “moves” each point in $X$ to its (approximate) nearest neighbor in $S$, resulting in a weighted set $(S,w_{S})$ where the weight $w_{S}(s)$ equal to the number of points moved to $s$.
+To realize this moving step in the dynamic setting, we need to maintain the weight $w_{S}$, which is efficiently supported by our cluster weight estimation structure introduced in Section 5.3.
 5.4.2 Augmented $k$-Means
@@
 There exists a data structure that handles two dynamic sets $X,S\subseteq[\Delta]^{d}$, where updates are generated by an adaptive adversary.
-For every query with an integer $a\geq 1$, it runs in time $\tilde{O}(a\cdot 2^{\epsilon d})$ to return a set $A\subseteq[\Delta]^{d}$ of $O(a\cdot\mathrm{poly}(\epsilon^{-1})\cdot d\log\Delta)$ points such that the following holds w.h.p.:
-$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|=a}\textsc{cost}(X,S+A^{\star}).$
+For every query with an integer $r\geq 1$, it runs in time $\tilde{O}(r\cdot 2^{\epsilon d})$ to return a set $A\subseteq[\Delta]^{d}$ of $O(r\cdot\mathrm{poly}(\epsilon^{-1})\cdot d\log\Delta)$ points such that the following holds w.h.p.:
+$\textsc{cost}(X,S+A)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\min_{A^{\star}\subseteq[\Delta]^{d}:|A^{\star}|=r}\textsc{cost}(X,S+A^{\star}).$
 Each update is handled in amortized time $\tilde{O}(2^{\epsilon d})$, and the algorithm succeeds w.h.p.
 Our approach is to iteratively select and add points to $S$ until a good augmentation is achieved.
-We borrow the idea from the classical $k$-means++ algorithm [AV07], which employs $D^{2}$-sampling – i.e., sampling a point from $X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ – to iteratively select centers and achieve an $O(\log k)$-approximation for the vanilla $k$-means problem.
-A follow-up work [ADK09] considers sampling $O(k)$ centers to achieve a bi-criteria $O(1)$-approximation for vanilla $k$-means, and we adapt their analysis to show that repeated $D^{2}$-sampling similarly yields a good approximate solution to the augmented $k$-means problem in our setting.
+We borrow the idea from the classical $k$-means++ algorithm [ArthurV07], which employs $D^{2}$-sampling – i.e., sampling a point from $X$ with probability proportional to $\mathrm{dist}^{2}(x,S)$ – to iteratively select centers and achieve an $O(\log k)$-approximation for the vanilla $k$-means problem.
+A follow-up work [DBLP:conf/approx/AggarwalDK09] considers sampling $O(k)$ centers to achieve a bi-criteria $O(1)$-approximation for vanilla $k$-means, and we adapt their analysis to show that repeated $D^{2}$-sampling similarly yields a good approximate solution to the augmented $k$-means problem in our setting.
 We employ an approximate $D^{2}$-sampling in Section 5.3 to implement this idea in dynamic setting.
@@
 Preprocessing.
-Throughout this section, we assume that the input space to our algorithm is $[\Delta]^{d}$, where the dimension $d$ is $O(\log n)$.
+Throughout this section, we assume that the input space to our algorithm is $\Delta]^{d}$, where the dimension $d$ is $O(\log n)$.
 This assumption can be achieved by the well-known Johnson-Lindenstrauss transformation [JL84].
@@
 We initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$.
-Whenever a point $x\in\mathbb{R}^{m}$ is inserted to the dynamic space, we first apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$.
+Whenever a point $x\in\mathbb{R}^{m}$ is inserted to the dynamic space, we first apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $\Delta]^{d}$.
 As a result, we can feed this new dynamic space into our algorithm.
@@
 […]
-approximation guarantee at most $\text{approximation guarantee at most }p_{\epsilon}\text{ and running time at most }\tilde{O}(n^{\epsilon}).$
+[…]
 First, we extend the building blocks of Definition 4.1 to work with approximate balls, because of the following reason: While using the fast data structures in Section 5, for every $x\in[\Delta]^{d}$ and $r>0$, some of the points in the space that has a distance less than $r$ to $x$ might be missing in the approximate ball, as well as some point that has distance more than $r$ to $x$ might be mistakenly included in the ball.
@@
 Definition 6.1.
-Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$.
+Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $\Delta]^{d}$.
 We call this sequence $t$-robust w.r.t.
@@
 We call a center set $S$, robust w.r.t. $X$ if and only if each center $u\in S$ satisfies the following
-$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\mathrm{dist}(u,S-u)/p_{\epsilon}^{20}.$ (9)
+[…]
 These definitions are relaxations of Definition 4.1 and Definition 4.2, which enable us to work with approximate balls while keeping the main properties of robust center (with a small overhead on the approximation ratio).
@@
 Algorithm 1 Computing an estimate $\widehat{\ell}$ of the value of $\ell^{\star}$.
-$i=0$to $\log_{2}k$
-[…]
-[…]
-[…]
-After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where ^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot p_{\epsilon}^{102}}\right\rfloor$.
+1: for $i=0$ to $\log_{2}k$ do
+2:   Using restricted $k$-means (Lemma 5.8), compute a subset $\widehat{S}_{i}\subseteq S_{\textnormal{{init}}}$ of $(k-s_{i})$ centers that is a $p_{\epsilon}$-approximation to $\mathrm{OPT}_{k-s_{i}}^{S_{\textnormal{{init}}}}\left(X^{(0)}\right)$.
+3:   if $\textsc{cost}(X^{(0)},\widehat{S}_{i})>p_{\epsilon}^{96}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})$ then
+4:    return $\widehat{\ell}:=s_{i-1}$.
+After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where $\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot p_{\epsilon}^{102}}\right\rfloor$.
 Implementing Step 2.
@@
 We consider the following condition.
-$t[u]\geq t,\text{ where }t\text{ is the smallest integer satisfying }p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$ (13)
+[…]
 Since $\widehat{\mathrm{dist}}(u,S-u)\geq\mathrm{dist}(u,S-u)$, it is obvious that if $u$ satisfies Condition (14), it definitely satisfies Condition (10).
@@
 Each center $u$ not detected in part 1, remains $t[u]$-robust w.r.t. the final space $X^{(\ell+1)}$.
-The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.10).
+The proof of this claim follows from [BCF24], but for the sake of completeness, we provide the proof in Part III (see Section 21.10).
 Partitioning $W^{\prime}$.
@@
 This claim shows that every update in the data set can contaminate at most one center in each scale $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$.
-The proof of this claim follows from [BCF25a], but for the sake of completeness, we provide the proof in Part III (see Section 20.11).
+The proof of this claim follows from [BCF24], but for the sake of completeness, we provide the proof in Part III (see Section 21.11).
 Partitioning According to $t[u]$ Value.
@@
 Each update in $S$ is handled in amortized $\tilde{O}(n^{\epsilon})$ time.
-This lemma follows from combining the ANN and [BGJ+24, Theorem 3.1].
-We have already implemented ANN in Section 4, and in [BGJ+24], it is explained how we can implement the above data structure by using an ANN.
+This lemma follows from combining the ANN and [BhattacharyaGJQ24, Theorem 3.1].
+We have already implemented ANN in Section 5, and in [BhattacharyaGJQ24], it is explained how we can implement the above data structure by using an ANN.
 Implementation Using Data Structure $D_{\gamma}(S)$.
@@
 This means that new centers might be inserted into $\mathcal{Y}$ and they might need a subsequent call to MakeRobust.
-Similar to [BCF25a], it is possible to show that if a call MakeRobust is made on a center in Robustify, it will continue to satisfy Conditions (13) and (9) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
+Similar to [BCF24], it is possible to show that if a call MakeRobust is made on a center in Robustify, it will continue to satisfy Conditions (14) and (10) until the end of the call to the current Robustify, and our procedure will not make another call on that center.
 Lemma 6.7.
-Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop.
-Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
-The proof of this lemma follows from [BCF25a] (although we have new definitions), and we provide the proof in Part III for the sake of completeness (see Section 20.9).
+[…]
+The proof of this lemma follows from [BCF24] (although we have new definitions), and we provide the proof in Part III for the sake of completeness (see Section 21.9).
 We summarize our implementation of Robustify in Algorithm 2.
@@
 Algorithm 2 Implementation of the call to $\texttt{Robustify}(W^{\prime})$ at the end of an epoch.
-[…]
-[…]
-[…]
-[…]
-[…]
-6:$u\leftarrow$pop$(\mathcal{Y})$
-[…]
-[…]
-[…]
+1: $W_{1}\leftarrow W^{\prime}\setminus S_{\textnormal{{init}}}$. $\triangleright$ All new centers
+2: for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ do
+3:   for each $i\in[0,\lceil\log_{p_{\epsilon}^{6}}(\sqrt{d}\Delta)\rceil]$ do
+4:    Let $u\leftarrow\widehat{\mathrm{dist}}(x,S[i]-x)$.
+5:    if $\mathrm{dist}(u,x)\leq p_{\epsilon}^{6i+3}$ then
+6:      $W_{2}\leftarrow W_{2}+u$. $\triangleright$ Mark $u$ as a contaminated center
+7: $S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$ $\triangleright$ This line updates data structures
+8: for each $u\in W_{1}\cup W_{2}$ do
+9:   $\texttt{MakeRobust}(u)$.
+10: while $\mathcal{Y}\neq\emptyset$ do $\triangleright$ Before this line, many centers might have been inserted to $\mathcal{Y}$
+11:   $u\leftarrowpop$$(\mathcal{Y})$
+12:   $t\leftarrow$ Smallest integer satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{20}$
+13:   if $t[u]<t$ then
+14:    $\texttt{MakeRobust}(u)$.
+15: return $S_{\textnormal{{final}}}\leftarrow S$ .
 6.5 Make-Robust
@@
 The goal is to follow the Definition 6.1.
-We define $x_{t}:=u$ and by iterating over $j=t$ down to 00, we find $x_{j-1}$ and $B_{j}$ satisfying the properties in Definition 5.1.
+We define $x_{t}:=u$ and by iterating over $j=t$ down to 0, we find $x_{j-1}$ and $B_{j}$ satisfying the properties in Definition 6.1.
 We use Lemma 5.5, and make a query on this data structure for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$.
@@
 Algorithm 3 Implementation of a call to $\texttt{MakeRobust}(u)$ at the end of an epoch.
-[…]
-[…]
-3:$x_{j-1}\leftarrow x_{j}$Else
-[…]
-5:$v\leftarrow x_{0}$
-6:$S\leftarrow S-u+v$
-7:Save $t[v]:=t$ together with $v$
+1: $t\leftarrow$ Smallest integer satisfying $p_{\epsilon}^{6t}\geq\widehat{\mathrm{dist}}(u,S-u)/p_{\epsilon}^{14}$,   $x_{t}\leftarrow u$
+2: for $j=t$ down to $1$ do
+3:   Invoke Lemma 5.5 for $x=x_{j}$ and $\rho=p_{\epsilon}^{6j}$ to get $c^{\star}\in[\Delta]^{d}$, $\widehat{\textsc{cost}}_{x_{j}}$, $\widehat{\textsc{cost}}_{c^{\star}}$ and $b_{x_{j}}$
+4:   if $\left(\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq p_{\epsilon}^{6j-3}\right)$ or $\left(\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{\star}}\right)$ then
+5:    $x_{j-1}\leftarrow x_{j}$
+6:   else
+7:    $x_{j-1}\leftarrow c^{\star}$
+8: $v\leftarrow x_{0}$
+9: $S\leftarrow S-u+v$
+10: Save $t[v]:=t$ together with $v$
 7 Analysis of Our Algorithm: Proof (Sketch) of Theorem 1.1
@@
 The approximation ratio of the algorithm is $\mathrm{poly}(1/\epsilon)$ w.h.p.
-The analysis of the approximation ratio of the algorithm follows from the prior work [BCF25a] with deliberate changes in the parameters inside the proofs.
+The analysis of the approximation ratio of the algorithm follows from the prior work [BCF24] with deliberate changes in the parameters inside the proofs.
 There are only two important considerations as follows.
-First, our new definition of robust centers still satisfy the main properties of robust centers as in [BCF25a] (see Section 15.1.1).
-Second, the analysis of [BCF25a] depends on the fact that the integrality gap of the standard LP-relaxation for $k$-median is constant.
-Since the same fact holds for $k$-means (see Theorem 1.1 in [ANFSW20]), we can extend the analysis for our purpose.
+First, our new definition of robust centers still satisfy the main properties of robust centers as in [BCF24] (see Section 16.1.1).
+Second, the analysis of [BCF24] depends on the fact that the integrality gap of the standard LP-relaxation for $k$-median is constant.
+Since the same fact holds for $k$-means (see Theorem 1.1 in [integrality-gap-k-means]), we can extend the analysis for our purpose.
 For the sake of completeness, we provide a thorough analysis in Part III.
@@
 The amortized recourse of the algorithm is at most $\tilde{O}(1)$.
-The recourse analysis of the algorithm also follows from the previous work [BCF25a] with deliberate changes.
+The recourse analysis of the algorithm also follows from the previous work [BCF24] with deliberate changes.
 There are only some considerations as follows.
 First, in augmented $k$-means (Lemma 5.11), instead of finding a set $A$ of size exactly $a$, we find a set of size $\tilde{O}(a)$.
-This increases the recourse of the algorithm by some $\log$ factors, but the same analysis works as in [BCF25a].
+This increases the recourse of the algorithm by some $\log$ factors, but the same analysis works as in [BCF24].
 Second, we changed the Condition (10) to a more restricted Condition (14), and we might call MakeRobust more often.
-But, the same analysis as in [BCF25a] can be applied here to show that the total amortized recourse of Robustify is still $\tilde{O}(1)$.
+But, the same analysis as in [BCF24] can be applied here to show that the total amortized recourse of Robustify is still $\tilde{O}(1)$.
 For the sake of completeness, we provide a thorough analysis in Part III.
@@
 In amortization, there are $\tilde{O}(n^{\epsilon})$ centers inserted into $\mathcal{Y}$.
-For each of these centers $u\in\mathcal{Y}$, during 6 to 7, the algorithm decides in $\tilde{O}(1)$ time whether or not a call to MakeRobust will be made on $u$ [^7] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 4.6.
+For each of these centers $u\in\mathcal{Y}$, during 11 to 13, the algorithm decides in $\tilde{O}(1)$ time whether or not a call to MakeRobust will be made on uu.
+[^7] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 5.6.
 As a result, the amortized time spend on part 2 is $\tilde{O}(n^{\epsilon})$.
@@
 The main for loop in the implementation consists of at most $O(\log_{p_{\epsilon}^{2}}(\sqrt{d}\Delta))$ many iterations.
-Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 4.5) used in 2 of Algorithm 3.
+Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 5.5) used in 3 of Algorithm 3.
 It is obvious that all other parts of the implementation take at most $\tilde{O}(1)$ time.
@@
 As a crucial foundation of our data structures, we design a time-efficient consistent hashing scheme in Section 9.
-Then, we elaborate on the construction and analysis of each data structure in Sections 9, 10, 11 and 12.
+Then, we elaborate on the construction and analysis of each data structure in Sections 10, 11, 12 and 13.
 Notation.
-Throughout this part, we fix a small constant $10<\epsilon<$ 1, denote by $X\subseteq[\Delta]^{d}$ and $w:X\to\mathbb{R}_{\geq 0}$ a dynamic weighted point set and its weight function, respectively, and denote by $S\subseteq[\Delta]^{d}$ a dynamic (unweighted) center set.
-Without loss of generality, we assume that both $X$ and $S$ are initially empty.
-We use $\tilde{O}(\cdot)$ to hide the $\mathrm{poly}(d\epsilon^{-1}\log(\Delta))$ factors.
+[…]
 8 Main Data Structures
@@
 All of them maintain both the dynamic weighted point set $X\subseteq[\Delta]^{d}$ and the dynamic unweighted center set $S\subseteq[\Delta]^{d}$; we thus omit these details from their statements for simplicity.
-The following Lemmas 7.1, 7.2, 7.3 and 7.4 are established in Section 9, where we design a unified range query structure for handling a class of “mergeable” query problems.
-Specifically, Lemmas 7.1, 7.2 and 7.3 address a set of approximate nearest neighbor (ANN) problems, while Lemma 7.4 concerns the $1$-means problem over approximate balls.
+The following Lemmas 8.1, 8.2, 8.3 and 8.4 are established in Section 10, where we design a unified range query structure for handling a class of “mergeable” query problems.
+Specifically, Lemmas 8.1, 8.2 and 8.3 address a set of approximate nearest neighbor (ANN) problems, while Lemma 8.4 concerns the $1$-means problem over approximate balls.
 Here, an “approximate ball” refers to a subset that is sandwiched between two balls with the same center and (up to a constant) roughly the same radii.
@@
 The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update and query time $\tilde{O}(2^{\epsilon d})$.
-The following Lemmas 7.5 and 7.6 address two crucial subproblems – restricted $k$-means and augmented $k$-means – respectively, which play a central role in our dynamic algorithm described in Part III.
-We present the data structures of Lemmas 7.5 and 7.6 in Sections 11 and 12, respectively, both of which are built on a common data structure that (implicitly) maintains an approximate assignment from the point set $X$ to the center set $S$, as described in Section 10.
+The following Lemmas 8.5 and 8.6 address two crucial subproblems – restricted $k$-means and augmented $k$-means – respectively, which play a central role in our dynamic algorithm described in Part III.
+We present the data structures of Lemmas 8.5 and 8.6 in Sections 12 and 13, respectively, both of which are built on a common data structure that (implicitly) maintains an approximate assignment from the point set $X$ to the center set $S$, as described in Section 11.
 Lemma 8.5 (Restricted $k$-Means; see Lemma 12.1).
 There exists a data structure that, on query of an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
-$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
+$\textsc{cost}(X,S-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
 where $\mathrm{OPT}_{|S|-r}^{S}(X):=\min_{S^{\prime}\subseteq S:|S^{\prime}|=|S|-r}\textsc{cost}(X,S^{\prime})$.
@@
 There exists a data structure that, on query of an integer $a\geq 1$, computes a size-$\Theta(a\cdot\epsilon^{-6}\cdot d\log\Delta)$ subset $A\subseteq[\Delta]^{d}$ in time $\tilde{O}(2^{\epsilon d}\cdot a)$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
-$\textsc{cost}(X,S+A)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
+$\textsc{cost}(X,S+A)\penalty 10000\ \leq\penalty 10000\ O(1)\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
 The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has amortized update time $\tilde{O}(2^{\epsilon d})$.
@@
   3. 3.
-    (efficiency) For every point $x\in[\Delta]^{d}$, both the hash value $\varphi(x)$ and a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ satisfying $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho)$ can be computed in time $\tilde{O}(\Lambda)$.
+    (efficiency) For every point $x\in[\Delta]^{d}$, both the hash value $\varphi(x)$ and a size-$(\leq\Lambda)$ subset $\Phi_{x}\subseteq\Phi$ satisfying $\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$ can be computed in time $\tilde{O}(\Lambda)$.
 Lemma 9.2 (Efficient Consistent Hashing).
-For any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\varphi$ on $[\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
+For any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\varphi$ on $\Delta]^{d}$, such that it can be sampled in $\mathrm{poly}(d\log\Delta)$ time, and with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash.
 9.1 Proof of Lemma 9.2
-Our construction of an efficient consistent hash builds on a previous construction from [FJL+25], which provides a weaker guarantee of consistency that holds in expectation.
+Our construction of an efficient consistent hash builds on a previous construction from [random-shift], which provides a weaker guarantee of consistency that holds in expectation.
 In addition, we leverage the special structure of their hash function – based on randomly shifted grids – and show that the set of hash values over any given ball can be computed efficiently (see Item 4 below).
 We summarize these results in the following lemma.
-Lemma 8.3 (Hashing with Weaker Consistency Guarantees [FJL+25]).
+Lemma 9.3 (Hashing with Weaker Consistency Guarantees [random-shift]).
 Given any $\Gamma\geq 2\sqrt{2\pi}$, $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$, and $\rho>0$, there exists a random hash $\widetilde{\varphi}:[\Delta]^{d}\to\mathbb{Z}^{d}$, such that
@@
 Proof.
-Items 1, 2 and 3 are explicitly stated in [FJL+25]. We now prove Item 4.
-We review the construction from [FJL+25]: Let $v\sim\operatorname{Unif}([0,r/\sqrt{d}]^{d})$ be a uniformly random real vector.
+Items 1, 2 and 3 are explicitly stated in [random-shift]. We now prove Item 4.
+We review the construction from [random-shift]: Let $v\sim\operatorname{Unif}([0,r/\sqrt{d}]^{d})$ be a uniformly random real vector.
 The hash function is defined as $\tilde{\varphi}:x\mapsto\lfloor\frac{x+v}{r/\sqrt{d}}\rfloor$ for $x\in[\Delta]^{d}$.
@@
 By construction, the hash preimages $\widetilde{\varphi}^{-1}(z)$ and $\widetilde{\varphi}^{-1}(z^{\prime})$ are hypercubes of the form
-$\widetilde{\varphi}^{-1}(z)$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z_{j}\cdot r/\sqrt{d}-v_{j},(z_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)},$
-$\widetilde{\varphi}^{-1}(z^{\prime})$ $\textstyle\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\leavevmode\nobreak\ =\leavevmode\nobreak\ \prod_{j\in[d]}\big{[}z^{\prime}_{j}\cdot r/\sqrt{d}-v_{j},(z^{\prime}_{j}+1)\cdot r/\sqrt{d}-v_{j}\big{)}.$
+$\widetilde{\varphi}^{-1}(z)$ $\textstyle\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\big[z_{j}\cdot r/\sqrt{d}-v_{j},(z_{j}+1)\cdot r/\sqrt{d}-v_{j}\big),$
+$\widetilde{\varphi}^{-1}(z^{\prime})$ $\textstyle\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\widetilde{\varphi}_{j}^{-1}(z)\penalty 10000\ =\penalty 10000\ \prod_{j\in[d]}\big[z^{\prime}_{j}\cdot r/\sqrt{d}-v_{j},(z^{\prime}_{j}+1)\cdot r/\sqrt{d}-v_{j}\big).$
 […]
@@
 Following Lemma 9.3 and using the sufficiently large integer $\Lambda=2^{{\Theta(d/\Gamma^{2/3})}}\cdot\mathrm{poly}(d\log\Delta)$ chosen in the statement of Lemma 9.2, every $\widetilde{\varphi}_{c}$ satisfies that:
-  • $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$. $\triangleright$ Item 1 of Lemma 8.3
-  • $\mathbb{E}\big{[}|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big{]}\leq\frac{\Lambda}{2C}$, for every point $x\in[\Delta]^{d}$. $\triangleright$ Item 2 of Lemma 8.3
+  • $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$, for every hash value $z\in\mathbb{Z}^{d}$.
+    $\triangleright$ Item 1 of Lemma 9.3
+  • $\mathbb{E}\big[|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\big]\leq\frac{\Lambda}{2C}$, for every point $x\in[\Delta]^{d}$.
+    $\triangleright$ Item 2 of Lemma 9.3
 The following Claim 9.4 measures the likelihood of a certain random event, which turns out to be the success probability of the construction of our consistent hash function $\varphi$.
@@
 Based on the weakly consistent hash functions $\{\widetilde{\varphi}_{c}\}_{c\in[C]}$, we define our consistent hash function $\varphi:[\Delta]^{d}\to\Phi$ as follows, where the output space $\Phi:=[C]\times\mathbb{Z}^{d}$; following Lemma 9.3, the entire construction takes time $C\cdot O(d)=\mathrm{poly}(d\log\Delta)$.
-$\varphi(x)\leavevmode\nobreak\ :=\leavevmode\nobreak\ (c_{x},\widetilde{\varphi}_{c_{x}}(x)),$ $\forall x\in[\Delta]^{d}.$ (16)
+$\varphi(x)\penalty 10000\ :=\penalty 10000\ (c_{x},\widetilde{\varphi}_{c_{x}}(x)),$ $\forall x\in[\Delta]^{d}.$ (17)
 The Performance Guarantees
 Below, we show that the constructed $\varphi$ is an efficient $(\Gamma,\Lambda,\rho)$-hash (Definition 9.1), conditioned on the high-probability event given in Equation 16.
-First, the diameter bound and the consistency properties (Items 1 and 2 of Definition 8.1) follow from Claims 8.5 and 8.6, respectively.
+First, the diameter bound and the consistency properties (Items 1 and 2 of Definition 9.1) follow from Claims 9.5 and 9.6, respectively.
 Claim 9.5 (Diameter).
@@
 Proof.
-By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ induces a natural partition of the input space $[\Delta]^{d}=\cup_{c\in[C]}G_{c}$, where $G_{c}:=\{x\in[\Delta]^{d}\mid c_{x}=c\}$, $\forall c\in[C]$.
+By construction, our consistent hash function $\varphi(x)=(c_{x},\widetilde{\varphi}_{c_{x}}(x))$ induces a natural partition of the input space $\Delta]^{d}=\cup_{c\in[C]}G_{c}$, where $G_{c}:=\{x\in[\Delta]^{d}\mid c_{x}=c\}$, $\forall c\in[C]$.
 In this manner, we have $|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\sum_{c\in[C]}|\varphi(\textsc{Ball}(x,\frac{\rho}{\Gamma})\cap G_{c})|$, so it suffices to show that
@@
 This finishes the proof. ∎
-Second, the efficiency guarantee (Item 3 of Definition 8.1) directly follows from a combination of Claims 8.7 and 8.8.
+Second, the efficiency guarantee (Item 3 of Definition 9.1) directly follows from a combination of Claims 9.7 and 9.8.
 Claim 9.7 (Hash Value Query).
@@
 If it answers $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))$ before stop – namely satisfaction $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|\leq\frac{\Lambda}{C}$ of Equation 16 – return $c_{x}\leftarrow c$ as the color (and terminate the for-loop $c=1,2,\dots,C$).
-Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, the depth-first search for every single $c=1,2,\dots,C$ stops in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$.
+Following Lemma 9.3 and the stop condition $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{2\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, the depth-first search for every single $c=1,2,\dots,C$ stops in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$.
 Next, provided the above color $c_{x}\in[C]$, we can simply compute $\widetilde{\varphi}_{c_{x}}(x)\in\mathbb{Z}^{d}$ using Lemma 9.3 in time $O(d)=\tilde{O}(1)$. This finishes the proof. ∎
@@
 If it answers $\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ before stop – namely $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|\leq\frac{\Lambda}{C}$ – augment the subset $\Phi_{x}\leftarrow\Phi_{x}\cup\{(c,z)\}_{z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))}$ (and move on to the next iteration of the for-loop $c=1,2,\dots,C$).
-Following Lemma 8.3 and the stop condition “$|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, every iteration $c=1,2,\dots,C$ will stop in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$ and augment the subset $\Phi_{x}$ by at most $\frac{\Lambda}{C}$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$ and the total size $|\Phi_{x}|\leq\Lambda$.
+Following Lemma 9.3 and the stop condition $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))|>\frac{\Lambda}{C}$”, every iteration $c=1,2,\dots,C$ will stop in time $\frac{\Lambda}{C}\cdot\mathrm{poly}(d)$ and augment the subset $\Phi_{x}$ by at most $\frac{\Lambda}{C}$, thus the total running time $\Lambda\cdot\mathrm{poly}(d)=\tilde{O}(\Lambda)$ and the total size $|\Phi_{x}|\leq\Lambda$.
 It remains to establish that $\varphi(\textsc{Ball}(x,\tfrac{\rho}{\Gamma}))\subseteq\Phi_{x}\cap\varphi([\Delta]^{d})\subseteq\varphi(\textsc{Ball}(x,2\rho))$.
@@
 (i) $|\widetilde{\varphi}_{c}(\textsc{Ball}(x,\tfrac{\rho}{\Gamma}))|\leq|\widetilde{\varphi}_{c}(\textsc{Ball}(y,\tfrac{2\rho}{\Gamma}))|\leq\tfrac{\Lambda}{C}$, given $\textsc{Ball}(x,\tfrac{\rho}{\Gamma})\subseteq\textsc{Ball}(y,\tfrac{2\rho}{\Gamma})$ and Equation 16.
-(ii) $z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$, given our construction of the hash function $\varphi$ (Equations 15 and 16).
+(ii) $z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$, given our construction of the hash function $\varphi$ (Equations 16 and 17).
 Thus, our subset $\Phi_{x}$ by construction will include the considered hash value $\Phi_{x}\ni(c,z)$.
@@
 Second, consider a specific hash value $(c,z)\in\Phi_{x}\cap\varphi([\Delta]^{d})$ – there exists a point $y\in[\Delta]^{d}$ such that $\varphi(y)=(c,z)\implies y\in\widetilde{\varphi}_{c}^{-1}(z)$ (Equation 17); we deduce from the triangle inequality that
-[…]
+$\mathrm{dist}(x,y)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\widetilde{\varphi}_{c}^{-1}(z))+\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\penalty 10000\ \leq\penalty 10000\ \tfrac{\rho}{\Gamma}+\rho\penalty 10000\ \leq\penalty 10000\ 2\rho.$
 Here, the second step applies $\mathrm{dist}(x,\widetilde{\varphi}_{c}^{-1}(z))\leq\frac{\rho}{\Gamma}\impliedby z\in\widetilde{\varphi}_{c}(\textsc{Ball}(x,\frac{\rho}{\Gamma}))$ (by our construction of the subset $\Phi_{x}$) and $\mathrm{diam}(\widetilde{\varphi}_{c}^{-1}(z))\leq\rho$ (Item 1 of Lemma 9.3).
@@
 Now, we explain how we answer queries. For a given $x\in[\Delta]^{d}$, we compute a set $\Phi_{x}$ such that
-$\varphi(\textsc{Ball}(x,r))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}\cap\varphi([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi(\textsc{Ball}(x,2\Gamma\cdot r)).$ (17)
+$\varphi(\textsc{Ball}(x,r))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{x}\cap\varphi([\Delta]^{d})\penalty 10000\ \subseteq\penalty 10000\ \varphi(\textsc{Ball}(x,2\Gamma\cdot r)).$ (18)
 Since $\varphi$ is an efficient $(\Gamma,\Lambda,r\cdot\Gamma)$-hash (Definition 9.1), the computation of $\Phi_{x}$ takes $\tilde{O}(\Lambda)$ time and we have $|\Phi_{x}|\leq\Lambda$.
@@
 As a warm-up application of Lemma 10.1, we describe how to design an approximate nearest neighbor (ANN) oracle – a data structure that, given a dynamic center set $S\subseteq[\Delta]^{d}$, returns for each query point $x$ a (distinct) approximate nearest center $s\in S-x$.
-We note that similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
+We note that similar guarantee of ANN has also been obtained in e.g., [DBLP:conf/nips/0001DFGHJL24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
 Without loss of generality, we assume hereafter that $S$ contains at least two centers, which guarantees the existence of a distinct nearest neighbor for any query point.
@@
 Lemma 10.2 (ANN Oracle).
-For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic center set $S\subseteq[\Delta]^{d}$, and for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$.
+For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic center set $S\subseteq[\Delta]^{d}$, and for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$.
 The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update and query time $\tilde{O}(2^{\epsilon d})$.
@@
 ∎
-An ANN oracle can be used to derive many useful data structures. By plugging Lemma 9.2 into the framework of [BGJ+24], we obtain the following ANN indicator data structure.
+An ANN oracle can be used to derive many useful data structures. By plugging Lemma 10.2 into the framework of [BhattacharyaGJQ24], we obtain the following ANN indicator data structure.
 Lemma 10.3 (ANN Indicators).
-For any $\gamma\geq 0$ and $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
+For any $\gamma\geq 0$ and $0<\epsilon<1$, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
   1. 1.
@@
 Proof.
-Follows from Lemma 9.2 and [BGJ+24, Theorem 3.1]
+Follows from Lemma 10.2 and [BhattacharyaGJQ24, Theorem 3.1]
 Furthermore, based on the ANN indicators maintained in Lemma 8.2, we derive the following result, which is a data structure that maintains the approximate distance from each center in $S$ to its nearest neighbor.
 Lemma 10.4 (ANN Distance).
-For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
+For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
 $\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(s,S-s).$
@@
 By construction, we have $b_{\ell_{s}}^{s}=1\implies\mathrm{dist}(s,S-s)\leq\epsilon^{-3/2}\cdot 2^{\ell_{s}}=\widehat{\mathrm{dist}}(s,S-s)$.
-When $\ell_{s}=1$, we have $\mathrm{dist}(s,S-s)\geq 1=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$, since two distinct points in the discrete space $[\Delta]^{d}$ have distance $\geq 1$.
+When $\ell_{s}=1$, we have $\mathrm{dist}(s,S-s)\geq 1=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$, since two distinct points in the discrete space $\Delta]^{d}$ have distance $\geq 1$.
 When $\ell_{s}\neq 1$, we have $b_{\ell_{s}-1}^{s}=0\implies\mathrm{dist}(s,S-s)>2^{\ell_{s}-1}=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$.
@@
 Here, we use Lemma 10.1 to devise a data structure that supports estimating both the $1$-means objective and its solution on any given approximate ball.
-This data structure is obtained by plugging a dynamic coreset algorithm from [HK20] (for the entire set $X$) into Lemma 9.1; hence, for each query, we first obtain a small-sized coreset and then compute the objective and/or solution based on it.
+This data structure is obtained by plugging a dynamic coreset algorithm from [HenzingerK20] (for the entire set $X$) into Lemma 10.1; hence, for each query, we first obtain a small-sized coreset and then compute the objective and/or solution based on it.
 Lemma 10.5 ($1$-Means over Approximate Balls).
-For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic weighted dataset $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{*}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
+For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic weighted dataset $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{*}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
   1. 1.
@@
 We first consider two extreme cases of the query radius $r$.
-In the case of a small radius $10\leq r<$ 1, we have $\textsc{Ball}(x,r)=\{x\}$ (recall that any two distinct points in the discrete space $[\Delta]^{d}$ have distance at least $1$), and we simply consider the point subset $B_{x}=\{x\}\cap X$.
+In the case of a small radius $0\leq r<1$, we have $\textsc{Ball}(x,r)=\{x\}$ (recall that any two distinct points in the discrete space $\Delta]^{d}$ have distance at least 11), and we simply consider the point subset $B_{x}=\{x\}\cap X$.
 Either subcase – $B_{x}=\{x\}$ or $B_{x}=\emptyset$ – is trivial.
@@
 Since there are at most $L+1$ such sub-data structures, the overall update time for Lemma 10.5 is $L+1=\tilde{O}(1)$ times that of a single sub-data structure, and hence asymptotically the same.
-The following claim provides a data structure that (1) maintains the total weight of the dataset (a trivial task), and (2) constructs a coreset for the entire dataset, following the approach of [HK20].101010[HK20] provides a general $(1+\epsilon)$-coreset construction for $k$-means.
+The following claim provides a data structure that (1) maintains the total weight of the dataset (a trivial task), and (2) constructs a coreset for the entire dataset, following the approach of [HenzingerK20].101010[HenzingerK20] provides a general $(1+\epsilon)$-coreset construction for $k$-means.
 In this paper, we only need a constant-factor coreset for $1$-means (e.g., $k=1$, $\epsilon=0.5$), so we state this special case.
@@
   • for every $j\in[t]$, $w_{j}=w(P_{j})$; and
-  • for every $j\in[t]$, the following holds with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
-[^11] While Lemma 9.1 does not explicitly consider success probability, its analysis can be easily adapted to handle data structures that succeed with high probability.
+  • for every $j\in[t]$, the following holds with probability at least $\frac{1}{\mathrm{poly}(\Delta^{d})}.11$1111While Lemma 10.1 does not explicitly consider success probability, its analysis can be easily adapted to handle data structures that succeed with high probability.
     $\forall c\in[\Delta]^{d},\qquad\textsc{cost}(Y_{j},c)=\Theta(\textsc{cost}(P_{j},c)).$ (20)
@@
 We first present static and dynamic objects that are explicitly maintained by our data structure (Definition 11.1).
-Then, we show how these explicit objects define an implicit partition of the non-center points $X-S$ (Claim 10.5).
+Then, we show how these explicit objects define an implicit partition of the non-center points XSX-S (Claim 11.5).
 Built on top of this implicit partition, we define an implicit assignment $\sigma:X-S\to S$ (Claim 11.7).
@@
     $S(i,z)$, a subset of close centers for every high-level bucket $(i,z)\in H$ that, if $z\in\varphi_{i}([\Delta]^{d})$, then [^13] For any hash value $z\notin\varphi_{i}([\Delta]^{d})$, we impose no requirement on the corresponding center subset $S(i,z)$.
-    $\big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\tfrac{\rho_{i}}{\Gamma}\big{\}}\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ S(i,z)\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\big{\}}.$ (21)
+    $\big\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\tfrac{\rho_{i}}{\Gamma}\big\}\penalty 10000\ \subseteq\penalty 10000\ S(i,z)\penalty 10000\ \subseteq\penalty 10000\ \big\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\big\}.$ (22)
   4. 6.
@@
 Proof.
-Initially, both $X$ and $S$ are empty =absent=$\emptyset$= ∅, so the initialization only needs to compute the $m+1$ hash functions $\{\varphi_{i}\}_{i\in[0,m]}$.
+Initially, both $X$ and $S$ are empty $\emptyset$, so the initialization only needs to compute the $m+1$ hash functions $\{\varphi_{i}\}_{i\in[0,m]}$.
 By Lemma 9.2, every $\varphi_{i}$ can be constructed in $\tilde{O}(1)$ time with a success probability of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
@@
 Implicit Partition and Assignment
-Finally, we show how the explicit objects (Definition 10.1) implicitly define a partition of the non-center points $X-S$ (Claim 10.5).
+Finally, we show how the explicit objects (Definition 11.1) implicitly define a partition of the non-center points XSX-S (Claim 11.5).
 This partition exhibits an approximate equidistant structure, where points in each part are roughly equally distant from the center set (Claim 11.6).
@@
 Therefore,
-$\mathrm{dist}(s,x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))+\mathrm{diam}(\varphi_{i-1}^{-1}(z_{i-1}))$ (Triangle inequality)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\rho_{i-1}+\rho_{i-1}$ (Item 2 of Definition 10.1)
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \rho_{i}/\Gamma.$ (Item 1 of Definition 10.1)
-Then, $\mathrm{dist}(s,\varphi_{i}^{-1}(z_{i}))\leq\mathrm{dist}(s,x)\leq\rho_{i}/\Gamma\implies s\in S(i,z_{i})=S(i,x)$ by Equations 21 and 22.
+$\mathrm{dist}(s,x)$ $\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))+\mathrm{diam}(\varphi_{i-1}^{-1}(z_{i-1}))$ (Triangle inequality)
+$\penalty 10000\ \leq\penalty 10000\ 2\rho_{i-1}+\rho_{i-1}$ (Item 2 of Definition 11.1)
+$\penalty 10000\ =\penalty 10000\ \rho_{i}/\Gamma.$ (Item 1 of Definition 11.1)
+Then, $\mathrm{dist}(s,\varphi_{i}^{-1}(z_{i}))\leq\mathrm{dist}(s,x)\leq\rho_{i}/\Gamma\implies s\in S(i,z_{i})=S(i,x)$ by Equations 22 and 23.
 The arbitrariness of $s\in S(i-1,x)$ implies that $S(i-1,x)\subseteq S(i,x)$.
@@
 Claim 11.5 (Partition).
-$\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$.
+$\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of XSX-S.
 Proof.
@@
 For every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$, it holds that
-$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
+$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,S)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\penalty 10000\ \leq\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
 Proof.
@@
 Since $S(i,z)\neq\emptyset$ and $z=\varphi_{i}(x)$, by Equation 22 and the fact that $\sigma^{\prime}(i,z)\in S(i,z)$ (Item 8 of Definition 11.1), we know
-[…]
+$\mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))\penalty 10000\ \leq\penalty 10000\ 2\rho_{i}\penalty 10000\ =\penalty 10000\ (3\Gamma)^{i}.$
 Together with the fact $\mathrm{diam}(\varphi_{i}^{-1}(z))\leq\rho_{i}=\frac{1}{2}\cdot(3\Gamma)^{i}$ (Item 1 of Definition 9.1), the distance from $x$ to $S$ can be bounded as:
-$\mathrm{dist}(x,S)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))+\mathrm{diam}(\varphi_{i}^{-1}(z))$ ($\varphi_{i}(x)=z$)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (3\Gamma)^{i}+\frac{1}{2}\cdot(3\Gamma)^{i}$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
+$\mathrm{dist}(x,S)$ $\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))$
+$\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))+\mathrm{diam}(\varphi_{i}^{-1}(z))$ ($\varphi_{i}(x)=z$)
+$\penalty 10000\ \leq\penalty 10000\ (3\Gamma)^{i}+\frac{1}{2}\cdot(3\Gamma)^{i}$
+$\penalty 10000\ =\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
 On the other hand, since $S(i-1,z^{\prime})=\emptyset$ and $z^{\prime}=\varphi_{i-1}(x)$, by Equation 22 again, we have
-$\min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\rho_{i}}{\Gamma}\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
+$\min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\penalty 10000\ =\penalty 10000\ \frac{\rho_{i}}{\Gamma}\penalty 10000\ >\penalty 10000\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
 Therefore,
-$\mathrm{dist}(x,S)\leavevmode\nobreak\ =\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(x,s)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
+$\mathrm{dist}(x,S)\penalty 10000\ =\penalty 10000\ \min_{s\in S}\mathrm{dist}(x,s)\penalty 10000\ \geq\penalty 10000\ \min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\penalty 10000\ >\penalty 10000\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
 Overall, we have
-$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
+$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,S)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\penalty 10000\ \leq\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
 The arbitrariness of $(i,z)\in H^{\prime}$ and $x\in\operatorname{bk}(i,z)$ finishes the proof. ∎
 Implicit Assignment.
-The following Claim 10.7 provides an implicit scheme for maintaining an approximate assignment, which follows directly from Claims 10.5 and 10.6.
+The following Claim 11.7 provides an implicit scheme for maintaining an approximate assignment, which follows directly from Claims 11.5 and 11.6.
 Claim 11.7 (Assignment).
@@
 To insert a new center $s\in[\Delta]^{d}-S$, we perform the following for every $i\in[0,m]$: compute a subset $\Phi_{i}\subseteq\Phi$ (Item 3 of Definition 9.1) such that
-$\varphi_{i}(\textsc{Ball}(s,\tfrac{\rho_{i}}{\Gamma}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{i}\cap\varphi_{i}([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(s,2\rho_{i})).$
+$\varphi_{i}(\textsc{Ball}(s,\tfrac{\rho_{i}}{\Gamma}))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{i}\cap\varphi_{i}([\Delta]^{d})\penalty 10000\ \subseteq\penalty 10000\ \varphi_{i}(\textsc{Ball}(s,2\rho_{i})).$
 Then, augment $S(i,z)\leftarrow S(i,z)+s$, for every $z\in\Phi_{i}$.
@@
 Then, to identify a to-update hash value $z\in\Phi_{x}$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, we can consider an arbitrary point $y\in\operatorname{bk}(i,z,z^{\prime})$; given that $\varphi_{i-1}(y)=z^{\prime}=\varphi_{i-1}(x)$ (Item 4 of Definition 11.1), we can deduce $\mathrm{dist}(x,y)\leq(3\Gamma)^{i-1}$ (Item 1 of Definition 9.1) and thus
-$z\leavevmode\nobreak\ =\leavevmode\nobreak\ \varphi_{i}(y)\leavevmode\nobreak\ \in\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,(3\Gamma)^{i-1}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,\tfrac{1}{\Gamma}\cdot(3\Gamma)^{i}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}.$
+$z\penalty 10000\ =\penalty 10000\ \varphi_{i}(y)\penalty 10000\ \in\penalty 10000\ \varphi_{i}(\textsc{Ball}(x,(3\Gamma)^{i-1}))\penalty 10000\ \subseteq\penalty 10000\ \varphi_{i}(\textsc{Ball}(x,\tfrac{1}{\Gamma}\cdot(3\Gamma)^{i}))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{x}.$
 Namely, to identify all to-update hash values $z\in\Phi_{x}$ with $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, it suffices to enumerate all hash values $z\in\Phi_{x}$ and check the conditions $\operatorname{bk}(i,z,z^{\prime})\neq\emptyset$, which runs in time $\tilde{O}(\Lambda)$.
@@
 Proof of Lemma 11.8.
-The total update time for the data structure of Definition 10.1 follows by combining the results from Claims 10.9, 10.10, 10.11, 10.12 and 10.13.
+The total update time for the data structure of Definition 11.1 follows by combining the results from Claims 11.9, 11.10, 11.11, 11.12 and 11.13.
 Among these, the most expensive one is maintaining the mappings $\{f(i,z)\}_{(i,z)\in H}$, which requires $\tilde{O}(\Lambda^{2})$ time.
@@
 The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update time $\tilde{O}(2^{\epsilon d})$.
-Before presenting our construction for Lemma 10.14, we state the following corollary, which follows as a consequence of Lemmas 10.14 and 9.4, and will be useful in the analysis of restricted $k$-means (Section 11).
+Before presenting our construction for Lemma 11.14, we state the following corollary, which follows as a consequence of Lemmas 11.14 and 10.4, and will be useful in the analysis of restricted $k$-means (Section 12).
 Corollary 11.15 (Importance of Centers).
-There exists a data structure that maintains an ordering $c_{1},c_{2},\ldots,c_{|S|}$ of centers $S$ such that for each $11\leq i\leq|S|-$ 1, the following holds,
+There exists a data structure that maintains an ordering $c_{1},c_{2},\ldots,c_{|S|}$ of centers $S$ such that for each $1\leq i\leq|S|-1$, the following holds,
 $w_{S}(c_{i})\cdot\widehat{\mathrm{dist}}^{2}(c_{i},S-c_{i})\leq w_{S}(c_{i+1})\cdot\widehat{\mathrm{dist}}^{2}(c_{i+1},S-c_{i+1}),$
@@
 Proof.
-Since both the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ and the weights $\{w_{S}(s)\}_{s\in S}$ are explicitly maintained by Lemmas 9.4 and 10.14, respectively, the desirable ordering $c_{1},c_{2},\ldots,c_{|S|}$ of the centers $S$ of centers $S$ can be maintained synchronously, with only a $\tilde{O}(1)$-factor increase in update time.
+Since both the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ and the weights $\{w_{S}(s)\}_{s\in S}$ are explicitly maintained by Lemmas 10.4 and 11.14, respectively, the desirable ordering $c_{1},c_{2},\ldots,c_{|S|}$ of the centers $S$ of centers $S$ can be maintained synchronously, with only a $\tilde{O}(1)$-factor increase in update time.
 ∎
@@
   • $w_{H}(i,z):=\sum_{z^{\prime}\in f(i,z)}w_{L}(i,z,z^{\prime})=\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$, for every $(i,z)\in H$.
-  • $w_{S}(s):=\sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(i,z)+w(s)$,141414 If $s\notin X$, we define $w(s)=0$. for every center $s\in S$.
+  • $w_{S}(s):=\sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(i,z)+w(s)$, [^14] If $s\notin X$, we define $w(s)=0$. for every center $s\in S$.
 By Lemma 11.2, the data structure of Definition 11.1 can be initialized with success probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$; we henceforth assume that this initialization succeeds.
@@
 where the last step follows from Equation 24. Now, for a specific center $s\in S$, we can write
-$w_{S}(s)$ $\leavevmode\nobreak\ \leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(s)+w(s)$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w(\operatorname{bk}(i,z))+w(s)$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}\sum_{x\in\operatorname{bk}(i,z)}w(s)+w(s).$
-Since $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$ (Claim 10.5) and $\sigma(x)=\sigma^{\prime}(i,z)$ for every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$ (Claim 10.7).
+$w_{S}(s)$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(s)+w(s)$
+$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w(\operatorname{bk}(i,z))+w(s)$
+$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}\sum_{x\in\operatorname{bk}(i,z)}w(s)+w(s).$
+Since $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of XSX-S (Claim 11.5) and $\sigma(x)=\sigma^{\prime}(i,z)$ for every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$ (Claim 11.7).
 It then follows that
-$w_{S}(s)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X-S:\sigma(x)=s}w(x)+w(s)=\sum_{x\in X:\sigma(x)=s}w(x).$
+$w_{S}(s)\penalty 10000\ =\penalty 10000\ \sum_{x\in X-S:\sigma(x)=s}w(x)+w(s)=\sum_{x\in X:\sigma(x)=s}w(x).$
 This finishes the proof of Item 2.
@@
 At any time, it can generate independent random points from $X$ such that each random point $y\in X$ is generated in time $\tilde{O}(1)$ and satisfies that
-$\forall x\in X:$ $\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{-6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$ (24)
+$\forall x\in X:$ $\Pr[y=x]\penalty 10000\ \geq\penalty 10000\ \Omega(\epsilon^{-6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$ (25)
 The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update time $\tilde{O}(2^{\epsilon d})$.
@@
 It remains to show the correctness of the procedure. Clearly, the final sampled point $y$ satisfies
-$y\leavevmode\nobreak\ \in\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\operatorname{bk}(i,z)\leavevmode\nobreak\ =\leavevmode\nobreak\ X-S.$ (25)
+$y\penalty 10000\ \in\penalty 10000\ \bigcup_{(i,z)\in H^{\prime}}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\penalty 10000\ =\penalty 10000\ \bigcup_{(i,z)\in H^{\prime}}\operatorname{bk}(i,z)\penalty 10000\ =\penalty 10000\ X-S.$ (26)
 Consider a specific point $x\in X$.
@@
 Since $\mathrm{dist}(x,S)=0$, Equation 25 trivially holds in this case.
-If $x\notin S$ is not a center, then there exists a unique pair $(i_{x},z_{x})\in H^{\prime}$ such that $x\in\operatorname{bk}(i_{x},z_{x})=\bigcup_{z^{\prime}\in f(i_{x},z_{x})}\operatorname{bk}(i,z,z^{\prime})$, since the collection $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ forms a partition of $X-S$ (Claim 10.5).
+If $x\notin S$ is not a center, then there exists a unique pair $(i_{x},z_{x})\in H^{\prime}$ such that $x\in\operatorname{bk}(i_{x},z_{x})=\bigcup_{z^{\prime}\in f(i_{x},z_{x})}\operatorname{bk}(i,z,z^{\prime})$, since the collection $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ forms a partition of XSX-S (Claim 11.5).
 Moreover, since the low-level buckets $\{\operatorname{bk}(i_{x},z_{x},z^{\prime})\}_{z^{\prime}\in f(i_{x},z_{x})}$ are disjoint (Item 4 of Definition 11.1), there exists a unique value $z_{x}^{\prime}\in f(i_{x},z_{x})$ such that $x\in\operatorname{bk}(i_{x},z_{x},z_{x}^{\prime})$.
 Thus, the point $x$ is sampled with probability
-$\Pr[y=x]$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{w_{H}(i_{x},z_{x})}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}\cdot\frac{w_{L}(i_{x},z_{x},z^{\prime}_{x})}{\sum_{z^{\prime}\in f(i_{x},z_{x})}w_{L}(i_{x},z_{x},z^{\prime})}\cdot\frac{w(x)}{w(\operatorname{bk}(i_{x},z_{x},z^{\prime}_{x}))}$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{(3\Gamma)^{2i_{x}}\cdot w(x)}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}.$
+$\Pr[y=x]$ $\penalty 10000\ =\penalty 10000\ \frac{w_{H}(i_{x},z_{x})}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}\cdot\frac{w_{L}(i_{x},z_{x},z^{\prime}_{x})}{\sum_{z^{\prime}\in f(i_{x},z_{x})}w_{L}(i_{x},z_{x},z^{\prime})}\cdot\frac{w(x)}{w(\operatorname{bk}(i_{x},z_{x},z^{\prime}_{x}))}$
+$\penalty 10000\ =\penalty 10000\ \frac{(3\Gamma)^{2i_{x}}\cdot w(x)}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}.$
 By Claim 11.6, the numerator $(3\Gamma)^{2i_{x}}\geq\Omega(1)\cdot\mathrm{dist}^{2}(x,S)$, and the denominator
-$\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}w(\operatorname{bk}(i,z))\cdot(3\Gamma)^{2i}$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}w(x)\cdot(3\Gamma)^{2i}$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}O(\Gamma^{4})\cdot w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.6)
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\sum_{x\in X-S}w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.5)
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\textsc{cost}(X,S).$
+$\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)$ $\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$
+$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}w(\operatorname{bk}(i,z))\cdot(3\Gamma)^{2i}$
+$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}w(x)\cdot(3\Gamma)^{2i}$
+$\penalty 10000\ \leq\penalty 10000\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}O(\Gamma^{4})\cdot w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 11.6)
+$\penalty 10000\ =\penalty 10000\ O(\Gamma^{4})\cdot\sum_{x\in X-S}w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 11.5)
+$\penalty 10000\ =\penalty 10000\ O(\Gamma^{4})\cdot\textsc{cost}(X,S).$
 Therefore, we conclude that
-$\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\Gamma^{4})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$
+$\Pr[y=x]\penalty 10000\ \geq\penalty 10000\ \Omega(\Gamma^{4})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}\penalty 10000\ =\penalty 10000\ \Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$
 This finishes the proof. ∎
@@
 In this section, we consider the restricted $k$-means problem: Given a weighted point set $X\subseteq[\Delta]^{d}$, a center set $S\subseteq[\Delta]^{d}$, and an integer $r\geq 1$ as input, the goal is to select a size-$r$ subset $R\subseteq S$ that minimizes the $k$-means objective $\textsc{cost}(X,S-R)$.
-Assuming that the weighted point set $X$ and the center set $S$ are maintained by the data structures previously discussed (specifically Corollaries 10.15, 9.2 and 10.14), we design a subroutine that computes an approximate restricted $k$-means solution in sublinear time.
+Assuming that the weighted point set $X$ and the center set $S$ are maintained by the data structures previously discussed (specifically Corollaries 11.15, 10.2 and 11.14), we design a subroutine that computes an approximate restricted $k$-means solution in sublinear time.
 Lemma 12.1 (Subroutine for Restricted $k$-Means).
-Given access to the data structures from Corollaries 10.15, 9.2 and 10.14, there exists a subroutine that, on input an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
-$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
+Given access to the data structures from Corollaries 11.15, 10.2 and 11.14, there exists a subroutine that, on input an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
+$\textsc{cost}(X,S-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
 where $\mathrm{OPT}_{|S|-r}^{S}(X):=\min_{S^{\prime}\subseteq S:|S^{\prime}|=|S|-r}\textsc{cost}(X,S^{\prime})$.
-Recall that all data structures from Corollaries 10.15, 9.2 and 10.14 support dynamic updates to both $X$ and $S$ against an adaptive adversary, each with update time $\tilde{O}(2^{\epsilon d})$.
+Recall that all data structures from Corollaries 11.15, 10.2 and 11.14 support dynamic updates to both $X$ and $S$ against an adaptive adversary, each with update time $\tilde{O}(2^{\epsilon d})$.
 By packing them into a single data structure together with the subroutine in Lemma 12.1, we obtain Lemma 8.5 stated in Section 8.
@@
 3: Compute a size-$r$ subset $\widehat{R}\subseteq T:=T_{1}+T_{2}$ such that $\triangleright$ $T\subseteq S$.
-$\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-\widehat{R})\leq O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-R),$
-where $w_{S}:S\to\mathbb{R}_{\geq 0}$ is the weight function maintained by Lemma 10.14.
-4:\$\widehat{R}$.
+$\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-\widehat{R})\leq O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-R),$ where $w_{S}:S\to\mathbb{R}_{\geq 0}$ is the weight function maintained by Lemma 11.14.
+4: return $\widehat{R}$.
 Running Time.
@@
 3 can be implemented in time $\tilde{O}(r^{1+\epsilon})$: Since the weight function $w_{S}:S\to\mathbb{R}_{\geq 0}$ is explicitly maintained by Lemma 11.14, we can construct the weighted set $T=T_{1}\cup T_{2}$, where every point $s\in T$ has weight $w_{S}(s)$, in time $\tilde{O}(r)$.
-Then, to compute $\widehat{R}$, it suffices to run an $(|T|-r)$-means algorithm on the weighted set $T$, with centers selected from $T$, and obtain a solution $\widehat{S}\subseteq T$ with $|\widehat{S}|=|T|-r$.
-Then, we return $\widehat{R}\leftarrow T-\widehat{S}$ as the desired output; for this, we can apply the static algorithm from [JJLL25], which takes $\tilde{O}(r^{1+\epsilon})$ time.
+Then, to compute $\widehat{R}$, it suffices to run an (|T|r)(|T|-r)-means algorithm on the weighted set $T$, with centers selected from $T$, and obtain a solution $\widehat{S}\subseteq T$ with $|\widehat{S}|=|T|-r$.
+Then, we return $\widehat{R}\leftarrow T-\widehat{S}$ as the desired output; for this, we can apply the static algorithm from [abs-2504-03513], which takes $\tilde{O}(r^{1+\epsilon})$ time.
 Notice that each step has a success probability (either from the data structures or the static algorithm) of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$; therefore, by the union bound, all steps succeed simultaneously with high probability, resulting in a total running time of $\tilde{O}(2^{\epsilon d}\cdot r+r^{1+\epsilon})$.
@@
 It remains to establish the correctness of Algorithm 4. Specifically, we aim to show that the subset $\widehat{R}\subseteq S$ computed in 3 satisfies
-$\textsc{cost}(X,S-\widehat{R})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$ (26)
+$\textsc{cost}(X,S-\widehat{R})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$ (27)
 We first have the following claim, which establishes a sufficient condition for Equation 27.
@@
 Equation 27 holds if
-$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$
+$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$
 We defer the proof of Claim 12.2 to Section 12.1.
@@
 Hence, the sufficient condition stated in Claim 12.2 becomes
-$\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R).$ (27)
+$\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R).$ (28)
 Let $R^{*}:=\mathrm{argmin}_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ denote the optimal solution to the restricted $k$-means on $S$.
@@
 There exists a size-$r$ subset $\widehat{R}^{*}\subseteq T_{1}$ such that
-$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
+$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
 Claim 12.4.
 For any size-$r$ set $R\subseteq T_{1}$, it holds that
-$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
-We defer the proof of Claims 11.3 and 11.4 to Sections 11.2 and 11.4, respectively.
+$\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
+We defer the proof of Claims 12.3 and 12.4 to Sections 12.2 and 12.4, respectively.
 We now show how these two claims together imply Equation 28.
@@
 Then, Equation 28 follows from the derivation below.
-LHS of Equation 27 $=\leavevmode\nobreak\ \min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
-$\leq\leavevmode\nobreak\ \textsc{cost}(T,T-\widehat{R}^{*})$ ($\widehat{R}^{*}\subseteq T_{1}\subseteq T$, $|\widehat{R}^{*}|=r$)
-$\leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-\widehat{R}^{*})$ (Claim 11.4)
-$\leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 11.3)
-$=\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Definition of $R^{*}$)
+LHS of Equation 28 $=\penalty 10000\ \min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
+$\leq\penalty 10000\ \textsc{cost}(T,T-\widehat{R}^{*})$ ($\widehat{R}^{*}\subseteq T_{1}\subseteq T$, $|\widehat{R}^{*}|=r$)
+$\leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-\widehat{R}^{*})$ (Claim 12.4)
+$\leq\penalty 10000\ O(\epsilon^{-6})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 12.3)
+$=\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Definition of $R^{*}$)
 […]
@@
 Equation 27 (i.e., $\textsc{cost}(X,S-\widehat{R})\leq O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)$) holds if
-$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$ (28)
+$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}(s,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$ (29)
 Proof.
@@
 Therefore,
-$\textsc{cost}(X,S-\widehat{R})$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot\mathrm{dist}^{2}(x,S-\widehat{R})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot(\mathrm{dist}(x,\sigma(x))+\mathrm{dist}(\sigma(x),S-\widehat{R}))^{2}$ (triangle inequality)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(\sigma(x),S-\widehat{R}))$ (AM-GM inequality)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(X,S)+2\cdot\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$ (Properties (i) and (ii))
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+2\cdot\textsc{cost}(S,S-\widehat{R}).$
-The last step applies $\textsc{cost}(X,S)\leq\mathrm{OPT}_{|S|-r}^{S}(X)$ and $\textsc{cost}(S,S-\widehat{R})=\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$.
-Since every center $s\in S-\widehat{R}$ has a zero distance to $S-\widehat{R}$, we have
-$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($T\subseteq S$)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($\widehat{R}\subseteq T$)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$ (3 of Algorithm 4)
+[…]
+[…]
+$\penalty 10000\ \leq\penalty 10000\ \sum_{x\in X}w(x)\cdot(\mathrm{dist}(x,\sigma(x))+\mathrm{dist}(\sigma(x),S-\widehat{R}))^{2}$ (triangle inequality)
+$)\displaystyle\penalty 10000\ \leq\penalty 10000\ \sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(\sigma(x),S-\widehat{R}$)) (AM-GM inequality)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$ (3 of Algorithm 4)
 Let $R^{*}_{X}\subseteq S$ be the optimal solution to the restricted $k$-means on $X$, i.e., $|R^{*}_{X}|=r$ and $\textsc{cost}(X,S-R^{*}_{X})=\mathrm{OPT}_{|S|-r}^{S}(X)$. It follows that
-$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Equation 28)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\textsc{cost}(S,S-R^{*}_{X})$ ($R^{*}_{X}\subseteq S$, $|R^{*}_{X}|=r$)
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,s)+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\textsc{cost}(X,S)+O(\epsilon^{-7})\cdot\textsc{cost}(X,S-R^{*}_{X})$ (Property (i))
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
+[…]
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Equation 29)
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\textsc{cost}(S,S-R^{*}_{X})$ ($R^{*}_{X}\subseteq S$, $|R^{*}_{X}|=r$)
+$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
+$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,s)+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
+$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\textsc{cost}(X,S)+O(\epsilon^{-7})\cdot\textsc{cost}(X,S-R^{*}_{X})$ (Property (i))
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
 Putting everything together, we obtain Equation 27:
-$\textsc{cost}(X,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
+[…]
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
 This finishes the proof. ∎
@@
 There exists a size-$r$ subset $\widehat{R}^{*}\subseteq T_{1}$ such that
-$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
+$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
 The Construction of $\widehat{R}^{*}$.
@@
 Hence, it remains to show that
-$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$ (29)
+$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$ (30)
 To this end, observe that
-$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$
+$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\$$=\penalty 10000\ \sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$
 $=\sum_{s\in\widehat{R}^{*}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$
@@
 Similarly, we can obtain
-$\textsc{cost}(S,S-R^{*})\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*})+\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$
+$\textsc{cost}(S,S-R^{*})\penalty 10000\ =\penalty 10000\ \sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*})+\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$
 As a result, we can further divide the proof of Equation 30 into the following two parts:
-$\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}),$ (30)
-$\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$ (31)
+$\sum_{s\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ \sum_{x\in R^{*}_{1}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}),$ (31)
+$\sum_{s\in\widehat{R}^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\sum_{s\in R^{*}_{2}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}).$ (32)
 Equation 31: Consider a specific center $s\in R^{*}_{1}$.
@@
 Therefore,
-$\mathrm{dist}(s,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,s^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathrm{dist}(s,S-R^{*}).$
+$\mathrm{dist}(s,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,s^{\prime})\penalty 10000\ =\penalty 10000\ \mathrm{dist}(s,S-R^{*}).$
 The arbitrariness of $s\in R^{*}_{1}$ implies Equation 31.
@@
 Since $\widehat{R}^{*}_{2}\subseteq T^{\prime}_{1}\subseteq T_{1}$ and $R^{*}_{2}=R^{*}-T_{1}\implies R^{*}_{2}\cap T_{1}=\emptyset$, it follows that
-$\forall i\in[t]:$ $w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i}).$
+$\forall i\in[t]:$ $w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})\penalty 10000\ \leq\penalty 10000\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i}).$
 Now consider a fixed $i\in[t]$. We can deduce from Corollary 11.15 that
-$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-r^{*}_{i})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
+$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$ $\penalty 10000\ \leq\penalty 10000\ w_{S}({\widehat{r}}^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
+$\penalty 10000\ \leq\penalty 10000\ w_{S}(r^{*}_{i})\cdot\widehat{\mathrm{dist}}^{2}(r^{*}_{i},S-r^{*}_{i})$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-r^{*}_{i})$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
 Since ${\widehat{r}}^{*}_{i}\in\widehat{R}^{*}_{2}\subseteq T^{\prime}_{1}$, by Item 3 of Claim 12.5, there exists $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ such that $\mathrm{dist}({\widehat{r}}^{*}_{i},s^{\prime})\leq 2\cdot\mathrm{dist}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$.
@@
 It follows that $s^{\prime}\in S-\widehat{R}^{*}$ and thus
-$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-\widehat{R}^{*})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},s^{\prime})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 4\cdot w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
+$w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-\widehat{R}^{*})$ $\penalty 10000\ \leq\penalty 10000\ w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},s^{\prime})$
+$\penalty 10000\ \leq\penalty 10000\ 4\cdot w_{S}({\widehat{r}}^{*}_{i})\cdot\mathrm{dist}^{2}({\widehat{r}}^{*}_{i},S-{\widehat{r}}^{*}_{i})$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot w_{S}(r^{*}_{i})\cdot\mathrm{dist}^{2}(r^{*}_{i},S-R^{*}).$
 The arbitrariness of $i\in[t]$ implies Equation 32.
@@
 This ensures that $s^{\prime}\in S-(R^{*}+T_{1})$, which implies Item 2, i.e., $s^{\prime}\in S-(R^{*}+T^{\prime}_{1})$ for any subset $T^{\prime}_{1}\subseteq T_{1}$.
-Clearly, the above process will remove at most $2\cdot|R^{*}|=2r$ centers from $T_{1}$, thus leaving at least $6r-2r=4r$ centers still in $T_{1}$ (1 of Algorithm 4).
+Clearly, the above process will remove at most $2\cdot|R^{*}|=2r$ centers from $T_{1}$, thus leaving at least 6r2r=4r6r-2r=4r centers still in $T_{1}$ (1 of Algorithm 4).
 Item 3.
@@
 We construct a desired subset $T^{\prime}_{1}\subseteq T_{1}$ as follows. First, consider the set $T^{\prime}_{1}=\bigcup_{y\in g(T_{1})-R^{*}}g^{-1}(y)$. This set has size
-[…]
-Further, for every $s\in T^{\prime}_{1}$, it holds that $g(s)\in g(T_{1})-R^{*}=S-R^{*}$ and $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$. We then apply the following claim from [Tho04].
-Claim 11.6 ([Tho04, Lemma 21]).
+$|T^{\prime}_{1}|\penalty 10000\ =\penalty 10000\ \sum_{y\in g(T_{1})-R^{*}}|g^{-1}(y)|\penalty 10000\ \geq\penalty 10000\ |g(T_{1})-R^{*}|\penalty 10000\ \geq\penalty 10000\ |g(T_{1})|-|R^{*}|\penalty 10000\ \geq\penalty 10000\ 3r-r\penalty 10000\ =\penalty 10000\ 2r.$
+Further, for every $s\in T^{\prime}_{1}$, it holds that $g(s)\in g(T_{1})-R^{*}=S-R^{*}$ and $\mathrm{dist}(s,g(s))=\mathrm{dist}(s,S-s)$. We then apply the following claim from [Thorup04].
+Claim 12.6 ([Thorup04, Lemma 21]).
 Given a set $A$, suppose that every element $a\in A$ has at most one successor $s(a)\in A$. Then, we can mark at least $|A|/2$ elements such that, if $a$ is marked, then $s(a)$ is unmarked.
@@
 Then, the subset $T^{\prime}_{1}$ is defined as
-$T^{\prime}_{1}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \bigcup_{i\in[t]}(g^{-1}(g_{i})-s^{\prime}_{i}).$ (32)
+$T^{\prime}_{1}\penalty 10000\ :=\penalty 10000\ \bigcup_{i\in[t]}(g^{-1}(g_{i})-s^{\prime}_{i}).$ (33)
 Clearly, we have $|T^{\prime}_{1}|\geq|T_{1}|-t>4r-3r=r$.
@@
 Hence, by applying the triangle inequality, we have
-[…]
+$\mathrm{dist}(s,s^{\prime}_{i})\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,g_{i})+\mathrm{dist}(s^{\prime}_{i},g_{i})\penalty 10000\ \leq\penalty 10000\ 2\cdot\mathrm{dist}(s,g_{i})\penalty 10000\ =\penalty 10000\ 2\cdot\mathrm{dist}(s,S-s).$
 Since $s^{\prime}_{i}\in T_{1}\subseteq S$, it remains to show that $s^{\prime}_{i}\notin R^{*}+T^{\prime}_{1}$: Recall that $T_{1}$ already satisfies (Item 1) $T_{1}\cap R^{*}=\emptyset$, and thus $s^{\prime}_{i}\in T_{1}\implies s^{\prime}_{i}\notin R^{*}$.
@@
 For any size-$r$ set $R\subseteq T_{1}$, it holds that
-$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
+$\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
 Proof.
 Consider a specific size-$r$ subset $R\subseteq T_{1}$. Since $T_{1}\cap T_{2}=\emptyset$ (1 and 2 of Algorithm 4), we have $T-R=(T_{1}-R)+T_{2}$ and $S-R=(S-T_{1})+(T_{1}-R)$. Therefore, we have
-$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\ \mathrm{dist}^{2}(s,T_{2})\}.$
+$\textsc{cost}(T,T-R)$ $\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)$
+$\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\ \mathrm{dist}^{2}(s,T_{2})\}.$
 Now, for every center $s\in R\subseteq T_{1}$, by 2 of Algorithm 4, there exists $s^{\prime}\in T_{2}$ such that $\mathrm{dist}(s,T_{2})\leq\mathrm{dist}(s,s^{\prime})\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(s,S-T_{1})$.
 Hence, it follows that
-$\textsc{cost}(T,T-R)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\mathrm{dist}^{2}(s,T_{2})\}$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),O(\epsilon^{-3})\cdot\mathrm{dist}^{2}(s,S-T_{1})\}$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R)$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
+$\textsc{cost}(T,T-R)$ $\penalty 10000\ =\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),\mathrm{dist}^{2}(s,T_{2})\}$
+$\penalty 10000\ \leq\penalty 10000\ \sum_{s\in R}w_{S}(s)\cdot\min\{\mathrm{dist}^{2}(s,T_{1}-R),O(\epsilon^{-3})\cdot\mathrm{dist}^{2}(s,S-T_{1})\}$
+$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\sum_{s\in R}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R)$
+$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
 This finishes the proof. ∎
@@
 Algorithm 5 Approximating augmented $k$-Means.
-[…]
-2:Sample $t:=\Theta(\epsilon^{-6}\cdot d\log\Delta)$ i.i.d.
+1: $A\leftarrow\emptyset$ $\triangleright$ The set of augmented centers
+2: for all $i\in[a+1]$ do
+3:   Sample $t:=\Theta(\epsilon^{-6}\cdot d\log\Delta)$ i.i.d.
 points $\{x_{i}\}_{i\in[t]}$ from $X$, such that $\Pr[x_{i}=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}$ for every $y\in X$.
-[…]
+4:   $A\leftarrow A+\{x_{i}\}_{i\in[t]}$ return AA
 Claim 13.2 (Fast Implementation of Algorithm 5).
@@
 Proof.
-We modify Algorithm 5 as follows: in each of the $a+1$ iterations, instead of only adding the points $\{x_{i}\}_{i\in[t]}$ (sampled in 2) to the set of augmented centers $A$, we also add them to the center set $S$ and update the data structure of Corollary 10.16.
+We modify Algorithm 5 as follows: in each of the $a+1$ iterations, instead of only adding the points $\{x_{i}\}_{i\in[t]}$ (sampled in 3) to the set of augmented centers $A$, we also add them to the center set $S$ and update the data structure of Corollary 11.16.
 This update takes a total of $\tilde{O}(2^{\epsilon d})$ time per iteration.
-With this modification, 2 becomes equivalent to sampling points $x\in X$ with probability at least $\Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}(x,S)}{\textsc{cost}(X,S)}$.
+With this modification, 3 becomes equivalent to sampling points $x\in X$ with probability at least $\Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}(x,S)}{\textsc{cost}(X,S)}$.
 By querying the data structure of Corollary 11.16 $t$ times, we obtain $t$ independent samples in total time $\tilde{O}(2^{\epsilon d})\cdot t=\tilde{O}(2^{\epsilon d})$.
@@
 $\textsc{cost}(X,S+A)\leq 16\cdot\min_{A^{\prime}\subseteq[\Delta]^{d}:|A^{\prime}|=a}\textsc{cost}(X,S+A^{\prime}).$
-The proof of Claim 12.3 follows almost directly from [ADK09, Theorem 1], where a similar sampling procedure is analyzed to obtain a bi-criteria approximation to the vanilla clustering problem.
+The proof of Claim 13.3 follows almost directly from [DBLP:conf/approx/AggarwalDK09, Theorem 1], where a similar sampling procedure is analyzed to obtain a bi-criteria approximation to the vanilla clustering problem.
 For completeness, we provide a proof in Section 13.1.
@@
 Effect of $D^{2}$-sampling on reducing bad clusters
-A crucial step in the analysis is to understand the behavior of the (distorted) $D^{2}$-sampling procedure, as described in 2 of Algorithm 5 – namely, sampling a random point $x\in X$ such that
+A crucial step in the analysis is to understand the behavior of the (distorted) $D^{2}$-sampling procedure, as described in 3 of Algorithm 5 – namely, sampling a random point $x\in X$ such that
 $\forall y\in X,\quad\Pr[x=y]\geq\Omega(\epsilon^{6})\cdot\frac{w(y)\cdot\mathrm{dist}^{2}(y,S+A)}{\textsc{cost}(X,S+A)}.$ (34)
@@
 Therefore,
-$\textsc{cost}(X,S+A)\leavevmode\nobreak$ $=\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A)+\sum_{j=1}^{a}\textsc{cost}(X_{j},S+A)$
-$\leq\leavevmode\nobreak\ \textsc{cost}(X_{0},S+A^{*})+\sum_{j=1}^{a}8\cdot\textsc{cost}(X_{j},S+A^{*})$
-$\leq\leavevmode\nobreak\ 8\cdot\textsc{cost}(X,S+A^{*}).$
+$\textsc{cost}(X,S+A)\penalty 10000\$$=\penalty 10000\ \textsc{cost}(X_{0},S+A)+\sum_{j=1}^{a}\textsc{cost}(X_{j},S+A)$
+$\leq\penalty 10000\ \textsc{cost}(X_{0},S+A^{*})+\sum_{j=1}^{a}8\cdot\textsc{cost}(X_{j},S+A^{*})$
+$\leq\penalty 10000\ 8\cdot\textsc{cost}(X,S+A^{*}).$
 The claim follows. ∎
@@
 For every bad cluster $X_{j}\in\mathcal{B}(A)$ – its center $c_{j}$ satisfies Equation 36 – we define
-$\widehat{X}_{j}\leavevmode\nobreak\ :=\leavevmode\nobreak\ \left\{y\in X_{j}:\ \mathrm{dist}^{2}(y,c_{j})\leq\frac{2}{w(X_{j})}\cdot\textsc{cost}(X_{j},c_{j})\right\}.$ (37)
-A random point $x\in X$ sampled via $D^{2}$-sampling (2 of Algorithm 5) turns out to satisfy:
+$\widehat{X}_{j}\penalty 10000\ :=\penalty 10000\ \left\{y\in X_{j}:\ \mathrm{dist}^{2}(y,c_{j})\leq\frac{2}{w(X_{j})}\cdot\textsc{cost}(X_{j},c_{j})\right\}.$ (38)
+A random point $x\in X$ sampled via $D^{2}$-sampling (3 of Algorithm 5) turns out to satisfy:
   1. (a)
-    Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]\geq\Omega(\epsilon^{6})$.
+    $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]\geq\Omega(\epsilon^{6})$.
   2. (b)
@@
 And since $\mathcal{B}(A+x)\subseteq\mathcal{B}(A)$ (Claim 13.4), we conclude with $|\mathcal{B}(A+x)|\leq|\mathcal{B}(A)|-1$.
-Item (a). According to the $D^{2}$-sampling procedure (2 of Algorithm 5), we have
-Pr[xX^j for some $\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]$ $\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\textsc{cost}(X,S+A)}$
-$\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}\Bigg{.}}_{(\dagger)}\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\Bigg{.}}_{(\ddagger)}.$
+Item (a). According to the $D^{2}$-sampling procedure (3 of Algorithm 5), we have
+$\Pr[\text{$x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$}]$ $\penalty 10000\ \geq\penalty 10000\ \Omega(\epsilon^{6})\cdot\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\textsc{cost}(X,S+A)}$
+$\penalty 10000\ =\penalty 10000\ \Omega(\epsilon^{6})\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}\Bigg.}_{(\dagger)}\cdot\underbrace{\frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\Bigg.}_{(\ddagger)}.$
 It remains to show that $(\dagger)=\Omega(1)$ and $(\ddagger)=\Omega(1)$.
-For the term ()($\dagger$)( † ), we deduce that
-(($\dagger$)( † ) = $\leavevmode\nobreak\ =\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A)+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}$
-$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 1-\frac{\textsc{cost}(X_{0},S+A^{*})+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}8\cdot\textsc{cost}(X_{j},S+A^{*})}{16\cdot\textsc{cost}(X,S+A^{*})}$
-$\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}.$
+For the term $(\dagger)$, we deduce that
+$(\dagger)$ = $\penalty 10000\ =\penalty 10000\ 1-\frac{\textsc{cost}(X_{0},S+A)+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}{\textsc{cost}(X,S+A)}$
+$\penalty 10000\ \geq\penalty 10000\ 1-\frac{\textsc{cost}(X_{0},S+A^{*})+\sum_{j\neq 0\land j\notin\mathcal{B}(A)}8\cdot\textsc{cost}(X_{j},S+A^{*})}{16\cdot\textsc{cost}(X,S+A^{*})}$
+$\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}.$
 Here the second step applies a combination of $\textsc{cost}(X_{0},S+A^{*})=\textsc{cost}(X_{0},S)\geq\textsc{cost}(X_{0},S+A)$ (Equation 35), $\textsc{cost}(X_{j},S+A)\leq 8\cdot\textsc{cost}(X_{j},S+A^{*})$ when $j\neq 0\land j\notin\mathcal{B}(A)$ (Equation 37), and $\textsc{cost}(X,S+A)>16\cdot\textsc{cost}(X,S+A^{*})$ (the premise of Claim 13.6).
-For the term ()($\ddagger$)( ‡ ), let us consider a specific bad cluster $j\in\mathcal{B}(A)$.
+For the term $(\ddagger)$, let us consider a specific bad cluster $j\in\mathcal{B}(A)$.
 Let $\alpha:=\sqrt{w(X_{j})}\cdot\mathrm{dist}(c_{j},S+A)$ and $\beta:=\sqrt{\textsc{cost}(X_{j},c_{j})}$ for brevity.
 We can upper-bound $\textsc{cost}(X_{j},S+A)$ using these parameters.
-$\textsc{cost}(X_{j},S+A)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}\mathrm{dist}(y,c_{j})+\mathrm{dist}(c_{j},S+A)\big{)}^{2}$ (triangle inequality)
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(c_{j},S+A)\big{)}$ (AM-GM inequality)
-= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 2\beta^{2}+2\alpha^{2}.$
+$\textsc{cost}(X_{j},S+A)$ $\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(\mathrm{dist}(y,c_{j})+\mathrm{dist}(c_{j},S+A)\big)^{2}$ (triangle inequality)
+$\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(c_{j},S+A)\big)$ (AM-GM inequality)
+= $\penalty 10000\ =\penalty 10000\ 2\beta^{2}+2\alpha^{2}.$
 This together with $\textsc{cost}(X_{j},S+A)\geq 8\cdot\textsc{cost}(X_{j},c_{j})=8\beta^{2}$ (Equation 37) also gives $\alpha\geq\sqrt{3}\beta$.
@@
 Altogether, we deduce that
-$\textsc{cost}(\widehat{X}_{j},S+A)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{y\in\widehat{X}_{j}}w(y)\cdot\mathrm{dist}^{2}(y,S+A)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{1}{2}\cdot(\alpha-\sqrt{2}\beta)^{2}.$
+$\textsc{cost}(\widehat{X}_{j},S+A)\penalty 10000\ =\penalty 10000\ \sum_{y\in\widehat{X}_{j}}w(y)\cdot\mathrm{dist}^{2}(y,S+A)\penalty 10000\ \geq\penalty 10000\ \frac{1}{2}\cdot(\alpha-\sqrt{2}\beta)^{2}.$
 These two observations in combination with arbitrariness of $j\in\mathcal{B}(A)$ implies that
-$(\ddagger)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{\left(\alpha-\sqrt{2}\beta\right)^{2}}{4\beta^{2}+4\alpha^{2}}\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \frac{(\sqrt{3}-\sqrt{2})^{2}}{4+4\cdot(\sqrt{3})^{2}}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{5-2\sqrt{6}}{16}.$
+$(\ddagger)\penalty 10000\ \geq\penalty 10000\ \frac{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(\widehat{X}_{j},S+A)}{\sum_{j\in\mathcal{B}(A)}\textsc{cost}(X_{j},S+A)}\penalty 10000\ \geq\penalty 10000\ \frac{\left(\alpha-\sqrt{2}\beta\right)^{2}}{4\beta^{2}+4\alpha^{2}}\penalty 10000\ \geq\penalty 10000\ \frac{(\sqrt{3}-\sqrt{2})^{2}}{4+4\cdot(\sqrt{3})^{2}}\penalty 10000\ =\penalty 10000\ \frac{5-2\sqrt{6}}{16}.$
 Here the last step follows from elementary algebra; specifically, the equality holds when $\alpha=\sqrt{3}\beta$.
 Item (b). If $x\in\widehat{X}_{j}$ for some $j\in\mathcal{B}(A)$, then the triangle and AM-GM inequalities together give
-$\textsc{cost}(X_{j},x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{y\in X_{j}}w(y)\cdot\big{(}2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(x,c_{j})\big{)}$
-$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\cdot\textsc{cost}(X_{j},c_{j})+4\cdot\textsc{cost}(X_{j},c_{j})$ (Equation 37)
-= $\leavevmode\nobreak\ =\leavevmode\nobreak\ 6\cdot\textsc{cost}(X_{j},S+A^{*}).$ (Equation 35)
+$\textsc{cost}(X_{j},x)$ $\penalty 10000\ \leq\penalty 10000\ \sum_{y\in X_{j}}w(y)\cdot\big(2\cdot\mathrm{dist}^{2}(y,c_{j})+2\cdot\mathrm{dist}^{2}(x,c_{j})\big)$
+$\penalty 10000\ \leq\penalty 10000\ 2\cdot\textsc{cost}(X_{j},c_{j})+4\cdot\textsc{cost}(X_{j},c_{j})$ (Equation 38)
+= $\penalty 10000\ =\penalty 10000\ 6\cdot\textsc{cost}(X_{j},S+A^{*}).$ (Equation 36)
 This finishes the proof of Claim 13.6. ∎
 Proof of Claim 13.3.
-For every $i\in[a+1]$, let $A_{i}$ denote the center set $A$ after the $i$-th iteration of the for loop (1, 2 and 3) in Algorithm 5.
+For every $i\in[a+1]$, let $A_{i}$ denote the center set $A$ after the $i$-th iteration of the for loop (2, 3 and 4) in Algorithm 5.
 Also, let $A_{0}:=\emptyset$.
@@
 where the last step applies the union bound. Then, consider a fixed $i\in[a+1]$. We have
-$\Pr\big{[}\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big{]}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \Pr\big{[}\neg\mathcal{F}_{i}\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Pr\big{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}.$
-Recall that $A_{i}=A_{i-1}+\{x_{1},\dots,x_{t}\}$, where $x_{1},\dots,x_{t}$ are the independent points sampled in the $i$-th iteration by 2. According to Claim 12.6, we have
-$\forall j\in[t],\quad\Pr\big{[}|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq 1-\Omega(\epsilon^{6}).$
+$\Pr\big[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}\big]\penalty 10000\ \leq\penalty 10000\ \Pr\big[\neg\mathcal{F}_{i}\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\penalty 10000\ =\penalty 10000\ \Pr\big[|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big].$
+Recall that $A_{i}=A_{i-1}+\{x_{1},\dots,x_{t}\}$, where $x_{1},\dots,x_{t}$ are the independent points sampled in the $i$-th iteration by 3. According to Claim 13.6, we have
+$\forall j\in[t],\quad\Pr\big[|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\leq 1-\Omega(\epsilon^{6}).$
 Let $t=O(\epsilon^{-6}\cdot d\log(\Delta))$ be sufficiently large, we have
-$\Pr\big{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\big{]}\leq\left(1-\Omega(\epsilon^{6})\right)^{t}\leq\frac{1}{\mathrm{poly}(\Delta^{d})}.$
+$\Pr\big[\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \big|\penalty 10000\ \neg\mathcal{E}_{i}\big]\leq\left(1-\Omega(\epsilon^{6})\right)^{t}\leq\frac{1}{\mathrm{poly}(\Delta^{d})}.$
 For every $j\in[t]$, by Claim 13.4, we have $\mathcal{B}(A_{i-1}+x_{j})\supseteq\mathcal{B}(A_{i})$.
@@
 As a result,
-$\Pr[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}]\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ \Pr\bigg{[}|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
-$\leq\leavevmode\nobreak\ \Pr\bigg{[}\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \neg\mathcal{E}_{i}\bigg{]}$
-$\leq\leavevmode\nobreak\ \frac{1}{\mathrm{poly}(\Delta^{d})}.$
+$\Pr[\neg\mathcal{E}_{i}\land\neg\mathcal{F}_{i}]\penalty 10000\$$\leq\penalty 10000\ \Pr\bigg[|\mathcal{B}(A_{i})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \bigg|\penalty 10000\ \neg\mathcal{E}_{i}\bigg]$
+$\leq\penalty 10000\ \Pr\bigg[\forall j\in[t],|\mathcal{B}(A_{i-1}+x_{j})|>|\mathcal{B}(A_{i-1})|-1\penalty 10000\ \bigg|\penalty 10000\ \neg\mathcal{E}_{i}\bigg]$
+$\leq\penalty 10000\ \frac{1}{\mathrm{poly}(\Delta^{d})}.$
 Finally, we apply the union bound over all $i\in[a+1]$, plug the result into Equation 39, and obtain $\Pr[\land_{i\in[a+1]}(\mathcal{E}_{i}\lor\mathcal{F}_{i})]\geq 1-\frac{1}{\mathrm{poly}(\Delta^{d})}$.
@@
 In this part of the paper, we provide our algorithm, its implementation, and its analysis.
-We start by explaining the framework of [BCF25a] in Section 13, with some major modifications that are necessary to ensure that we can implement it using our geometric data structures (see Part II).
+We start by explaining the framework of [BCF24] in Section 14, with some major modifications that are necessary to ensure that we can implement it using our geometric data structures (see Part II).
 Next, in Section 15, we show how to implement this modified framework in Euclidean spaces with an update time of $\tilde{O}(n^{\epsilon})$.
-Finally, in Sections 15, 16, 17 and 18, we provide the analysis of the algorithm.
+Finally, in Sections 16, 17, 18 and 19, we provide the analysis of the algorithm.
 Theorem 13.7.
@@
 $d$ is the dimension of the space.
-$n$ is the maximum size of the space, and $[0,\Delta]$ is the range of the coordinates of the points.
-Note that the aspect ratio of the space would be $\sqrt{d}\Delta$ since we only consider points in the grid $[\Delta]^{d}$.
+$n$ is the maximum size of the space, and $0,\Delta]$ is the range of the coordinates of the points.
+Note that the aspect ratio of the space would be $\sqrt{d}\Delta$ since we only consider points in the grid $\Delta]^{d}$.
 Preprocessing.
-Throughout Sections 14, 15, 16, 17 and 18, we assume that the dimension of the input space $d$ is $O(\log n)$.
+Throughout Sections 15, 16, 17, 18 and 19, we assume that the dimension of the input space $d$ is $O(\log n)$.
 Hence, the running time performance of all the data structures in Part II becomes $\tilde{O}(n^{\epsilon})$.
@@
 Assume that the input space is $\mathbb{R}^{m}$.
-We can initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$, and whenever we have an insertion of a point $x\in\mathbb{R}^{m}$, we apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $[\Delta]^{d}$.
+We can initialize a random matrix $A$ of size $m\times d$ where $d=O(\log n)$, and whenever we have an insertion of a point $x\in\mathbb{R}^{m}$, we apply the transformation (compute $Ax\in\mathbb{R}^{d}$) and then round it to a discrete point in $\Delta]^{d}$.
 Then, we consider feeding this new dynamic space into our algorithm.
 Note that according to this transformation, we will have an additive $\tilde{O}(m)$ in the update time of the algorithm, but we omit to write this down from now on since it only appears because of this preprocessing step, and the rest of the algorithm will not depend on that.
-13 The Algorithmic Framework of [BCF25a]
+14 The Algorithmic Framework of [BCF24]
 Roadmap.
-In this section, we explain how the algorithmic framework of [BCF25a] works, with some significant modifications that are necessary for our purpose.
+In this section, we explain how the algorithmic framework of [BCF24] works, with some significant modifications that are necessary for our purpose.
 We start by defining the modified version of the main tool used in the algorithm in Section 14.1, which is called the robust solution.
@@
 Remark 14.1.
-Although the algorithm of [BCF25a] is primarily designed for $k$-median, it also works for $k$-means.
-The only consideration is that the analysis of [BCF25a] depends on the fact that the integrality gap of the LP-relaxation for $k$-median is constant.
-Since this fact is also correct for $k$-means (see Theorem 1.3 in [ANFSW20]), we can extend the analysis of [BCF25a] for $k$-means.
+Although the algorithm of [BCF24] is primarily designed for $k$-median, it also works for $k$-means.
+The only consideration is that the analysis of [BCF24] depends on the fact that the integrality gap of the LP-relaxation for $k$-median is constant.
+Since this fact is also correct for $k$-means (see Theorem 1.3 in [integrality-gap-k-means]), we can extend the analysis of [BCF24] for $k$-means.
 The constant factors in the analysis might change, and here we provide a thorough analysis of the algorithm specifically for Euclidean $k$-means.
 14.1 Robust Solution
-Before we explain the algorithm of [BCF25a], we start with the necessary definitions for describing the algorithm.
-Since our goal is to achieve an update time of $\tilde{O}(n^{\epsilon})$, we slightly changed some of the structural definitions of [BCF25a] that enable us to work with approximate balls around points instead of exact balls.
+Before we explain the algorithm of [BCF24], we start with the necessary definitions for describing the algorithm.
+Since our goal is to achieve an update time of $\tilde{O}(n^{\epsilon})$, we slightly changed some of the structural definitions of [BCF24] that enable us to work with approximate balls around points instead of exact balls.
 Definition 14.2.
-Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $[\Delta]^{d}$.
+Let $t\geq 0$ be an integer and $(x_{0},x_{1},\ldots,x_{t})$ be a sequence of $t+1$ points of $\Delta]^{d}$.
 We call this sequence $t$-robust w.r.t.
@@
 We call a center set $S$, robust w.r.t. $X$ if and only if each center $u\in S$ satisfies the following
-$u\text{ is at least }t\text{-robust, where }t\text{ is the smallest integer satisfying }\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}.$ (42)
+[…]
 14.2 Description of The Algorithm
-In this section, we describe the algorithm framework of [BCF25a].
+In this section, we describe the algorithm framework of [BCF24].
 For simplicity.
@@
 Algorithm 6 Computing an estimate $\widehat{\ell}$ of the value of $\ell^{\star}$.
-[…]
-[…]
-[…]
+1: for $i=0$ to $\log_{2}k$ do Using restricted $k$-means (Lemma 8.5), in $\tilde{O}(n^{\epsilon}s_{i})$ time compute a subset $\widehat{S}_{i}\subseteq S_{\textnormal{{init}}}$ of $(k-s_{i})$ centers, that is a $p_{\epsilon}$-approximation to $\mathrm{OPT}_{k-s_{i}}^{S_{\textnormal{{init}}}}\left(X^{(0)}\right)$.
+2:   if $\textsc{cost}(X^{(0)},\widehat{S}_{i})>(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})$ then
+3:    return $\widehat{\ell}:=s_{i-1}$.
 After finding $\widehat{\ell}$, we set the length of the epoch to be $\ell+1$ where
-^$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot\lambda^{51}}\right\rfloor.$
-With some extra calculations, with the same argument as in [BCF25a], we can show that $\ell+1=\tilde{\Omega}(\widehat{\ell}+1)=\tilde{\Omega}(\ell^{\star}+1)$ and
+$\ell\leftarrow\left\lfloor\frac{\widehat{\ell}}{36\cdot\lambda^{51}}\right\rfloor.$
+With some extra calculations, with the same argument as in [BCF24], we can show that $\ell+1=\tilde{\Omega}(\widehat{\ell}+1)=\tilde{\Omega}(\ell^{\star}+1)$ and
 $\frac{\mathrm{OPT}_{k-{\ell}}(X^{(0)})}{\lambda^{51}}\leq\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+{\ell}}(X^{(0)}).$
@@
 we want to maintain the variant that $t[u]$ is always greater than or equal to the smallest integer $t$ satisfying $\lambda^{3t}\geq\mathrm{dist}(u,S-u)/\lambda^{10}$ (see Condition (43)).
-Hence, to check whether a center $u$ satisfies Condition (42), it suffices to check whether $t[u]\geq t$ [^16] Note that there might be a center $u$ such that $t[u]<t$, but $u$ is still $t$-robust.
+Hence, to check whether a center $u$ satisfies Condition (43), it suffices to check whether $t[u]\geq t$.
+[^16] Note that there might be a center $u$ such that $t[u]<t$, but $u$ is still $t$-robust.
 More precisely, we maintain an approximation of $\mathrm{dist}(u,S-u)$ using Lemma 8.3 and find the smallest integer $t$ satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$.
@@
 Lemma 15.4.
-Consider any call to Robustify$(W^{\prime})$, and suppose that it sets $w_{0}\leftarrow\texttt{MakeRobust}(w)$ during some iteration of the while loop.
-Then in subsequent iterations of the while loop in the same call to Robustify$(W^{\prime})$, we will not make any call to MakeRobust$(w_{0})$.
+[…]
 We provide the proof of this lemma in Section 21.9.
@@
 Algorithm 7 Implementation of the call to $\texttt{Robustify}(W^{\prime})$ at the end of an epoch.
-[…]
-[…]
-3:$S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$Foreach $u\in W_{1}\cup W_{2}$
-[…]
-5:$u\leftarrow$pop$(\mathcal{Y})$
-[…]
-[…]
-[…]
-[…]
+1: $W_{1}\leftarrow W^{\prime}\setminus S_{\textnormal{{init}}}$ $\triangleright$ All new centers
+2: for each $x\in X^{(0)}\oplus X^{(\ell+1)}$ do
+3:   for each $i\in[0,\lceil\log_{\lambda^{3}}(\sqrt{d}\Delta)\rceil]$ do
+4:    Let $u\leftarrow\textsc{NN}(x,S[i])$
+5:    if $\mathrm{dist}(u,x)\leq\lambda^{3i+1.5}$ then $W_{2}\leftarrow W_{2}+u$ $\triangleright$ Mark $u$ as a contaminated center
+6: $S\leftarrow S_{\textnormal{{init}}}-(S_{\textnormal{{init}}}\setminus W^{\prime})+(W^{\prime}\setminus S_{\textnormal{{init}}})$
+7: for each $u\in W_{1}\cup W_{2}$ do
+8:   $\texttt{MakeRobust}(u)$.
+9: while $\mathcal{Y}\neq\emptyset$ do
+10:   $u\leftarrowpop$$(\mathcal{Y})$
+11:   $t\leftarrow$ Smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{10}$
+12:   if $t[u]\geq t$ then
+13:    go back to 9
+14:   else
+15:    $\texttt{MakeRobust}(u)$.
+16: return $S_{\textnormal{{final}}}\leftarrow S$ .
 15.2.2 Make-Robust
@@
 1: $t\leftarrow$ Smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(u,S-u)/\lambda^{7}$
-[…]
-[…]
-4:$x_{j-1}\leftarrow x_{j}$Else
-[…]
-6:$u_{0}\leftarrow x_{0}$
-7:$S\leftarrow S-u+u_{0}$
-8:Save $t[u_{0}]\leftarrow t$ together with $u_{0}$
+2: $x_{t}\leftarrow u$
+3: for $j=t$ down to $1$ do
+4:   Invoke Lemma 8.4 for $x=x_{j}$ and $r=\lambda^{3j}$ to get $c^{*}\in[\Delta]^{d}$, $\widehat{\textsc{cost}}_{x_{j}}$, $\widehat{\textsc{cost}}_{c^{*}}$ and $b_{x_{j}}$
+5:   if $\left(\widehat{\textsc{cost}}_{x_{j}}/b_{x_{j}}\geq\lambda^{3j-1}/p_{\epsilon}\right)$ or $\left(\widehat{\textsc{cost}}_{x_{j}}/p_{\epsilon}\leq\widehat{\textsc{cost}}_{c^{*}}\right)$ then
+6:    $x_{j-1}\leftarrow x_{j}$
+7:   else
+8:    $x_{j-1}\leftarrow c^{*}$
+9: $u_{0}\leftarrow x_{0}$
+10: $S\leftarrow S-u+u_{0}$
+11: Save $t[u_{0}]\leftarrow t$ together with $u_{0}$
 16 Analysis of Our Dynamic Algorithm: Proof of Theorem 13.7
@@
 where $C_{v}$ is the cluster of $v$ w.r.t. the center set $V$, i.e. $C_{v}=\{x\in X\mid\mathrm{dist}(x,v)=\mathrm{dist}(x,V)\}$.
-Lemma 15.6 ($k$-Means Version of Lemma 4.2 [BCF25a]).
+Lemma 16.6 ($k$-Means Version of Lemma 4.2 [BCF24]).
 Let $r\geq 0$ and $m\in[0,k]$.
@@
 If the number of well-separated pairs w.r.t.
-$(S,V)$ is $k-m$, then there exists a subset $\tilde{S}\subseteq S$ of size at most $k-\lfloor(m-r)/4\rfloor$ such that
+$(S,V)$ is kmk-m, then there exists a subset $\tilde{S}\subseteq S$ of size at most $k-\lfloor(m-r)/4\rfloor$ such that
 $\textsc{cost}(X,\tilde{S})\leq 81\lambda^{40}\cdot\left(\textsc{cost}(X,S)+\textsc{cost}(X,V)\right).$
@@
 Then for every $k\geq 1$, we have $\mathrm{OPT}_{k+s}(X^{\prime})\leq\mathrm{OPT}_{k}(X)$.
-Lemma 15.9 ($k$-Means Version of Double-Sided Stability (Lemma 4.1 [BCF25a])).
+Lemma 16.9 ($k$-Means Version of Double-Sided Stability (Lemma 4.1 [BCF24])).
 Consider any $r\in[0,k-1]$ and any $\eta\geq 1$.
@@
 the current dataset.
-According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Step 1 of our algorithm (see Algorithm 6, Algorithm 6), we infer that
+According to the definition of $\widehat{\ell}=s_{i-1}$ in Step 1 of our algorithm (see Algorithm 6, 2), we infer that
 $\mathrm{OPT}_{k-\widehat{\ell}}(X^{(0)})\leq\textsc{cost}(X^{(0)},\widehat{S}_{i-1})\leq(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}).$
@@
 $\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor}(X^{(0)}).$
-Finally, since =^$\ell=\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor\leq\widehat{\ell}$, we have
+Finally, since $\ell=\lfloor\widehat{\ell}/(36\cdot\lambda^{51})\rfloor\leq\widehat{\ell}$, we have
 $\frac{\mathrm{OPT}_{k-\ell}(X^{(0)})}{\lambda^{51}}\leq\mathrm{OPT}_{k}(X^{(0)})\leq 12\cdot\mathrm{OPT}_{k+\ell}(X^{(0)}).$ (54)
@@
 Let $V\subseteq[\Delta]^{d}$ be an optimal $(k+\ell+1)$-median solution for the dataset $X^{(0)}$, i.e., $\left|V\right|=k+\ell+1$ and $\textsc{cost}(X^{(0)},V)=\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)$.
-Let $m\in[0,k]$ be the unique integer such that there are $(k-m)$ well-separated pairs w.r.t.
+Let $m\in[0,k]$ be the unique integer such that there are (km)(k-m) well-separated pairs w.r.t.
 $\left(S_{\textnormal{{init}}},T\right)$.
-Let $\{(u_{1},v_{1}),(u_{2},v_{2}),\ldots,(u_{k-m},v_{k-m})\}\subseteq S_{\textnormal{{init}}}\times V$ denote the collection of $(k-m)$ well-separated pairs w.r.t.
+Let $\{(u_{1},v_{1}),(u_{2},v_{2}),\ldots,(u_{k-m},v_{k-m})\}\subseteq S_{\textnormal{{init}}}\times V$ denote the collection of (km)(k-m) well-separated pairs w.r.t.
 $\left(S_{\textnormal{{init}}},V\right)$.
@@
 $\textsc{cost}(X^{(0)},S^{\prime})$ $\leq p_{\epsilon}\cdot\min\limits_{A\subseteq[\Delta]^{d}:|A|\leq s}\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+A)$
-$\leq p_{\epsilon}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}}+\tilde{A})$
+[…]
 $\leq 4\cdot(p_{\epsilon})^{4}\cdot\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right),$
@@
 We apply Lemma 16.6, by setting $r=\ell+1$, $S=S_{\textnormal{{init}}}$, $X=X^{(0)}$.
-This implies the existence of a set $\tilde{S}\subseteq S_{\textnormal{{init}}}$ of at most $(k-b)$ centers, with $b=\lfloor(m-\ell-1)/4\rfloor$, such that
-$\textsc{cost}(X^{(0)},\tilde{S})$$\leq$ $81\lambda^{40}\cdot\left(\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})+\textsc{cost}(X^{(0)},V)\right)$
+This implies the existence of a set $\tilde{S}\subseteq S_{\textnormal{{init}}}$ of at most (kb)(k-b) centers, with $b=\lfloor(m-\ell-1)/4\rfloor$, such that
+[…]
 $\leq$ $81\lambda^{40}\cdot\left((p_{\epsilon})^{6}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right)+\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)\right)\leq\lambda^{44}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right).$
-In the above derivation, the second inequality follows from the initial assumption at the beginning of the epoch (52), and the last inequality holds because $\mathrm{OPT}_{k+\ell+1}\left(X^{(0)}\right)\leq\mathrm{OPT}_{k}(X^{(0)})$ and $\lambda=(p_{\epsilon})^{2}$ is a large number (see Part III).
-Since $\mathrm{OPT}_{k-b}\left(X^{(0)}\right)\leq\textsc{cost}(X^{(0)},\tilde{S})$, we get
+[…]
 $\mathrm{OPT}_{k-b}\left(X^{(0)}\right)\leq\lambda^{44}\cdot\mathrm{OPT}_{k}\left(X^{(0)}\right).$ (65)
@@
 From (65), it follows that $b\leq\ell^{\star}$.
-According to the definition of ^$\widehat{\ell}=s_{i-1}$ in Algorithm 6, for $s_{i}=2s_{i-1}=2\widehat{\ell}$, from (Algorithm 6), we have
+According to the definition of $\widehat{\ell}=s_{i-1}$ in Algorithm 6, for $s_{i}=2s_{i-1}=2\widehat{\ell}$, from (2), we have
 $(p_{\epsilon})^{8}\cdot\lambda^{44}\cdot\textsc{cost}(X^{(0)},S_{\textnormal{{init}}})<\textsc{cost}(X^{(0)},\widehat{S}_{i})\leq p_{\epsilon}\cdot\mathrm{OPT}^{S_{\textnormal{{init}}}}_{k-2\widehat{\ell}}(X^{(0)})$
@@
 This shows that $2\widehat{\ell}\geq\ell^{\star}$ by the definition of $\ell^{\star}$.
-Now, we have $b\geq\frac{m-\ell-1}{4}-1=\frac{m-\ell-5}{4}$ and ^$\ell\geq\frac{\widehat{\ell}}{36\cdot\lambda^{51}}-1$, we get
+Now, we have $b\geq\frac{m-\ell-1}{4}-1=\frac{m-\ell-5}{4}$ and $\ell\geq\frac{\widehat{\ell}}{36\cdot\lambda^{51}}-1$, we get
 $\frac{m-\ell-5}{4}\leq b\leq\ell^{\star}\leq 2\widehat{\ell}\leq 2\cdot 36\cdot\lambda^{51}\cdot(\ell+1)$
@@
 Lemma 18.1.
-The amortized recourse of the algorithm is at most $\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$, where $d$ is the dimension of the space, $n$ is the upper bound on the number of the points in the data-set, and $[0,\Delta]$ is the range of the coordinates of the points.
+The amortized recourse of the algorithm is at most $\mathrm{poly}(d\cdot\log(n\Delta)/\epsilon)$, where $d$ is the dimension of the space, $n$ is the upper bound on the number of the points in the data-set, and $0,\Delta]$ is the range of the coordinates of the points.
 Roadmap.
@@
 $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|\leq\tilde{O}(\ell+1).$ (68)
-Since the epoch lasts for $\ell+1$ updates, the $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ and $24\ell+$ 2 terms in the right hand side of (66) contributes an amortized recourse of $\tilde{O}(1)$.
+Since the epoch lasts for $\ell+1$ updates, the $\left|S_{\textnormal{{init}}}\oplus W^{\prime}\right|$ and $4\ell+2$ terms in the right hand side of (67) contributes an amortized recourse of $\tilde{O}(1)$.
 Moreover, the term $\left|W^{\prime}\oplus S_{\textnormal{{final}}}\right|$ is proportional to the number of MakeRobust calls made while computing $S_{\textnormal{{final}}}\leftarrow\texttt{Robustify}\left(W^{\prime}\right)$.
@@
 Consider one Type III call $w_{i}\leftarrow\texttt{MakeRobust}(w_{i-1})$.
-This call is made on 8 in Algorithm 7, which means that $t[w_{i-1}]$ satisfies $t[w_{i-1}]<t$ where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}$.
+This call is made on 15 in Algorithm 7, which means that $t[w_{i-1}]$ satisfies $t[w_{i-1}]<t$ where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{i-1},W^{\prime}-w_{i-1})/\lambda^{10}$.
 Hence,
@@
 As a result, $t[w_{0}]<t[w_{1}]<\ldots<t[w_{j}]$.
-Since the aspect ratio of the space is $(\sqrt{d}\Delta)$, all of the integers $t[w_{i}]$ are between 00 and $\log_{\lambda^{3}}(\sqrt{d}\Delta)<\log(\sqrt{d}\Delta)$ which concludes $j\leq\log(\sqrt{d}\Delta)$.
+Since the aspect ratio of the space is $(\sqrt{d}\Delta)$, all of the integers $t[w_{i}]$ are between 0 and $\log_{\lambda^{3}}(\sqrt{d}\Delta)<\log(\sqrt{d}\Delta)$ which concludes $j\leq\log(\sqrt{d}\Delta)$.
 ∎
@@
 But, considering the entire sequence of updates, we show that the amortized update time of this part of the Robustify subroutine is at most $\tilde{O}(n^{\epsilon})$.
-According to our algorithm, for each center $u$ in the yellow set $\mathcal{Y}$, in $\tilde{O}(1)$ time (5 to 6) [^20] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 7.3 ., our algorithm decides whether or not a call to MakeRobust will be made on $u$.
+According to our algorithm, for each center $u$ in the yellow set $\mathcal{Y}$, in $\tilde{O}(1)$ time (10 to 12) [^20] Note that the number $\widehat{\mathrm{dist}}(u,S-u)$ is maintained explicitly by the data structure in Lemma 8.3 ., our algorithm decides whether or not a call to MakeRobust will be made on $u$.
 Each call to MakeRobust also takes $\tilde{O}(n^{\epsilon})$ time according to Section 19.3.
@@
 The main for loop in the implementation consists of at most $O(\log_{\lambda}(\sqrt{d}\Delta))$ many iterations.
-Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 7.4) used in 3 of the implementation.
+Each iteration is done in $\tilde{O}(n^{\epsilon})$ time according to the data structure (Lemma 8.4) used in 4 of the implementation.
 It is obvious that all other parts of the algorithm take at most $\tilde{O}(1)$ time.
@@
 In this section, we show how to improve the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$.
-We do this by using dynamic sparsification as a black box, in a similar manner to previous dynamic clustering algorithms [BCG+24, BCF25a].
-In [BCF25a], the authors give an algorithm that maintains a $O(1)$-approximation with $\tilde{O}(1)$ recourse and $\tilde{O}(n)$ update time.
-Using the dynamic $k$-means algorithm of [BCLP23] as a dynamic sparsifier in a black-box manner, they can then improve the update time of their algorithm to $\tilde{O}(k)$, while only incurring $O(1)$ overhead in the approximation and $\tilde{O}(1)$ overhead in the recourse.
-We speed up the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ using a completely analogous approach, except that we use the sparsifier of [DHS24], which is more well-suited to Euclidean spaces.
+We do this by using dynamic sparsification as a black box, in a similar manner to previous dynamic clustering algorithms [BhattacharyaCGLP24, BCF24].
+In [BCF24], the authors give an algorithm that maintains a $O(1)$-approximation with $\tilde{O}(1)$ recourse and $\tilde{O}(n)$ update time.
+Using the dynamic $k$-means algorithm of [nips/BhattacharyaCLP23] as a dynamic sparsifier in a black-box manner, they can then improve the update time of their algorithm to $\tilde{O}(k)$, while only incurring $O(1)$ overhead in the approximation and $\tilde{O}(1)$ overhead in the recourse.
+We speed up the update time of our algorithm from $\tilde{O}(n^{\epsilon})$ to $\tilde{O}(k^{\epsilon})$ using a completely analogous approach, except that we use the sparsifier of [esa/TourHS24], which is more well-suited to Euclidean spaces.
 Our Sparsified Dynamic Algorithm.
-The following theorem summarises the relevant properties of the dynamic algorithm of [DHS24], which we call Sparsifier.
-[^21] The dynamic algorithm of [DHS24] dynamically maintains a coreset of the input space.
+The following theorem summarises the relevant properties of the dynamic algorithm of [esa/TourHS24], which we call Sparsifier.
+[^21] The dynamic algorithm of [esa/TourHS24] dynamically maintains a coreset of the input space.
 For the sake of simplicity, we do not define the notion of a coreset and instead focus on the fact that it can be used as a sparsifier.
 We use this algorithm as a black-box.
-Theorem 19.1 ([DHS24]).
+Theorem 20.1 ([esa/TourHS24]).
 There is an algorithm Sparsifier that, given a dynamic weighted $d$-dimensional Euclidean space $X$, maintains a weighted subspace $U\subseteq X$ of size $\tilde{O}(k)$ with $\tilde{O}(1)$ amortized update time and $\tilde{O}(1)$ amortized recourse.
@@
 Proof.
-By Theorem 19.2, at each point in time, we have that $\textsc{cost}(U,S_{i}^{\prime})\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $21-1/\mathrm{poly}(k)\geq 1/$ 2 for each $i\in[L]$.
+By Theorem 20.2, at each point in time, we have that $\textsc{cost}(U,S_{i}^{\prime})\leq\alpha\cdot\mathrm{OPT}_{k}(U)$ with probability at least $1-1/\mathrm{poly}(k)\geq 1/2$ for each $i\in[L]$.
 Since each of the algorithms $\mathrm{\textsc{ALG}}_{1}^{\prime},\dots,\mathrm{\textsc{ALG}}_{L}^{\prime}$ uses independent randomness, the probability that $\textsc{cost}(U,S_{i}^{\prime})>\alpha\cdot\mathrm{OPT}_{k}(U)$ for all $i\in[L]$ is at most $(1/2)^{L}=1/\mathrm{poly}(n)$.
@@
 ∎
-Now, let $p(n)=\mathrm{poly}(n)=\Omega(n)$ be the polynomial such that the approximation guarantee of ALG holds with probability at least $1-1/p(n)$ after each update, as described in Theorem 19.2.
+Now, let $p(n)=\mathrm{poly}(n)=\Omega(n)$ be the polynomial such that the approximation guarantee of ALG holds with probability at least 11/p(n)1-1/p(n) after each update, as described in Theorem 20.2.
 Consider some subsequence of updates $\sigma_{\ell+1},\dots,\sigma_{\ell+q}$, where $q\leq p(k)/2$.
@@
 if we restart the algorithm $\mathrm{\textsc{ALG}}^{\star}$ with fresh randomness at the end of the update.
-Suppose that $r^{\prime}\geq i-1$ and let $\sigma_{j_{1}}^{\prime},\dots,\sigma_{j_{i-1}}^{\prime}$ be the first $i-1$ bad updates in $\sigma^{\prime}_{1},\dots,\sigma^{\prime}_{q}$.
+Suppose that $r^{\prime}\geq i-1$ and let $\sigma_{j_{1}}^{\prime},\dots,\sigma_{j_{i-1}}^{\prime}$ be the first i1i-1 bad updates in $\sigma^{\prime}_{1},\dots,\sigma^{\prime}_{q}$.
 Since we restart $\mathrm{\textsc{ALG}}^{\star}$ after update $\sigma^{\prime}_{j_{i-1}}$, the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ that effect whether any of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ are bad are independent of the random bits used by $\mathrm{\textsc{ALG}}^{\star}$ to handle the previous updates.
 Thus, conditioned on this event, each of the updates $\sigma_{j_{i-1}+1}^{\prime},\dots,\sigma_{q}^{\prime}$ is bad with probability at most $1/p(k)$ by Claim 20.7.
-Taking a union bound, we get that none of them is bad with probability at least $21-q/p(k)\geq 1/$ 2.
+Taking a union bound, we get that none of them is bad with probability at least $1-q/p(k)\geq 1/2$.
 ∎
@@
 ∎
-It follows from Lemmas 19.5 and 19.9 that $R=\tilde{O}(T)+k\cdot\tilde{O}(T/k)=\tilde{O}(T)$ with probability at least $1-1/\mathrm{poly}(n)$.
+It follows from Lemmas 20.5 and 20.9 that $R=\tilde{O}(T)+k\cdot\tilde{O}(T/k)=\tilde{O}(T)$ with probability at least $1-1/\mathrm{poly}(n)$.
 Thus, the amortized recourse of our algorithm is $\tilde{O}(1)$ with high probability.
@@
 Since the algorithm uses fresh randomness, the probability that we restart it again (i.e.
-that $r_{i}^{\star}\geq j+1$) is at most $21/p(k)\leq 1/$ 2.
+that $r_{i}^{\star}\geq j+1$) is at most $1/p(k)\leq 1/2$.
 Thus, it follows that
@@
 ∎
-It follows from Lemmas 19.11, 19.10 and 19.9 that the total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot T$ with probability at least $1-1/\mathrm{poly}(n)$.
+It follows from Lemmas 20.11, 20.10 and 20.9 that the total update time of our algorithm is $\tilde{O}(k^{\epsilon})\cdot T$ with probability at least $1-1/\mathrm{poly}(n)$.
 Thus, the amortized update time of our algorithm is $\tilde{O}(k^{\epsilon})$ with high probability.
@@
 $=\textsc{cost}(B_{i-1},x_{i-2})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{i-1},x_{0})$
-[…]
+$\ \ \!\vdots$
 $\geq\textsc{cost}(B_{1},x_{0})+\frac{1}{4}\cdot\textsc{cost}(U\setminus B_{1},x_{0})\geq\frac{1}{4}\cdot\textsc{cost}(U,x_{0}).$
@@
 $\leq$ $\mathrm{dist}(w_{i},w_{j})+4\cdot\lambda^{3t[w_{i}^{\prime}]-1}\leq(1+4/\lambda^{5})\cdot\mathrm{dist}(w_{i},w_{j}).$
-The first inequality holds by Equation 70, the third one holds by Lemma 15.1 (note that $w_{i}^{\prime}=$ MakeRobust$(w_{i})$) and the last one holds by Equation 71 (by exchanging $i$ and $j$ in this equation since $i<j$).
-Since $\lambda$ is a large number, we conclude that for each $i<j$, the following holds
+[…]
 $\mathrm{dist}(w_{i},w_{j})>\lambda^{4}/(1+4/\lambda^{5})\cdot\lambda^{3t[w_{j}^{\prime}]}>2\cdot\lambda^{3t[w_{j}^{\prime}]}.$
@@
 Equation 42 holds for $i=t^{*}$. So,
-[…]
+$\textsc{cost}(B_{t^{*}},x_{t^{*}})/w(B_{t^{*}})\leq\lambda^{6t^{*}-2}\text{ and }\textsc{cost}(B_{t^{*}},x_{t^{*}-1})\leq(p_{\epsilon})^{3}\cdot\mathrm{OPT}_{1}(B_{t^{*}}).$
 In this case, for every $y\in C_{v}-B_{t^{*}}$,
@@
 Consider the standard LP relaxation for the weighted $k$-means problem.
-We consider the set of potential centers to open as $S$, and the goal is to open at most $k-(m-r)/4$ many centers, i.e., find $\tilde{S}\subseteq S$.
+We consider the set of potential centers to open as $S$, and the goal is to open at most k(mr)/4k-(m-r)/4 many centers, i.e., find $\tilde{S}\subseteq S$.
 $\min$ $\sum_{p\in X}\sum_{u\in S}w(p)\cdot\mathrm{dist}(u,p)^{2}\cdot x_{up}$
@@
 Each center $u\in S$ that forms a well-separated pair with a center $v\in V$ has $|\pi^{-1}_{S}(u)|\geq 1$ since $u$ must be the closest center to $v$ in $S$.
-Since the number of well-separated pairs is $k-m$, we have
+Since the number of well-separated pairs is kmk-m, we have
 $k+r\geq|V|=\sum_{u\in S}|\pi^{-1}(u)|\geq\sum_{u\in S_{I}}|\pi^{-1}(u)|\geq k-m+2\cdot\left(|S_{I}|-(k-m)\right).$
@@
 The third inequality follows by Equation 81.
-Finally, since the integrality gap of the LP relaxation is known to be at most $9$ [ANFSW20].
-[^23] Note that although it is not explicitly mentioned, the result of [ANFSW20] works for the weighted case.
+Finally, since the integrality gap of the LP relaxation is known to be at most $9$ [integrality-gap-k-means].
+[^23] Note that although it is not explicitly mentioned, the result of [integrality-gap-k-means] works for the weighted case.
 One way to show this, which only uses the unweighted case as a black box is as follows.
@@
 Since the space $X$ is fixed here, we denote $\mathrm{OPT}_{k}(X)$ by $\mathrm{OPT}_{k}$.
-It is known that the integrality gap of this relaxation is at most $9$ [ANFSW20].
+It is known that the integrality gap of this relaxation is at most $9$ [integrality-gap-k-means].
 So, for every $k$ we have
@@
 The first inequality holds since $w^{\prime}$ is present in $W^{\text{new}}$, the second inequality is triangle inequality, the third inequality holds by Equation 85 ,and the last inequality holds since $\lambda=(p_{\epsilon})^{2}$ is a large number.
-According to 6 in Algorithm 7, we must have $t[w_{0}]<t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}$, as otherwise we should not made a call to MakeRobust on $w_{0}$.
+According to 12 in Algorithm 7, we must have $t[w_{0}]<t$, where $t$ is the smallest integer satisfying $\lambda^{3t}\geq\widehat{\mathrm{dist}}(w_{0},W^{\text{new}}-w_{0})/\lambda^{10}$, as otherwise we should not made a call to MakeRobust on $w_{0}$.
 Hence,
@@
 Hence, the distance of any center $v$ in $S[i]$ to $x$ is strictly greater than $\lambda^{3i+2}/p_{\epsilon}=\lambda^{3i+1.5}$ (since $\lambda=(p_{\epsilon})^{2}$ by Part III), which concludes there is no center $v\in S[i]$ contaminated by $x$.
-Appendix A Barriers Towards Adapting the Algorithm of [BCG+24]
-In [BCG+24], the authors give a dynamic algorithm for $k$-means on general metric spaces, with $O(1/\epsilon)$-approximation ratio and $\tilde{O}(k^{\epsilon})$ update time.
+Appendix A Barriers Towards Adapting the Algorithm of [BhattacharyaCGLP24]
+In [BhattacharyaCGLP24], the authors give a dynamic algorithm for $k$-means on general metric spaces, with $O(1/\epsilon)$-approximation ratio and $\tilde{O}(k^{\epsilon})$ update time.
 With the current state-of-the-art tools for designing geometric algorithms in Euclidean spaces, we cannot avoid a tradeoff between a $\mathrm{poly}(1/\epsilon)$ term in the approximation and a $k^{\epsilon}$ term in the update time for this problem.
@@
 Bypassing them, even in the static setting, is a major open problem.
-Thus, it might appear that building on top of the work of [BCG+24] might be a simpler and more natural approach than building on [BCF25a], since we will eventually incur such a tradeoff in our final algorithm.
-Unfortunately, the algorithm of [BCG+24] is not as well suited to implementation in high-dimensional Euclidean spaces, and it is not clear how to use this algorithm without avoiding an exponential blowup in the approximation ratio, leading to a $\exp(1/\epsilon)$-approximation.
-The Algorithm of [BCG+24].
-The high level approach taken by [BCG+24] is to maintain a hierarchy of nested solutions $X\supseteq S_{0}\supseteq\dots\supseteq S_{\ell}$, where $S_{\ell}$ has size $k$ and is the output of the algorithm.
+Thus, it might appear that building on top of the work of [BhattacharyaCGLP24] might be a simpler and more natural approach than building on [BCF24], since we will eventually incur such a tradeoff in our final algorithm.
+Unfortunately, the algorithm of [BhattacharyaCGLP24] is not as well suited to implementation in high-dimensional Euclidean spaces, and it is not clear how to use this algorithm without avoiding an exponential blowup in the approximation ratio, leading to a $\exp(1/\epsilon)$-approximation.
+The Algorithm of [BhattacharyaCGLP24].
+The high level approach taken by [BhattacharyaCGLP24] is to maintain a hierarchy of nested solutions $X\supseteq S_{0}\supseteq\dots\supseteq S_{\ell}$, where $S_{\ell}$ has size $k$ and is the output of the algorithm.
 Using a new randomized variant of local search, they can maintain this hierarchy by periodically reconstructing these solutions and maintaining them lazily between reconstructions, leading to an update time of $\tilde{O}(k^{1+1/\ell})$ and a recourse of $\tilde{O}(k^{1/\ell})$.
@@
 Applying this argument $O(\ell)$ times, they show that $\textsc{cost}(X,S_{\ell})\leq O(\ell)\cdot\mathrm{OPT}_{k}(X)$.
-The Barrier to Using [BCG+24].
-If we want to implement [BCG+24] efficiently in high-dimensional Euclidean spaces, the most natural approach is to modify some components within the algorithm to use approximate nearest neighbor data structures, which can be implemented efficiently in this setting.
+The Barrier to Using [BhattacharyaCGLP24].
+If we want to implement [BhattacharyaCGLP24] efficiently in high-dimensional Euclidean spaces, the most natural approach is to modify some components within the algorithm to use approximate nearest neighbor data structures, which can be implemented efficiently in this setting.
 However, the hierarchical nature of this algorithm requires us to have very strong guarantees on the quality of the solutions in order for the analysis to go through.